User’s Guide

to

PARI-GP

C. Batut, K. Belabas, D. Bernardi, H. Cohen, M. Olivier

Laboratoire A2X, U.M.R. 9936 du C.N.R.S. Université Bordeaux I, 351 Cours de la Libération

33405 TALENCE Cedex, FRANCE

        e-mail:  pari@math.u-bordeaux.fr

Home Page:

http://hasse.mathematik.tu-muenchen.de/ntsw/pari/

Primary ftp site:

   ftp://megrez.math.u-bordeaux.fr/pub/pari/

last updated 20 June 1999 for version 2.0.17

Table of Contents

Chapter 1: Overview of the PARI system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.2 The PARI types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.3 Operations and functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

Chapter 2: Specific Use of the GP Calculator . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.1 Defaults and output formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 2.2 Simple metacommands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.3 Input formats for the PARI types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.4 GP operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.5 The general GP input line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.6 The GP/PARI programming language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 2.7 Interfacing GP with other languages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 2.8 The preferences file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 2.9 Using GP under GNU Emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 2.10 Using GP with readline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

Chapter 3: Functions and Operations Available in PARI and GP . . . . . . . . . . . . . . . . 45

3.1 Standard monadic or dyadic operators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 3.2 Conversions and similar elementary functions or commands . . . . . . . . . . . . . . . . . . . . . . . 48 3.3 Transcendental functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 3.4 Arithmetic functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 3.5 Functions related to elliptic curves . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 3.6 Functions related to general number fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 3.7 Polynomials and power series . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 3.8 Vectors, matrices, linear algebra and sets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 3.9 Sums, products, integrals and similar functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 3.10 Plotting functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 3.11 Programming under GP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137

Chapter 4: Programming PARI in Library Mode . . . . . . . . . . . . . . . . . . . . . . . . . 145

4.1 Introduction: initializations, universal objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 4.2 Important technical notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 4.3 Creation of PARI objects, assignments, conversions . . . . . . . . . . . . . . . . . . . . . . . . . . 149 4.4 Garbage collection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 4.5 Implementation of the PARI types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 4.6 PARI variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 4.7 Input and output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 4.8 A complete program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 4.9 Adding functions to PARI . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

Chapter 5: Technical Reference Guide for Low-Level Functions . . . . . . . . . . . . . . . . . 179

5.1 Level 0 kernel (operations on unsigned longs) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 179 5.2 Level 1 kernel (operations on longs, integers and reals) . . . . . . . . . . . . . . . . . . . . . . . . 180 5.3 Level 2 kernel (operations on general PARI objects) . . . . . . . . . . . . . . . . . . . . . . . . . . 185

Appendix A: Installation Guide for the UNIX Versions . . . . . . . . . . . . . . . . . . . . . 191

1. Required tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 2. Compiling the library and the GP calculator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 192 3. Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 4. Getting Started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194

Appendix B: A Sample program and Makefile . . . . . . . . . . . . . . . . . . . . . . . . . . 197 Appendix C: Summary of Available Constants . . . . . . . . . . . . . . . . . . . . . . . . . . 199 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 4

Chapter 1: Overview of the PARI system

1.1 Introduction.

The PARI system is a package which is capable of doing formal computations on recursive types at high speed; it is primarily aimed at number theorists, but can be used by anybody whose primary need is speed.

Although quite an amount of symbolic manipulation is possible in PARI, this system does very badly compared to much more sophisticated systems like Axiom, Macsyma, Maple, Mathematica or Reduce on such manipulations (e.g. multivariate polynomials, formal integration, etc…). On the other hand, the three main advantages of the system are its speed (which can be between 5 and 100 times better on many computations), the possibility of using directly data types which are familiar to mathematicians, and its extensive algebraic number theory module which has no equivalent in the above-mentioned systems.

It is possible to use PARI in two different ways:

1) as a library, which can be called from an upper-level language application (for instance written in C, C++, Pascal or Fortran);

2) as a sophisticated programmable calculator, named GP, which contains most of the control instructions of a standard language like C.

The use of GP is explained in chapters 2 and 3, and the programming in library mode is explained in chapters 3, 4 and 5. In the present Chapter 1, we give an overview of the system.

Important note: A tutorial for GP is provided in the standard distribution (tutorial.dvi) and you should read this first (at least the beginning of it, you can skip the specialized topics you’re not interested in). You can then start over and read the more boring stuff which lies ahead. But you should do that eventually, at the very least the various Chapter headings. You can have a quick idea of what is available by looking at the GP reference card (refcard.dvi or refcard.ps). In case of need, you can then refer to the complete function description in Chapter 3.

How to get the latest version?

This package can be obtained by anonymous ftp from quite a number of sites (ask archie or your favourite Web search engine for the site nearest to you). But, if you want the very latest version (including development versions), you should use the anonymous ftp address

    ftp://megrez.math.u-bordeaux.fr/pub/pari

where you will find all the different ports and possibly some binaries. A lot of version information, mailing list archives, and various tips can be found on PARI’s (fledgling) home page:

    http://hasse.mathematik.tu-muenchen.de/ntsw/pari/

Implementation notes: (You can skip this section and switch to Section 1.2 if you’re not interested in hardware technicalities. You won’t miss anything that would be mentioned here.)

Version 2.0.17 of the PARI package contains essentially three versions. The first one is a specific implementation for 680x0 based computers which contains a kernel (for the elementary arithmetic operations on multiprecise integers and real numbers, and binary/decimal conversion routines) entirely written in MC68020 assembly language (around 6000 lines), the rest being at present entirely written in ANSI C with a C++-compatible syntax. The system runs on SUN-3/xx, Sony News, NeXT cubes and on 680x0 based Macs with x≥2. It should be very easy to port on any other 680x0 based machine like for instance the Apollo Domain workstations.

Note that the assembly language source code uses the SUN syntax, which for some strange reason differs from the Motorola standard used by most other 680x0 machines in the world. In the Mac distribution, we have included a program which automatically converts from the SUN syntax into the standard one, at least for the needed PARI assembly file. On the Unix distribution, we have included other versions of the assembly file, using different syntaxes. This version is not really maintained anymore since we lack the hardware to update/test it.

The second version is a version where most of the kernel routines are written in C, but the time-critical parts are written in a few hundred lines of assembler at most. At present there exist three versions for the Sparc architecture: one for Sparc version 7 (e.g. Sparcstation 1, 1+, IPC, IPX or 2), one for Sparc version 8 with supersparc processors (e.g. Sparcstation 10 and 20) and one for Sparc version 8 with microsparc I or II processors (e.g. Sparcclassic or Sparcstation 4 and 5). No specific version is written for the Ultrasparc since it can use the microsparc II version. In addition, versions exist for the HP-PA architecture, for the PowerPC architecture (only for the 601), for the Intel family starting at the 386 (under Linux, OS/2, MSDOS, or Windows), and finally for the DEC Alpha 64-bit processors.

Finally, a third version is written entirely in C, and should be portable without much trouble to any 32 or 64-bit computer having no real memory constraints. It is about 2 times slower than versions with a small assembly kernel. This version has been tested for example on MIPS based DECstations 3100 and 5000 and SGI computers.

In addition to Unix workstations and Macs, PARI has been ported to a considerable number of smaller and larger machines, for example the VAX, 68000-based machines like the Atari, Mac Classic or Amiga 500, 68020 machines such as the Amiga 2500 or 3000, and even to MS-DOS 386 or better machines, using the EMX port of the GNU C compiler and DOS-extender.

1.2 The PARI types.

The crucial word in PARI is recursiveness: most of the types it knows about are recursive. For example, the basic type Complex exists (actually called t COMPLEX). However, the components (i.e. the real and imaginary part) of such a “complex number” can be of any type. The only sensible ones are integers (we are then in Z[i]), rational numbers (Q[i]), real numbers (R[i] = C), or even elements of Z/nZ ((Z/nZ)[i] when this makes sense), or p-adic numbers when p ≡ 3 mod 4 (Q_{p}[i]).

This feature must of course not be used too rashly: for example you are in principle allowed to create objects which are “complex numbers of complex numbers”, but don’t expect PARI to make sensible use of such objects: you will mainly get nonsense.

On the other hand, one thing which is allowed is to have components of different, but compatible, types. For example, taking again complex numbers, the real part could be of type integer, and the imaginary part of type rational.

By compatible, we mean types which can be freely mixed in operations like + or ×. For example if the real part is of type real, the imaginary part cannot be of type integermod (integers modulo a given number n).

Let us now describe the types. As explained above, they are built recursively from basic types which are as follows. We use the letter T to designate any type; the symbolic names correspond to the internal representations of the types.

type t INT :  Z  Integers (with arbitrary precision)
type t REAL:  R  Real numbers (with arbitrary precision)
type t INTMOD: Z/nZ  Integermods (integers modulo n)
type t FRAC:  Q  Rational numbers (in irreducible form)
type t FRACN:  Q  Rational numbers (not necessarily in irreducible form)
type t COMPLEX: T [i]  Complex numbers
type t PADIC:  Q_{p}  p-adic numbers
type t QUAD:  Q[w]  Quadratic Numbers (where [Z[w] : Z] = 2)
type t POLMOD: T [X]/P (X)T [X] Polmods (polynomials modulo P)
type t POL:  T [X]  Polynomials
type t SER:  T ((X))  Power series (finite Laurent series)
type t RFRAC:  T (X)  Rational functions (in irreducible form)
type t RFRACN: T (X)  Rational functions (not necessarily in irreducible form)
type t VEC:  T^{n}  Row (i.e. horizontal) vectors
type t COL:  T^{n}  Column (i.e. vertical) vectors
type t MAT:  M_{m,n}(T)  Matrices
type t LIST:  T^{n}  Lists
type t STR:  Character strings

and where the types T in recursive types can be different in each component.

In addition, there exist types t QFR and t QFI for binary quadratic forms of respectively positive and negative discriminants, which can be used in specific operations, but which may disappear in future versions.

Every PARI object (called GEN in the sequel) belongs to one of these basic types. Let us have a closer look.

1.2.1 Integers and reals: they are of arbitrary and varying length (each number carrying in its internal representation its own length or precision) with the following mild restrictions (given for 32-bit machines, the restrictions for 64-bit machines being so weak as to be considered inexistent): integers must be in absolute value less than 2^{268435456} (i.e. roughly 80807124 digits). The precision of real numbers is also at most 80807124 significant decimal digits, and the binary exponent must be in absolute value less than 2^{23} = 8388608.

Note that PARI has been optimized so that it works as fast as possible on numbers with at most a few thousand decimal digits. In particular, not too much effort has been put into fancy multiplication techniques (only the Karatsuba algorithm is implemented). Hence, although it is possible to use PARI to do computations with 80000000 decimal digits, much better programs can be written for such huge numbers.

Integers and real numbers are completely non-recursive types and are sometimes called the leaves.

1.2.2 Integermods, rational numbers (irreducible or not), p-adic numbers, polmods, and rational functions: these are recursive, but in a restricted way.

For integermods or polmods, there are two components: the modulus, which must be of type integer (resp. polynomial), and the representative number (resp. polynomial).

For rational numbers or rational functions, there are also only two components: the numerator and the denominator, which must both be of type integer (resp. polynomial).

Finally, p-adic numbers have three components: the prime p, the “modulus” p^{k}, and an approximation to the p-adic number. Here Z_{p} is considered as \underleftarrow{lim}Z/p^{k}Z, and Q_{p} as its field of fractions. Like real numbers, the codewords contain an exponent (giving essentially the p-adic valuation of the number) and also the information on the precision of the number (which is in fact redundant with p^{k}, but is included for the sake of efficiency).

1.2.3 Complex numbers and quadratic numbers: quadratic numbers are numbers of the form a + bw, where w is such that [Z[w] : Z] = 2, and more precisely w = \sqrt{d}/2 when d ≡ 0 mod 4, and w = (1 + \sqrt{d})/2 when d ≡ 1 mod 4, where d is the discriminant of a quadratic order. Complex numbers correspond to the very important special case w = \sqrt{−1}.

Complex and quadratic numbers are partially recursive: the two components a and b can be of type integer, real, rational, integermod or p-adic, and can be mixed, subject to the limitations mentioned above. For example, a + bi with a and b p-adic is in Q_{p}[i], but this is equal to Q_{p} when p ≡ 1 mod 4, hence we must exclude these p when one explicitly uses a complex p-adic type.

1.2.4 Polynomials, power series, vectors, matrices and lists: they are completely recursive: their components can be of any type, and types can be mixed (however beware when doing operations). Note in particular that a polynomial in two variables is simply a polynomial with polynomial coefficients.

Note that in the present version 2.0.17 of PARI, there is a bug in the handling of power series of power series (i.e. power series in several variables). However power series of polynomials (which are power series in several variables of a special type) are OK. The reason for this bug is known, but it is difficult to correct because the mathematical problem itself contains some amount of imprecision.

1.2.5 Strings: These contain objects just as they would be printed by the GP calculator.

1.2.6 Notes:

1.2.6.1 Exact and imprecise objects: we have already said that integers and reals are called the leaves because they are ultimately at the end of every branch of a tree representing a PARI object. Another important notion is that of an exact object: by definition, numbers of basic type real, p-adic or power series are imprecise, and we will say that a PARI object having one of these imprecise types anywhere in its tree is not exact. All other PARI objects will be called exact. This is a very important notion since no numerical analysis is involved when dealing with exact objects.

1.2.6.2 Scalar types: the first nine basic types, from t INT to t POLMOD, will be called scalar types because they essentially occur as coefficients of other more complicated objects. Note that type t POLMOD is used to define algebraic extensions of a base ring, and as such is a scalar type.

1.2.6.3 What is zero? This is a crucial question in all computer systems. The answer we give in PARI is the following. For exact types, all zeros are equivalent and are exact, and thus are usually represented as an integer zero. The problem becomes non-trivial for imprecise types. For p-adics the answer is as follows: every p-adic number (including 0) has an exponent e and a “mantissa” (a purist would say a “significand”) u which is a p-adic unit, except when the number is zero (in which case u is zero), the significand having a certain “precision” k (i.e. being defined modulo p^{k}). Then this p-adic zero is understood to be equal to O(p^{e}), i.e. there are infinitely many distinct p-adic zeros. The number k is thus irrelevant.

For power series the situation is similar, with p replaced by X, i.e. a power series zero will be O(X^{e}), the number k (here the length of the power series) being also irrelevant.

For real numbers, the precision k is also irrelevant, and a real zero will in fact be O(2^{e}) where e is now usually a negative binary exponent. This of course will be printed as usual for a real number (0.0000 ⋯ in f format or 0.Exx in e format) and not with a O() symbol as with p-adics or power series.

1.3 Operations and functions.

1.3.1 The PARI philosophy. The basic philosophy which governs PARI is that operations and functions should, firstly, give as exact a result as possible, and secondly, be permitted if they make any kind of sense.

More specifically, if you do an operation (not a transcendental one) between exact objects, you will get an exact object. For example, dividing 1 by 3 does not give 0.33333 ⋯ as you might expect, but simply the rational number (1/3). If you really want the result in type real, evaluate 1./3 or add 0. to (1/3).

The result of operations between imprecise objects will be as precise as possible. Consider for example one of the most difficult cases, that is the addition of two real numbers x and y. The accuracy of the result is a priori unpredictable; it depends on the precisions of x and y, on their sizes (i.e. their exponents), and also on the size of x + y. PARI works out automatically the right precision for the result, even when it is working in calculator mode GP where there is a default precision.

In particular, this means that if an operation involves objects of different accuracies, some digits will be disregarded by PARI. It is a common source of errors to forget, for instance, that a real number is given as r + 2^{e}ε where r is a rational approximation, e a binary exponent and ε is a nondescript real number less than 1 in absolute value*. Hence, any number less than 2^{e} may be treated as an exact zero:

    ?  0.E-28 + 1.E-100
    %1 = 0.E-28

As an exercise, if a = 2^-100, why do a + 0. and a * 1. differ ?

The second part of the PARI philosophy is that PARI operations are in general quite permissive. For instance taking the exponential of a vector should not make sense. However, it frequently happens that a computation comes out with a result which is a vector with many components, and one wants to get the exponential of each one. This could easily be done either under GP or in

* this is actually not quite true: internally, the format is 2^{b}(a + ε), where a and b are integers library mode, but in fact PARI assumes that this is exactly what you want to do when you take the exponential of a vector, so no work is necessary. Most transcendental functions work in the same way (see Chapter 3 for details).

An ambiguity would arise with square matrices. PARI always considers that you want to do componentwise function evaluation, hence to get for example the exponential of a square matrix you would need to use a function with a different name, matexp for instance. In the present version 2.0.17, this is not yet implemented. See however the program in Appendix C, which is a first attempt for this particular function.

The available operations and functions in PARI are described in detail in Chapter 3. Here is a brief summary:

1.3.2 Standard operations.

Of course, the four standard operators +, -, *, / exist. It should once more be emphasized that division is, as far as possible, an exact operation: 4 divided by 3 gives (4/3). In addition to this, operations on integers or polynomials, like \ (Euclidean division), % (Euclidean remainder) exist (and for integers, \/ computes the quotient such that the remainder has smallest possible absolute value). There is also the exponentiation operator ^, when the exponent is of type integer. Otherwise, it is considered as a transcendental function. Finally, the logical operators ! (not prefix operator), && (and operator), || (or operator) exist, giving as results 1 (true) or 0 (false). Note that & and | are also accepted as synonyms respectively for && and ||. However, there is no bitwise

and or or.

1.3.3 Conversions and similar functions.

Many conversion functions are available to convert between different types. For example floor, ceiling, rounding, truncation, etc. . .. Other simple functions are included like real and imaginary part, conjugation, norm, absolute value, changing precision or creating an integermod or a polmod.

1.3.4 Transcendental functions.

They usually operate on any object in C, and some also on p-adics. The list is everexpanding and of course contains all the elementary functions, plus already a number of others. Recall that by extension, PARI usually allows a transcendental function to operate componentwise on vectors or matrices.

1.3.5 Arithmetic functions.

Apart from a few like the factorial function or the Fibonacci numbers, these are functions which explicitly use the prime factor decomposition of integers. The standard functions are included. In the present version 2.0.17, a primitive, but useful version of Lenstra’s Elliptic Curve Method (ECM) has been implemented.

There is now a very large package which enables the number theorist to work with ease in algebraic number fields. All the usual operations on elements, ideals, prime ideals, etc… are available.

More sophisticated functions are also implemented, like solving Thue equations, finding integral bases and discriminants of number fields, computing class groups and fundamental units, computing in relative number field extensions (including explicit class field theory), and also many functions dealing with elliptic curves over Q or over local fields.

1.3.6 Other functions.

Quite a number of other functions dealing with polynomials (e.g. finding complex or p-adic roots, factoring, etc), power series (e.g. substitution, reversion), linear algebra (e.g. determinant, characteristic polynomial, linear systems), and different kinds of recursions are also included. In addition, standard numerical analysis routines like Romberg integration (open or closed, on a finite or infinite interval), real root finding (when the root is bracketed), polynomial interpolation, infinite series evaluation, and plotting are included. See the last sections of Chapter 3 for details.

12

Specific Use of the GP Calculator

Originally, GP was designed as a debugging tool for the PARI system library, and hence not much thought had been given to making it user-friendly. The situation has now changed somewhat, and GP is very useful as a stand-alone tool. The operations and functions available in PARI and GP will be described in the next chapter. In the present one, we describe the specific use of the GP programmable calculator.

For starting the calculator, the general commandline syntax is:

    gp [-s stacksize] [-p primelimit]

where items within brackets are optional*. These correspond to some internal parameters of GP, or defaults. See Section 2.1 below for a list and explanation of all defaults, there are many more than just those two. These defaults can be changed by adding parameters to the input line as above, or interactively during a GP session or in a preferences file (also known as gprc).

UNIX: Some new features were developed on UNIX platforms, and depend heavily on the operating system in use. It is possible that some of these will be ported to other operating systems (BeOS, MacOS, DOS, OS/2, Windows, etc.) in future versions (most of them should be easy tasks for anybody acquainted with those). As for now, most of them were not. So, whenever a specific feature of the UNIX version is discussed in a paragraph, a UNIX sign sticks out in the left margin, like here. Just skip these if you’re stranded on a different operating system: the core GP functions (i.e. at least everything which is even faintly mathematical in nature) will still be available to you. It may also be possible (and then definitely advisable) to install Linux or FreeBSD on your machine.

Note (added in version 2.0.12): All the UNIX goodies are now available for DOS, OS/2 and Windows 3.1, thanks to the EMX runtime package (install excluded under DOS, since DLLs are not supported by the OS). They’ve been reported to be available under Windows 95/98 and NT using the Cygwin package (untested by us, but supposedly supported by Configure).

EMACS: If you have GNU Emacs, you can work in a special Emacs shell (see Section 2.9), which is started by typing M-x gp (where as usual M is the Meta key) if you accept the default stack, prime and buffer sizes, or C-u M-x gp which will ask you for the name of the gp executable, the stack size, the prime limit and the buffer size. Specific features of this Emacs shell will be indicated by an EMACS sign.

If a preferences file (or gprc, to be discussed in Section 2.8) can be found, GP will then read it and execute the commands it contains. This provides an easy way to customize GP without having to delve into the code to hardwire it to your likings.

A copyright message then appears which includes the version number. Please note this number, so as to be sure to have the most recent version if you wish to have updates of PARI. The present manual is written for version 2.0.17, and has undergone major changes since the 1.39.xx versions.

* On the Macintosh, even after clicking on the gp icon, once in the MPW Shell, you still need to type explicitly a command of the above form.

After the copyright, the computer works for a few seconds (it is in fact computing and storing a table of primes), writes the top-level help information, some initial defaults, and then waits after printing its prompt (initially: ?).

Note that at any point the user can type Ctrl-C (that is press simultaneously the Control and C keys): the current computation will be interrupted and control given back to the user at the GP prompt.

The top-level help information tells you that (as in many systems) to get help, you should type a ?. When you do this and hit return, a menu appears, describing the eleven main categories of available functions and what to do to get more detailed help. If you now type ?n with 1 ≤ n ≤ 11, you will get the list of commands corresponding to category n and simultaneously to Section 3.n of this manual.

If you type ?functionname where functionname is the name of a PARI function, you will get a short explanation of this function.

UNIX: If extended help (see Section 2.2.1) is available on your system, you can double or triple the ? sign to get much more: respectively the complete description of the function (e.g. ?? sqrt), or a list of GP functions relevant to your query (e.g. ??? "elliptic curve" or ??? "quadratic

field").

If GP was compiled with the right options (see Appendix A), a line editor will be available to correct the command line, get automatic completions, and so on. See Section 2.10.1 for a short summary of available commands. This might not be available for all architectures.

Whether extended on-line help and line editing are available or not is indicated in the GP banner, between the version number and the copyright message.

If you type ?\ you will get a short description of the metacommands (keyboard shortcuts).

Finally, typing ?. will return the list of available (pre-defined) member functions. These are functions attached to specific kind of objects, used to retrieve easily some information from complicated structures (you can define your own but they won’t be shown here). We will soon describe these commands in more detail.

As a general rule, under GP, commands starting with \ or with some other symbols like ? or #, are not computing commands, but are metacommands which allow the user to exchange information with GP. The available metacommands can be divided into default setting commands (explained below) and simple commands (or keyboard shortcuts, to be dealt with in Section 2.2).

2.1 Defaults and output formats.

There are many internal variables in GP, defining how the system will behave in certain situations, unless a specific override has been given. Most of them are a matter of basic customization (colors, prompt) and will be set once and for all in your preferences file (see Section 2.8), but some of them are useful interactively (set timer on, increase precision, etc.).

The function used to manipulate these values is called default, which is described in Section 3.11.2.4. The basic syntax is

    default(def , value),

which sets the default def to value. In interactive use, most of these can be abbreviated using historic GP metacommands (mostly, starting with \), which we shall describe in the next section.

Here we will only describe the available defaults and how they are used. Just be aware that typing default by itself will list all of them, as well as their current values (see \d). Just after the default name, we give between parentheses the initial value when GP starts (assuming you did not tamper with it using command-line switches or a gprc).

(somewhat technical) Note: As we will see in Section 2.6.5, the second argument to default will be subject to string context expansion, which means you can use run-time values. In other words, something like a = 3; default(logfile, "some filename" a ".log") will work.

For the user’s convenience, some defaults will be expanded further when the values are used (after the above expansion has been performed):

• time expansion: the string is sent through the library function strftime. This means that %char combinations have a special meaning, usually related to the time and date. For instance, %H = hour (24-hour clock) and %M = minute [00,59] (on a Unix system, you can try man strftime at your shell prompt to get a complete list). This is applied to prompt, psfile, and logfile. For instance,

    default(prompt,"(%R) ? ")

will prepend the time of day, in the form (hh:mm) to GP’s usual prompt.

UNIX: • environment expansion: When the string contains a sequence of the form $SOMEVAR (e.g. $HOME) the environment is searched and if SOMEVAR is defined, the sequence is replaced by the corresponding value. Also the ~ symbol has the same meaning as in the C and bash shells — ~ by itself stands for your home directory, and ~user is expanded to user’s home directory. This is applied to all filenames.

2.1.1 buffersize (default 30000): GP input is buffered, which means only so many bytes of data can be read at a time before a command is executed. This used to be a very important variable, to allow for very large input files to be read into GP, for example large matrices, without it complaining about “unused characters”. Currently, buffersize is automatically adjusted to the size of the data that are to be read. It will never go down by itself though. Thus this option may come in handy to decrease the buffer size after some unusually large read, when you don’t need to keep gigantic buffers around anymore.

UNIX: 2.1.2 colors (default ""): this default is only usable if GP is running within certain color-capable terminals. For instance rxvt, color xterm and modern versions of xterm under X Windows, or standard Linux/DOS text consoles. It causes GP to use a small palette of colors for its output. With xterms, the colormap used corresponds to the resources Xterm*colorn where n ranges from 0 to 15 (see the file misc/color.dft for an example). Legal values for this default are strings "a_{1},…,a_{k}" where k ≤ 7 and each a_{i} is either

• the keyword no (use the default color, usually black)

• an integer between 0 and 15 corresponding to the aforementioned colormap

• a triple [c_{0}, c_{1}, c_{2}] where c_{0} stands for foreground color, c_{1} for background color, and c_{2} for attributes (0 is default, 1 is bold, 4 is underline).

The output objects thus affected are respectively error messages, history numbers, prompt, input line, output, help messages, timer (that’s seven of them). If k < 7, the remaining a_{i} are assumed to be no. For instance

    default(colors,"9, 5, no, no, 4")

typesets error messages in color 9, history numbers in color 5, output in color 4, and does not affect the rest.

EMACS: In the present version, this default is incompatible with Emacs. Changing it will just fail silently (the alternative would be to display escape sequences as is, since Emacs will refuse to interpret them). On the other hand, you can customize highlighting in your .emacs so as to mimic exactly this behaviour. See emacs/pariemacs.txt.

If you use an old readline library (version number less than 2.0), you should do as in the example above and leave a_{3} and a_{4} (prompt and input line) strictly alone. Since old versions of readline did not handle escape characters correctly (or more accurately, treated them in the only sensible way since they did not care to check all your terminal capabilities: it just ignored them), changing them would result in many annoying display bugs.

The hacker’s way to check if this is the case would be to look in the readline.h include file (wherever your readline include files are) for the string RL PROMPT START IGNORE. If it’s there, you are safe.

A more sensible way is to make some experiments, and get a more recent readline if yours doesn’t work the way you’d like it to. See the file misc/gprc.dft for some examples.

2.1.3 compatible (default 0): The GP function names and syntax have changed tremendously between versions 1.xx and 2.00. To help you cope with this we provide some kind of backward compatibility, depending on the value of this default:

compatible = 0: no backward compatibility. In this mode, a very handy function, to be described in Section 3.11.2.26, is whatnow, which tells you what has become of your favourite functions, which GP suddenly can’t seem to remember.

compatible = 1: warn when using obsolete functions, but otherwise accept them. The output uses the new conventions though, and there may be subtle incompatibilities between the behaviour of former and current functions, even when they share the same name (the current function is used in such cases, of course!). We thought of this one as a transitory help for GP oldtimers. Thus, to encourage switching to compatible=0, it is not possible to disable the warning.

compatible = 2: use only the old function naming scheme (as used up to version 1.39.15), but taking case into account. Thus I ( = \sqrt{−1}) is not the same as i (user variable, unbound by default), and you won’t get an error message using i as a loop index as used to be the case.

compatible = 3: try to mimic exactly the former behaviour. This is not always possible when functions have changed in a fundamental way. But these differences are usually for the better (they were meant to, anyway), and will probably not be discovered by the casual user.

One adverse side effect is that any user functions and aliases that have been defined before changing compatible will get erased if this change modifies the function list, i.e. if you move between groups {0, 1} and {2, 3} (variables are unaffected). We of course strongly encourage you to try and get used to the setting compatible=0.

2.1.4 debug (default 0): debugging level. If it is non-zero, some extra messages may be printed (some of it in French), according to what is going on (see \g).

2.1.5 debugfiles (default 0): file usage debugging level. If it is non-zero, GP will print information on file descriptors in use, from PARI’s point of view (see \gf).

2.1.6 debugmem (default 0): memory debugging level. If it is non-zero, GP will regularly print information on memory usage. If it’s greater than 2, it will indicate any important garbage collecting and the function it is taking place in (see \gm).

Important Note: As it noticeably slows down the performance (and triggers bugs in a popular compiler), the first functionality (memory usage) is disabled if you’re not running a version compiled for debugging (see Appendix A).

2.1.7 echo (default 0): this is a toggle, which can be either 1 (on) or 0 (off). When echo mode is on, each command is reprinted before being executed. This can be useful when reading a file with the \r or read commands. For example, it is turned on at the beginning of the test files used to check whether GP has been built correctly (see \e).

2.1.8 format (default "g0.28" and "g0.38" on 32-bit and 64-bit machines, respectively): of the form xm.n, where x is a letter in {e, f, g}, and n, m are integers. If x is f, real numbers will be printed in fixed floating point format with no explicit exponent (e.g. 0.000033); if the letter is e, they will be printed in scientific format, always with an explicit exponent (e.g. 3.3e-5). If the letter is g, real numbers will be printed in f format, except when their absolute value is less than 2^{−32}, in which case they are printed in e format.

The number n is the number of significant digits printed for real numbers, except if n < 0 where all the significant digits will be printed (initial default 28, or 38 for 64-bit machines), and the number m is the number of characters to be used for printing integers, but is ignored if equal to 0 (which is the default). This is a feeble attempt at formatting.

UNIX: 2.1.9 help (default: the location of the gphelp script): the name of the external help program which will be used from within GP when extended help is invoked, usually through a ?? or ??? request (see Section 2.2.1), or M-H under readline (see Section 2.10.1).

2.1.10 histsize (default 5000): GP keeps a history of the last histsize results computed so far, which you can recover using the % notation (see Section 2.2.4). When this number is exceeded, the oldest values are erased. Tampering with this default is the only way to get rid of the ones you don’t need anymore.

2.1.11 lines (default 0): if set to a positive value, GP prints at most that many lines from each result, terminating the last line shown with [+++] if further material has been suppressed. The various print commands (see Section 3.11.2) are unaffected, so you can always type print(%), \a, or \b to view the full result. If the actual screen width cannot be determined, a “line” is assumed to be 80 characters long.

2.1.12 log (default 0): this is a toggle, which can be either 1 (on) or 0 (off). When logging mode is turned on, GP opens a log file, whose exact name is determined by the logfile default. Subsequently, all the commands and results will be written to that file (see \l). In case a file with this precise name already existed, it will not be erased: your data will be appended at the end.

2.1.13 logfile (default "pari.log"): name of the log file to be used when the log toggle is on. Tilde and time expansion are performed.

2.1.14 output (default 1): this can take any of the following three values: 0 (= raw), 1 (= prettymatrix), or 2 (= prettyprint). This means that, independently of the default format for reals which we explained above, you can print results in three ways: either in raw format, i.e. a format which is equivalent to what you input, including explicit multiplication signs, and everything typed on a line instead of two dimensional boxes. This can have several advantages, for instance it allows you to pick the result with a mouse or an editor, and to put it somewhere else.

The second format is the prettymatrix format. The only difference to raw format is that matrices are printed as boxes instead of horizontally. This is prettier, but takes more space and cannot be used for input. Column vectors are still printed horizontally.

The third format is the prettyprint or beautified format. In the present version 2.0.17, this is not beautiful at all.

Independently of the setting of this default, an object can be printed in any of the three formats at any time using the commands \a, \m and \b respectively (see below).

2.1.15 parisize (default, 1000000 bytes on the Mac, 4000000 otherwise): GP, and in fact any program using the PARI library, needs a stack in which to do its computations. parisize is the stack size, in bytes. It is strongly recommended you increase this default (using the -s commandline switch, or a gprc) if you can afford it. In case of emergency, you can use the allocatemem function to increase parisize, once the session is started. GP will try to double the stack size by itself when memory runs low during a computation, but this very computation will then be lost, and you will have to type the command again.

2.1.16 path (default ".:~:~/gp" on UNIX systems, ".;C:\;C:\GP on DOS, OS/2 and Windows, and "." otherwise): This is a list of directories, separated by colons ’:’ (semicolons ’;’ in the DOS world, since colons are pre-empted for drive names). When asked to read a file whose name does not contain / (i.e. no explicit path was given), GP will look for it in these directories, in the order they were written in path. Here, as usual, ’.’ means the current directory, and ’. .’ its immediate parent. Tilde expansion is performed.

2.1.17 primelimit (default 200000 on the Mac, and 500000 otherwise): GP precomputes a list of all primes less than primelimit at initialization time. These are used by many arithmetical functions. If you don’t plan to invoke any of them, you can just set this to 1.

2.1.18 prompt (default "? "): a string that will be printed as prompt. Note that most usual escape sequences are available there: \e for Esc, \n for Newline, …, \\ for \. Time expansion is performed.

This string is sent through the library function strftime (on a Unix system, you can try man strftime at your shell prompt). This means that % constructs have a special meaning, usually related to the time and date. For instance, %H = hour (24-hour clock) and %M = minute [00,59] (use %% to get a real %).

If you use readline, escape sequences in your prompt will result in display bugs. If you have a relatively recent readline (see the comment at the end of Section 2.1.2), you can brace them with special sequences (\[ and \]), and you will be safe. If these just result in extra spaces in your prompt, then you’ll have to get a more recent readline. See the file misc/gprc.dft for an example.

EMACS: Caution: Emacs needs to know about the prompt pattern to separate your input from previous GP results, without ambiguity. It’s not a trivial problem to adapt automatically this regular expression to an arbitrary prompt (which can be self-modifying!). Thus, in this version 2.0.17, Emacs relies on the prompt being the default one. So, do not tamper with the prompt variable unless you modify it simultaneously in your .emacs file (see emacs/pariemacs.txt and misc/gprc.dft for examples).

2.1.19 psfile (default "pari.ps"): name of the default file where GP is to dump its PostScript drawings (these will always be appended, so that no previous data are lost). Tilde and time expansion are performed.

2.1.20 realprecision (default 28 and 38 on 32-bit and 64-bit machines respectively): the number of significant digits and, at the same time, the number of printed digits of real numbers (see \p). Note that PARI internal precision works on a word basis (32 or 64 bits), hence may not coincide with the number of decimal digits you input. For instance to get 2 decimal digits you need one word of precision which, on a 32-bit machine, actually gives you 9 digits (9 < log_{10}(2^{32}) < 10):

    ?  default(realprecision, 2)
      realprecision = 9 significant digits (2 digits displayed)

2.1.21 secure (default 0): this is a toggle which can be either 1 (on) or 0 (off). If on, the system and extern command are disabled. These two commands are potentially dangerous when you execute foreign scripts since they let GP execute arbitrary UNIX commands. GP will ask for confirmation before letting you (or a script) unset this toggle.

2.1.22 seriesprecision (default 16): precision of power series (see \ps).

2.1.23 simplify (default 1): this is a toggle which can be either 1 (on) or 0 (off). When the PARI library computes something, the type of the result is not always the simplest possible. The only type conversions which the PARI library does automatically are rational numbers to integers (when they are of type t FRAC and equal to integers), and similarly rational functions to polynomials (when they are of type t RFRAC and equal to polynomials). This feature is useful in many cases, and saves time, but can be annoying at times. Hence you can disable this and, whenever you feel like it, use the function simplify (see Chapter 3) which allows you to simplify objects to the simplest possible types recursively (see \y).

2.1.24 strictmatch (default 1): this is a toggle which can be either 1 (on) or 0 (off). If on, unused characters after a sequence has been processed will produce an error. Otherwise just a warning is printed. This can be useful when you’re not sure how many parentheses you have to close after complicated nested loops.

2.1.25 timer (default 0): this is a toggle which can be either 1 (on) or 0 (off). If on, every instruction sequence (anything ended by a newline in your input) is timed, to some accuracy depending on the hardware and operating system. The time measured is the user CPU time, not including the time for printing the results (see # and ##).

2.1.26 Note on output formats.

A zero real number is printed in e format as 0.Exx where xx is the (usually negative) decimal exponent of the number (cf. Section 1.2.6.3). This allows the user to check the accuracy of the zero in question (this could also be done using \x, but that would be more technical).

When the integer part of a real number x is not known exactly because the exponent of x is greater than the internal precision, the real number is printed in e format (note that in versions before 1.38.93, this was instead printed with a ∗ at the end).

Note also that in beautified format, a number of type integer or real is written without enclosing parentheses, while most other types have them. Hence, if you see the expression (3.14), it is not of type real, but probably of type complex with zero imaginary part (if you want to be sure, type \x or use the function type).

2.2 Simple metacommands.

Simple metacommands are meant as shortcuts and should not be used in GP scripts (see Section 3.11). Beware that these, as all of GP input, are now case sensitive. For example, \Q is no longer identical to \q. In the following list, braces are used to denote optional arguments, with their default values when applicable, e.g. {n = 0} means that if n is not there, it is assumed to be 0. Whitespace (or spaces) between the metacommand and its arguments and within arguments is optional. (This can cause problems only with \w, when you insist on having a filename whose first character is a digit, and with \r or \w, if the filename itself contains a space. In such cases, just use the underlying read or write function; see Section 3.11.2.27).

2.2.1 ? {command}: GP on-line help interface. As already mentioned, if you type ?n where n is a number from 1 to 11, you will get the list of functions in Section 3.n of the manual (the list of sections being obtained by simply typing ?).

These names are in general not informative enough. More details can be obtained by typing ?function, which gives a short explanation of the function’s calling convention and effects. Of course, to have complete information, read Chapter 3 of this manual (the source code is at your disposal as well, though a trifle less readable!). Much better help can be obtained through the extended help system (see below).

You then get the function description exactly as it stands in Chapter 3. All operators (e.g. +, &&, etc.) are accepted by this extended help, as well as a few other keywords describing key GP concepts, e.g. readline (the line editor), integer, nf (“number field” as used in most algebraic number theory computations), ell (elliptic curves), etc.

UNIX: If the line before the copyright message indicates that extended help is available (this means perl is installed on your system, GP was told about it at compile time, and the whole PARI distribution was correctly installed), you can add more ? signs for extended functionalities:

?? keyword yields the functions description as it stands in this manual, usually in Chapter 2 or 3. If you’re not satisfied with the default chapter chosen, you can impose a given chapter by ending the keyword with @ followed by the chapter number, e.g. ?? Hello@2 will look in Chapter 2 for section heading Hello (which doesn’t exist, by the way).

??? pattern produces a list of sections in Chapter 3 of the manual related to your query. As before, if pattern ends by @ ifollowed by a chapter number, that chapter is searched instead; you also have the option to append a simple @ (without a chapter number) to browse through the whole manual.

If your query contains dangerous characters (e.g ? or blanks) it is advisable to enclose it within double quotes, as for GP strings (e.g ??? "elliptic curve").

Note that extended help is much more powerful than the short help, since it knows about operators as well: you can type ?? * or ?? &&, whereas a single ? would just yield a not too helpful

    *** unknown identifier.

message. Also, you can ask for extended help on section number n in Chapter 3, just by typing ?? n (where ?n would yield merely a list of functions). Finally, a few key concepts in GP are documented in this way: metacommands (e.g ?? "??"), defaults (e.g ?? psfile) and type names (e.g t INT or integer), as well as various miscellaneous keywords such as edit (short summary of

line editor commands), operator, member, "user defined", nf, ell, . . .

Last but not least : ?? without argument will open a dvi previewer (xdvi by default, $GPXDVI if it is defined in your environment) containing the full user’s manual. ??tutorial and ??refcard do the same with the tutorial and reference card respectively.

Technical note: these functionalities are provided by an external perl script that you are free to use outside any GP session (and modify to your liking, if you are perl-knowledgeable). It is called gphelp, lies in the doc subdirectory of your distribution (just make sure you run Configure first, see Appendix A) and is really two programs in one. The one which is used from within GP is gphelp which runs TEX on a selected part of this manual, then opens a previewer. gphelp -detex is a text mode equivalent, which looks often nicer especially on a colour-capable terminal (see misc/gprc.dft for examples). The default help selects which help program will be used from within GP. You are welcome to improve this help script, or write new ones (and we really would like to know about it so that we may include them in future distributions). By the way, outside of GP you can give more than one keyword as argument to gphelp.

2.2.2 /*...*/: comment. Everything between the stars is ignored by GP. These comments can span any number of lines.

2.2.3 \\: one-line comment. The rest of the line is ignored by GP.

2.2.4 \a {n}: prints the object number n (%n) in raw format. If the number n is omitted, print the latest computed object (%).

2.2.5 \b {n}: Same as \a, in prettyprint (i.e. beautified) format.

2.2.6 \c: prints the list of all available hardcoded functions under GP, not including operators written as special symbols (see Section 2.4). More information can be obtained using the ? metacommand (see above). For user-defined functions / member functions, see \u and \um.

2.2.7 \d: prints the defaults as described in the previous section (shortcut for default(), see Section 3.11.2.4).

2.2.8 \e {n}: switches the echo mode on (1) or off (0). If n is explicitly given, set echo to n.

2.2.9 \g {n}: sets the debugging level debug to the non-negative integer n.

2.2.10 \gf {n}: sets the file usage debugging level debufiles to the non-negative integer n.

2.2.11 \gm {n}: sets the memory debugging level debugmem to the non-negative integer n.

2.2.12 \h {m-n}: outputs some debugging info about the hashtable. If the argument is a number n, outputs the contents of cell n. Ranges can be given in the form m-n (from cell m to cell n, $ = last cell). If a function name is given instead of a number or range, outputs info on the internal structure of the hash cell this function occupies (a struct entree in C). If the range is reduced to a dash (’-’), outputs statistics about hash cell usage.

2.2.13 \l {logfile/}: switches log mode on and off. If a logfile argument is given, change the default logfile name to logfile and switch log mode on.

2.2.14 \m: as \a, but using prettymatrix format.

2.2.15 \p {n}: sets realprecision to n decimal digits. Prints its current value if n is omitted.

2.2.16 \ps {n}: sets seriesprecision to n significant terms. Prints its current value if n is omitted.

2.2.17 \q: quits the GP session and returns to the system. Shortcut for the function quit (see Section 3.11.2.20).

2.2.18 \r {filename}: reads into GP all the commands contained in the named file as if they had been typed from the keyboard, one line after the other. Can be used in combination with the \w command (see below). Related but not equivalent to the function read (see Section 3.11.2.21); in particular, if the file contains more than one line of input, there will be one history entry for each of them, whereas read would only record the last one. If filename is omitted, re-read the previously used input file (fails if no file has ever been successfully read in the current session).

UNIX: This command accepts compressed files in compressed (.Z) or gzipped (.gz or .z) format. They will be uncompressed on the fly as GP reads them, without changing the files themselves.

2.2.19 \s: prints the state of the PARI stack and heap. This is used primarily as a debugging device for PARI, and is not intended for the casual user.

2.2.20 \t: prints the internal longword format of all the PARI types. The detailed bit or byte format of the initial codeword(s) is explained in Chapter 4, but its knowledge is not necessary for a GP user.

2.2.21 \u: prints the definitions of all user-defined functions.

2.2.22 \um: prints the definitions of all user-defined member functions.

2.2.23 \v: prints the version number and implementation architecture (680x0, Sparc, Alpha, other) of the GP executable you are using. In library mode, you can use instead the two character strings PARIVERSION and PARIINFO, which correspond to the first two lines printed by GP just before the Copyright message.

2.2.24 \w {n} {filename}: writes the object number n ( %n ) into the named file, in raw format. If the number n is omitted, writes the latest computed object ( % ). If filename is omitted, appends to logfile (the GP function write is a trifle more powerful, as you can have filenames whose first character is a digit).

2.2.25 \x: prints the complete tree with addresses and contents (in hexadecimal) of the internal representation of the latest computed object in GP. As for \s, this is used primarily as a debugging device for PARI. However, used on a PARI integer, it can be used as a decimal→hexadecimal converter.

2.2.26 \y {n}: switches simplify on (1) or off (0). If n is explicitly given, set simplify to n.

2.2.27 #: switches the timer on or off.

2.2.28 ##: prints the time taken by the latest computation. Useful when you forgot to turn on

the timer.

2.3 Input formats for the PARI types.

Before describing more sophisticated functions in the next section, let us see here how to input values of the different data types known to PARI. Recall that blanks are ignored in any expression which is not a string (see below).

2.3.1 Integers (type t INT): type the integer (with an initial + or -, if desired) with no decimal point.

2.3.2 Real numbers (type t REAL): type the number with a decimal point. The internal precision of the real number will be the supremum of the input precision and the default precision. For example, if the default precision is 28 digits, typing 2. will give a number with internal precision 28, but typing a 45 significant digit real number will give a number with internal precision at least 45 (although less may be printed).

You can also use scientific notation with the letter E or e (like 6.02 E 23 or 1e-5).

2.3.3 Integermods (type t INTMOD): to enter n mod m, type Mod(n,m), not n%m (see Chapter 3).

2.3.4 Rational numbers (types t FRAC and t FRACN): under GP, all fractions are automatically reduced to lowest terms, so it is in principle impossible to work with reducible fractions (of type t FRACN), although of course in library mode this is easy. To enter n/m just type it as written. As explained in Section 3.1.4, division will not be performed, only reduction to lowest terms.

If you really want a reducible fraction under GP, you must use the type function (see Section 3.11.2.25), by typing type(x,FRACN). Be warned however that this function must be used with extreme care.

2.3.5 Complex numbers (type t COMPLEX): to enter x + iy, type x + I*y (not x+i*y). The letter I stands for \sqrt{−1}. Recall from Chapter 1 that x and y can be of type t INT, t REAL,

t INTMOD, t FRAC/t FRACN, or t PADIC.

2.3.6 p-adic numbers (type t PADIC): to enter a p-adic number, simply write a rational or integer expression and add to it O(p^k), where p and k are integers. This last expression indicates three things to GP: first that it is dealing with a t PADIC type (the fact that p is an integer, and not a polynomial, which would be used to enter a series, see Section 2.3.10), secondly the “prime” p (note that it is not checked whether p is indeed prime; you can work on 10-adics if you want, but beware of disasters as soon as you do something non-trivial like taking a square root), and finally the number of significant p-adic digits k. Note that O(25) is not the same as O(5^2); you probably want the latter!

For example, you can type in the 7-adic number

    2*7^(-1) + 3 + 4*7 + 2*7^2 + O(7^3)

exactly as shown, or equivalently as 905/7 + O(7^3).

2.3.7 Quadratic numbers (type t QUAD): first, you must define the default quadratic order or field in which you want to work. This is done using the quadgen function, in the following way. Write something like

    w = quadgen(d)

where d is the discriminant of the quadratic order in which you want to work (hence d is congruent to 0 or 1 modulo 4). The name w is of course just a suggestion, but corresponds to traditional usage. You can of course use any variable name that you like. However, quadratic numbers are always printed with a w, regardless of the discriminant. So beware, two numbers can be printed in the same way and not be equal. However GP will refuse to add or multiply them for example.

Now (1, w) will be the “canonical” integral basis of the quadratic order (i.e. w = \sqrt{d}/2 if d ≡ 0 mod 4, and w = (1 + \sqrt{d})/2 if d ≡ 1 mod 4, where d is the discriminant), and to enter x + yw you just type x + y*w.

2.3.8 Polmods (type t POLMOD): exactly as for integermods, to enter x mod y (where x and y are polynomials), type Mod(x,y), not x%y (see Section 3.2.3). Note that when y is an irreducible polynomial in one variable, polmods whose modulus is y are simply algebraic numbers in the finite extension defined by the polynomial y. This allows us to work easily in number fields, finite extensions of the p-adic field Q_{p}, or finite fields.

Important remark. Since the variables occurring in a polmod are not free variables, it is essential in order to avoid inconsistencies that polmods use the same variable in internal operations (i.e. between polmods) and variables of lower priority (which have been introduced later in the GP session) for external operations (typically between a polynomial and a polmod). For example, PARI will not recognize that Mod(y, y^2 + 1) is the same as Mod(x, x^2 + 1). Hopefully, this problem will pass away when type “element of a number field” is eventually introduced.

    On the other hand, Mod(x, x^2 + 1) + Mod(x, x^2 + 1) (which gives Mod(2*x, x^2 + 1))

and x + Mod(y, y^2 + 1) (which gives a result mathematically equivalent to x + i with i^{2} = −1) are completely correct, while y + Mod(x, x^2 + 1) gives Mod(x + y, x^2 + 1), which may not be what you want (y is treated here as a numerical parameter, not as a polynomial variable).

Note (added in version 2.0.16) As long as the main variables are the same, it is allowed to mix t POL and t POLMODs. The result will be the expected t POLMOD. For instance x + Mod(x, x^2 + 1) is equal to Mod(2*x, x^2 + 1). This wasn’t the case prior to version 2.0.16: it returned a polynomial in x equivalent to x + i, which was in fact an invalid object (you couldn’t lift it).

2.3.9 Polynomials (type t POL): type the polynomial in a natural way, not forgetting to put a “∗” between a coefficient and a formal variable (this ∗ does not appear in beautified output). Any variable name can be used except for the reserved names I (used exclusively for the square root of −1), Pi (3.14 …), Euler (Euler’s constant), and all the function names: predefined functions, as described in Chapter 3 (use \c to get the complete list of them) and user-defined functions, which you ought to know about (use \u if you are subject to memory lapses). The total number of different variable names is limited to 16384 and 65536 on 32-bit and 64-bit machines respectively, which should be enough. If you ever need hundreds of variables, you should probably be using vectors instead.

2.3.10 Power series (type t SER): type a rational function or polynomial expression and add to it O(expr ^k), where expr is an expression which has non-zero valuation (it can be a polynomial, power series, or a rational function; the most common case being simply a variable name). This indicates to GP that it is dealing with a power series, and the desired precision is k times the valuation of expr with respect to the main variable of expr (to check the ordering of the variables, or to modify it, use the function reorder; see Section 3.11.2.22).

2.3.11 Rational functions (types t RFRAC and t RFRACN): as for fractions, all rational functions are automatically reduced to lowest terms under GP. All that was said about fractions in Section 2.3.4 remains valid here.

2.3.12 Binary quadratic forms of positive or negative discriminant (type t QFR and t QFI): these are input using the function Qfb (see Chapter 3). For example Qfb(1,2,3) will create the binary form x^{2} + 2xy + 3y^{2}. It will be imaginary (of internal type t QFI) since 2^{2} − 4 ∗ 3 = −8 is negative.

In the case of forms with positive discriminant (type t QFR), you may add an optional fourth component (related to the regulator, more precisely to Shanks and Lenstra’s “distance”), which must be a real number. See also the function qfbprimeform which directly creates a prime form of given discriminant (see Chapter 3).

2.3.13 Row and column vectors (types t VEC and t COL): to enter a row vector, type the components separated by commas “,”, and enclosed between brackets “[ ” and “ ]”, e.g. [1,2,3]. To enter a column vector, type the vector horizontally, and add a tilde “˜ ” to transpose. [ ] yields the empty (row) vector. The function Vec can be used to transform any object into a vector (see Chapter 3).

2.3.14 Matrices (type t MAT): to enter a matrix, type the components line by line, the components being separated by commas “,”, the lines by semicolons “;”, and everything enclosed in brackets “[ ” and “ ]”, e.g. [x,y; z,t; u,v]. [ ; ] yields the empty (0x0) matrix. The function Mat can be used to transform any object into a matrix (see Chapter 3).

Note that although the internal representation is essentially the same (only the type number is different), a row vector of column vectors is not a matrix; for example, multiplication will not work in the same way.

Note also that it is possible to create matrices (by conversion of empty column vectors and concatenation, or using the matrix function) with a given positive number of columns, each of which has zero rows. It is not possible to create or represent matrices with zero columns and a nonzero number of rows.

2.3.15 Lists (type t LIST): lists cannot be input directly; you have to use the function listcreate first, then listput each time you want to append a new element (but you can access the elements directly as with the vector types described above). The function List can be used to transform (row or column) vectors into lists (see Chapter 3).

2.3.16 Strings (type t STR): to enter a string, just enclose it between double quotes ", like this: "this is a string". The function Str can be used to transform any object into a string (see Chapter 3).

2.4 GP operators.

Loosely speaking, an operator is a function (usually associated to basic arithmetic operations) whose name contains only non-alphanumeric characters. In practice, most of these are simple functions, which take arguments, and return a value; assignment operators also have side effects. Each of these has some fixed and unchangeable priority, which means that, in a given expression, the operations with the highest priority will be performed first. Operations at the same priority level will always be performed in the order they were written, i.e. from left to right. Anything enclosed between parenthesis is considered a complete subexpression, and will be resolved independently of the surrounding context. For instance, assuming that op_{1}, op_{2}, op_{3} are standard binary operators with increasing priorities (think of +, *, ^for instance),

x op_{1} y op_{2} z op_{2} x op_{3} y

is equivalent to

x op_{1} ((y op_{2} z) op_{2} (x op_{3} y)).

GP knows quite a lot of different operators, some of them unary (having only one argument), some binary. Unary operators are defined for either prefix (preceding their single argument: op x) or postfix (following the argument: x op) position, never both (some are syntactically correct in both positions, but with different meanings). Binary operators all use the syntax x op y. Most of them are well known, some are borrowed from C syntax, and a few are specific to GP. Beware that some GP operators may differ slightly from their C counterparts. For instance, GP’s postfix ++ returns the new value, like the prefix ++ of C, and the binary shifts <<, >> have a priority which is different from (higher than) that of their C counterparts. When in doubt, just surround everything by parentheses (besides, your code will probably be more legible).

Here is the complete list (in order of decreasing priority, binary unless mentioned otherwise):

• Priority 9 ++ and -- (unary, postfix): x++ assigns the value x + 1 to x, then returns the new value of

x. This corresponds to the C statement ++x (there is no prefix ++ operator in GP). x-- does the same with x − 1.

• Priority 8 op=, where op is any simple binary operator (i.e. a binary operator with no side effects, i.e. one

of those defined below) which is not a boolean operator (comparison or logical). x op= y assigns (x op y) to x, and returns the new value of x, not a reference to the variable x. (Thus an assignment cannot occur on the lefthand side of another assignment.)

• Priority 7 = is the assignment operator. The result of x = y is the value of the expression y, which is

also assigned to the variable x. This is not the equality test operator. Beware that a statement like x = 1 is always true (i.e. non-zero), and sets x to 1.

• Priority 6 ! (unary, prefix): logical not. !x return 1 if x is equal to 0 (specifically, if gcmp0(x)==1), and

0 otherwise.

' (unary, prefix): quote its argument without evaluating it.

    ?  a = x + 1; x = 1;
    ?  subst(a,x,1)
      ***   variable name expected:  subst(a,x,1)
                                            ^---
    ?  subst(a,'x,1)
    %1 = 2

• Priority 5 ^: powering.

' (unary, postfix): derivative with respect to the main variable.

~ (unary, postfix): vector/matrix transpose.

! (unary, postfix): factorial. x! = x(x − 1) ⋯ 1.

.: x.b extracts member b from structure x.

• Priority 4 +, - (unary, prefix): - toggles the sign of its argument, + has no effect whatsoever.

• Priority 3 *: multiplication.

/: exact division (3/2=3/2, not 1.5).

\, %: euclidean quotient and remainder, i.e. if x = qy + r, with 0 ≤ r < y (if x and y are polynomials, assume instead that deg r < deg y and that the leading terms of r and x have the same sign), then x\y = q, x%y = r.

\/: rounded euclidean quotient for integers (rounded towards +∞ when the exact quotient would be a half-integer).

<<, >>: left and right binary shift: x<<n = x ∗ 2^{n} if n > 0, and x\/2^{−n} otherwise; and

x>>n = x<<(-n).

• Priority 2 +, -: addition/subtraction.

• Priority 1 <, >, <=, >=: the usual comparison operators, returning 1 for true and 0 for false. For

instance, x<=1 returns 1 if x ≤ 1 and 0 otherwise.

<>, !=: test for (exact) inequality.

==: test for (exact) equality.

• Priority 0 &, &&: logical and.

|, ||: logical (inclusive) or. Any sequence of logical or and and operations is evaluated from left to right, and aborted as soon as the final truth value is known. Thus, for instance, (x && 1/x) or (type(p) == "t INT" && isprime(p)) will never produce an error since the second argument need not (and will not) be processed when the first is already zero (false).

Remark: For the sake of efficiency, you should use the ++, -- and op= operators whenever possible:

?  a=200000;
?  i=0; while(i<a, i=i+1)
time = 4,919 ms.
?  i=0; while(i<a, i+=1)
time = 4,478 ms.
?  i=0; while(i<a, i++)
time = 3,639 ms.

The shift operators should be preferred to multiplication whenever possible:

?  a=1<<20000;
?  i=1; while(i<a, i=i*2);
time = 5,255 ms.
?  i=1; while(i<a, i<<=1);
time = 988 ms.

2.5 The general GP input line.

2.5.1 Generalities. User interaction with a GP session proceeds as follows: a sequence of characters is typed by the user at the GP prompt. This can be either a \ command, a function definition, an expression, or a sequence of expressions (i.e. a program). In the latter two cases, after the last expression has been computed its result is put into an internal (“history”) array, and printed. The successive elements of this array are called %1, %2, … As a shortcut, the latest computed expression can also be called %, the previous one %`, the one before that %`` and so on.

If you want to suppress the printing of the result, for example because it is a long unimportant intermediate result, end the expression with a ; sign. This same sign is used as an instruction separator when several instructions are written on the same line (note that for the pleasure of BASIC addicts, the : sign can also be used, but we will try to stick to C-style conventions in this manual). The final expression computed, even if not printed, will still be assigned to the history array, so you may have to pay close attention when you intend to refer back to it by number since this number does not appear explicitly. Of course, if you just want to use it on the next line, use % as usual.

Any legal expression can be typed in, and is evaluated using the conventions about operator priorities and left to right associativity (see the previous section), using the available operator symbols, function names (including user-defined functions and member functions see Section 2.6.3), and special variables. Please note that, from version 1.900 on, there is a distinction between lowercase and uppercase. Also, note that, outside of constant strings, blanks are completely ignored in the input to GP.

The special variablevariable (special) names known to GP are Euler (Euler’s constant γ = 0.577 …), I (the square root of −1), Pi (3.14…) — which could be thought of as functions with no arguments, and which may therefore be invoked without parentheses —, and O which obeys the following syntax:

    O(expr^k)

When expr is an integer or a rational number, this creates an expr-adic number (zero in fact) of precision k. When expr is a polynomial, a power series or a rational function whose main variable is X, say, this creates a power series (also zero) of precision v ∗ k where v is the X-adic valuation of expr (see 2.3.6 and 2.3.9).

2.5.2 Special editing characters. A GP program can of course have more than one line. Since GP executes your commands as soon as you have finished typing them, there must be a way to tell it to wait for the next line or lines of input before doing anything. There are three ways of doing this.

The first one is simply to use the backslash character \ at the end of the line that you are typing, just before hitting <Return>. This tells GP that what you will write on the next line is the physical continuation of what you have just written. In other words, it makes GP forget your newline character. For example if you use this while defining a function, and if you ask for the definition of the function using ?name, you will see that your backslash has disappeared and that everything is on the same line. You can type a \ anywhere. It will be interpreted as above only if (apart from ignored whitespace characters) it is immediately followed by a newline. For example, you can type

    ?  3 + \
    4

instead of typing 3 + 4.

The second one is a slight variation on the first, and is mostly useful when defining a user function (see Section 2.6.3): since an equal sign can never end a valid expression, GP will disregard a newline immediately following an =.

    ?  a =
    123
    %1 = 123

The third one cannot be used everywhere, but is in general much more useful. It is the use of braces { and }. When GP sees an opening brace ({) at the beginning of a line (modulo spaces as usual), it understands that you are typing a multi-line command, and newlines will be ignored until you type a closing brace }. However, there is an important (but easily obeyed) restriction: inside an open brace-close brace pair, all your input lines will be concatenated, suppressing any newlines. Thus, all newlines should occur after a semicolon (;), a comma (,) or an operator (for clarity’s sake, we don’t recommend splitting an identifier over two lines in this way). For instance, the following program

    {
      a = b
      b = c
    }

would silently produce garbage, since what GP will really see is a=bb=c which will assign the value of c to both bb and a (if this really is what you intended, you’re a hopeless case).

2.6 The GP/PARI programming language.

The GP calculator uses a purely interpreted language. The structure of this language is reminiscent of LISP with a functional notation, f(x,y) rather than (f x y): all programming constructs, such as if, while, etc... are functions * (see Section 3.11 for a complete list), and the main loop does not really execute, but rather evaluates (sequences of) expressions. Of course, it is by no means a true LISP.

2.6.1 Variables and symbolic expressions.

In GP you can use up to 16383 variable names (up to 65535 on 64-bit machines). These names can be any standard identifier names, i.e. they must start with a letter and contain only valid keyword characters: or alphanumeric characters ([ A-Za-z0-9]). To avoid confusion with other symbols, you must not use other non-alphanumeric symbols like $, or ’.’. In addition to the function names which you must not use (see the list with \c), there are exactly three special variable names which you are not allowed to use: Pi and Euler, which represent well known constants, and I = \sqrt{−1}.

Note that GP names are case sensitive since version 1.900. This means for instance that the symbol i is perfectly safe to use, and will not be mistaken for \sqrt{−1}, and that o is not synonymous anymore to O. If you grew addicted to the previous behaviour, you can have it back by setting the default compatible to 3.

* Not exactly, since not all their arguments need be evaluated. For instance it would be stupid to evaluate both branches of an if statement: since only one will apply, GP only expands this one.

Now the main thing to understand is that PARI/GP is not a symbolic manipulation package, although it shares some of the functionalities. One of the main consequences of this fact is that all expressions are evaluated as soon as they are written, they never stay in a purely abstract form**. As an important example, consider what happens when you use a variable name before assigning a value into it. This is perfectly acceptable to GP, which considers this variable in fact as a polynomial of degree 1, with coefficients 1 in degree 1, 0 in degree 0, whose variable is the variable name you used.

If later you assign a value to that variable, the objects which you have created before will still be considered as polynomials. If you want to obtain their value, use the function eval (see Section 3.7.3).

Finally, note that if the variable x contains a vector or list, you can assign a result to x[m] (i.e. write something like x[k] = expr). If x is a matrix, you can assign a result to x[m, n], but not to x[m]. If you want to assign an expression to the m-th column of a matrix x, use x[, m] = expr instead. Similarly, use x[m,] = expr to assign an expression to the m-th row of x. This process is recursive, so if x is a matrix of matrices of …, an expression such as x[1, 1][, 3][4] = 1 would be perfectly valid (assuming of course that all matrices along the way have the correct dimensions).

Note: We’ll see in Section 2.6.3 that it is possible to restrict the use of a given variable by declaring it to be global or local. This can be useful to enforce clean programming style, but is in no way mandatory.

(Technical) Note: Variables are numbered in the order that they appear since the beginning of the session, and the main variable of an expression is always the lowest numbered variable. Hence if you are working with expressions involving several variables and want to have them ordered in a specific manner in the internal representation, the simplest is just to write down the variables one after the other under GP before starting any real computations. If you already have started working and want to change the names of the variables in an object, use the function changevar. If you only want to have them ordered when the result is printed, you can also use the function reorder, but this won’t change anything to the internal representation.

(Very technical) Note: Each variable has a stack of values, implemented as a linked list. When a new scope is entered (during a function call which uses it as a parameter, or if the variable is used as a loop index, see Section 2.6.3 and Section 3.11), the value of the actual parameter is pushed on the stack. If the parameter is not supplied, a special 0 value called gnil is pushed on the stack (this value is not printed if it is returned as the result of a GP expression sequence). Upon exit, the stack decreases. You can kill a variable, decreasing the stack yourself. This should be used only at the top level of GP, to undo the effect of an assignment, not from a function. However, the stack has a bottom: the value of a variable is the monomial of degree 1 in this variable, as is natural for a mathematician.

** An obvious but important exception are character strings which are evaluated… essentially to themselves (type t STR). Not exactly so though, since we do some work to treat the quoted characters correctly (those preceded by a \).

2.6.2 Expressions and expression sequences.

An expression is formed by combining the GP operators, functions (including user-defined functions, see below) and control statements. It may be preceded by an assignment statement ’=’ into a variable. It always has a value, which can be any PARI object.

Several expressions can be combined on a single line by separating them with semicolons (’;’) and also with colons (’:’) for those who are used to BASIC. Such an expression sequence will be called simply a seq. A seq also has a value, which is the value of the last non-empty expression in the sequence. Under GP, the value of the seq, and only this last value, is always put on the stack (i.e. it will become the next object %n). The values of the other expressions in the seq are discarded after the execution of the seq is complete, except of course if they were assigned into variables. In addition, the value of the seq (or of course of an expression if there is only one) is printed if the line does not end with a semicolon (’;’).

2.6.3 User defined functions.

It is very easy to define a new function under GP, which can then be used like any other function. The syntax is as follows:

    name(list of formal variables) = local(list of local variables); seq

which looks better written on consecutive lines:

    {
      local(t_{0}, t_{1}, . . .);
      local(. . .);

…

    }

(note that the first newline is disregarded due to the preceding = sign, and the others because of the enclosing braces). Both lists of variables are comma-separated and allowed to be empty. The local statements can be omitted; as usual seq is any expression sequence.

name is the name given to the function and is subject to the same restrictions as variable names. In addition, variable names are not valid function names, you have to kill the variable first (the converse is true: function names can’t be used as variables, see Section 3.11.2.14). Previously used function names can be recycled: you are just redefining the function (the previous definition is lost of course).

list of formal variables is the list of variables corresponding to those which you will actually use when calling your function. The number of actual parameters supplied when calling the function has to be less than the number of formal variables.

Uninitialized formal variables will be given a default value. An equal (=) sign following a variable name in the function definition, followed by any expression, gives the variable a default value. The expression gets evaluated the moment the function is defined, and is fixed afterward. A variable for which you supply no default value will be initialized to zero.

list of local variables is the list of the additional local variables which are used in the function body. Note that if you omit some or all of these local variable declarations, the nondeclared variables will become global, hence known outside of the function, and this may have undesirable side-effects. On the other hand, in some cases it may also be what you want. Local variables can be given a default value as the formal variables.

Example: For instance foo(x=1,y=2,z=3) = print(x ":" y ":" z), defines a function which prints its arguments (at most three of them), separated by colons. This then follows the rules of default arguments generation as explained at the beginning of Section 3.0.0.

    ?  foo(6,7)
    6:7:3
    ?  foo(,5)
    1:5:3
    ?  foo
    1:2:3

Once the function is defined using the above syntax, you can use it like any other function. In addition, you can also recall its definition exactly as you do for predefined functions, that is by writing ?name. This will print the list of arguments, as well as their default values, the text of seq, and a short help text if one was provided using the addhelp function (see Section 3.11.2.1). One small difference to predefined functions is that you can never redefine the built-in functions, while you can redefine a user-defined function as many times as you want.

Typing \u will output the list of user-defined functions.

An amusing example of a user-defined function is the following. It is intended to illustrate both the use of user-defined functions and the power of the sumalt function. Although the Riemann zeta-function is included in the standard functions, let us assume that this is not the case (or that we want another implementation). One way to define it, which is probably the simplest (but certainly not the most efficient), is as follows:

    zet(s) =
    {
      local(j); /* not needed, and possibly confusing (see below) */
      sumalt(j=1, (-1)^(j-1)*j^(-s)) / (1 - 2^(1-s))
    }

This gives reasonably good accuracy and speed as long as you are not too far from the domain of convergence. Try it for s integral between −5 and 5, say, or for s = 0.5 + i ∗ t where t = 14.134 …

The iterative constructs which use a variable name (forxxx, prodxxx, sumxxx, vector, matrix, plot, etc.) also consider the given variable to be local to the construct. A value is pushed on entry and pulled on exit. So, it is not necessary for a function using such a construct to declare the variable as a dummy formal parameter.

In particular, since loop variables are not visible outside their loops, the variable j need not be declared in the protoype of our zet function above.

    zet(s) = sumalt(j=1, (-1)^(j-1)*j^(-s)) / (1 - 2^(1-s))

would be a perfectly sensible (and in fact better) definition. Since local/global scope is a very tricky point, here’s one more example. What’s wrong with the following definition?

    ?  first prime div(x) =
    {
      local(p);
      forprime(p=2, x, if (x%p == 0, break));
      p
    }

    ?  first prime div(10)
    %1 = 0

Answer: the index p in the forprime loop is local to the loop and is not visible to the outside world. Hence, it doesn’t survive the break statement. More precisely, at this point the loop index is restored to its preceding value, which is 0 (local variables are initialized to 0 by default). To sum up, the routine returns the p declared local to it, not the one which was local to forprime and ran through consecutive prime numbers. Here’s a corrected version:

    ?  first prime div(x) = forprime(p=2, x, if (x%p == 0, return(p)))

Again, it is strongly recommended to declare all other local variables that are used inside a function: if a function accesses a variable which is not one of its formal parameters, the value used will be the one which happens to be on top of the stack at the time of the call. This could be a “global” value, or a local value belonging to any function higher in the call chain. So, be warned.

There’s no problem with recursive functions as long as one pays proper attention to variable scope. Here’s a last example, used to retrieve the coefficient array of a multivariate polynomial (a non-trivial task due to PARI’s unsophisticated representation for those objects) :

    coeffs(P, nbvar) =
    {
      local(v);
      if (type(P) != "t POL",
        for (i=0, nbvar-1, P = [P]);
        return (P)
      );
      v = vector(poldegree(P)+1, i, polcoeff(P,i-1));
      vector(length(v), i, coeffs(v[i], nbvar-1))
    }

If P is a polynomial in k variables, show that after the assignment v = coeffs(P,k), the coefficient of x_{1}^{n_{1}} … x_{k}^{n_{k}} in P is given by v[n_{1}+1][…][n_{k}+1]. What would happen if the declaration local(v) had been omitted ?

Function which take functions as parameters ? This is easy in GP using the following trick (neat example due to Bill Daly):

    calc(f, x) = eval(Str( f "(x)"))

If you call this with calc("sin", 1), it will return sin(1) (evaluated!).

Restrictions on variable use: it is not allowed to use the same variable name for different parameters of your function. Or to use a given variable both as a formal parameter and a local variable in a given function. Hence

    ?  f(x,x) = 1
      *** user function f:  variable x declared twice.

Also, the statement global(x, y, z, t) (see Section 3.11.2.11) declares the corresponding variables to be global. It is then forbidden to use them as formal parameters or loop indexes as described above, since the parameter would “shadow” the variable.

Implementation note. For the curious reader, here is how these stacks are handled: a hashing function is computed from the identifier, and used as an index in hashtable, a table of pointers. Each of these pointers begins a linked list of structures (type entree). The linked list is searched linearly for the identifier (each list will typically have less than 7 components or so). When the correct entree is found, it points to the top of the stack of values for that identifier if it is a variable, to the function itself if it is a predefined function, and to a copy of the text of the function if it is a user-defined function. When an error occurs, all of this maze (rather a tree, in fact) is searched and (hopefully) restored to the state preceding the last call of the main evaluator.

Note: The above syntax (using the local keyword) was introduced in version 2.0.13. The old syntax

    name(list of true formal variables, list of local variables) = seq

is still recognized but is deprecated since genuine arguments and local variables become undistinguishable.

2.6.4 Member functions.

Member functions use the ‘dot’ notation to retrieve information from complicated structures (by default: types ell, nf, bnf, bnr and prime ideals). The syntax structure.member is taken to mean: retrieve member from structure, e.g. ell.j returns the j-invariant of the elliptic curve ell (or outputs an error message if ell doesn’t have the correct type).

You can define your own member functions using the syntax:

    structure.member = function text

where function text is written as the seq in a standard user function (without local variables), whose only argument would be structure. For instance, the current implementation of the ell type is simply an horizontal vector, the j-invariant being the thirteenth component. This could be implemented as

    x.j =
    {
      if (type(x) != "t VEC" || length(x) < 14,
        error("this is not a proper elliptic curve:  " x)
      );
      x[13]
    }

You can redefine one of your own member functions simply by typing a new definition for it. On the other hand, as a safety measure, you can’t redefine the built-in member functions, so typing the above text would in fact produce an error (you’d have to call it e.g. x.j2 in order for GP to accept it).

Typing \um will output the list of user-defined member functions.

Note: Member functions were not meant to be too complicated or to depend on any data that wouldn’t be global. Hence they do no have parameters (besides the implicit structure) or local variables. Of course, if you need some preprocessing work in there, there’s nothing to prevent you from calling your own functions (using freely their local variables) from a member function. For instance, one could implement (a dreadful idea as far as efficiency goes):

    correct ell if needed(x) =
    {
      local(tmp);
      if (type(x) != "t VEC", tmp = ellinit(x))

\\some further checks

      tmp
    }
    x.j = correct ell if needed(x)[13];

2.6.5 Strings and Keywords

GP variables can now hold values of type character string (internal type t STR). This section describes how they are actually used, as well as some convenient tricks (automatic concatenation and expansion, keywords) valid in string context.

As explained above, the general way to input a string is to enclose characters between quotes ". This is the only input construct where whitespace characters are significant: the string will contain the exact number of spaces you typed in. Besides, you can “escape” characters by putting a \ just before them; this has the following effects:

\e: the <Escape> character.

\n: the <Newline> character.

\t: the <Tab> character.

      \any-other-char: the any-other-char character.

In particular, the only way to put a " into a string is to escape it. Thus, for instance, "\"a\"" would produce the string whose content is “a”. This is definitely not the same thing as typing "a", whose content is merely the one-letter string a.

You can concatenate two strings using the concat function. If either argument is a string, the other is automatically converted to a string if necessary (it will be evaluated first).

    ?  concat("ex", 1+1)
    %1 = "ex2"
    ?  a = 2; b = "ex"; concat(b, a)
    %2 = "ex2"
    ?  concat(a, b)
    %3 = "2ex"

Some functions expect strings for some of their arguments: print would be an obvious example, Str is a less obvious but very useful one (see the end of this section for a complete list). While typing in such an argument, you will be said to be in string context. The rest of this section is devoted to special syntactical tricks which can be used with such arguments (and only here; you will get an error message if you try these outside of string context):

• Writing two strings alongside one another will just concatenate them, producing a longer string. Thus it is equivalent to type in "a " "b" or "a b". A little tricky point in the first expression: the first whitespace is enclosed between quotes, and so is part of a string; while the second (before the "b") is completely optional and GP actually suppresses it, as it would with any number of whitespace characters at this point (i.e. outside of any string).

• If you insert an expression without quotes when GP expects a string, it gets “expanded”: it is evaluated as a standard GP expression, and the final result (as would have been printed if you had typed it by itself) is then converted to a string, as if you had typed it directly. For instance "a" 1+1 "b" is equivalent to "a2b": three strings get created, the middle one being the expansion of 1+1, and these are then concatenated according to the rule described above. Another tricky point here: assume you did not assign a value to aaa in a GP expression before. Then typing aaa by itself in a string context will actually produce the correct output (i.e. the string whose content is aaa), but in a fortuitous way. This aaa gets expanded to the monomial of degree one in the variable aaa, which is of course printed as aaa, and thus will expand to the three letters you were expecting. But you will have defined a variable as a side effect.

• Since there are cases where expansion is not really desirable, we now distinguish between “Keywords” and “Strings”. String is what has been described so far. Keywords are special relatives of Strings which are automatically assumed to be quoted, whether you actually type in the quotes or not. Thus expansion is never performed on them. They get concatenated, though. The analyzer supplies automatically the quotes you have “forgotten” and treats Keywords just as normal strings otherwise. For instance, if you type "a"b+b in Keyword context, you will get the string whose contents are ab+b. In String context, on the other hand, you would get a2*b (and you would have created the variable b in the process if it didn’t exist before, but not the variable a).

All GP functions have prototypes (described in Chapter 3 below) which specify the types of arguments they expect: either generic PARI objects (GEN), or strings, or keywords, or unevaluated expression sequences. In the keyword case, only a very small set of words will actually be meaningful (the default function is a prominent example).

Let’s now try some not-so-stupid exercises to get the hang of it. Try to guess the results of the following commands without actually typing them, assuming that the print command evaluates and prints its (string) arguments in left-to-right order, ending with a newline (and returns 0 as an unprinted result):

      print()
      print(1+3"a,3" ,4)
      print(a=3, (1 + ((a-3)==print())) (a = (a == 5\/2)))

To round this up, here is a less artificial example, used to create generic matrices:

    ?  genmat(u,v,s="x") = matrix(u,v,i,j, eval(Str(s "" i "" j)))
    ?  genmat(2,3) + genmat(2,3,m)
    %1 =
    [x11 + m11 x12 + m12 x13 + m13]
    [x21 + m21 x22 + m22 x23 + m23]

Note that the argument of Str is evaluated in string context, and really consists of 5 pieces (exercise: why are the empty strings necessary?). This part could also have been written as con-

cat(concat(Str(s), i), j) (but not as concat(Str(s), concat(i,j))!). In practice, Str will

often be easier to use than concat, if slightly more cryptic.

The arguments of the following functions are processed in string context:

    Str

addhelp (second argument) default (second argument)

    error
    extern

plotstring (second argument) plotterm (first argument) all the printxxx functions

    read
    system

all the writexxx functions

The arguments of the following functions are processed as keywords:

    alias

default (first argument)

install (all arguments but the last)

type (second argument)

    whatnow

2.7 Interfacing GP with other languages.

The PARI library was meant to be interfaced with C programs. This specific use will be dealt with extensively in Chapter 4. GP itself provides a convenient, if simple-minded, interpreter, which enables you to execute rather intricate scripts (see Section 3.11).

Scripts, when properly written, tend to be shorter and much clearer than C programs, and are certainly easier to write, maintain or debug. You don’t need to deal with memory management, garbage collection, pointers, declarations, and so on. Because of their intrinsic simplicity, they are more robust as well. They are unfortunately somewhat slower. Thus their use will remain complementary: it is suggested that you test and debug your algorithms using scripts, before actually coding them in C for the sake of speed.

UNIX: Note that the install command enables you to concentrate on critical parts of your programs only (which can of course be written with the help of other mathematical libraries than PARI!), and to easily and efficiently import foreign functions for use under GP (see Section 3.11.2.13).

We are aware of three PARI-related public domain libraries. We neither endorse nor support any of them. You might want to give them a try if you are familiar with the languages they are based on. First, there are PariPerl*, written by Ilya Zakharevich (ilya@math.ohio-state.edu), and PariPython**, by Stéfane Fermigier (fermigie@math.jussieu.fr). Finaly, Michael Stoll (Michael Stoll@math.uni-bonn.de) has integrated PARI into CLISP, which is a Common Lisp implementation by Bruno Haible, Marcus Daniels and others. These provide interfaces to GP functions for use in perl, python or Lisp programs. To our knowledge, only the python and perl interfaces have been upgraded to version 2.0 of PARI, the CLISP one being still based on version 1.39.xx.

  * see http://nswt.tuwien.ac.at:8000/htdocs/internet/unix/perl/math-pari.html
 ** see http://www.math.jussieu.fr/~fermigie/PariPython/readme.html

2.8 The preferences file.

When GP is started, it looks for a customization file, or gprc in the following places (in this order, only the first one found will be read):

• On the Macintosh (only), GP looks in the directory which contains the GP executable itself for a file called gprc. No other places are examined.

• If the operating system supports environment variables (essentially, anything but MacOS), GP checks whether the environment variable GPRC is set. Under DOS, you can set it in AUTOEXEC.BAT. On Unix, this can be done with something like:

    GPRC=/my/dir/anyname; export GPRC in sh syntax (for instance in your .profile),
    setenv GPRC /my/dir/anyname  in csh syntax (in your .login or .cshrc file).

If so, the file named by $GPRC is the gprc.

• If GPRC is not set, and if the environment variable HOME is defined, GP then tries $HOME/.gprc on a Unix system $HOME\ gprc on a DOS, OS/2, or Windows system.

• If HOME also leaves us clueless, we try ~/.gprc on a Unix system (where as usual ~ stands for your home directory), or \ gprc on a DOS, OS/2, or Windows system.

• Finally, if no gprc was found among the user files mentioned above we look for /etc/gprc (\etc\gprc) for a system-wide gprc file (you’ll need root privileges to set up such a file yourself).

Note that on Unix systems, the gprc’s default name starts with a ’.’ and thus is hidden to regular ls commands; you need to type ls -a to see whether it’s already there without your knowing about it.

In any case, GP will open the corresponding file and process the commands in there, before doing anything else, e.g. creating the PARI stack. If the file doesn’t exist or cannot be read, GP will proceed to the initialization phase at once, eventually emitting a prompt. If any explicit commandline switches are given, they will override the values read from the gprc file.

The syntax in this file (and valid in this file only, at this very precise moment!) is simpleminded, but should be sufficient for most purposes. It is read line by line, white space being optional as usual (unless surrounded by quotes). Two types of lines are first dealt with by a preprocessor:

• comments are removed. This applies to all text surrounded by /* … */ as well as everything following \\ on a given line.

• lines starting with #if keyword are treated as comments if keyword is not defined, and read normally otherwise. The condition can be negated using either #if not (or #if !). Only two keywords are recognized:

EMACS: defined if GP is running in an Emacs shell (see Section 2.9).

READL: defined if GP is compiled with readline support (see Section 2.10.1). For instance you could set your prompt in the following portable way:

    \\ self modifying prompt looking like (18:03) gp >

    prompt = "(%R) \e[1mgp\e[m > "
    \\ readline wants non-printing characters to be braced between ^A/^B pairs
    #if READL prompt = "(%R) ^A\e[1m^Bgp^A\e[m^B > "
    \\ escape sequences not supported under emacs
    #if EMACS prompt = "(%R) gp > "

After the preprocessing there remain two types of lines:

• lines of the form default = value, where default is one of the available defaults (see Section 2.1), which will be set to value on actual startup. Don’t forget the quotes around strings (e.g. for prompt

or help).

• lines of the form read "some GP file" where some GP file is a regular GP script this time, which will be read just before GP prompts you for commands, but after initializing the defaults. This is the right place to input files containing alias commands, or your favorite macros.

A sample gprc file called gprc.dft is provided in the standard distribution (in directory lib). It’s a good idea to have a look at it and customize it to your needs.

2.9 Using GP under GNU Emacs.

Thanks to the initial help of Annette Hoffman from the University of Saarbrücken, and David Carlisle from the University of Manchester, it is possible to use GP as a subprocess of GNU Emacs. (Of course, you need GNU Emacs to be installed on your machine!). To use this, you should include in your .emacs file the following lines:

    (autoload 'gp-mode "../../../OpenXM/lib/pari/pari" nil t)
    (autoload 'gp-script-mode "../../../OpenXM/lib/pari/pari" nil t)
    (autoload 'gp "../../../OpenXM/lib/pari/pari" nil t)
    (autoload 'gpman "../../../OpenXM/lib/pari/pari" nil t)
    (setq auto-mode-alist
      (cons '("\\.gp$" .  gp-script-mode) auto-mode-alist))

where pari.el is the name of the file that will have to be loaded by GNU Emacs (if you have changed the name, or if you have the file in a different directory, you must of course supply the correct name). This file is included in the PARI distribution and probably has been installed at the same time as GP.

Once this is done, under GNU Emacs if you type M-x gp (where as usual M is the Meta key, i.e. Escape, or on SUN keyboards, the Left key), a special shell will be started, which in particular launches GP with the default stack size, prime limit and input buffer size. If you type instead C-u M-x gp, you will be asked for the name of the GP executable, the stack size, the prime limit and the input buffer size before the execution of GP begins. If for any of these you simply type return, the default value will be used. On UNIX machines it will be the place you told Configure (usually /usr/local/bin/gp) for the executable, 4000000 for the stack, 500000 for the prime limit and 30000 for the buffer size.

You can then work as usual under GP, but with two notable advantages (which don’t really matter if readline is available to you, see below). First and foremost, you have at your disposal all the facilities of a text editor like Emacs, in particular for correcting or copying blocks. Second, you can have an on-line help which is much more complete than what you obtain by typing ?name. This is done by typing M-?. In the minibuffer, Emacs asks what function you want to describe, and after your reply you obtain the description which is in the users manual, including the description of functions (such as \, %) which use special symbols.

This help system can also be menu-driven, by using the command M-\c which opens a help menu window which enables you to choose the category of commands for which you want an explanation.

Nevertheless, if extended help is available on your system (see Section 2.2.1), you should use it instead of the above, since it’s nicer (it ran through TEX) and understands many more keywords.

Finally you can use command completion in the following way. After the prompt, type the first few letters of the command, then <TAB> where <TAB> is the TAB key. If there exists a unique command starting with the letters you have typed, the command name will be completed. If not, either the list of commands starting with the letters you typed will be displayed in a separate window (which you can then kill by typing as usual C-x 1 or by typing in more letters), or “no match found” will be displayed in the Emacs command line. If your GP was linked with the readline library, read the section on completion in the section below (the paragraph on online help is not relevant).

Note that if for some reason the session crashes (due to a bug in your program or in the PARI system), you will usually stay under Emacs, but the GP buffer will be killed. To recover it, simply type again M-x gp (or C-u M-x gp), and a new session of GP will be started after the old one, so you can recover what you have typed. Note that this will of course not work if for some reason you exited Emacs before coming back (except for the C-z temporary stopping command).

You also have at your disposal a few other commands and many possible customizations (colours, prompt). Read the file emacs/pariemacs.txt in standard distribution for details.

2.10 Using GP with readline.

Thanks to the initial help of Ilya Zakharevich, there is a possibility of line editing and command name completion outside of an Emacs buffer if you have compiled GP with the GNU readline library. If you don’t have Emacs available, or can’t stand using it, we really advise you to make sure you get this very useful library before configuring or compiling GP. In fact, with readline, even line editing becomes more powerful outside an Emacs buffer!

2.10.1 A (too) short introduction to readline: The basics are as follows (read the readline user manual !), assume that C- stands for “the Control key combined with another” and the same for M- with the Meta key (generally C- combinations act on characters, while the M- ones operate on words). The Meta key might be called Alt on some keyboards, will display a black diamond on most others, and can safely be replaced by Esc in any case. Typing any ordinary key inserts text where the cursor stands, the arrow keys enabling you to move in the line. There are many more movement commands, which will be familiar to the Emacs user, for instance C-a/C-e will take you to the start/end of the line, M-b/M-f move the cursor backward/forward by a word, etc. Just press the Return key at any point to send your command to GP.

All the commands you type in are stored in a history (with multiline commands being saved as single concatenated lines). The Up and Down arrows (or C-p/C-n) will move you through it, M-</M- > sending you to the start/end of the history. C-r/C-s will start an incremental backward/forward search. You can kill text (C-k kills till the end of line, M-d to the end of current word) which you can then yank back using the C-y key (M-y will rotate the kill-ring). C- will undo your last changes incrementally (M-r undoes all changes made to the current line). C-t and M-t will transpose the character (word) preceding the cursor and the one under the cursor.

Keeping the M- key down while you enter an integer (a minus sign meaning reverse behaviour) gives an argument to your next readline command (for instance M-- C-k will kill text back to the start of line). If you prefer Vi–style editing, M-C-j will toggle you to Vi mode.

Of course you can change all these default bindings. For that you need to create a file named .inputrc in your home directory. For instance (notice the embedding conditional in case you would want specific bindings for GP):

    $if Pari-GP
      set show-all-if-ambiguous
      "\C-h":  backward-delete-char
      "\e\C-h":  backward-kill-word
      "\C-xd":  dump-functions
      (:  "\C-v()\C-b"     # can be annoying when copy-pasting !
      [:  "\C-v[]\C-b"
    $endif

C-x C-r will re-read this init file, incorporating any changes made to it during the current session.

Note: By default, ( and [ are bound to the function pari-matched-insert which, if “electric parentheses” are enabled (default: off) will automatically insert the matching closure (respectively ) and ]). This behaviour can be toggled on and off by giving the numeric argument −2 to ( (M--2(), which is useful if you want, e.g to copy-paste some text into the calculator. If you don’t want a toggle, you can use M--0 / M--1 to specifically switch it on or off).

Note: In recent versions of readline (2.1 for instance), the Alt or Meta key can give funny results (output 8-bit accented characters for instance). If you don’t want to fall back to the Esc combination, put the following two lines in your .inputrc:

    set convert-meta on
    set output-meta off

2.10.2 Command completion and online help

As in the Emacs shell, <TAB> will complete words for you. But, under readline, this mechanism will be context-dependent: GP will strive to only give you meaningful completions in a given context (it will fail sometimes, but only under rare and restricted conditions).

For instance, shortly after a ~, we expect a user name, then a path to some file. Directly after default( has been typed, we would expect one of the default keywords. After whatnow( , we expect the name of an old function, which may well have disappeared from this version. After a ’.’, we expect a member keyword. And generally of course, we expect any GP symbol which may be found in the hashing lists: functions (both yours and GP’s), and variables.

If, at any time, only one completion is meaningful, GP will provide it together with

• an ending comma if we’re completing a default,

• a pair of parentheses if we’re completing a function name. In that case hitting <TAB> again will provide the argument list as given by the online help*.

* recall that you can always undo the effect of the preceding keys by hitting C-

Otherwise, hitting <TAB> once more will give you the list of possible completions. Just experiment with this mechanism as often as possible, you’ll probably find it very convenient. For instance,

you can obtain default(seriesprecision,10), just by hitting def<TAB>se<TAB>10, which saves

18 keystrokes (out of 27).

Hitting M-h will give you the usual short online help concerning the word directly beneath the cursor, M-H will yield the extended help corresponding to the help default program (usually opens a dvi previewer, or runs a primitive tex-to-ASCII program). None of these disturb the line you were editing.

44

Functions and Operations Available in PARI and GP

The functions and operators available in PARI and in the GP/PARI calculator are numerous and everexpanding. Here is a description of the ones available in version 2.0.17. It should be noted that many of these functions accept quite different types as arguments, but others are more restricted. The list of acceptable types will be given for each function or class of functions. Except when stated otherwise, it is understood that a function or operation which should make natural sense is legal. In this chapter, we will describe the functions according to a rough classification. For the functions in alphabetical order, see the general index. The general entry looks something like:

foo(x, {flag = 0}): short description.

The library syntax is foo(x, flag).

This means that the GP function foo has one mandatory argument x, and an optional one, flag, whose default value is 0 (the {} should never be typed, it is just a convenient notation we will use throughout to denote optional arguments). That is, you can type foo(x,2), or foo(x), which is then understood to mean foo(x,0). As well, a comma or closing parenthesis, where an optional argument should have been, signals to GP it should use the default. Thus, the syntax foo(x,) is also accepted as a synonym for our last expression. When a function has more than one optional argument, the argument list is filled with user supplied values, in order. And when none are left, the defaults are used instead. Thus, assuming that foo’s prototype had been

foo({x = 1}, {y = 2}, {z = 3}),

typing in foo(6,4) would give you foo(6,4,3). In the rare case when you want to set some far away flag, and leave the defaults in between as they stand, you can use the “empty arg” trick alluded to above: foo(6,,1) would yield foo(6,2,1). By the way, foo() by itself yields foo(1,2,3) as was to be expected. In this rather special case of a function having no mandatory argument, you can even omit the (): a standalone foo would be enough (though we don’t really recommend it for your scripts, for the sake of clarity). In defining GP syntax, we strove to put optional arguments at the end of the argument list (of course, since they would not make sense otherwise), and in order of decreasing usefulness so that, most of the time, you will be able to ignore them.

For some of these optional flags, we adopted the customary binary notation as a compact way to represent many toggles with just one number. Letting (p_{0}, …, p_{n}) be a list of switches (i.e. of properties which can be assumed to take either the value 0 or 1), the number 2^{3} + 2^{5} = 40 means that p_{3} and p_{5} have been set (that is, set to 1), and none of the others were (that is, they were set to 0). This will usually be announced as “The binary digits of flag mean 1: p_{0}, 2: p_{1}, 4: p_{2}”, and so on, using the available consecutive powers of 2.

To finish with our generic simple-minded example, the library function foo, as defined above, is seen to have two mandatory arguments, x and flag(no PARI mathematical function has been implemented so as to accept a variable number of arguments). When not mentioned otherwise, the result and arguments of a function are assumed implicitly to be of type GEN. Most other functions return an object of type long integer in C (see Chapter 4). The variable or parameter names prec and flag always denote long integers.

Pointers. If a parameter in the function prototype is prefixed with a & sign, as in

foo(x, &e)

it means that, besides the normal return value, the variable named e may be set as a side effect. When passing the argument, the & sign has to be typed in explicitly. As of version 2.0.17, this pointer argument is optional for all documented functions, hence the & will always appear between brackets as in issquare(x, {&e}).

3.1 Standard monadic or dyadic operators.

3.1.1 +/-: The expressions +x and -x refer to monadic operators (the first does nothing, the second negates x).

The library syntax is gneg(x) for -x.

3.1.2 +, -: The expression x + y is the sum and x - y is the difference of x and y. Among the prominent impossibilities are addition/subtraction between a scalar type and a vector or a matrix, between vector/matrices of incompatible sizes and between an integermod and a real number.

The library syntax is gadd(x, y) x + y, gsub(x, y) for x - y.

3.1.3 *: The expression x * y is the product of x and y. Among the prominent impossibilities are multiplication between vector/matrices of incompatible sizes, between an integermod and a real number. Note that because of vector and matrix operations, * is not necessarily commutative. Note also that since multiplication between two column or two row vectors is not allowed, to obtain the scalar product of two vectors of the same length, you must multiply a line vector by a column vector, if necessary by transposing one of the vectors (using the operator ~ or the function mattranspose, see Section 3.8).

If x and y are binary quadratic forms, compose them. See also qfbnucomp and qfbnupow.

The library syntax is gmul(x, y) for x * y. Also available is gsqr(x) for x * x (faster of course!).

3.1.4 /: The expression x / y is the quotient of x and y. In addition to the impossibilities for multiplication, note that if the divisor is a matrix, it must be an invertible square matrix, and in that case the result is x∗y^{−1}. Furthermore note that the result is as exact as possible: in particular, division of two integers always gives a rational number (which may be an integer if the quotient is exact) and not the Euclidean quotient (see x \ y for that), and similarly the quotient of two polynomials is a rational function in general. To obtain the approximate real value of the quotient of two integers, add 0. to the result; to obtain the approximate p-adic value of the quotient of two integers, add O(p^k) to the result; finally, to obtain the Taylor series expansion of the quotient of two polynomials, add O(X^k) to the result or use the taylor function (see Section 3.7.31).

The library syntax is gdiv(x, y)for x / y.

3.1.5 \: The expression x \ y is the Euclidean quotient of x and y. The types must be either both integer or both polynomials. The result is the Euclidean quotient. In the case of integer division, the quotient is such that the corresponding remainder is non-negative.

The library syntax is gdivent(x, y) for x \ y.

3.1.6 \/: The expression x \/ y is the Euclidean quotient of x and y. The types must be either both integer or both polynomials. The result is the rounded Euclidean quotient. In the case of integer division, the quotient is such that the corresponding remainder is smallest in absolute value and in case of a tie the quotient closest to +∞ is chosen.

The library syntax is gdivround(x, y) for x \/ y.

3.1.7 %: The expression x % y is the Euclidean remainder of x and y. The modulus y must be of type integer or polynomial. The result is the remainder, always non-negative in the case of integers. Allowed dividend types are scalar exact types when the modulus is an integer, and polynomials, polmods and rational functions when the modulus is a polynomial.

The library syntax is gmod(x, y) for x % y.

3.1.8 divrem(x, y): creates a column vector with two components, the first being the Euclidean quotient, the second the Euclidean remainder, of the division of x by y. This avoids the need to do two divisions if one needs both the quotient and the remainder. The arguments must be both integers or both polynomials; in the case of integers, the remainder is non-negative.

The library syntax is gdiventres(x, y).

3.1.9 ^: The expression x^y is powering. If the exponent is an integer, then exact operations are performed using binary (left-shift) powering techniques. In particular, in this case the first argument cannot be a vector or matrix unless it is a square matrix (and moreover invertible if the exponent is negative). If the exponent is not of type integer, this is treated as a transcendental function (see Section 3.3), and in particular has the effect of componentwise powering on vector or matrices.

The library syntax is gpow(x, y, prec) for x^y.

3.1.10 shift(x, n) or x << n (= x >> (−n)): shifts x componentwise left by n bits if n ≥ 0 and right by |n| bits if n < 0. A left shift by n corresponds to multiplication by 2^{n}. A right shift of an integer x by |n| corresponds to a Euclidean division of x by 2^{|n|} with a remainder of the same sign as x, hence is not the same (in general) as x\2^{n}.

The library syntax is gshift(x, n) where n is a long.

3.1.11 shiftmul(x, n): multiplies x by 2^{n}. The difference with shift is that when n < 0, ordinary division takes place, hence for example if x is an integer the result may be a fraction, while for shift Euclidean division takes place when n < 0 hence if x is an integer the result is still an integer.

The library syntax is gmul2n(x, n) where n is a long.

3.1.12 Comparison and boolean operators. The six standard comparison operators <=, <, >=, >, ==, != are available in GP, and in library mode under the names gle, glt, gge, ggt, geq, gne respectively. The library syntax is co(x, y), where co is the comparison operator. The result is 1 (as a GEN) if the comparison is true, 0 (as a GEN) if it is false.

The standard boolean functions || (inclusive or), && (and) and ! (not) are also available, and the library syntax is gor(x, y), gand(x, y) and gnot(x) respectively.

In library mode, it is in fact usually preferable to use the two basic functions which are gcmp(x, y) which gives the sign (1, 0, or -1) of x − y, where x and y must be in R, and gegal(x, y) which can be applied to any two PARI objects x and y and gives 1 (i.e. true) if they are equal (but not necessarily identical), 0 (i.e. false) otherwise. Particular cases of gegal which should be used are gcmp0(x) (x == 0 ?), gcmp1(x) (x == 1 ?), and gcmp 1(x) (x == −1 ?).

Note that gcmp0(x) tests whether x is equal to zero, even if x is not an exact object. To test whether x is an exact object which is equal to zero, one must use isexactzero.

Also note that the gcmp and gegal functions return a C-integer, and not a GEN like gle etc.

GP accepts the following synonyms for some of the above functions: since there is no bitwise and or bitwise or, | and & are accepted as synonyms of || and && respectively. Also, <> is accepted as a synonym for !=. On the other hand, = is definitely not a synonym for == since it is the assignment statement.

3.1.13 lex(x, y): gives the result of a lexicographic comparison between x and y. This is to be interpreted in quite a wide sense. For example, the vector [1, 3] will be considered smaller than the longer vector [1, 3, −1] (but of course larger than [1, 2, 5]), i.e. lex([1,3], [1,3,-1]) will return −1.

The library syntax is lexcmp(x, y).

3.1.14 sign(x): sign (0, 1 or −1) of x, which must be of type integer, real or fraction.

The library syntax is gsigne(x). The result is a long.

3.1.15 max(x, y) and min(x, y): creates the maximum and minimum of x and y when they can be compared.

The library syntax is gmax(x, y) and gmin(x, y).

3.1.16 vecmax(x): if x is a vector or a matrix, returns the maximum of the elements of x, otherwise returns a copy of x. Returns −∞ in the form of −(2^{31} − 1) (or −(2^{63} − 1) for 64-bit machines) if x is empty.

The library syntax is vecmax(x).

3.1.17 vecmin(x): if x is a vector or a matrix, returns the minimum of the elements of x, otherwise returns a copy of x. Returns +∞ in the form of 2^{31} − 1 (or 2^{63} − 1 for 64-bit machines) if x is empty.

The library syntax is vecmin(x).

3.2 Conversions and similar elementary functions or commands.

Many of the conversion functions are rounding or truncating operations. In this case, if the argument is a rational function, the result is the Euclidean quotient of the numerator by the denominator, and if the argument is a vector or a matrix, the operation is done componentwise. This will not be restated for every function.

3.2.1 List(x = [ ]): transforms a (row or column) vector x into a list. The only other way to create a t LIST is to use the function listcreate.

This is useless in library mode.

3.2.2 Mat(x = [ ]): transforms the object x into a matrix. If x is not a vector or a matrix, this creates a 1 × 1 matrix. If x is a row (resp. column) vector, this creates a 1-row (resp. 1-column) matrix. If x is already a matrix, a copy of x is created.

This function can be useful in connection with the function concat (see there).

The library syntax is gtomat(x).

3.2.3 Mod(x, y, {flag = 0}): creates the PARI object (x mod y), i.e. an integermod or a polmod. y must be an integer or a polynomial. If y is an integer, x must be an integer. If y is a polynomial, x must be a scalar or a polynomial. The result is put on the PARI stack.

This function is not the same as x % y, the result of which is an integer or a polynomial.

If flag is equal to 1, the modulus of the created result is put on the heap and not on the stack, and hence becomes a permanent copy which cannot be erased later by garbage collecting (see Section 4.4). In particular, care should be taken to avoid creating too many such objects, since the heap is very small (typically a few thousand objects at most).

The library syntax is Mod0(x, y, flag). Also available are

• for flag = 1: gmodulo(x, y).

• for flag = 0: gmodulcp(x, y).

3.2.4 Pol(x, {v = x}): transforms the object x into a polynomial with main variable v. If x is a scalar, this gives a constant polynomial. If x is a power series, the effect is identical to truncate (see there), i.e. it chops off the O(X^{k}). If x is a vector, this function creates the polynomial whose coefficients are given in x, with x[1] being the leading coefficient (which can be zero).

Warning: this is not a substitution function. It is intended to be quick and dirty. So if you try Pol(a,y) on the polynomial a = x+y, you will get y+y, which is not a valid PARI object.

The library syntax is gtopoly(x, v), where v is a variable number.

3.2.5 Polrev(x, {v = x}): transform the object x into a polynomial with main variable v. If x is a scalar, this gives a constant polynomial. If x is a power series, the effect is identical to truncate (see there), i.e. it chops off the O(X^{k}). If x is a vector, this function creates the polynomial whose coefficients are given in x, with x[1] being the constant term. Note that this is the reverse of Pol if x is a vector, otherwise it is identical to Pol.

The library syntax is gtopolyrev(x, v), where v is a variable number.

3.2.6 Ser(x, {v = x}): transforms the object x into a power series with main variable v (x by default). If x is a scalar, this gives a constant power series with precision given by the default serieslength (corresponding to the C global variable precdl). If x is a polynomial, the precision is the greatest of precdl and the degree of the polynomial. If x is a vector, the precision is similarly given, and the coefficients of the vector are understood to be the coefficients of the power series starting from the constant term (i.e. the reverse of the function Pol).

The warning given for Pol applies here: this is not a substitution function.

The library syntax is gtoser(x, v), where v is a variable number (i.e. a C integer).

3.2.7 Set({x = [ ]}): converts x into a set, i.e. into a row vector with strictly increasing entries. x can be of any type, but is most useful when x is already a vector. The components of x are put in canonical form (type t STR) so as to be easily sorted. To recover an ordinary GEN from such an element, you can apply eval to it.

The library syntax is gtoset(x).

3.2.8 Str({x = ""}, {flag = 0}): converts x into a character string (type t STR, the empty string if x is omitted). To recover an ordinary GEN from a string, apply eval to it. The arguments of Str are evaluated in string context (see Section 2.6.5). If flag is set, treat x as a filename and perform environment expansion on the string. This feature can be used to read environment variable values.

    ?  i = 1; Str("x" i)
    %1 = "x1"
    ?  eval(%)
    %2 = x1;
    ?  Str("$HOME", 1)
    %2 = "/home/pari"

The library syntax is strtoGENstr(x, flag). This function is mostly useless in library mode. Use the pair strtoGEN/GENtostr to convert between char* and GEN.

3.2.9 Vec(x = [ ]): transforms the object x into a row vector. The vector will be with one component only, except when x is a vector/matrix or a quadratic form (in which case the resulting vector is simply the initial object considered as a row vector), but more importantly when x is a polynomial or a power series. In the case of a polynomial, the coefficients of the vector start with the leading coefficient of the polynomial, while for power series only the significant coefficients are taken into account, but this time by increasing order of degree.

The library syntax is gtovec(x).

3.2.10 binary(x): outputs the vector of the binary digits of |x|. Here x can be an integer, a real number (in which case the result has two components, one for the integer part, one for the fractional part) or a vector/matrix.

The library syntax is binaire(x).

3.2.11 bittest(x, n): outputs the n^{th} bit of |x| starting from the right (i.e. the coefficient of 2^{n} in the binary expansion of x). The result is 0 or 1. To extract several bits at once as a vector, pass a vector for n.

The library syntax is bittest(x, n), where n and the result are longs.

3.2.12 ceil(x): ceiling of x. When x is in R, the result is the smallest integer greater than or equal to x.

The library syntax is gceil(x).

3.2.13 centerlift(x, {v}): lifts an element x = a mod n of Z/nZ to a in Z, and similarly lifts a polmod to a polynomial. This is the same as lift except that in the particular case of elements of Z/nZ, the lift y is such that −n/2 < y ≤ n/2. If x is of type fraction, complex, quadratic, polynomial, power series, rational function, vector or matrix, the lift is done for each coefficient. Real and p-adics are forbidden.

The library syntax is centerlift0(x, v), where v is a long and an omitted v is coded as −1. Also available is centerlift(x) = centerlift0(x,-1).

3.2.14 changevar(x, y): creates a copy of the object x where its variables are modified according to the permutation specified by the vector y. For example, assume that the variables have been introduced in the order x, a, b, c. Then, if y is the vector [x,c,a,b], the variable a will be replaced by c, b by a, and c by b, x being unchanged. Note that the permutation must be completely specified, e.g. [c,a,b] would not work, since this would replace x by c, and leave a and b unchanged (as well as c which is the fourth variable of the initial list). In particular, the new variable names must be distinct.

The library syntax is changevar(x, y).

3.2.15 components of a PARI object:

There are essentially three ways to extract the components from a PARI object.

The first and most general, is the function component(x, n) which extracts the n^{th}-component of x. This is to be understood as follows: every PARI type has one or two initial code words. The components are counted, starting at 1, after these code words. In particular if x is a vector, this is indeed the n^{th}-component of x, if x is a matrix, the n^{th} column, if x is a polynomial, the n^{th} coefficient (i.e. of degree n − 1), and for power series, the n^{th} significant coefficient. The use of the function component implies the knowledge of the structure of the different PARI types, which can be recalled by typing \t under GP.

The library syntax is compo(x, n), where n is a long.

The two other methods are more natural but more restricted. First, the function polcoeff(x, n) gives the coefficient of degree n of the polynomial or power series x, with respect to the main variable of x (to see the order of the variables or to change it, use the function reorder, see Section 3.11.2.22). In particular if n is less than the valuation of x or in the case of a polynomial, greater than the degree, the result is zero (contrary to compo which would send an error message). If x is a power series and n is greater than the largest significant degree, then an error message is issued.

For greater flexibility, vector or matrix types are also accepted for x, and the meaning is then identical with that of compo.

Finally note that a scalar type is considered by polcoeff as a polynomial of degree zero.

The library syntax is truecoeff(x, n).

The third method is specific to vectors or matrices under GP. If x is a (row or column) vector, then x[n] represents the n^{th} component of x, i.e. compo(x,n). It is more natural and shorter to write. If x is a matrix, x[m,n] represents the coefficient of row m and column n of the matrix, x[m,] represents the m^{th} row of x, and x[,n] represents the n^{th} column of x.

Finally note that in library mode, the macros coeff and mael are available to deal with the non-recursivity of the GEN type from the compiler’s point of view. See the discussion on typecasts in Chapter 4.

3.2.16 conj(x): conjugate of x. The meaning of this is clear, except that for real quadratic numbers, it means conjugation in the real quadratic field. This function has no effect on integers, reals, integermods, fractions or p-adics. The only forbidden type is polmod (see conjvec for this).

The library syntax is gconj(x).

3.2.17 conjvec(x): conjugate vector representation of x. If x is a polmod, equal to Mod(a, q), this gives a vector of length degree(q) containing the complex embeddings of the polmod if q has integral or rational coefficients, and the conjugates of the polmod if q has some integermod coefficients. The order is the same as that of the polroots functions. If x is an integer or a rational number, the result is x. If x is a (row or column) vector, the result is a matrix whose columns are the conjugate vectors of the individual elements of x.

The library syntax is conjvec(x, prec).

3.2.18 denominator(x): lowest denominator of x. The meaning of this is clear when x is a rational number or function. When x is an integer or a polynomial, the result is equal to 1. When x is a vector or a matrix, the lowest common denominator of the components of x is computed. All other types are forbidden.

The library syntax is denom(x).

3.2.19 floor(x): floor of x. When x is in R, the result is the largest integer smaller than or equal to x.

The library syntax is gfloor(x).

3.2.20 frac(x): fractional part of x. Identical to x − floor(x). If x is real, the result is in [0, 1[.

The library syntax is gfrac(x).

3.2.21 imag(x): imaginary part of x. When x is a quadratic number, this is the coefficient of ω in the “canonical” integral basis (1, ω).

The library syntax is gimag(x).

3.2.22 length(x): number of non-code words in x really used (i.e. the effective length minus 2 for integers and polynomials). In particular, the degree of a polynomial is equal to its length minus 1. If x has type t STR, output number of letters.

The library syntax is glength(x) and the result is a C long.

3.2.23 lift(x, {v}): lifts an element x = a mod n of Z/nZ to a in Z, and similarly lifts a polmod to a polynomial if v is omitted. Otherwise, lifts only polmods with main variable v (if v does not occur in x, lifts only intmods). If x is of type fraction, complex, quadratic, polynomial, power series, rational function, vector or matrix, the lift is done for each coefficient. Forbidden types for x are reals and p-adics.

The library syntax is lift0(x, v), where v is a long and an omitted v is coded as −1. Also available is lift(x) = lift0(x,-1).

3.2.24 norm(x): algebraic norm of x, i.e. the product of x with its conjugate (no square roots are taken), or conjugates for polmods. For vectors and matrices, the norm is taken componentwise and hence is not the L^{2}-norm (see norml2). Note that the norm of an element of R is its square, so as to be compatible with the complex norm.

The library syntax is gnorm(x).

3.2.25 norml2(x): square of the L^{2}-norm of x. x must be a (row or column) vector.

The library syntax is gnorml2(x).

3.2.26 numerator(x): numerator of x. When x is a rational number or function, the meaning is clear. When x is an integer or a polynomial, the result is x itself. When x is a vector or a matrix, then numerator(x) is defined to be denominator(x)*x. All other types are forbidden.

The library syntax is numer(x).

3.2.27 numtoperm(n, k): generates the k-th permutation (as a row vector of length n) of the numbers 1 to n. The number k is taken modulo n! , i.e. inverse function of permtonum.

The library syntax is permute(n, k), where n is a long.

3.2.28 padicprec(x, p): absolute p-adic precision of the object x. This is the minimum precision of the components of x. The result is VERYBIGINT (2^{31} − 1 for 32-bit machines or 2^{63} − 1 for 64-bit machines) if x is an exact object.

The library syntax is padicprec(x, p) and the result is a long integer.

3.2.29 permtonum(x): given a permutation x on n elements, gives the number k such that x = numtoperm(n, k), i.e. inverse function of numtoperm.

The library syntax is permuteInv(x).

3.2.30 precision(x, {n}): gives the precision in decimal digits of the PARI object x. If x is an exact object, the largest single precision integer is returned. If n is not omitted, creates a new object equal to x with a new precision n. This is to be understood as follows:

For exact types, no change. For x a vector or a matrix, the operation is done componentwise.

For real x, n is the number of desired significant decimal digits. If n is smaller than the precision of x, x is truncated, otherwise x is extended with zeros.

For x a p-adic or a power series, n is the desired number of significant p-adic or X-adic digits, where X is the main variable of x.

Note that the function precision never changes the type of the result. In particular it is not possible to use it to obtain a polynomial from a power series. For that, see truncate.

The library syntax is precision0(x, n), where n is a long. Also available are ggprecision(x) (result is a GEN) and gprec(x, n), where n is a long.

3.2.31 random({N = 2^{31}}): gives a random integer between 0 and N − 1. N can be arbitrary large. This is an internal PARI function and does not depend on the system’s random number generator. Note that the resulting integer is obtained by means of linear congruences and will not be well distributed in arithmetic progressions.

The library syntax is genrand(N).

3.2.32 real(x): real part of x. In the case where x is a quadratic number, this is the coefficient of 1 in the “canonical” integral basis (1, ω).

The library syntax is greal(x).

3.2.33 round(x, {&e}): If x is in R, rounds x to the nearest integer and set e to the number of error bits, that is the binary exponent of the difference between the original and the rounded value (the “fractional part”). If the exponent of x is too large compared to its precision (i.e. e > 0), the result is undefined and an error occurs if e was not given.

Important remark: note that, contrary to the other truncation functions, this function operates on every coefficient at every level of a PARI object. For example

truncate (\frac{2.4 ∗ X^{2} − 1.7}{X}) = 2.4 ∗ X,

whereas

round (\frac{2.4 ∗ X^{2} − 1.7}{X}) = \frac{2 ∗ X^{2} − 2}{X}.

An important use of round is to get exact results after a long approximate computation, when theory tells you that the coefficients must be integers.

The library syntax is grndtoi(x, &e), where e is a long integer. Also available is ground(x).

3.2.34 simplify(x): this function tries to simplify the object x as much as it can. The simplifications do not concern rational functions (which PARI automatically tries to simplify), but type changes. Specifically, a complex or quadratic number whose imaginary part is exactly equal to 0 (i.e. not a real zero) is converted to its real part, and a polynomial of degree zero is converted to its constant term. For all types, this of course occurs recursively. This function is useful in any case, but in particular before the use of arithmetic functions which expect integer arguments, and not for example a complex number of 0 imaginary part and integer real part (which is however printed as an integer).

The library syntax is simplify(x).

3.2.35 sizebyte(x): outputs the total number of bytes occupied by the tree representing the PARI object x.

The library syntax is taille2(x) which returns a long. The function taille returns the number of words instead.

3.2.36 sizedigit(x): outputs a quick bound for the number of decimal digits of (the components of) x, off by at most 1. If you want the exact value, you an use length(Str(x)), which is much slower.

The library syntax is gsize(x) which returns a long.

3.2.37 truncate(x, {&e}): truncate x and set e to the number of error bits. When x is in R, this means that the part after the decimal point is chopped away, integer and set e to the number of error bits that is the binary exponent of the difference between the original and the truncated value (the “fractional part”). If the exponent of x is too large compared to its precision (i.e. e > 0), the result is undefined and an error occurs if e was not given.

Note a very special use of truncate: when applied to a power series, it transforms it into a polynomial or a rational function with denominator a power of X, by chopping away the O(X^{k}). Similarly, when applied to a p-adic number, it transforms it into an integer or a rational number by chopping away the O(p^{k}).

The library syntax is gcvtoi(x, &e), where e is a long integer. Also available is gtrunc(x).

3.2.38 valuation(x, p): computes the highest exponent of p dividing x. If p is of type integer, x must be an integer, an integermod whose modulus is divisible by p, a fraction, a q-adic number with q = p, or a polynomial or power series in which case the valuation is the minimum of the valuation of the coefficients.

If p is of type polynomial, x must be of type polynomial or rational function, and also a power series if x is a monomial. Finally, the valuation of a vector, complex or quadratic number is the minimum of the component valuations.

If x = 0, the result is VERYBIGINT (2^{31} − 1 for 32-bit machines or 2^{63} − 1 for 64-bit machines) if x is an exact object. If x is a p-adic numbers or power series, the result is the exponent of the zero. Any other type combinations gives an error.

The library syntax is ggval(x, p), and the result is a long.

3.2.39 variable(x): gives the main variable of the object x, and p if x is a p-adic number. Gives an error if x has no variable associated to it. Note that this function is useful only in GP, since in library mode the function gvar is more appropriate.

The library syntax is gpolvar(x). However, in library mode, this function should not be used. Instead, test whether x is a p-adic (type t PADIC), in which case p is in x[2], or call the function gvar(x) which returns the variable number of x if it exists, BIGINT otherwise.

3.3 Transcendental functions.

As a general rule, which of course in some cases may have exceptions, transcendental functions operate in the following way:

• If the argument is either an integer, a real, a rational, a complex or a quadratic number, it is, if necessary, first converted to a real (or complex) number using the current precision held in the default realprecision. Note that only exact arguments are converted, while inexact arguments such as reals are not.

Under GP this is transparent to the user, but when programming in library mode, care must be taken to supply a meaningful parameter prec as the last argument of the function if the first argument is an exact object. This parameter is ignored if the argument is inexact.

Note that in library mode the precision argument prec is a word count including codewords, i.e. represents the length in words of a real number, while under GP the precision (which is changed by the metacommand \p or using default(realprecision,...)) is the number of significant decimal digits.

Note that some accuracies attainable on 32-bit machines cannot be attained on 64-bit machines for parity reasons. For example the default GP accuracy is 28 decimal digits on 32-bit machines, corresponding to prec having the value 5, but this cannot be attained on 64-bit machines.

After possible conversion, the function is computed. Note that even if the argument is real, the result may be complex (e.g. acos(2.0) or acosh(0.0)). Note also that the principal branch is always chosen.

• If the argument is an integermod or a p-adic, at present only a few functions like sqrt (square root), sqr (square), log, exp, powering, teichmuller (Teichmüller character) and agm (arithmetic-geometric mean) are implemented. Note that in the case of a 2-adic number, sqr(x) is not identical to x∗x: for example if x = 1+O(2^{5}) then x∗x = 1+O(2^{5}) while sqr(x) = 1+O(2^{6}). (Remark: note that if we wanted to be strictly consistent with the PARI philosophy, we should have x ∗ y = (4 mod 8) when both x and y are congruent to 2 modulo 4, or sqr(x) = (4 mod 32) when x is congruent to 2 modulo 4. However, since an integermod is an exact object, PARI assumes that the modulus must not change, and the result is hence 0 mod 4 in both cases. On the other hand, p-adics are not exact objects, hence are treated differently.)

• If the argument is a polynomial, power series or rational function, it is, if necessary, first converted to a power series using the current precision held in the variable precdl. Under GP this again is transparent to the user. When programming in library mode, however, the global variable precdl must be set before calling the function if the argument has an exact type (i.e. not a power series). Here precdl is not an argument of the function, but a global variable.

Then the Taylor series expansion of the function around X = 0 (where X is the main variable) is computed to a number of terms depending on the number of terms of the argument and the function being computed.

• If the argument is a vector or a matrix, the result is the componentwise evaluation of the function. In particular, transcendental functions on square matrices, which are not implemented in the present version 2.0.17 (see Appendix B however), will have a slightly different name if they are implemented some day.

3.3.1 ^: If y is not of type integer, x^y has the same effect as exp(y*ln(x)). It can be applied to p-adic numbers as well as to the more usual types.

The library syntax is gpow(x, y, prec).

3.3.2 Euler: Euler’s constant 0.57721 · · ·. Note that Euler is one of the few special reserved names which cannot be used for variables (the others are I and Pi, as well as all function names).

The library syntax is mpeuler(prec) where prec must be given. Note that this creates γ on the PARI stack. If one does not want to create it on the stack but stash it for later use under the global name geuler (with no parentheses), use instead consteuler(prec).

3.3.3 I: the complex number \sqrt{−1}.

The library syntax is the global variable gi (of type GEN).

3.3.4 Pi: the constant π (3.14159 ⋯).

The library syntax is mppi(prec) where prec must be given. Note that this creates π on the PARI stack. If one does not want to create it on the stack but stash it for later use under the global name gpi (with no parentheses), use instead constpi(prec).

3.3.5 abs(x): absolute value of x (modulus if x is complex). Polynomials, power series and rational functions are not allowed. Contrary to most transcendental functions, an integer is not converted to a real number before applying abs.

The library syntax is gabs(x, prec).

3.3.6 acos(x): principal branch of cos^{−1}(x), i.e. such that Re(acos(x)) ∈ [0, π]. If x ∈ R and |x| > 1, then acos(x) is complex.

The library syntax is gacos(x, prec).

3.3.7 acosh(x): principal branch of cosh^{−1}(x), i.e. such that Im(acosh(x)) ∈ [0, π]. If x ∈ R and x < 1, then acosh(x) is complex.

The library syntax is gach(x, prec).

3.3.8 agm(x, y): arithmetic-geometric mean of x and y. In the case of complex or negative numbers, the principal square root is always chosen. p-adic or power series arguments are also allowed. Note that a p-adic agm exists only if x/y is congruent to 1 modulo p (modulo 16 for p = 2). x and y cannot both be vectors or matrices.

The library syntax is agm(x, y, prec).

3.3.9 arg(x): argument of the complex number x, such that −π < arg(x) ≤ π.

The library syntax is garg(x, prec).

3.3.10 asin(x): principal branch of sin^{−1}(x), i.e. such that Re(asin(x)) ∈ [−π/2, π/2]. If x ∈ R and |x| > 1 then asin(x) is complex.

The library syntax is gasin(x, prec).

3.3.11 asinh(x): principal branch of sinh^{−1}(x), i.e. such that Im(asinh(x)) ∈ [−π/2, π/2].

The library syntax is gash(x, prec).

3.3.12 atan(x): principal branch of tan^{−1}(x), i.e. such that Re(atan(x)) ∈ ] − π/2, π/2[.

The library syntax is gatan(x, prec).

3.3.13 atanh(x): principal branch of tanh^{−1}(x), i.e. such that Im(atanh(x)) ∈ ] − π/2, π/2]. If x ∈ R and |x| > 1 then atanh(x) is complex.

The library syntax is gath(x, prec).

3.3.14 bernfrac(x): Bernoulli number B_{x}, where B_{0} = 1, B_{1} = −1/2, B_{2} = 1/6,…, expressed as a rational number. The argument x should be of type integer.

The library syntax is bernfrac(x).

3.3.15 bernreal(x): Bernoulli number B_{x}, as bernfrac, but B_{x} is returned as a real number (with the current precision).

The library syntax is bernreal(x, prec).

3.3.16 bernvec(x): creates a vector containing, as rational numbers, the Bernoulli numbers B_{0}, B_{2},…, B_{2x}. These Bernoulli numbers can then be used as follows. Assume that this vector has been put into a variable, say bernint. Then you can define under GP:

    bern(x) =
    {
      if (x==1, return(-1/2));
      if ((x<0) || (x%2), return(0));
      bernint[x/2+1]
    }

and then bern(k) gives the Bernoulli number of index k as a rational number, exactly as bernreal(k) gives it as a real number. If you need only a few values, calling bernfrac(k) each time will be much more efficient than computing the huge vector above.

The library syntax is bernvec(x).

3.3.17 besseljh(n, x): J-Bessel function of half integral index. More precisely, besseljh(n, x) computes J_{n+1/2}(x) where n must be of type integer, and x is any element of C. In the present version 2.0.17, this function is not very accurate when x is small.

The library syntax is jbesselh(n, x, prec).

3.3.18 besselk(nu, x, {flag = 0}): K-Bessel function of index nu (which can be complex) and argument x. Only real and positive arguments x are allowed in the present version 2.0.17. If flag is equal to 1, uses another implementation of this function which is often faster.

The library syntax is kbessel(nu, x, prec) and kbessel2(nu, x, prec) respectively.

3.3.19 cos(x): cosine of x.

The library syntax is gcos(x, prec).

3.3.20 cosh(x): hyperbolic cosine of x.

The library syntax is gch(x, prec).

3.3.21 cotan(x): cotangent of x.

The library syntax is gcotan(x, prec).

3.3.22 dilog(x): principal branch of the dilogarithm of x, i.e. analytic continuation of the power series log_{2}(x) = ∑_{n≥1} x^{n}/n^{2}.

The library syntax is dilog(x, prec).

3.3.23 eint1(x, \{n\}): exponential integral ∫_{x}^{∞} \frac{e^{−t}}{t} dt (x ∈ R)

If n is present, outputs the n-dimensional vector [eint1(x), …, eint1(nx)] (x ≥ 0). This is faster than repeatedly calling eint1(i * x).

The library syntax is veceint1(x, n, prec). Also available is eint1(x, prec).

3.3.24 erfc(x): complementary error function (2/\sqrt{π}) ∫_{x}^{∞} e^{−t^{2}} dt.

The library syntax is erfc(x, prec).

3.3.25 eta(x, {flag = 0}): Dedekind’s η function, without the q^{1/24}. This means the following: if x is a complex number with positive imaginary part, the result is ∏_{n=1}^{∞}(1 − q^{n}), where q = e^{2iπx}. If x is a power series (or can be converted to a power series) with positive valuation, the result is ∏_{n=1}^{∞}(1 − x^{n}).

If flag = 1 and x can be converted to a complex number (i.e. is not a power series), computes the true η function, including the leading q^{1/24}.

The library syntax is eta(x, prec).

3.3.26 exp(x): exponential of x. p-adic arguments with positive valuation are accepted.

The library syntax is gexp(x, prec).

3.3.27 gammah(x): gamma function evaluated at the argument x + 1/2. When x is an integer, this is much faster than using gamma(x + 1/2).

The library syntax is ggamd(x, prec).

3.3.28 gamma(x): gamma function of x. In the present version 2.0.17 the p-adic gamma function is not implemented.

The library syntax is ggamma(x, prec).

3.3.29 hyperu(a, b, x): U-confluent hypergeometric function with parameters a and b.

The library syntax is hyperu(a, b, x, prec).

3.3.30 incgam(s, x, y): incomplete gamma function.

The arguments s and x must be positive. The result returned is ∫_{x}^{∞} e^{−t}t^{s−1} dt. When y is given, assume (of course without checking!) that y = Γ(s). For small x, this will tremendously speed up the computation.

The library syntax is incgam(s, x, prec) and incgam4(s, x, y, prec), respectively. There exist also the functions incgam1 and incgam2 which are used for internal purposes.

3.3.31 incgamc(s, x): complementary incomplete gamma function.

The arguments s and x must be positive. The result returned is ∫_{0}^{x} e^{−t}t^{s−1} dt, when x is not too large.

The library syntax is incgam3(s, x, prec).

3.3.32 log(x, {flag = 0}): principal branch of the natural logarithm of x, i.e. such that Im(ln(x)) ∈ ] − π, π]. The result is complex (with imaginary part equal to π) if x ∈ R and x < 0.

p-adic arguments are also accepted for x, with the convention that ln(p) = 0. Hence in particular exp(ln(x))/x will not in general be equal to 1 but to a (p − 1)-th root of unity (or ±1 if p = 2) times a power of p.

If flag is equal to 1, use an agm formula suggested by Mestre, when x is real, otherwise identical

to log.

The library syntax is glog(x, prec) or glogagm(x, prec).

3.3.33 lngamma(x): principal branch of the logarithm of the gamma function of x. Can have much larger arguments than gamma itself. In the present version 2.0.17, the p-adic lngamma function is not implemented.

The library syntax is glngamma(x, prec).

3.3.34 polylog(m, x, flag = 0): one of the different polylogarithms, depending on flag:

If flag = 0 or is omitted: m^{th} polylogarithm of x, i.e. analytic continuation of the power series Li_{m}(x) = ∑_{n≥1} x^{n}/n^{m}. The program uses the power series when |x|^{2} ≤ 1/2, and the power series expansion in log(x) otherwise. It is valid in a large domain (at least |x| < 230), but should not be used too far away from the unit circle since it is then better to use the functional equation linking the value at x to the value at 1/x, which takes a trivial form for the variant below. Power series, polynomial, rational and vector/matrix arguments are allowed.

For the variants to follow we need a notation: let ℜ_{m} denotes ℜ or ℑ depending whether m is odd or even.

If flag = 1: modified m^{th} polylogarithm of x, called D̃_{m}(x) in Zagier, defined for |x| ≤ 1 by

ℜ_{m} ( ∑_{k=0}^{m−1} \frac{(− log |x|)^{k}}{k!} Li_{m−k}(x) + \frac{(− log |x|)^{m−1}}{m!} log |1 − x|) .

If flag = 2: modified m^{th} polylogarithm of x, called D_{m}(x) in Zagier, defined for |x| ≤ 1 by

ℜ_{m} ( ∑_{k=0}^{m−1} \frac{(− log |x|)^{k}}{k!} Li_{m−k}(x) − \frac{1}{2} \frac{(− log |x|)^{m}}{m!} ) .

If flag = 3: another modified m^{th} polylogarithm of x, called P_{m}(x) in Zagier, defined for |x| ≤ 1 by

ℜ_{m} ( ∑_{k=0}^{m−1} \frac{2^{k}B_{k}}{k!} (log |x|)^{k}Li_{m−k}(x) − \frac{2^{m−1}B_{m}}{m!} (log |x|)^{m}) .

These three functions satisfy the functional equation f_{m}(1/x) = (−1)^{m−1}f_{m}(x).

The library syntax is polylog0(m, x, flag, prec).

3.3.35 psi(x): the ψ-function of x, i.e. the logarithmic derivative Γ′(x)/Γ(x). The library syntax is gpsi(x, prec).

3.3.36 sin(x): sine of x.

The library syntax is gsin(x, prec).

3.3.37 sinh(x): hyperbolic sine of x.

The library syntax is gsh(x, prec).

3.3.38 sqr(x): square of x. Not identical to x ∗ x in the case of 2-adics, where it returns a more precise result.

The library syntax is gsqr(x).

3.3.39 sqrt(x): principal branch of the square root of x, i.e. such that Arg(sqrt(x)) ∈ ]−π/2, π/2], or in other words such that ℜ(sqrt(x)) > 0 or ℜ(sqrt(x)) = 0 and ℑ(sqrt(x)) ≥ 0. If x ∈ R and x < 0, then the result is complex with positive imaginary part.

Integermod a prime and p-adics are allowed as arguments. In that case, the square root (if it exists) which is returned is the one whose first p-adic digit (or its unique p-adic digit in the case of integermods) is in the interval [0, p/2]. When the argument is an integermod a non-prime (or a non-prime-adic), the result is undefined (and the function may not even return).

The library syntax is gsqrt(x, prec).

3.3.40 tan(x): tangent of x.

The library syntax is gtan(x, prec).

3.3.41 tanh(x): hyperbolic tangent of x.

The library syntax is gth(x, prec).

3.3.42 teichmuller(x): Teichmüller character of the p-adic number x. The library syntax is teich(x).

3.3.43 theta(q, z): Jacobi sine theta-function.

The library syntax is theta(q, z, prec).

3.3.44 thetanullk(q, k): k-th derivative at z = 0 of theta(q, z).

The library syntax is thetanullk(q, k, prec), where k is a long.

3.3.45 weber(x, {flag = 0}): one of Weber’s three f functions. If flag = 0, returns

f(x) = exp(−iπ/24) · η((x + 1)/2) / η(x) such that j = (f^{24} − 16)^{3}/f^{24} ,

where j is the elliptic j-invariant (see the function ellj). If flag = 1, returns

f_{1}(x) = η(x/2) / η(x) such that j = (f_{1}^{24} + 16)^{3}/f_{1}^{24} .

Finally, if flag = 2, returns

f_{2}(x) = \sqrt{2}η(2x) / η(x) such that j = (f_{2}^{24} + 16)^{3}/f_{2}^{24}.

Note the identities f^{8} = f_{1}^{8} + f_{2}^{8} and ff_{1}f_{2} = \sqrt{2}.

The library syntax is weber0(x, flag, prec), or wf(x, prec), wf1(x, prec) or wf2(x, prec).

3.3.46 zeta(s): Riemann’s zeta function ζ(s) = ∑_{n≥1} n^{−s}, computed using the Euler-Maclaurin summation formula, except when s is of type integer, in which case it is computed using Bernoulli numbers for s ≤ 0 or s > 0 and even, and using modular forms for s > 0 and odd.

The library syntax is gzeta(s, prec).

3.4 Arithmetic functions.

These functions are by definition functions whose natural domain of definition is either Z (or Z_{>0}), or sometimes polynomials over a base ring. Functions which concern polynomials exclusively will be explained in the next section. The way these functions are used is completely different from transcendental functions: in general only the types integer and polynomial are accepted as arguments. If a vector or matrix type is given, the function will be applied on each coefficient independently.

In the present version 2.0.17, all arithmetic functions in the narrow sense of the word — Euler’s totient function, the Möbius function, the sums over divisors or powers of divisors etc.— call, after trial division by small primes, the same versatile factoring machinery described under factorint. It includes Pollard Rho, ECM and MPQS stages, and has an early exit option for the functions moebius and (the integer function underlying) issquarefree. Note that it relies on a (fairly strong) probabilistic primality test: numbers found to be strong pseudo-primes after 10 successful trials of the Rabin-Miller test are declared primes.

3.4.1 Qfb(a, b, c, {D = 0.}): creates the binary quadratic form ax^{2} + bxy + cy^{2}. If b^{2} − 4ac > 0, initialize Shanks’ distance function to D.

The library syntax is Qfb0(a, b, c, D, prec). Also available are qfi(a, b, c) (when b^{2} − 4ac < 0), and qfr(a, b, c, d) (when b^{2} − 4ac > 0).

3.4.2 addprimes({x = [ ]}): adds the primes contained in the vector x (or the single integer x) to the table computed upon GP initialization (by pari init in library mode), and returns a row vector whose first entries contain all primes added by the user and whose last entries have been filled up with 1’s. In total the returned row vector has 100 components. Whenever factor or smallfact is subsequently called, first the primes in the table computed by pari init will be checked, and then the additional primes in this table. If x is empty or omitted, just returns the current list of extra primes.

The entries in x are not checked for primality. They need only be positive integers not divisible by any of the pre-computed primes. It’s in fact a nice trick to add composite numbers, which for example the function factor(x, 0) was not able to factor. In case the message “impossible inverse modulo ⟨some integermod⟩” shows up afterwards, you have just stumbled over a non-trivial factor. Note that the arithmetic functions in the narrow sense, like eulerphi, do not use this extra table.

The present PARI version 2.0.17 allows up to 100 user-specified primes to be appended to the table. This limit may be changed by altering NUMPRTBELT in file init.c. To remove primes from

the list use removeprimes.

The library syntax is addprimes(x).

3.4.3 bestappr(x, k): if x ∈ R, finds the best rational approximation to x with denominator at most equal to k using continued fractions.

The library syntax is bestappr(x, k).

3.4.4 bezout(x, y): finds u and v minimal in a natural sense such that x ∗ u + y ∗ v = gcd(x, y). The arguments must be both integers or both polynomials, and the result is a row vector with three components u, v, and gcd(x, y).

The library syntax is vecbezout(x, y) to get the vector, or gbezout(x, y, &u, &v) which gives as result the address of the created gcd, and puts the addresses of the corresponding created objects into u and v.

3.4.5 bezoutres(x, y): as bezout, with the resultant of x and y replacing the gcd.

The library syntax is vecbezoutres(x, y) to get the vector, or subresext(x, y, &u, &v) which gives as result the address of the created gcd, and puts the addresses of the corresponding created objects into u and v.

3.4.6 bigomega(x): number of prime divisors of x counted with multiplicity. x must be an integer.

The library syntax is bigomega(x), the result is a long.

3.4.7 binomial(x, y): binomial coefficient ( \begin{matrix}x \\ y \end{matrix} ). Here y must be an integer, but x can be any

PARI object.

The library syntax is binome(x, y), where y must be a long.

3.4.8 chinese(x, y): if x and y are both integermods or both polmods, creates (with the same type) a z in the same residue class as x and in the same residue class as y, if it is possible.

This function also allows vector and matrix arguments, in which case the operation is recursively applied to each component of the vector or matrix. For polynomial arguments, it is applied to each coefficient. Finally chinese(x, x) = x regardless of the type of x; this allows vector arguments to contain other data, so long as they are identical in both vectors.

The library syntax is chinois(x, y).

3.4.9 content(x): computes the gcd of all the coefficients of x, when this gcd makes sense. If x is a scalar, this simply returns x. If x is a polynomial (and by extension a power series), it gives the usual content of x. If x is a rational function, it gives the ratio of the contents of the numerator and the denominator. Finally, if x is a vector or a matrix, it gives the gcd of all the entries.

The library syntax is content(x).

3.4.10 contfrac(x, {b}, {lmax}): creates the row vector whose components are the partial quotients of the continued fraction expansion of x, the number of partial quotients being limited to lmax. If x is a real number, the expansion stops at the last significant partial quotient if lmax is omitted. x can also be a rational function or a power series.

If a vector b is supplied, the numerators will be equal to the coefficients of b. The length of the result is then equal to the length of b, unless a partial remainder is encountered which is equal to zero. In which case the expansion stops. In the case of real numbers, the stopping criterion is thus different from the one mentioned above since, if b is too long, some partial quotients may not be significant.

The library syntax is contfrac0(x, b, lmax). Also available are gboundcf(x, lmax), gcf(x), or gcf2(b, x), where lmax is a C integer.

3.4.11 contfracpnqn(x): when x is a vector or a one-row matrix, x is considered as the list of partial quotients [a_{0}, a_{1}, …, a_{n}] of a rational number, and the result is the 2 by 2 matrix [p_{n}, p_{n−1}; q_{n}, q_{n−1}] in the standard notation of continued fractions, so p_{n}/q_{n} = a_{0} + 1/(a_{1} + … + 1/a_{n}) …). If x is a matrix with two rows [b_{0}, b_{1}, …, b_{n}] and [a_{0}, a_{1}, …, a_{n}], this is then considered as a generalized continued fraction and we have similarly p_{n}/q_{n} = 1/b_{0}(a_{0}+b_{1}/(a_{1}+…+b_{n}/a_{n}) …). Note that in this case one usually has b_{0} = 1.

The library syntax is pnqn(x).

3.4.12 core(n, {flag = 0}): if n is a non-zero integer written as n = df^{2} with d squarefree, returns d. If flag is non-zero, returns the two-element row vector [d, f].

The library syntax is core0(n, flag). Also available are core(n) (= core(n, 0)) and core2(n) (= core(n, 1)).

3.4.13 coredisc(n, {flag}): if n is a non-zero integer written as n = df^{2} with d fundamental discriminant (including 1), returns d. If flag is non-zero, returns the two-element row vector [d, f]. Note that if n is not congruent to 0 or 1 modulo 4, f will be a half integer and not an integer.

The library syntax is coredisc0(n, flag). Also available are coredisc(n) (= coredisc(n, 0)) and coredisc2(n) (= coredisc(n, 1)).

3.4.14 dirdiv(x, y): x and y being vectors of perhaps different lengths but with y[1] ≠ 0 considered as Dirichlet series, computes the quotient of x by y, again as a vector.

The library syntax is dirdiv(x, y).

3.4.15 direuler(p = a, b, expr): computes the Dirichlet series to b terms of the Euler product of expression expr as p ranges through the primes from a to b. expr must be a polynomial or rational function in another variable than p (say X) and expr(X) is understood as the Dirichlet series (or more precisely the local factor) expr(p^{−s}).

The library syntax is direuler(entree *ep, GEN a, GEN b, char *expr) (see the section on sums and products for explanations of this).

3.4.16 dirmul(x, y): x and y being vectors of perhaps different lengths considered as Dirichlet series, computes the product of x by y, again as a vector.

The library syntax is dirmul(x, y).

3.4.17 divisors(x): creates a row vector whose components are the positive divisors of the integer x in increasing order. The factorization of x (as output by factor) can be used instead.

The library syntax is divisors(x).

3.4.18 eulerphi(x): Euler’s ϕ (totient) function of x. x must be of type integer.

The library syntax is phi(x).

3.4.19 factor(x, {lim = −1}): general factorization function. If x is of type integer, rational, polynomial or rational function, the result is a two-column matrix, the first column being the irreducibles dividing x (prime numbers or polynomials), and the second the exponents. If x is a vector or a matrix, the factoring is done componentwise (hence the result is a vector or matrix of two-column matrices).

If x is of type integer or rational, an argument lim can be added, meaning that we look only for factors up to lim, or to primelimit, whichever is lowest (except when lim = 0 where the effect is identical to setting lim = primelimit). Hence in this case, the remaining part is not necessarily prime. See factorint for more information about the algorithms used.

The polynomials or rational functions to be factored must have scalar coefficients. In particular PARI does not know how to factor multivariate polynomials.

Note that PARI tries to guess in a sensible way over which ring you want to factor. Note also that factorization of polynomials is done up to multiplication by a constant. In particular, the factors of rational polynomials will have integer coefficients, and the content of a polynomial or rational function is discarded and not included in the factorization. If you need to, you can always ask for the content explicitly:

    ?  factor(t^2 + 5/2*t + 1)
    %1 =
    [2*t + 1 1]
    [t + 2 1]
    ?  content(t^2 + 5/2*t + 1)
    %2 = 1/2

See also factornf.

The library syntax is factor0(x, lim), where lim is a C integer. Also available are factor(x) (= factor0(x, −1)), smallfact(x) (= factor0(x, 0)).

3.4.20 factorback(f, {nf}): f being any factorization, gives back the factored object. If a second argument nf is supplied, f is assumed to be a prime ideal factorization in the number field nf . The resulting ideal is given in HNF form.

The library syntax is factorback(f, nf ), where an omitted nf is entered as NULL.

3.4.21 factorcantor(x, p): factors the polynomial x modulo the prime p, using distinct degree plus Cantor-Zassenhaus. The coefficients of x must be operation-compatible with Z/pZ. The result is a two-column matrix, the first column being the irreducible polynomials dividing x, and the second the exponents. If you want only the degrees of the irreducible polynomials (for example for computing an L-function), use factormod(x, p, 1). Note that the factormod algorithm is usually

faster than factorcantor.

The library syntax is factcantor(x, p).

3.4.22 factorff(x, p, a): factors the polynomial x in the field F_{q} defined by the irreducible polynomial a over F_{p}. The coefficients of x must be operation-compatible with Z/pZ. The result is a two-column matrix, the first column being the irreducible polynomials dividing x, and the second the exponents. It is recommended to use for the variable of a (which will be used as variable of a polmod) a name distinct from the other variables used, so that a lift() of the result will be legible.

The library syntax is factmod9(x, p, a).

3.4.23 factorial(x) or x!: factorial of x. The expression x! gives a result which is an integer, while fact(x) gives a real number.

The library syntax is mpfact(x) for x! and mpfactr(x, prec) for fact(x). x must be a long integer and not a PARI integer.

3.4.24 factorint(n, {flag = 0}): factors the integer n using a combination of the Pollard Rho method (with modifications due to Brent), Lenstra’s ECM (with modifications by Montgomery), and MPQS (the latter adapted from the LiDIA code with the kind permission of the LiDIA maintainers), as well as a search for pure powers with exponents≤ 10. The output is a two-column matrix as for factor.

This gives direct access to the integer factoring engine called by most arithmetical functions. flag is optional; its binary digits mean 1: avoid MPQS, 2: skip first stage ECM (we may still fall back to it later), 4: avoid Rho, 8: don’t run final ECM (as a result, a huge composite may be declared to be prime). Note that a (strong) probabilistic primality test is used; thus composites might (very rarely) not be detected.

The machinery underlying this function is still in a somewhat experimental state, but should be much faster on average than pure ECM as used by all PARI versions up to 2.0.8, at the expense of heavier memory use. You are invited to play with the flag settings and watch the internals at work by using GP’s debuglevel default parameter (level 3 shows just the outline, 4 turns on time keeping, 5 and above show an increasing amount of internal details). If you see anything funny happening, please let us know.

The library syntax is factorint(n, flag).

3.4.25 factormod(x, p, {flag = 0}): factors the polynomial x modulo the prime integer p, using Berlekamp. The coefficients of x must be operation-compatible with Z/pZ. The result is a twocolumn matrix, the first column being the irreducible polynomials dividing x, and the second the exponents. If flag is non-zero, outputs only the degrees of the irreducible polynomials (for example, for computing an L-function). A different algorithm for computing the mod p factorization is factorcantor which is sometimes faster.

The library syntax is factormod(x, p, flag). Also available are factmod(x, p) (which is equivalent to factormod(x, p, 0)) and simplefactmod(x, p) (= factormod(x, p, 1)).

3.4.26 fibonacci(x): x^{th} Fibonacci number.

The library syntax is fibo(x). x must be a long.

3.4.27 gcd(x, y, {flag = 0}): creates the greatest common divisor of x and y. x and y can be of quite general types, for instance both rational numbers. Vector/matrix types are also accepted, in which case the GCD is taken recursively on each component. Note that for these types, gcd is not commutative.

If flag = 0, use Euclid’s algorithm.

If flag = 1, use the modular gcd algorithm (x and y have to be polynomials, with integer coefficients).

If flag = 2, use the subresultant algorithm.

The library syntax is gcd0(x, y, flag). Also available are ggcd(x, y), modulargcd(x, y), and srgcd(x, y) corresponding to flag = 0, 1 and 2 respectively.

3.4.28 hilbert(x, y, {p}): Hilbert symbol of x and y modulo p. If x and y are of type integer or fraction, an explicit third parameter p must be supplied, p = 0 meaning the place at infinity. Otherwise, p needs not be given, and x and y can be of compatible types integer, fraction, real, integermod or p-adic.

The library syntax is hil(x, y, p).

3.4.29 isfundamental(x): true (1) if x is equal to 1 or to the discriminant of a quadratic field, false (0) otherwise.

The library syntax is gisfundamental(x), but the simpler function isfundamental(x) which returns a long should be used if x is known to be of type integer.

3.4.30 isprime(x): true (1) if x is a strong pseudo-prime for 10 randomly chosen bases, false (0) otherwise.

The library syntax is gisprime(x), but the simpler function isprime(x) which returns a long should be used if x is known to be of type integer.

3.4.31 ispseudoprime(x): true (1) if x is a strong pseudo-prime for a randomly chosen base, false (0) otherwise.

The library syntax is gispsp(x), but the simpler function ispsp(x) which returns a long should be used if x is known to be of type integer.

3.4.32 issquare(x, {&n}): true (1) if x is square, false (0) if not. x can be of any type. If n is given and an exact square root had to be computed in the checking process, puts that square root in n. This is in particular the case when x is an integer or a polynomial. This is not the case for intmods (use quadratic reciprocity) or series (only check the leading coefficient).

The library syntax is gcarrecomplet(x, &n). Also available is gcarreparfait(x).

3.4.33 issquarefree(x): true (1) if x is squarefree, false (0) if not. Here x can be an integer or a polynomial.

The library syntax is gissquarefree(x), but the simpler function issquarefree(x) which returns a long should be used if x is known to be of type integer. This issquarefree is just the square of the Möbius function, and is computed as a multiplicative arithmetic function much like the latter.

3.4.34 kronecker(x, y): Kronecker (i.e. generalized Legendre) symbol (\frac{x}{y}). x and y must be of type integer.

The library syntax is kronecker(x, y), the result (0 or ±1) is a long.

3.4.35 lcm(x, y): least common multiple of x and y, i.e. such that lcm(x, y)∗gcd(x, y) = abs(x∗y).

The library syntax is glcm(x, y).

3.4.36 moebius(x): Möbius μ-function of x. x must be of type integer.

The library syntax is mu(x), the result (0 or ±1) is a long.

3.4.37 nextprime(x): finds the smallest prime greater than or equal to x. x can be of any real type. Note that if x is a prime, this function returns x and not the smallest prime strictly larger than x.

The library syntax is nextprime(x).

3.4.38 numdiv(x): number of divisors of x. x must be of type integer, and the result is a long.

The library syntax is numbdiv(x).

3.4.39 omega(x): number of distinct prime divisors of x. x must be of type integer.

The library syntax is omega(x), the result is a long.

3.4.40 precprime(x): finds the largest prime less than or equal to x. x can be of any real type. Returns 0 if x ≤ 1. Note that if x is a prime, this function returns x and not the largest prime strictly smaller than x.

The library syntax is precprime(x).

3.4.41 prime(x): the x^{th} prime number, which must be among the precalculated primes.

The library syntax is prime(x). x must be a long.

3.4.42 primes(x): creates a row vector whose components are the first x prime numbers, which must be among the precalculated primes.

The library syntax is primes(x). x must be a long.

3.4.43 qfbclassno(x, {flag = 0}): class number of the quadratic field of discriminant x. In the present version 2.0.17, a simple algorithm is used for x > 0, so x should not be too large (say x < 10^{7}) for the time to be reasonable. On the other hand, for x < 0 one can reasonably compute classno(x) for |x| < 10^{25}, since the method used is Shanks’ method which is in O(|x|^{1/4}). For larger values of |D|, see quadclassunit.

If flag = 1, compute the class number using Euler products and the functional equation. However, it is in O(|x|^{1/2}).

Important warning. For D < 0, this function often gives incorrect results when the class group is non-cyclic, because the authors were too lazy to implement Shanks’ method completely. It is therefore strongly recommended to use either the version with flag = 1, the function qfhclassno(−x) if x is known to be a fundamental discriminant, or the function quadclassunit.

The library syntax is qfbclassno0(x, flag). Also available are classno(x) (= qfbclassno(x)), classno2(x) (= qfbclassno(x, 1)), and finally there exists the function hclassno(x) which computes the class number of an imaginary quadratic field by counting reduced forms, an O(|x|) algorithm. See also qfbhclassno.

3.4.44 qfbcompraw(x, y) composition of the binary quadratic forms x and y, without reduction of the result. This is useful e.g. to compute a generating element of an ideal.

The library syntax is compraw(x, y).

3.4.45 qfbhclassno(x): Hurwitz class number of x, where x is non-negative and congruent to 0 or 3 modulo 4. See also qfbclassno.

The library syntax is hclassno(x).

3.4.46 qfbnucomp(x, y, l): composition of the primitive positive definite binary quadratic forms x and y using the NUCOMP and NUDUPL algorithms of Shanks (à la Atkin). l is any positive constant, but for optimal speed, one should take l = |D|^{1/4}, where D is the common discriminant of x and y.

The library syntax is nucomp(x, y, l). The auxiliary function nudupl(x, l) should be used instead for speed when x = y.

3.4.47 qfbnupow(x, n): n-th power of the primitive positive definite binary quadratic form x using the NUCOMP and NUDUPL algorithms (see qfbnucomp).

The library syntax is nupow(x, n).

3.4.48 qfbpowraw(x, n): n-th power of the binary quadratic form x, computed without doing any reduction (i.e. using qfbcompraw). Here n must be non-negative and n < 2^{31}.

The library syntax is powraw(x, n) where n must be a long integer.

3.4.49 qfbprimeform(x, p): prime binary quadratic form of discriminant x whose first coefficient is the prime number p. Returns an error if x is not a quadratic residue mod p. In the case where x > 0, the “distance” component of the form is set equal to zero according to the current precision.

3.4.50 qfbred(x, {flag = 0}, {D}, {isqrtD}, {sqrtD}): reduces the binary quadratic form x. flag can be any of 0: default behaviour, uses Shanks’ distance function d, 1: uses d, but performs only a single reduction step, 2: does not compute the distance function d, or 3: does not use d, single reduction step.

D, isqrtD, sqrtD, if present, supply the values of the discriminant, ⌊\sqrt{D}⌋, and \sqrt{D} respectively (no checking is done of these facts). If D < 0 these values are useless, and all references to Shanks’s distance are irrelevant.

The library syntax is qfbred0(x, flag, D, isqrtD, sqrtD). Use NULL to omit any of D, isqrtD, sqrtD.

Also available are

redimag(x) (= qfbred(x) where x is definite),

and for indefinite forms:

redreal(x) (= qfbred(x)),

rhoreal(x) (= qfbred(x, 1)),

redrealnod(x, sq) (= qfbred(x, 2, , isqrtD)),

rhorealnod(x, sq) (= qfbred(x, 3, , isqrtD)).

The library syntax is primeform(x, p, prec), where the third variable prec is a long, but is only taken into account when x > 0.

3.4.51 quadclassunit(D, {flag = 0}, {tech = []}): Buchmann-McCurley’s sub-exponential algorithm for computing the class group of a quadratic field of discriminant D. If D is not fundamental, the function may or may not be defined, but usually is, and often gives the right answer (a warning is issued). The more general function bnrinit should be used to compute the class group of an order.

This function should be used instead of qfbclassno or quadregula when D < −10^{25}, D > 10^{10}, or when the structure is wanted.

If flag is non-zero and D > 0, computes the narrow class group and regulator, instead of the ordinary (or wide) ones. In the current version 2.0.17, this doesn’t work at all : use the general

function bnfnarrow.

tech

is a row vector of the form [c_{1}, c_{2}], where c_{1} and c_{2} are positive real numbers which control the execution time and the stack size. To get maximum speed, set c_{2} = c. To get a rigorous result (under GRH) you must take c_{2} = 6. Reasonable values for c are between 0.1 and 2.

The result of this function is a vector v with 4 components if D < 0, and 5 otherwise. The correspond respectively to

• v[1] : the class number

• v[2] : a vector giving the structure of the class group as a product of cyclic groups; • v[3] : a vector giving generators of those cyclic groups (as binary quadratic forms).

• v[4] : (omitted if D < 0) the regulator, computed to an accuracy which is the maximum of an internal accuracy determined by the program and the current default (note that once the regulator is known to a small accuracy it is trivial to compute it to very high accuracy, see the tutorial).

• v[5] : a measure of the correctness of the result. If it is close to 1, the result is correct (under GRH). If it is close to a larger integer, this shows that the class number is off by a factor equal to this integer, and you must start again with a larger value for c_{1} or a different random seed. In this case, a warning message is printed.

The library syntax is quadclassunit0(D, flag, tech). Also available are buchimag(D, c_{1}, c_{2}) and buchreal(D, flag, c_{1}, c_{2}).

3.4.52 quaddisc(x): discriminant of the quadratic field Q(\sqrt{x}), where x ∈ Q.

The library syntax is quaddisc(x).

3.4.53 quadhilbert(D, {flag = 0}): relative equation defining the Hilbert class field of the quadratic field of discriminant D. If flag is non-zero and D < 0, outputs [form, root(form)] (to be used for constructing subfields). Uses complex multiplication in the imaginary case and Stark units in the real case.

The library syntax is quadhilbert(D, flag, prec).

3.4.54 quadgen(x): creates the quadratic number ω = (a + \sqrt{x})/2 where a = 0 if x ≡ 0 mod 4, a = 1 if x ≡ 1 mod 4, so that (1, ω) is an integral basis for the quadratic order of discriminant x. x must be an integer congruent to 0 or 1 modulo 4.

The library syntax is quadgen(x).

3.4.55 quadpoly(D, {v = x}): creates the “canonical” quadratic polynomial (in the variable v) corresponding to the discriminant D, i.e. the minimal polynomial of quadgen(x). D must be an integer congruent to 0 or 1 modulo 4.

The library syntax is quadpoly0(x, v).

3.4.56 quadray(D, f, {flag = 0}): relative equation for the ray class field of conductor f for the quadratic field of discriminant D (which can also be a bnf). flag is only meaningful when D < 0. If it’s an odd integer, outputs instead the vector of [ideal, corresponding root].

If flag = 0 or 1, uses the σ function, while if flag > 1, uses the Weierstrass ℘ function, which is less efficient and may disappear in future versions (not all special cases have been implemented in this case). Finally, flag can also be a two-component vector [λ, flag], where flag is as above and λ is the technical element of bnf necessary for Schertz’s method using σ. In that case, returns 0 if λ is not suitable.

If D > 0, the function may fail with the following message

    "Cannot find a suitable modulus in FindModulus"

See the comments in bnrstark about this problem.

The library syntax is quadray(D, f, flag).

3.4.57 quadregulator(x): regulator of the quadratic field of positive discriminant x. Returns an error if x is not a discriminant (fundamental or not) or if x is a square. See also quadclassunit if x is large.

The library syntax is regula(x, prec).

3.4.58 quadunit(x): fundamental unit of the real quadratic field Q(\sqrt{x}) where x is the positive discriminant of the field. If x is not a fundamental discriminant, this probably gives the fundamental unit of the corresponding order. x must be of type integer, and the result is a quadratic number.

The library syntax is fundunit(x).

3.4.59 removeprimes({x = [ ]}): removes the primes listed in x from the prime number table. x can also be a single integer. List the current extra primes if x is omitted.

The library syntax is removeprimes(x).

3.4.60 sigma(x, {k = 1}): sum of the k^{th} powers of the positive divisors of x. x must be of type integer.

The library syntax is sumdiv(x) (= sigma(x)) or gsumdivk(x, k) (= sigma(x, k)), where k is a C long integer.

3.4.61 sqrtint(x): integer square root of x, which must be of PARI type integer. The result is non-negative and rounded towards zero. A negative x is allowed, and the result in that case is

I*sqrtint(-x).

The library syntax is racine(x).

3.4.62 znlog(x, g): g must be a primitive root mod a prime p, and the result is the discrete log of x in the multiplicative group (Z/pZ)^{∗}. This function using a simple-minded baby-step/giant-step approach and requires O(\sqrt{p}) storage, hence it cannot be used for p greater than about 10^{1}3.

The library syntax is znlog(x, g).

3.4.63 znorder(x): x must be an integer mod n, and the result is the order of x in the multiplicative group (Z/nZ)^{∗}. Returns an error if x is not invertible.

The library syntax is order(x).

3.4.64 znprimroot(x): returns a primitive root of x, where x is a prime power.

The library syntax is gener(x).

3.4.65 znstar(n): gives the structure of the multiplicative group (Z/nZ)^{∗} as a 3-component row vector v, where v[1] = ϕ(n) is the order of that group, v[2] is a k-component row-vector d of integers d[i] such that d[i] > 1 and d[i] | d[i − 1] for i ≥ 2 and (Z/nZ)^{∗} ≃ ∏_{i=1}^{k}(Z/d[i]Z), and v[3] is a k-component row vector giving generators of the image of the cyclic groups Z/d[i]Z.

The library syntax is znstar(n).

3.5 Functions related to elliptic curves.

We have implemented a number of functions which are useful for number theorists working on elliptic curves. We always use Tate’s notations. The functions assume that the curve is given by a general Weierstrass model

y^{2} + a_{1}xy + a_{3}y = x^{3} + a_{2}x^{2} + a_{4}x + a_{6},

where a priori the a_{i} can be of any scalar type. This curve can be considered as a five-component vector E=[a1,a2,a3,a4,a6]. Points on E are represented as two-component vectors [x,y], except for the point at infinity, i.e. the identity element of the group law, represented by the one-component vector [0].

It is useful to have at one’s disposal more information. This is given by the function ellinit (see there), which usually gives a 19 component vector (which we will call a long vector in this section). If a specific flag is added, a vector with only 13 component will be output (which we will call a medium vector). A medium vector just gives the first 13 components of the long vector corresponding to the same curve, but is of course faster to compute. The following member functions are available to deal with the output of ellinit:

a1–a6, b2–b8, c4–c6 : coefficients of the elliptic curve.
area  : volume of the complex lattice defining E.
disc  : discriminant of the curve.
j  : j-invariant of the curve.
omega  : [ω_{1}, ω_{2}], periods forming a basis of the complex lattice defining E (ω_{1} is the

real period, and ω_{2}/ω_{1} belongs to Poincaré’s half-plane).

eta  : quasi-periods [η_{1}, η_{2}], such that η_{1}ω_{2} − η_{2}ω_{1} = iπ.
roots  : roots of the associated Weierstrass equation.
tate  : [u^{2}, u, v] in the notation of Tate.
w  : Mestre’s w (this is technical).

Their use is best described by an example: assume that E was output by ellinit, then typing E.disc will retrieve the curve’s discriminant. The member functions area, eta and omega are only available for curves over Q. Conversely, tate and w are only available for curves defined over Q_{p}.

Some functions, in particular those relative to height computations (see ellheight) require also that the curve be in minimal Weierstrass form. This is achieved by the function ellglobalred.

All functions related to elliptic curves share the prefix ell, and the precise curve we are interested in is always the first argument, in either one of the three formats discussed above, unless otherwise specified. For instance, in functions which do not use the extra information given by long vectors, the curve can be given either as a five-component vector, or by one of the longer vectors computed by ellinit.

3.5.1 elladd(E, z1, z2): sum of the points z1 and z2 on the elliptic curve corresponding to the vector E.

The library syntax is addell(E, z1, z2).

3.5.2 ellak(E, n): computes the coefficient a_{n} of the L-function of the elliptic curve E, i.e. in principle coefficients of a newform of weight 2 assuming Taniyama-Weil. E must be a medium or long vector of the type given by ellinit. For this function to work for every n and not just those prime to the conductor, E must be a minimal Weierstrass equation. If this is not the case, use the function ellglobalred first before using ellak.

The library syntax is akell(E, n).

3.5.3 ellan(E, n): computes the vector of the first n a_{k} corresponding to the elliptic curve E. All comments in ellak description remain valid.

The library syntax is anell(E, n), where n is a C integer.

3.5.4 ellap(E, p, {flag = 0}): computes the a_{p} corresponding to the elliptic curve E and the prime number p. These are defined by the equation #E(F_{p}) = p + 1 − a_{p}, where #E(F_{p}) stands for the number of points of the curve E over the finite field F_{p}. When flag is 0, this uses the baby-step giant-step method and a trick due to Mestre. This runs in time O(p^{1/4}) and requires O(p^{1/4}) storage, hence becomes unreasonable when p has about 30 digits.

If flag is 1, computes the a_{p} as a sum of Legendre symbols. This is slower than the previous method as soon as p is greater than 100, say.

No checking is done that p is indeed prime. E must be a medium or long vector of the type given by ellinit, defined over Q, F_{p} or Q_{p}.

The library syntax is ellap0(E, p, flag). Also available are apell(E, p), corresponding to flag = 0, and apell2(E, p) (flag = 1).

3.5.5 ellbil(E, z1, z2): if z1 and z2 are points on the elliptic curve E, this function computes the value of the canonical bilinear form on z1, z2:

               ellheight(E, z1+z2) − ellheight(E, z1) − ellheight(E, z2)

where + denotes of course addition on E. In addition, z1 or z2 (but not both) can be vectors or matrices. Note that this is equal to twice some normalizations. E is assumed to be integral, given by a minimal model.

The library syntax is bilhell(E, z1, z2, prec).

3.5.6 ellchangecurve(E, v): changes the data for the elliptic curve E by changing the coordinates using the vector v=[u,r,s,t], i.e. if x′ and y′ are the new coordinates, then x = u^{2}x′ + r, y = u^{3}y′ + su^{2}x′ + t. The vector E must be a medium or long vector of the type given by ellinit.

The library syntax is coordch(E, v).

3.5.7 ellchangepoint(x, v): changes the coordinates of the point or vector of points x using the vector v=[u,r,s,t], i.e. if x′ and y′ are the new coordinates, then x = u^{2}x′ +r, y = u^{3}y′ +su^{2}x′ +t

(see also ellchangecurve).

The library syntax is pointch(x, v).

3.5.8 elleisnum(E, k, {flag = 0}): E being an elliptic curve as output by ellinit (or, alternatively, given by a 2-component vector [ω_{1}, ω_{2}]), and k being an even positive integer, computes the numerical value of the Eisenstein series of weight k at E. When flag is non-zero and k = 4 or 6, returns g_{2} or g_{3} with the correct normalization.

The library syntax is elleisnum(E, k, flag).

3.5.9 elleta(om): returns the two-component row vector [η_{1}, η_{2}] of quasi-periods associated to om = [ω_{1}, ω_{2}]

The library syntax is elleta(om, prec)

3.5.10 ellglobalred(E): calculates the arithmetic conductor, the global minimal model of E and the global Tamagawa number c. Here E is an elliptic curve given by a medium or long vector of the type given by ellinit, and is supposed to have all its coefficients a_{i} in Q. The result is a 3 component vector [N, v, c]. N is the arithmetic conductor of the curve, v is itself a vector [u, r, s, t] with rational components. It gives a coordinate change for E over Q such that the resulting model has integral coefficients, is everywhere minimal, a_{1} is 0 or 1, a_{2} is 0, 1 or −1 and a_{3} is 0 or 1. Such a model is unique, and the vector v is unique if we specify that u is positive. To get the new model, simply type ellchangecurve(E,v). Finally c is the product of the local Tamagawa numbers c_{p}, a quantity which enters in the Birch and Swinnerton-Dyer conjecture.

The library syntax is globalreduction(E).

3.5.11 ellheight(E, z, {flag = 0}): global Néron-Tate height of the point z on the elliptic curve E. The vector E must be a long vector of the type given by ellinit, with flag = 1. If flag = 0, this computation is done using sigma and theta-functions and a trick due to J. Silverman. If flag = 1, use Tate’s 4^{n} algorithm, which is much slower.

The library syntax is ellheight0(E, z, flag, prec). The Archimedean contribution alone is given by the library function hell(E, z, prec). Also available are ghell(E, z, prec) (flag = 0) and ghell2(E, z, prec) (flag = 1).

3.5.12 ellheightmatrix(E, x): x being a vector of points, this function outputs the Gram matrix of x with respect to the Néron-Tate height, in other words, the (i, j) component of the matrix is equal to ellbil(E,x[i],x[j]). The rank of this matrix, at least in some approximate sense, gives the rank of the set of points, and if x is a basis of the Mordell-Weil group of E, its determinant is equal to the regulator of E. Note that this matrix should be divided by 2 to be in accordance with certain normalizations. E is assumed to be integral, given by a minimal model.

The library syntax is mathell(E, x, prec).

3.5.13 ellinit(E, {flag = 0}): computes some fixed data concerning the elliptic curve given by the five-component vector E, which will be essential for most further computations on the curve. The result is a 19-component vector E (called a long vector in this section), shortened to 13 components (medium vector) if flag = 1. Both contain the following information in the first 13 components:

a_{1}, a_{2}, a_{3}, a_{4}, a_{6}, b_{2}, b_{4}, b_{6}, b_{8}, c_{4}, c_{6}, Δ, j.

In particular, the discriminant is E[12] (or E.disc), and the j-invariant is E[13] (or E.j).

The other six components are only present if flag is 0 (or omitted!). Their content depends on whether the curve is defined over R or not:

• When E is defined over R, E[14] (E.roots) is a vector whose three components contain the roots of the associated Weierstrass equation. If the roots are all real, then they are ordered by decreasing value. If only one is real, it is the first component of E[14].

E[15] (E.omega[1]) is the real period of E (integral of dx/(2y + a_{1}x + a_{3}) over the connected component of the identity element of the real points of the curve), and E[16] (E.omega[2]) is a complex period. In other words, ω_{1} = E[15] and ω_{2} = E[16] form a basis of the complex lattice defining E (E.omega), with τ = \frac{ω_{2}}{ω_{1}} having positive imaginary part.

E[17] and E[18] are the corresponding values η_{1} and η_{2} such that η_{1}ω_{2} − η_{2}ω_{1} = iπ, and both can be retrieved by typing E.eta (as a row vector whose components are the η_{i}).

Finally, E[19] (E.area) is the volume of the complex lattice defining E.

• When E is defined over Q_{p}, the p-adic valuation of j must be negative. Then E[14] (E.roots) is the vector with a single component equal to the p-adic root of the associated Weierstrass equation corresponding to −1 under the Tate parametrization.

E[15] is equal to the square of the u-value, in the notation of Tate.

E[16] is the u-value itself, if it belongs to Q_{p}, otherwise zero.

E[17] is the value of Tate’s q for the curve E.

E.tate will yield the three-component vector [u^{2}, u, q].

E[18] (E.w) is the value of Mestre’s w (this is technical), and E[19] is arbitrarily set equal to zero.

For all other base fields or rings, the last six components are arbitrarily set equal to zero (see also the description of member functions related to elliptic curves at the beginning of this section).

The library syntax is ellinit0(E, flag, prec). Also available are initell(E, prec) (flag = 0) and smallinitell(E, prec) (flag = 1).

3.5.14 ellisoncurve(E, z): gives 1 (i.e. true) if the point z is on the elliptic curve E, 0 otherwise. If E or z have imprecise coefficients, an attempt is made to take this into account, i.e. an imprecise equality is checked, not a precise one.

The library syntax is oncurve(E, z), and the result is a long.

3.5.15 ellj(x): elliptic j-invariant. x must be a complex number with positive imaginary part, or convertible into a power series or a p-adic number with positive valuation.

The library syntax is jell(x, prec).

3.5.16 elllocalred(E, p): calculates the Kodaira type of the local fiber of the elliptic curve E at the prime p. E must be given by a medium or long vector of the type given by ellinit, and is assumed to have all its coefficients a_{i} in Z. The result is a 4-component vector [f, kod, v, c]. Here f is the exponent of p in the arithmetic conductor of E, and kod is the Kodaira type which is coded as follows:

1 means good reduction (type I_{0}), 2, 3 and 4 mean types II, III and IV respectively, 4 + ν with ν > 0 means type I_{ν}; finally the opposite values −1, −2, etc. refer to the starred types I_{0}^{∗}, II^{∗}, etc. The third component v is itself a vector [u, r, s, t] giving the coordinate changes done during the local reduction. Normally, this has no use if u is 1, that is, if the given equation was already minimal. Finally, the last component c is the local Tamagawa number c_{p}.

The library syntax is localreduction(E, p).

3.5.17 elllseries(E, s, {A = 1}): E being a medium or long vector given by ellinit, this computes the value of the L-series of E at s. It is assumed that E is a minimal model over Z and that the curve is a modular elliptic curve. The optional parameter A is a cutoff point for the integral, which must be chosen close to 1 for best speed. The result must be independent of A, so this allows some internal checking of the function.

Note that if the conductor of the curve is large, say greater than 10^{12}, this function will take an unreasonable amount of time since it uses an O(N^{1/2}) algorithm.

The library syntax is lseriesell(E, s, A, prec) where prec is a long and an omitted A is coded

as NULL.

3.5.18 ellorder(E, z): gives the order of the point z on the elliptic curve E if it is a torsion point, zero otherwise. In the present version 2.0.17, this is implemented only for elliptic curves defined over Q.

The library syntax is orderell(E, z).

3.5.19 ellordinate(E, x): gives a 0, 1 or 2-component vector containing the y-coordinates of the points of the curve E having x as x-coordinate.

The library syntax is ordell(E, x).

3.5.20 ellpointtoz(E, z): if E is an elliptic curve with coefficients in R, this computes a complex number t (modulo the lattice defining E) corresponding to the point z, i.e. such that, in the standard Weierstrass model, ℘(t) = z[1], ℘′(t) = z[2]. In other words, this is the inverse function

of ellztopoint.

If E has coefficients in Q_{p}, then either Tate’s u is in Q_{p}, in which case the output is a p-adic number t corresponding to the point z under the Tate parametrization, or only its square is, in which case the output is t + 1/t. E must be a long vector output by ellinit.

The library syntax is zell(E, z, prec).

3.5.21 ellpow(E, z, n): computes n times the point z for the group law on the elliptic curve E. Here, n can be in Z, or n can be a complex quadratic integer if the curve E has complex multiplication by n (if not, an error message is issued).

The library syntax is powell(E, z, n).

3.5.22 ellrootno(E, {p = 1}): E being a medium or long vector given by ellinit, this computes the local (if p ≠ 1) or global (if p = 1) root number of the L-series of the elliptic curve E. Note that the global root number is the sign of the functional equation and conjecturally is the parity of the rank of the Mordell-Weil group. The equation for E must have coefficients in Q but need not be minimal.

The library syntax is ellrootno(E, p) and the result (equal to ±1) is a long.

3.5.23 ellsigma(E, z, {flag = 0}): value of the Weierstrass σ function of the lattice associated to E as given by ellinit (alternatively, E can be given as a lattice [ω_{1}, ω_{2}]).

If flag = 1, computes an (arbitrary) determination of log(σ(z)).

If flag = 2, 3, same using the product expansion instead of theta series. The library syntax is ellsigma(E, z, flag)

3.5.24 ellsub(E, z1, z2): difference of the points z1 and z2 on the elliptic curve corresponding to the vector E.

The library syntax is subell(E, z1, z2).

3.5.25 elltaniyama(E): computes the modular parametrization of the elliptic curve E, where E is given in the (long or medium) format output by ellinit, in the form of a two-component vector [u, v] of power series, given to the current default series precision. This vector is characterized by the following two properties. First the point (x, y) = (u, v) satisfies the equation of the elliptic curve. Second, the differential du/(2v+a_{1}u+a_{3}) is equal to f(z)dz, a differential form on H/Γ_{0}(N) where N is the conductor of the curve. The variable used in the power series for u and v is x, which is implicitly understood to be equal to exp(2iπz). It is assumed that the curve is a strong Weil curve, and the Manin constant is equal to 1. The equation of the curve E must be minimal (use ellglobalred to get a minimal equation).

The library syntax is taniyama(E), and the precision of the result is determined by the global variable precdl.

3.5.26 elltors(E, {flag = 0}): if E is an elliptic curve defined over Q, outputs the torsion subgroup of E as a 3-component vector [t,v1,v2], where t is the order of the torsion group, v1 gives the structure of the torsion group as a product of cyclic groups (sorted by decreasing order), and v2 gives generators for these cyclic groups. E must be a long vector as output by ellinit.

    ?  E = ellinit([0,0,0,-1,0]);
    ?  elltors(E)
    %1 = [4, [2, 2], [[0, 0], [1, 0]]]

Here, the torsion subgroup is isomorphic to Z/2Z × Z/2Z, with generators [0, 0] and [1, 0].

If flag = 0, use Doud’s algorithm : bound torsion by computing #E(F_{p}) for small primes of good reduction, then look for torsion points using Weierstrass parametrization (and Mazur’s classification).

If flag = 1, use Lutz–Nagell (much slower), E is allowed to be a medium vector.

The library syntax is elltors0(E, flag).

3.5.27 ellwp(E, {z = x}, {flag = 0}):

Computes the value at z of the Weierstrass ℘ function attached to the elliptic curve E as given by ellinit (alternatively, E can be given as a lattice [ω_{1}, ω_{2}]).

If z is omitted or is a simple variable, computes the power series expansion in z (starting z^{−2} + O(z^{2})). The number of terms to an even power in the expansion is the default serieslength in GP, and the second argument (C long integer) in library mode.

Optional flag is (for now) only taken into account when z is numeric, and means 0: compute only ℘(z), 1: compute [℘(z), ℘′(z)].

The library syntax is ellwp0(E, z, flag, prec, precdl). Also available is weipell(E, precdl) for the power series (in x = polx[0]).

3.5.28 ellzeta(E, z): value of the Weierstrass ζ function of the lattice associated to E as given by ellinit (alternatively, E can be given as a lattice [ω_{1}, ω_{2}]).

The library syntax is ellzeta(E, z).

3.5.29 ellztopoint(E, z): E being a long vector, computes the coordinates [x, y] on the curve E corresponding to the complex number z. Hence this is the inverse function of ellpointtoz. In other words, if the curve is put in Weierstrass form, [x, y] represents the Weierstrass ℘-function and its derivative. If z is in the lattice defining E over C, the result is the point at infinity [0].

The library syntax is pointell(E, z, prec).

3.6 Functions related to general number fields.

In this section can be found functions which are used almost exclusively for working in general number fields. Other less specific functions can be found in the next section on polynomials. Functions related to quadratic number fields can be found in the section Section 3.4 (Arithmetic functions).

We shall use the following conventions:

• nf denotes a number field, i.e. a 9-component vector in the format output by nfinit. This contains the basic arithmetic data associated to the number field: signature, maximal order, discriminant, etc.

• bnf denotes a big number field, i.e. a 10-component vector in the format output by bnfinit. This contains nf and the deeper invariants of the field: units, class groups, as well as a lot of technical data necessary for some complex fonctions like bnfisprincipal.

• bnr denotes a big “ray number field”, i.e. some data structure output by bnrinit, even more complicated than bnf , corresponding to the ray class group structure of the field, for some modulus.

• rnf denotes a relative number field (see below).

• ideal can mean any of the following:

– a Z-basis, in Hermite normal form (HNF) or not. In this case x is a square matrix.

– an idele, i.e. a 2-component vector, the first being an ideal given as a Z–basis, the second being a r_{1} + r_{2}-component row vector giving the complex logarithmic Archimedean information.

– a Z_{K}-generating system for an ideal.

– a column vector x expressing an element of the number field on the integral basis, in which case the ideal is treated as being the principal idele (or ideal) generated by x.

– a prime ideal, i.e. a 5-component vector in the format output by idealprimedec.

– a polmod x, i.e. an algebraic integer, in which case the ideal is treated as being the principal idele generated by x.

– an integer or a rational number, also treated as a principal idele.

• a character on the Abelian group ⨁(Z/N_{i}Z)g_{i} is given by a row vector χ = [a_{1}, …, a_{n}] such that χ(∏ g_{i}^{n_{i}}) = exp(2iπ ∑ a_{i}n_{i}/N_{i}).

Warnings:

1) An element in nf can be expressed either as a polmod or as a vector of components on the integral basis nf .zk. It is absolutely essential that all such vectors be column vectors.

2) When giving an ideal by a Z_{K} generating system to a function expecting an ideal, it must be ensured that the function understands that it is a Z_{K}-generating system and not a Z-generating system. When the number of generators is strictly less than the degree of the field, there is no ambiguity and the program assumes that one is giving a Z_{K}-generating set. When the number of generators is greater than or equal to the degree of the field, however, the program assumes on the contrary that you are giving a Z-generating set. If this is not the case, you must absolutely change it into a Z-generating set, the simplest manner being to use idealhnf(nf ,x).

Concerning relative extensions, some additional definitions are necessary.

• A relative matrix will be a matrix whose entries are elements of a (given) number field nf , always expressed as column vectors on the integral basis nf .zk. Hence it is a matrix of vectors.

• An ideal list will be a row vector of (fractional) ideals of the number field nf .

• A pseudo-matrix will be a pair (A, I) where A is a relative matrix and I an ideal list whose length is the same as the number of columns of A. This pair will be represented by a 2-component row vector.

• The module generated by a pseudo-matrix (A, I) is the sum ∑_{i} a_{j}A_{j} where the a_{j} are the ideals of I and A_{j} is the j-th column of A.

• A pseudo-matrix (A, I) is a pseudo-basis of the module it generates if A is a square matrix with non-zero determinant and all the ideals of I are non-zero. We say that it is in Hermite Normal Form (HNF) if it is upper triangular and all the elements of the diagonal are equal to 1.

• The determinant of a pseudo-basis (A, I) is the ideal equal to the product of the determinant of A by all the ideals of I. The determinant of a pseudo-matrix is the determinant of any pseudobasis of the module it generates.

Finally, when defining a relative extension, the base field should be defined by a variable having a lower priority (i.e. a higher number) than the variable defining the extension. For example, under GP you can use the variable name y (or t) to define the base field, and the variable name x to define the relative extension.

Now a last set of definitions concerning the way big ray number fields (or bnr) are input, using class field theory. These are defined by a triple a1, a2, a3, where the defining set [a1, a2, a3] can have any of the following forms: [bnr], [bnr, subgroup], [bnf , module], [bnf , module, subgroup], where:

• bnf is as output by bnfclassunit or bnfinit, where units are mandatory unless the ideal is trivial; bnr by bnrclass (with flag > 0) or bnrinit. This is the ground field.

• module is either an ideal in any form (see above) or a two-component row vector containing an ideal and an r_{1}-component row vector of flags indicating which real Archimedean embeddings to take in the module.

• subgroup is the HNF matrix of a subgroup of the ray class group of the ground field for the modulus module. This is input as a square matrix expressing generators of a subgroup of the ray class group bnr.clgp on the given generators.

The corresponding bnr is then the subfield of the ray class field of the ground field for the given modulus, associated to the given subgroup.

All the functions which are specific to relative extensions, number fields, big number fields, big number rays, share the prefix rnf, nf, bnf, bnr respectively. They are meant to take as first argument a number field of that precise type, respectively output by rnfinit, nfinit, bnfinit,

and bnrinit.

However, and even though it may not be specified in the descriptions of the functions below, it is permissible, if the function expects a nf , to use a bnf instead (which contains much more information). The program will make the effort of converting to what it needs. On the other hand, if the program requires a big number field, the program will not launch bnfinit for you, which can be a costly operation. Instead, it will give you a specific error message.

The data types corresponding to the structures described above are rather complicated. Thus, as we already have seen it with elliptic curves, GP provides you with some “member functions” to retrieve the data you need from these structures (once they have been initialized of course). The relevant types of number fields are indicated between parentheses:

bnf (bnr, bnf ) : big number field.
clgp (bnr, bnf  ) : classgroup. This one admits the following three subclasses:
cyc  : cyclic decomposition (SNF).
gen  : generators.
no  : number of elements.
diff (bnr, bnf , nf ) : the different ideal.
codiff (bnr, bnf , nf ) : the codifferent (inverse of the different in the ideal group).
disc (bnr, bnf , nf ) : discriminant.
fu  (bnr, bnf , nf ) : fundamental units.
futu (bnr, bnf  ) : [u, w], u is a vector of fundamental units, w generates the torsion.
nf  (bnr, bnf , nf ) : number field.
reg  (bnr, bnf , ) : regulator.
roots (bnr, bnf , nf ) : roots of the polnomial generating the field.
sign (bnr, bnf , nf ) : [r_{1}, r_{2}] the signature of the field. This means that the field has r_{1} real

embeddings, 2r_{2} complex ones.

t2  (bnr, bnf , nf ) : the T2 matrix (see nfinit).
tu  (bnr, bnf , ) : a generator for the torsion units.
tufu (bnr, bnf , ) : as futu, but outputs [w, u].
zk  (bnr, bnf , nf ) : integral basis, i.e. a Z-basis of the maximal order.
zkst (bnr  ) : structure of (Z_{K}/m)^{∗} (can be extracted also from an idealstar).

For instance, assume that bnf = bnfinit(pol), for some polynomial. Then bnf .clgp retrieves the class group, and bnf .clgp.no the class number. If we had set bnf = nfinit(pol), both would have output an error message. All these functions are completely recursive, thus for instance bnr.bnf.nf.zk will yield the maximal order of bnr (which you could get directly with a simple bnr.zk of course).

The following functions, starting with buch in library mode, and with bnf under GP, are implementations of the sub-exponential algorithms for finding class and unit groups under GRH, due to Hafner-McCurley, Buchmann and Cohen-Diaz-Olivier.

The general call to the functions concerning class groups of general number fields (i.e. excluding quadclassunit) involves a polynomial P and a technical vector

tech = [c, c2, nrel, borne, nrpid, minsfb], where the parameters are to be understood as follows:

P is the defining polynomial for the number field, which must be in Z[X], irreducible and, preferably, monic. In fact, if you supply a non-monic polynomial at this point, GP will issue a warning, then transform your polynomial so that it becomes monic. Instead of the normal result, say res, you then get a vector [res,Mod(a,Q)], where Mod(a,Q)=Mod(X,P) gives the change of variables.

The numbers c and c2 are positive real numbers which control the execution time and the stack size. To get maximum speed, set c2 = c. To get a rigorous result (under GRH) you must take c2 = 12 (or c2 = 6 in the quadratic case, but then you should use the much faster function quadclassunit). Reasonable values for c are between 0.1 and 2. (The defaults are c = c2 = 0.3).

nrel is the number of initial extra relations requested in computing the relation matrix. Reasonable values are between 5 and 20. (The default is 5).

borne is a multiplicative coefficient of the Minkowski bound which controls the search for small norm relations. If this parameter is set equal to 0, the program does not search for small norm relations. Otherwise reasonable values are between 0.5 and 2.0. (The default is 1.0).

nrpid is the maximal number of small norm relations associated to each ideal in the factor base. Irrelevant when borne = 0. Otherwise, reasonable values are between 4 and 20. (The default is 4).

minsfb is the minimal number of elements in the “sub-factorbase”. If the program does not seem to succeed in finding a full rank matrix (which you can see in GP by typing \g 2), increase this number. Reasonable values are between 2 and 5. (The default is 3).

Remarks.

Apart from the polynomial P, you don’t need to supply any of the technical parameters (under the library you still need to send at least an empty vector, cgetg(1,t VEC)). However, should you choose to set some of them, they must be given in the requested order. For example, if you want to specify a given value of nrel, you must give some values as well for c and c2, and provide a vector [c, c2, nrel].

Note also that you can use an nf instead of P, which avoids recomputing the integral basis and analogous quantities.

3.6.1 bnfcertify(bnf ): bnf being a big number field as output by bnfinit or bnfclassunit, checks whether the result is correct, i.e. whether it is possible to remove the assumption of the Generalized Riemann Hypothesis. If it is correct, the answer is 1. If not, the program may output some error message, but more probably will loop indefinitely. In no occasion can the program give a wrong answer (barring bugs of course): if the program answers 1, the answer is certified.

The library syntax is certifybuchall(bnf ), and the result is a C long.

3.6.2 bnfclassunit(P, {flag = 0}, {tech = [ ]}): Buchmann’s sub-exponential algorithm for computing the class group, the regulator and a system of fundamental units of the general algebraic number field K defined by the irreducible polynomial P with integer coefficients.

The result of this function is a vector v with 10 components (it is not a bnf , you need bnfinit for that), which for ease of presentation is in fact output as a one column matrix. First we describe the default behaviour (flag = 0):

v[1] is equal to the polynomial P. Note that for optimum performance, P should have gone through polred or nfinit(x, 2).

v[2] is the 2-component vector [r1, r2], where r1 and r2 are as usual the number of real and half the number of complex embeddings of the number field K.

v[3] is the 2-component vector containing the field discriminant and the index.

v[4] is an integral basis in Hermite normal form.

v[5] (v.clgp) is a 3-component vector containing the class number (v.clgp.no), the structure of the class group as a product of cyclic groups of order n_{i} (v.clgp.cyc), and the corresponding generators of the class group of respective orders n_{i} (v.clgp.gen).

v[6] (v.reg) is the regulator computed to an accuracy which is the maximum of an internally determined accuracy and of the default.

v[7] is a measure of the correctness of the result. If it is close to 1, the results are correct (under GRH). If it is close to a larger integer, this shows that the product of the class number by the regulator is off by a factor equal to this integer, and you must start again with a larger value for c or a different random seed, i.e. use the function setrand. (Since the computation involves a random process, starting again with exactly the same parameters may give the correct result.) In this case a warning message is printed.

v[8] (v.tu) a vector with 2 components, the first being the number w of roots of unity in K and the second a primitive w-th root of unity expressed as a polynomial.

v[9] (v.fu) is a system of fundamental units also expressed as polynomials.

v[10] gives a measure of the correctness of the computations of the fundamental units (not of the regulator), expressed as a number of bits. If this number is greater than 20, say, everything is OK. If v[10] ≤ 0, then we have lost all accuracy in computing the units (usually an error message will be printed and the units not given). In the intermediate cases, one must proceed with caution (for example by increasing the current precision).

If flag = 1, and the precision happens to be insufficient for obtaining the fundamental units exactly, the internal precision is doubled and the computation redone, until the exact results are obtained. The user should be warned that this can take a very long time when the coefficients of the fundamental units on the integral basis are very large, for example in the case of large real quadratic fields. In that case, there are alternate methods for representing algebraic numbers which are not implemented in PARI.

If flag = 2, the fundamental units and roots of unity are not computed. Hence the result has only 7 components, the first seven ones.

tech is a technical vector (empty by default) containing c, c2, nrel, borne, nbpid, minsfb, in this order (see the beginning of the section or the keyword bnf). You can supply any number of these provided you give an actual value to each of them (the “empty arg” trick won’t work here). Careful use of these parameters may speed up your computations considerably.

The library syntax is bnfclassunit0(P, flag, tech, prec).

3.6.3 bnfclgp(P, {tech = [ ]}): as bnfclassunit, but only outputs v[5], i.e. the class group.

The library syntax is bnfclassgrouponly(P, tech, prec), where tech is as described under

bnfclassunit.

3.6.4 bnfdecodemodule(nf , m): if m is a module as output in the first component of an extension given by bnrdisclist, outputs the true module.

The library syntax is decodemodule(nf , m).

3.6.5 bnfinit(P, {flag = 0}, {tech = [ ]}): essentially identical to bnfclassunit except that the output contains a lot of technical data, and should not be printed out explicitly in general. The

result of bnfinit is used in programs such as bnfisprincipal, bnfisunit or bnfnarrow. The

result is a 10-component vector bnf .

• The first 6 and last 2 components are technical and in principle are not used by the casual user. However, for the sake of completeness, their description is as follows. We use the notations explained in the book by H. Cohen, A Course in Computational Algebraic Number Theory, Graduate Texts in Maths 138, Springer-Verlag, 1993, Section 6.5, and subsection 6.5.5 in particular.

bnf [1] contains the matrix W, i.e. the matrix in Hermite normal form giving relations for the class group on prime ideal generators (p_{i})_{1≤i≤r}.

bnf [2] contains the matrix B, i.e. the matrix containing the expressions of the prime ideal factorbase in terms of the p_{i}. It is an r × c matrix.

bnf [3] contains the complex logarithmic embeddings of the system of fundamental units which has been found. It is an (r_{1} + r_{2}) × (r_{1} + r_{2} − 1) matrix.

bnf [4] contains the matrix M′′_{C} of Archimedean components of the relations of the matrix M′′, except that the first r_{1} + r_{2} − 1 columns are suppressed since they are already in bnf [3].

bnf [5] contains the prime factor base, i.e. the list of k prime ideals used in finding the relations.

bnf [6] contains the permutation of the prime factor base which was necessary to reduce the relation matrix to the form explained in subsection 6.5.5 of GTM 138 (i.e. with a big c × c identity matrix on the lower right). Note that in the above mentioned book, the need to permute the rows of the relation matrices which occur was not emphasized.

bnf [9] is a 3-element row vector obtained as follows. Let b = u_{1}^{−1}bnf [1]u_{2} obtained by applying the Smith normal form algorithm to the matrix W (= bnf [1]). Then bnf [9] = [u_{1}, u_{2}, b]. Note that the final class group generators given by bnfinit or bnfclassunit are obtained by LLL-reducing the generators whose list is b.

Finally, bnf [10] is unused and set equal to 0, but it is essential that this component be present, because PARI distinguishes a number field nf from a big number field bnf by the number of its components.

• The less technical components are as follows:

bnf [7] or bnf .nf is equal to the number field data nf as would be given by nfinit.

bnf [8] is a vector containing the last 6 components of bnfclassunit[,1], i.e. the classgroup bnf .clgp, the regulator bnf .reg, the general “check” number which should be close to 1, the number of roots of unity and a generator bnf .tu, the fundamental units bnf .fu, and finally the check on their computation. If the precision becomes insufficient, GP outputs a warning (fundamental units too large, not given) and does not strive to compute the units by default (flag = 0).

When flag = 1, GP insists on finding the fundamental units exactly, the internal precision being doubled and the computation redone, until the exact results are obtained. The user should be warned that this can take a very long time when the coefficients of the fundamental units on the integral basis are very large.

When flag = 2, on the contrary, it is initially agreed that GP will not compute units.

When flag = 3, computes a very small version of bnfinit, a “small big number field” (or sbnf for short) which contains enough information to recover the full bnf vector very rapidly, but which is much smaller and hence easy to store and print. It is supposed to be used in conjunction with bnfmake. The output is a 12 component vector v, as follows. Let bnf be the result of a full bnfinit, complete with units. Then v[1] is the polynomial P, v[2] is the number of real embeddings r_{1}, v[3] is the field discriminant, v[4] is the integral basis, v[5] is the list of roots as in the sixth component of nfinit, v[6] is the matrix MD of nfinit giving a Z-basis of the different, v[7] is the matrix W = bnf [1], v[8] is the matrix matalpha = bnf [2], v[9] is the prime ideal factor base bnf [5] coded in a compact way, and ordered according to the permutation bnf [6], v[10] is the 2-component vector giving the number of roots of unity and a generator, expressed on the integral basis, v[11] is the list of fundamental units, expressed on the integral basis, v[12] is a vector containing the algebraic numbers alpha corresponding to the columns of the matrix matalpha, expressed on the integral basis.

Note that all the components are exact (integral or rational), except for the roots in v[5]. In practice, this is the only component which a user is allowed to modify, by recomputing the roots to a higher accuracy if desired. Note also that the member functions will not work on sbnf , you have to use bnfmake explicitly first.

The library syntax is bnfinit0(P, flag, tech, prec).

3.6.6 bnfisintnorm(bnf , x): computes a complete system of solutions (modulo units of positive norm) of the absolute norm equation Norm(a) = x, where a is an integer in bnf . If bnf has not been certified, the correctness of the result depends on the validity of GRH.

The library syntax is bnfisintnorm(bnf , x).

3.6.7 bnfisnorm(bnf , x, {flag = 1}): tries to tell whether the rational number x is the norm of some element y in bnf . Returns a vector [a, b] where x = Norm(a) ∗ b. Looks for a solution which is an S-unit, with S a certain set of prime ideals containing (among others) all primes dividing x. If bnf is known to be Galois, set flag = 0 (in this case, x is a norm iff b = 1). If flag is non zero the program adds to S the following prime ideals, depending on the sign of flag. If flag > 0, the ideals of norm less than flag. And if flag < 0 the ideals dividing flag.

If you are willing to assume GRH, the answer is guaranteed (i.e. x is a norm iff b = 1), if S contains all primes less than 12 log(disc(Bnf ))^{2}, where Bnf is the Galois closure of bnf .

The library syntax is bnfisnorm(bnf , x, flag, prec), where flag and prec are longs.

3.6.8 bnfissunit(bnf , sfu, x): bnf being output by bnfinit, sfu by bnfsunit, gives the column vector of exponents of x on the fundamental S-units and the roots of unity. If x is not a unit, outputs an empty vector.

The library syntax is bnfissunit(bnf , sfu, x).

3.6.9 bnfisprincipal(bnf , x, {flag = 1}): bnf being the number field data output by bnfinit, and x being either a Z-basis of an ideal in the number field (not necessarily in HNF) or a prime ideal in the format output by the function idealprimedec, this function tests whether the ideal is principal or not. The result is more complete than a simple true/false answer: it gives a row vector [v_{1}, v_{2}, check], where

v_{1} is the vector of components c_{i} of the class of the ideal x in the class group, expressed on the generators g_{i} given by bnfinit (specifically bnf .clgp.gen which is the same as bnf [8][1][3]). The c_{i} are chosen so that 0 ≤ c_{i} < n_{i} where n_{i} is the order of g_{i} (the vector of n_{i} being

bnf .clgp.cyc, that is bnf [8][1][2]).

v_{2} gives on the integral basis the components of α such that x = α ∏_{i} g_{i}^{c_{i}}. In particular, x is principal if and only if v_{1} is equal to the zero vector, and if this the case x = αZ_{K} where α is given by v_{2}. Note that if α is too large to be given, a warning message will be printed and v_{2} will be set equal to the empty vector.

Finally the third component check is analogous to the last component of bnfclassunit: it gives a check on the accuracy of the result, in bits. check should be at least 10, and preferably much more. In any case, the result is checked for correctness.

If flag = 0, outputs only v_{1}, which is much easier to compute.

If flag = 2, does as if flag were 0, but doubles the precision until a result is obtained.

If flag = 3, as in the default behaviour (flag = 1), but doubles the precision until a result is obtained.

The user is warned that these two last setting may induce very lengthy computations.

The library syntax is isprincipalall(bnf , x, flag).

3.6.10 bnfisunit(bnf , x): bnf being the number field data output by bnfinit and x being an algebraic number (type integer, rational or polmod), this outputs the decomposition of x on the fundamental units and the roots of unity if x is a unit, the empty vector otherwise. More precisely, if u_{1},…,u_{r} are the fundamental units, and ζ is the generator of the group of roots of unity (found by bnfclassunit or bnfinit), the output is a vector [x_{1}, …, x_{r}, x_{r+1}] such that x = u_{1}^{x_{1}} ⋯ u_{r}^{x_{r}} ·ζ^{x_{r+1}}. The x_{i} are integers for i ≤ r and is an integer modulo the order of ζ for i = r + 1.

The library syntax is isunit(bnf , x).

3.6.11 bnfmake(sbnf ): sbnf being a “small bnf ” as output by bnfinit(x, 3), computes the complete bnfinit information. The result is not identical to what bnfinit would yield, but is functionally identical. The execution time is very small compared to a complete bnfinit. Note that if the default precision in GP (or prec in library mode) is greater than the precision of the roots sbnf [5], these are recomputed so as to get a result with greater accuracy.

Note that the member functions are not available for sbnf , you have to use bnfmake explicitly first.

The library syntax is makebigbnf(sbnf , prec), where prec is a C long integer.

3.6.12 bnfnarrow(bnf ): bnf being a big number field as output by bnfinit, computes the narrow class group of bnf . The output is a 3-component row vector v analogous to the corresponding class group component bnf .clgp (bnf [8][1]): the first component is the narrow class number v.no, the second component is a vector containing the SNF cyclic components v.cyc of the narrow class group, and the third is a vector giving the generators of the corresponding v.gen cyclic groups. Note that this function is a special case of bnrclass.

The library syntax is buchnarrow(bnf ).

3.6.13 bnfsignunit(bnf ): bnf being a big number field output by bnfinit, this computes an r_{1} × (r_{1} + r_{2} − 1) matrix having ±1 components, giving the signs of the real embeddings of the fundamental units.

The library syntax is signunits(bnf ).

3.6.14 bnfreg(bnf ): bnf being a big number field output by bnfinit, computes its regulator.

The library syntax is regulator(bnf , tech, prec), where tech is as in bnfclassunit.

3.6.15 bnfsunit(bnf , S): computes the fundamental S-units of the number field bnf (output by bnfinit), where S is a list of prime ideals (output by idealprimedec). The output is a vector v with 6 components.

v[1] gives a minimal system of (integral) generators of the S-unit group modulo the unit group.

v[2] contains technical data needed by bnfissunit.

v[3] is an empty vector (used to give the logarithmic embeddings of the generators in v[1] in version 2.0.16).

v[4] is the S-regulator (this is the product of the regulator, the determinant of v[2] and the natural logarithms of the norms of the ideals in S).

v[5] gives the S-class group structure, in the usual format (a row vector whose three components give in order the S-class number, the cyclic components and the generators).

v[6] is a copy of S.

The library syntax is bnfsunit(bnf , S, prec).

3.6.16 bnfunit(bnf ): bnf being a big number field as output by bnfinit, outputs a twocomponent row vector giving in the first component the vector of fundamental units of the number field, and in the second component the number of bit of accuracy which remained in the computation (which is always correct, otherwise an error message is printed). This function is mainly for people who used the wrong flag in bnfinit and would like to skip part of a lengthy bnfinit computation.

The library syntax is buchfu(bnf ).

3.6.17 bnrL1(bnr, {flag = 0}): bnr being the number field data which is output by bnrinit(,,1), returns for each character χ of the corresponding ray class group, the value at s = 1 (or s = 0) of the abelian L-functions associated to χ. For the value at s = 0, the function returns in fact for each character χ a vector [r_{χ}, c_{χ}] where r_{χ} is the order of L(s, χ) at s = 0 and c_{χ} the first non-zero term in the expansion of L(s, χ) at s = 0; in other words

L(s, χ) = c_{χ} · s^{r_{χ}} + O(s^{r_{χ}+1})

near 0. flag is optional, default value is 0; its binary digits mean 1: compute at s = 1 if set to 1 or s = 0 if set to 0, 2: compute the primitive L-functions associated to χ if set to 0 or the L-function with Euler factors at prime ideals dividing the modulus of bnr removed if set to 1 (this is the so-called L_{S}(s, χ) function where S is the set of infinite places of the number field together with the finite prime ideals dividing the modulus of bnr, see the example below), 3: returns also the character.

Example:

    bnf = bnfinit(x^2-229);
    bnr = bnrinit(bnf,1,1);
    bnrL1(bnr)

returns the order and the first non-zero term of the abelian L-functions L(s, χ) at s = 0 where χ runs through the characters of the class group of Q(\sqrt{229}). Then

    bnr2 = bnrinit(bnf,2,1);
    bnrL1(bnr2,2)

returns the order and the first non-zero terms of the abelian L-functions L_{S}(s, χ) at s = 0 where χ runs through the characters of the class group of Q(\sqrt{229}) and S is the set of infinite places of Q(\sqrt{229}) together with the finite prime 2 (note that the ray class group modulo 2 is in fact the class group, so bnrL1(bnr2) returns exactly the same answer as bnrL1(bnr)!).

The library syntax is bnrL1(bnr, flag, prec)

3.6.18 bnrclass(bnf , ideal, {flag = 0}): bnf being a big number field as output by bnfinit (the units are mandatory unless the ideal is trivial), and ideal being either an ideal in any form or a twocomponent row vector containing an ideal and an r_{1}-component row vector of flags indicating which real Archimedean embeddings to take in the module, computes the ray class group of the number field for the module ideal, as a 3-component vector as all other finite Abelian groups (cardinality, vector of cyclic components, corresponding generators).

If flag = 2, the output is different. It is a 6-component vector w. w[1] is bnf . w[2] is the result of applying idealstar(bnf , I, 2). w[3], w[4] and w[6] are technical components used only by the function bnrisprincipal. w[5] is the structure of the ray class group as would have been output with flag = 0.

If flag = 1, as above, except that the generators of the ray class group are not computed, which saves time.

The library syntax is bnrclass0(bnf , ideal, flag, prec).

3.6.19 bnrclassno(bnf , I): bnf being a big number field as output by bnfinit (units are mandatory unless the ideal is trivial), and I being either an ideal in any form or a two-component row vector containing an ideal and an r_{1}-component row vector of flags indicating which real Archimedean embeddings to take in the modulus, computes the ray class number of the number field for the modulus I. This is faster than bnrclass and should be used if only the ray class number is desired.

The library syntax is rayclassno(bnf , I).

3.6.20 bnrclassnolist(bnf , list): bnf being a big number field as output by bnfinit (units are mandatory unless the ideal is trivial), and list being a list of modules as output by ideallist of ideallistarch, outputs the list of the class numbers of the corresponding ray class groups.

The library syntax is rayclassnolist(bnf , list).

3.6.21 bnrconductor(a_{1}, {a_{2}}, {a_{3}}, {flag = 0}): conductor of the subfield of a ray class field as defined by [a_{1}, a_{2}, a_{3}] (see bnr at the beginning of this section).

The library syntax is bnrconductor(a_{1}, a_{2}, a_{3}, flag, prec), where an omitted argument among the a_{i} is input as gzero, and flag is a C long.

3.6.22 bnrconductorofchar(bnr, chi): bnr being a big ray number field as output by bnrclass, and chi being a row vector representing a character as expressed on the generators of the ray class group, gives the conductor of this character as a modulus.

The library syntax is bnrconductorofchar(bnr, chi, prec) where prec is a long.

3.6.23 bnrdisc(a1, {a2}, {a3}, {flag = 0}): a1, a2, a3 defining a big ray number field L over a groud field K (see bnr at the beginning of this section for the meaning of a1, a2, a3), outputs a 3-component row vector [N, R_{1}, D], where N is the (absolute) degree of L, R_{1} the number of real places of L, and D the discriminant of L/Q, including sign (if flag = 0).

If flag = 1, as above but outputs relative data. N is now the degree of L/K, R_{1} is the number of real places of K unramified in L (so that the number of real places of L is equal to R_{1} times the relative degree N), and D is the relative discriminant ideal of L/K.

If flag = 2, does as in case 0, except that if the modulus is not the exact conductor corresponding to the L, no data is computed and the result is 0 (gzero).

If flag = 3, as case 2, outputting relative data.

The library syntax is bnrdisc0(a1, a2, a3, flag, prec).

3.6.24 bnrdisclist(bnf , bound, {arch}, {flag = 0}): bnf being a big number field as output by bnfinit (the units are mandatory), computes a list of discriminants of Abelian extensions of the number field by increasing modulus norm up to bound bound, where the ramified Archimedean places are given by arch (unramified at infinity if arch is void or omitted). If flag is non-zero, give arch all the possible values. (See bnr at the beginning of this section for the meaning of a1, a2, a3.)

The alternative syntax bnrdisclist(bnf , list) is supported, where list is as output by ideal-

list or ideallistarch (with units).

The output format is as follows. The output v is a row vector of row vectors, allowing the bound to be greater than 2^{16} for 32-bit machines, and v[i][j] is understood to be in fact V [2^{15}(i − 1) + j] of a unique big vector V (note that 2^{15} is hardwired and can be increased in the source code only on 64-bit machines and higher).

Such a component V [k] is itself a vector W (maybe of length 0) whose components correspond to each possible ideal of norm k. Each component W [i] corresponds to an Abelian extension L of bnf whose modulus is an ideal of norm k and no Archimedean components (hence the extension is unramified at infinity). The extension W [i] is represented by a 4-component row vector [m, d, r, D] with the following meaning. m is the prime ideal factorization of the modulus, d = [L : Q] is the absolute degree of L, r is the number of real places of L, and D is the factorization of the absolute discriminant. Each prime ideal pr = [p, α, e, f, β] in the prime factorization m is coded as p · n^{2} + (f − 1) · n + (j − 1), where n is the degree of the base field and j is such that

    pr=idealprimedec(nf ,p)[j].

m can be decoded using bnfdecodemodule.

The library syntax is bnrdisclist0(a1, a2, a3, bound, arch, flag).

3.6.25 bnrinit(bnf , ideal, {flag = 0}): bnf is as output by bnfinit, ideal is a valid ideal (or a module), initializes data linked to the ray class group structure corresponding to this module. This is the same as bnrclass(bnf , ideal, flag + 1).

The library syntax is bnrinit0(bnf , ideal, flag, prec).

3.6.26 bnrisconductor(a1, {a2}, {a3}): a1, a2, a3 represent an extension of the base field, given by class field theory for some modulus encoded in the parameters. Outputs 1 if this modulus is the conductor, and 0 otherwise. This is slightly faster than bnrconductor.

The library syntax is bnrisconductor(a1, a2, a3) and the result is a long.

3.6.27 bnrisprincipal(bnr, x, {flag = 1}): bnr being the number field data which is output by bnrinit and x being an ideal in any form, outputs the components of x on the ray class group generators in a way similar to bnfisprincipal. That is a 3-component vector v where v[1] is the vector of components of x on the ray class group generators, v[2] gives on the integral basis an element α such that x = α ∏_{i} g_{i}^{x_{i}}. Finally v[3] indicates the number of bits of accuracy left in the result. In any case the result is checked for correctness, but v[3] is included to see if it is necessary to increase the accuracy in other computations.

If flag = 0, outputs only v_{1}.

The settings flag = 2 or 3 are not available in this case.

The library syntax is isprincipalrayall(bnr, x, flag).

3.6.28 bnrrootnumber(bnr, chi, {flag = 0}): if χ = chi is a (not necessarily primitive) character over bnr, let L(s, χ) = ∑_{id} χ(id)N(id)^{−s} be the associated Artin L-function. Returns the so-called Artin root number, i.e. the complex number W (χ) of modulus 1 such that

Λ(1 − s, χ) = W (χ)Λ(s, \overline{χ})

where Λ(s, χ) = A(χ)^{s/2}γ_{χ}(s)L(s, χ) is the enlarged L-function associated to L.

The generators of the ray class group are needed, and you can set flag = 1 if the character is known to be primitive. Example:

    bnf = bnfinit(x^2-145);
    bnr = bnrinit(bnf,7,1);
    bnrrootnumber(bnr, [5])

returns the root number of the character χ of Cl_{7}(Q(\sqrt{145})) such that χ(g) = ζ^{5}, where g is the generator of the ray-class field and ζ = e^{2iπ/N} where N is the order of g (N = 12 as bnr.cyc readily tells us).

The library syntax is bnrrootnumber(bnf , chi, flag)

3.6.29 bnrstark(bnr, subgroup, {flag = 0}): bnr being as output by bnrinit(,,1), finds a relative equation for the class field corresponding to the modulus in bnr and the given congruence subgroup using Stark units (set subgroup = 0 if you want the whole ray class group). The main variable of bnr must not be x, and the ground field and the class field must be totally real and not isomorphic to Q. flag is optional and may be set to 0 to obtain a reduced relative polynomial, 1 to be satisfied with any relative polynomial, 2 to obtain an absolute polynomial and 3 to obtain the irreducible relative polynomial of the Stark unit, 0 being default. Example:

    bnf = bnfinit(y^2-3);
    bnr = bnrinit(bnf,5,1);
    bnrstark(bnr,0)

returns the ray class field of Q(\sqrt{3}) modulo 5.

Remark. The function may fail, returning the error message

    "Cannot find a suitable modulus in FindModule".

In this case, the corresponding congruence group is a product of cyclic groups and, for the time being, the class field has to be obtained by splitting this group into its cyclic components.

The library syntax is bnrstark(bnr, subgroup, flag).

3.6.30 dirzetak(nf , b): gives as a vector the first b coefficients of the Dedekind zeta function of the number field nf considered as a Dirichlet series.

The library syntax is dirzetak(nf , b).

3.6.31 factornf(x, t): factorization of the univariate polynomial x over the number field defined by the (univariate) polynomial t. x may have coefficients in Q or in the number field. The main variable of t must be of lower priority than that of x (in other words the variable number of t must be greater than that of x). However if the coefficients of the number field occur explicitly (as polmods) as coefficients of x, the variable of these polmods must be the same as the main variable of

t. For example factornf(x^2 + Mod(y,y^2+1), y^2+1) and factornf(x^2+1, y^2+1) are legal
but factornf(x^2 + Mod(z,z^2+1), y^2+1) is not.

The library syntax is polfnf(x, t).

3.6.32 ffinit(p, n, {v = x}): computes a monic polynomial of degree n which is irreducible over F_{p}. For instance if P = ffinit(3,2,y), you can represent elements in F_{3^{2}} as polmods modulo P. This function is rather crude and expects p to be relatively small (p < 2^{3}1).

The library syntax is ffinit(p, n, v), where v is a variable number.

3.6.33 idealadd(nf , x, y): sum of the two ideals x and y in the number field nf . When x and y are given by Z-bases, this does not depend on nf and can be used to compute the sum of any two Z-modules. The result is given in HNF.

The library syntax is idealadd(nf , x, y).

3.6.34 idealaddtoone(nf , x, {y}): x and y being two co-prime integral ideals (given in any form), this gives a two-component row vector [a, b] such that a ∈ x, b ∈ y and a + b = 1.

The alternative syntax idealaddtoone(nf , v), is supported, where v is a k-component vector of ideals (given in any form) which sum to Z_{K}. This outputs a k-component vector e such that e[i] ∈ x[i] for 1 ≤ i ≤ k and ∑_{1≤i≤k} e[i] = 1.

The library syntax is idealaddtoone0(nf , x, y), where an omitted y is coded as NULL.

3.6.35 idealappr(nf , x, {flag = 0}): if x is a fractional ideal (given in any form), gives an element α in nf such that for all prime ideals p such that the valuation of x at p is non-zero, we have v_{p}(α) = v_{p}(x), and. v_{p}(α) ≥ 0 for all other p.

If flag is non-zero, x must be given as a prime ideal factorization, as output by idealfactor, but possibly with zero or negative exponents. This yields an element α such that for all prime ideals p occurring in x, v_{p}(α) is equal to the exponent of p in x, and for all other prime ideals, v_{p}(α) ≥ 0. This generalizes idealappr(nf , x, 0) since zero exponents are allowed. Note that the algorithm used is slightly different, so that idealapp(nf ,idealfactor(nf ,x)) may not be the

same as idealappr(nf ,x,1).

The library syntax is idealappr0(nf , x, flag).

3.6.36 idealchinese(nf , x, y): x being a prime ideal factorization (i.e. a 2 by 2 matrix whose first column contain prime ideals, and the second column integral exponents), y a vector of elements in nf indexed by the ideals in x, computes an element b such that

v_{p}(b − y_{p}) ≥ v_{p}(x) for all prime ideals in x and v_{p}(b) ≥ 0 for all other p.

The library syntax is idealchinese(nf , x, y).

3.6.37 idealcoprime(nf , x, y): given two integral ideals x and y in the number field nf , finds a β in the field, expressed on the integral basis nf [7], such that β · y is an integral ideal coprime to x.

The library syntax is idealcoprime(nf , x).

3.6.38 idealdiv(nf , x, y, {flag = 0}): quotient x · y^{−1} of the two ideals x and y in the number field nf . The result is given in HNF.

If flag is non-zero, the quotient x · y^{−1} is assumed to be an integral ideal. This can be much faster when the norm of the quotient is small even though the norms of x and y are large.

The library syntax is idealdiv0(nf , x, y, flag). Also available are idealdiv(nf , x, y) (flag = 0) and idealdivexact(nf , x, y) (flag = 1).

3.6.39 idealfactor(nf , x): factors into prime ideal powers the ideal x in the number field nf . The output format is similar to the factor function, and the prime ideals are represented in the form output by the idealprimedec function, i.e. as 5-element vectors.

The library syntax is idealfactor(nf , x).

3.6.40 idealhnf(nf , a, {b}): gives the Hermite normal form matrix of the ideal a. The ideal can be given in any form whatsoever (typically by an algebraic number if it is principal, by a Z_{K}-system of generators, as a prime ideal as given by idealprimedec, or by a Z-basis).

If b is not omitted, assume the ideal given was aZ_{K} + bZ_{K}, where a and b are elements of K given either as vectors on the integral basis nf [7] or as algebraic numbers.

The library syntax is idealhnf0(nf , a, b) where an omitted b is coded as NULL. Also available is idealhermite(nf , a) (b omitted).

3.6.41 idealintersect(nf , x, y): intersection of the two ideals x and y in the number field nf . When x and y are given by Z-bases, this does not depend on nf and can be used to compute the intersection of any two Z-modules. The result is given in HNF.

The library syntax is idealintersect(nf , x, y).

3.6.42 idealinv(nf , x, {flag = 0}): inverse of the ideal x in the number field nf . The result is the Hermite normal form of the inverse of the ideal, together with the opposite of the Archimedean information if it is given.

If flag = 1, uses the different. This is usually slower.

The library syntax is idealinv0(nf , x, flag). Also available is idealinv(nf , x) (flag = 0).

3.6.43 ideallist(nf , bound, {flag = 4}): computes the list of all ideals of norm less or equal to bound in the number field nf . The result is a row vector with exactly bound components. Each component is itself a row vector containing the information about ideals of a given norm, in no specific order. This information can be either the HNF of the ideal or the idealstar with possibly some additional information.

If flag is present, its binary digits are toggles meaning

1: give also the generators in the idealstar.

2: output [L, U], where L is as before and U is a vector of zinternallogs of the units.

4: give only the ideals and not the idealstar or the ideallog of the units.

The library syntax is ideallist0(nf , bound, flag), where bound must be a C long integer. Also available is ideallist(nf , bound), corresponding to the case flag = 0.

3.6.44 ideallistarch(nf , list, {arch = [ ]}, {flag = 0}): vector of vectors of all idealstarinit (see idealstar) of all modules in list, with Archimedean part arch added (void if omitted). list is a vector of big ideals, as output by ideallist(…, flag) for instance. flag is optional; its binary digits are toggles meaning: 1: give generators as well, 2: list format is [L, U] (see ideallist).

The library syntax is ideallistarch0(nf , list, arch, flag), where an omitted arch is coded as

NULL.

3.6.45 ideallog(nf , x, bid): nf being a number field, bid being a “big ideal” as output by idealstar and x being a non-necessarily integral element of nf which must have valuation equal to 0 at all prime ideals dividing I = bid[1], computes the “discrete logarithm” of x on the generators given in bid[2]. In other words, if g_{i} are these generators, of orders d_{i} respectively, the result is a column vector of integers (x_{i}) such that 0 ≤ x_{i} < d_{i} and

x ≡ ∏_{i} g_{i}^{x_{i}} (mod ^{∗}I) .

Note that when I is a module, this implies also sign conditions on the embeddings.

The library syntax is zideallog(nf , x, bid).

3.6.46 idealmin(nf , x, vdir): computes a minimum of the ideal x in the direction vdir in the number field nf .

The library syntax is minideal(nf , x, vdir, prec).

3.6.47 idealmul(nf , x, y, {flag = 0}): ideal multiplication of the ideals x and y in the number field nf . The result is a generating set for the ideal product with at most n elements, and is in Hermite normal form if either x or y is in HNF or is a prime ideal as output by idealprimedec, and this is given together with the sum of the Archimedean information in x and y if both are given.

If flag is non-zero, reduce the result using idealred.

The library syntax is idealmul(nf , x, y) (flag = 0) or idealmulred(nf , x, y, prec) (flag ≠ 0), where as usual, prec is a C long integer representing the precision.

3.6.48 idealnorm(nf , x): computes the norm of the ideal x in the number field nf .

The library syntax is idealnorm(nf , x).

3.6.49 idealpow(nf , x, k, {flag = 0}): computes the k-th power of the ideal x in the number field nf . k can be positive, negative or zero. The result is NOT reduced, it is really the k-th ideal power, and is given in HNF.

If flag is non-zero, reduce the result using idealred. Note however that this is NOT the same as as idealpow(nf , x, k) followed by reduction, since the reduction is performed throughout the powering process.

The library syntax corresponding to flag = 0 is idealpow(nf , x, k). If k is a long, you can use idealpows(nf , x, k). Corresponding to flag = 1 is idealpowred(nf , vp, k, prec), where prec is

a long.

3.6.50 idealprimedec(nf , p): computes the prime ideal decomposition of the prime number p in the number field nf . p must be a (positive) prime number. Note that the fact that p is prime is not checked, so if a non-prime number p is given it may lead to unpredictable results.

The result is a vector of 5-component vectors, each representing one of the prime ideals above p in the number field nf . The representation vp = [p, a, e, f, b] of a prime ideal means the following. The prime ideal is equal to pZ_{K} +αZ_{K} where Z_{K} is the ring of integers of the field and α = ∑_{i} a_{i}ω_{i} where the ω_{i} form the integral basis nf .zk, e is the ramification index, f is the residual index, and b is an n-component column vector representing a β ∈ Z_{K} such that vp^{−1} = Z_{K} + β/pZ_{K} which will be useful for computing valuations, but which the user can ignore. The number α is guaranteed to have a valuation equal to 1 at the prime ideal (this is automatic if e > 1).

The library syntax is idealprimedec(nf , p).

3.6.51 idealprincipal(nf , x): creates the principal ideal generated by the algebraic number x (which must be of type integer, rational or polmod) in the number field nf . The result is a onecolumn matrix.

The library syntax is principalideal(nf , x).

3.6.52 idealred(nf , x, {vdir = 0}): LLL reduction of the ideal x in the number field nf , along the direction vdir. Here vdir must be either an r1 + r2-component vector (r1 and r2 number of real and complex places of nf as usual), or the PARI zero, in which case vdir is assumed to be equal to the vector having only components equal to 1. The notion of reduction along a direction is technical and cannot be explained here. Note that this is not the same as the LLL reduction of the lattice x since ideal operations are involved. The result is the Hermite normal form of the LLL-reduced ideal, which is usually, but not always, a reduced ideal. x may also be a 2-component vector, the first being as above, and the second containing a matrix of Archimedean information. In that case, this matrix is suitably updated.

The library syntax is ideallllred(nf , x, vdir, prec).

3.6.53 idealstar(nf , I, {flag = 1}): nf being a number field, and I either and ideal in any form, or a row vector whose first component is an ideal and whose second component is a row vector of r_{1} 0 or 1, outputs necessary data for computing in the group (Z_{K}/I)^{∗}.

If flag = 2, the result is a 5-component vector w. w[1] is the ideal or module I itself. w[2] is the structure of the group. The other components are difficult to describe and are used only in conjunction with the function ideallog.

If flag = 1 (default), as flag = 2, but do not compute explicit generators for the cyclic components, which saves time.

If flag = 0, computes the structure of (Z_{K}/I)^{∗} as a 3-component vector v. v[1] is the order, v[2] is the vector of SNF cyclic components and v[3] the corresponding generators. When the row vector is explicitly included, the non-zero elements of this vector are considered as real embeddings of nf in the order given by polroots, i.e. in nf [6] (nf .roots), and then I is a module with components at infinity.

To solve discrete logarithms (using ideallog), you have to choose flag = 2.

The library syntax is idealstar0(nf , I, flag).

3.6.54 idealtwoelt(nf , x, {a}): computes a two-element representation of the ideal x in the number field nf , using a straightforward (exponential time) search. x can be an ideal in any form, (including perhaps an Archimedean part, which is ignored) and the result is a row vector [a, α] with two components such that x = aZ_{K} + αZ_{K} and a ∈ Z, where a is the one passed as argument if any. If x is given by at least two generators, a is chosen to be the positive generator of x ∩ Z.

Note that when an explicit a is given, we use an asymptotically faster method, however in practice it is usually slower.

The library syntax is ideal two elt0(nf , x, a), where an omitted a is entered as NULL.

3.6.55 idealval(nf , x, vp): gives the valuation of the ideal x at the prime ideal vp in the number field nf , where vp must be a 5-component vector as given by idealprimedec.

The library syntax is idealval(nf , x, vp), and the result is a long integer.

3.6.56 ideleprincipal(nf , x): creates the principal idele generated by the algebraic number x (which must be of type integer, rational or polmod) in the number field nf . The result is a twocomponent vector, the first being a one-column matrix representing the corresponding principal ideal, and the second being the vector with r_{1} + r_{2} components giving the complex logarithmic embedding of x.

The library syntax is principalidele(nf , x).

3.6.57 matalgtobasis(nf , x): nf being a number field in nfinit format, and x a matrix whose coefficients are expressed as polmods in nf , transforms this matrix into a matrix whose coefficients are expressed on the integral basis of nf . This is the same as applying nfalgtobasis to each entry, but it would be dangerous to use the same name.

The library syntax is matalgtobasis(nf , x).

3.6.58 matbasistoalg(nf , x): nf being a number field in nfinit format, and x a matrix whose coefficients are expressed as column vectors on the integral basis of nf , transforms this matrix into a matrix whose coefficients are algebraic numbers expressed as polmods. This is the same as applying nfbasistoalg to each entry, but it would be dangerous to use the same name.

The library syntax is matbasistoalg(nf , x).

3.6.59 modreverse(a): a being a polmod A(X) modulo T (X), finds the “reverse polmod” B(X) modulo Q(X), where Q is the minimal polynomial of a, which must be equal to the degree of T, and such that if θ is a root of T then θ = B(α) for a certain root α of Q.

This is very useful when one changes the generating element in algebraic extensions.

The library syntax is polmodrecip(x).

3.6.60 newtonpoly(x, p): gives the vector of the slopes of the Newton polygon of the polynomial x with respect to the prime number p. The n components of the vector are in decreasing order, where n is equal to the degree of x. Vertical slopes occur iff the constant coefficient of x is zero and are denoted by VERYBIGINT, the biggest single precision integer representable on the machine (2^{31} − 1 (resp. 2^{63} − 1) on 32-bit (resp. 64-bit) machines), see Section 3.2.38.

The library syntax is newtonpoly(x, p).

3.6.61 nfalgtobasis(nf , x): this is the inverse function of nfbasistoalg. Given an object x whose entries are expressed as algebraic numbers in the number field nf , transforms it so that the entries are expressed as a column vector on the integral basis nf .zk.

The library syntax is algtobasis(nf , x).

3.6.62 nfbasis(x, {flag = 0}, {p}): integral basis of the number field defined by the irreducible, preferably monic, polynomial x, using the round 4 algorithm by default. (This program is the translation into C by Pascal Letard of a program written by David Ford in Maple.) The binary digits of flag have the following meaning:

1: assume that no square of a prime greater than the default primelimit divides the discriminant of x, i.e. that the index of x has only small prime divisors.

2: use round 2 algorithm. For small degrees and coefficient size, this is sometimes a little faster. (This program is the translation into C of a program written by David Ford in Algeb.)

Thus for instance, if flag = 3, this uses the round 2 algorithm and outputs an order which will be maximal at all the small primes.

If p is present, we assume (without checking!) that it is the two-column matrix of the factorization of the discriminant of the polynomial x. Note that it does not have to be a complete factorization. This is especially useful if only a local integral basis for some small set of places is desired: only factors with exponents greater or equal to 2 will be considered.

The library syntax is nfbasis0(x, flag, p). An extended version is nfbasis(x, &d, flag, p), where d will receive the discriminant of the number field (not of the polynomial x), and an omitted p should be input as gzero. Also available are base(x, &d) (flag = 0), base2(x, &d) (flag = 2) and factoredbase(x, p, &d).

3.6.63 nfbasistoalg(nf , x): this is the inverse function of nfalgtobasis. Given an object x whose entries are expressed on the integral basis nf .zk, transforms it into an object whose entries are algebraic numbers (i.e. polmods).

The library syntax is basistoalg(nf , x).

3.6.64 nfdetint(nf , x): given a pseudo-matrix x, computes a non-zero ideal contained in (i.e. multiple of) the determinant of x. This is particularly useful in conjunction with nfhnfmod.

The library syntax is nfdetint(nf , x).

3.6.65 nfdisc(x, {flag = 0}, {p}): field discriminant of the number field defined by the integral, preferably monic, irreducible polynomial x. flag and p are exactly as in nfbasis. That is, p provides the matrix of a partial factorization of the discriminant of x, and binary digits of flag are as follows:

1: assume that no square of a prime greater than primelimit divides the discriminant.

2: use the round 2 algorithm, instead of the default round 4. This should be slower except maybe for polynomials of small degree and coefficients.

The library syntax is nfdiscf0(x, flag, p) where, to omit p, you should input gzero. You can also use discf(x) (flag = 0).

3.6.66 nfeltdiv(nf , x, y): given two elements x and y in nf , computes their quotient x/y in the number field nf .

The library syntax is element div(nf , x, y).

3.6.67 nfeltdiveuc(nf , x, y): given two elements x and y in nf , computes an algebraic integer q in the number field nf such that the components of x − qy are reasonably small. In fact, this is functionally identical to round(nfeltdiv(nf ,x,y)).

The library syntax is nfdiveuc(nf , x, y).

3.6.68 nfeltdivmodpr(nf , x, y, pr): given two elements x and y in nf and pr a prime ideal in modpr format (see nfmodprinit), computes their quotient x/y modulo the prime ideal pr.

The library syntax is element divmodpr(nf , x, y, pr).

3.6.69 nfeltdivrem(nf , x, y): given two elements x and y in nf , gives a two-element row vector [q, r] such that x = qy + r, q is an algebraic integer in nf , and the components of r are reasonably small.

The library syntax is nfdivres(nf , x, y).

3.6.70 nfeltmod(nf , x, y): given two elements x and y in nf , computes an element r of nf of the form r = x−qy with q and algebraic integer, and such that r is small. This is functionally identical to

                     x − nfeltmul(nf , round(nfeltdiv(nf , x, y)), y).

The library syntax is nfmod(nf , x, y).

3.6.71 nfeltmul(nf , x, y): given two elements x and y in nf , computes their product x ∗ y in the number field nf .

The library syntax is element mul(nf , x, y).

3.6.72 nfeltmulmodpr(nf , x, y, pr): given two elements x and y in nf and pr a prime ideal in modpr format (see nfmodprinit), computes their product x ∗ y modulo the prime ideal pr.

The library syntax is element mulmodpr(nf , x, y, pr).

3.6.73 nfeltpow(nf , x, k): given an element x in nf , and a positive or negative integer k, computes x^{k} in the number field nf .

The library syntax is element pow(nf , x, k).

3.6.74 nfeltpowmodpr(nf , x, k, pr): given an element x in nf , an integer k and a prime ideal pr in modpr format (see nfmodprinit), computes x^{k} modulo the prime ideal pr.

The library syntax is element powmodpr(nf , x, k, pr).

3.6.75 nfeltreduce(nf , x, ideal): given an ideal in Hermite normal form and an element x of the number field nf , finds an element r in nf such that x − r belongs to the ideal and r is small.

The library syntax is element reduce(nf , x, ideal).

3.6.76 nfeltreducemodpr(nf , x, pr): given an element x of the number field nf and a prime ideal pr in modpr format compute a canonical representative for the class of x modulo pr.

The library syntax is nfreducemodpr2(nf , x, pr).

3.6.77 nfeltval(nf , x, pr): given an element x in nf and a prime ideal pr in the format output by idealprimedec, computes their the valuation at pr of the element x. The same result could be obtained using idealval(nf ,x,pr) (since x would then be converted to a principal ideal), but it would be less efficient.

The library syntax is element val(nf , x, pr), and the result is a long.

3.6.78 nffactor(nf , x): factorization of the univariate polynomial x over the number field nf given by nfinit. x has coefficients in nf (i.e. either scalar, polmod, polynomial or column vector). The main variable of nf must be of lower priority than that of x (in other words, the variable number of nf must be greater than that of x). However if the polynomial defining the number field occurs explicitly in the coefficients of x (as modulus of a t POLMOD), its main variable must be the same as the

main variable of x. For example, if nf = nfinit(y^2+1) then nffactor(nf ,x^2+Mod(y,y^2+1))
and nffactor(nf ,x^2+1) are both legal but nffactor(nf ,x^2+Mod(z,z^2+1)) is not.

The library syntax is nffactor(nf , x).

3.6.79 nffactormod(nf , x, pr): factorization of the univariate polynomial x modulo the prime ideal pr in the number field nf . x can have coefficients in the number field (scalar, polmod, polynomial, column vector) or modulo the prime ideal (integermod modulo the rational prime under pr, polmod or polynomial with integermod coefficients, column vector of integermod). The prime ideal pr must be in the format output by idealprimedec. The main variable of nf must be of lower priority than that of x (in other words the variable number of nf must be greater than that of x). However if the coefficients of the number field occur explicitly (as polmods) as coefficients of x, the variable of these polmods must be the same as the main variable of t (see nffactor).

The library syntax is nffactormod(nf , x, pr).

3.6.80 nfgaloisapply(nf , aut, x): nf being a number field as output by nfinit, and aut being a Galois automorphism of nf expressed either as a polynomial or a polmod (such automorphisms being found using for example one of the variants of nfgaloisconj), computes the action of the automorphism aut on the object x in the number field. x can be an element (scalar, polmod, polynomial or column vector) of the number field, an ideal (either given by Z_{K}-generators or by a Z-basis), a prime ideal (given as a 5-element row vector) or an idele (given as a 2-element row vector). Because of possible confusion with elements and ideals, other vector or matrix arguments are forbidden.

The library syntax is galoisapply(nf , aut, x).

3.6.81 nfgaloisconj(nf , {flag = 0}, {d}): nf being a number field as output by nfinit, computes the conjugates of a root r of the non-constant polynomial x = nf [1] expressed as polynomials in r. This can be used even if the number field nf is not Galois since some conjugates may lie in the field. As a note to old-timers of PARI, starting with version 2.0.17 this function works much better than in earlier versions.

nf can simply be a polynomial if flag ≠ 1.

If no flags or flag = 0, if nf is a number field use a combination of flag 4 and 1 and the result is always complete, else use a combination of flag 4 and 2 and the result is subject to the restriction of flag = 2.

If flag = 1, use nfroots (require a number field).

If flag = 2, use complex approximations to the roots and an integral LLL. The result is not guaranteed to be complete: some conjugates may be missing (especially so if the corresponding polynomial has a huge index). In that case, increasing the default precision may help.

If flag = 4, use Allombert’s algorithm and permutation testing. If the field is Galois with “weakly” super solvable Galois group, return the complete list of automorphisms, else only the identity element. If present, d is assumed to be a multiple of the index of the power basis in the maximal order.

A group G is “weakly” super solvable if it contains a super solvable normal subgroup H such that G = H , or G/H ≃ A_{4} , or G/H ≃ S_{4}. Abelian and nilpotent groups are “weakly” super solvable. In practice, almost all groups of small order are weakly super solvable, the exceptions having order 36(1 exception), 48(2), 56(1), 60(1), 72(5), 75(1), 80(1), 96(10) and ≥ 108.

Hence flag = 4 permits to quickly check whether a polynomial of order strictly less than 36 is Galois or not. This method is much faster than nfroots and be applied to polynomial of degree more than 50.

The library syntax is galoisconj0(nf , flag, d, prec). Also available are galoisconj(nf ) for flag = 0, galoisconj2(nf , n, prec) for flag = 2 where n is a bound on the number of conjugates, and galoisconj4(nf , d) corresponding to flag = 4.

3.6.82 nfhilbert(nf , a, b, {pr}): if pr is omitted, compute the global Hilbert symbol (a, b) in nf , that is 1 if x^{2} − ay^{2} − bz^{2} has a non trivial solution (x, y, z) in nf , and −1 otherwise. Otherwise compute the local symbol modulo the prime ideal pr (as output by idealprimedec).

The library syntax is nfhilbert(nf , a, b, pr), where an omitted pr is coded as NULL.

3.6.83 nfhnf(nf , x): given a pseudo-matrix (A, I), finds a pseudo-basis in Hermite normal form of the module it generates.

The library syntax is nfhermite(nf , x).

3.6.84 nfhnfmod(nf , x, detx): given a pseudo-matrix (A, I) and an ideal detx which is contained in (read integral multiple of) the determinant of (A, I), finds a pseudo-basis in Hermite normal form of the module generated by (A, I). This avoids coefficient explosion. detx can be computed using the function nfdetint.

The library syntax is nfhermitemod(nf , x, detx).

3.6.85 nfinit(pol, {flag = 0}): pol being a non-constant, preferably monic, irreducible polynomial in Z[X], computes a 9-component vector nf useful in working in the number field K defined by pol.

nf [1] contains the polynomial pol (nf .pol).

nf [2] contains [r1, r2] (nf .sign), the number of real and complex places of K.

nf [3] contains the discriminant d(K) (nf .disc) of the number field K.

nf [4] contains the index of nf [1], i.e. [Z_{K} : Z[θ]], where θ is any root of nf [1].

nf [5] is a vector containing 7 matrices M, MC, T 2, T, MD, T I, MDI useful for certain computations in the number field K.

• M is the (r1+r2)×n matrix whose columns represent the numerical values of the conjugates of the elements of the integral basis.

• MC is essentially the conjugate of the transpose of M, except that the last r2 columns are also multiplied by 2.

• T 2 is an n×n matrix equal to the real part of the product MC ·M (which is a real positive definite symmetric matrix), the so-called T_{2}-matrix (nf .t2).

• T is the n × n matrix whose coefficients are Tr(ω_{i}ω_{j}) where the ω_{i} are the elements of the integral basis. Note that T = \overline{MC} · M and in particular that T = T_{2} if the field is totally real (in practice T_{2} will have real approximate entries and T will have integer entries). Note also that det(T) is equal to the discriminant of the field K.

• The columns of MD (nf .diff) express a Z-basis of the different of K on the integral basis.

• T I is equal to d(K)T^{−1}, which has integral coefficients.

• Finally, MDI has the form [x, y, n], where (x, y) expresses a Z_{K}-basis of d(K) times the codifferent ideal (nf .disc∗nf .codiff, which is an integral ideal) and n is its norm (this ideal is

used in idealinv).

nf [6] is the vector containing the r1+r2 roots (nf .roots) of nf [1] corresponding to the r1+r2 embeddings of the number field into C (the first r1 components are real, the next r2 have positive imaginary part).

nf [7] is an integral basis in Hermite normal form for Z_{K} (nf .zk) expressed on the powers of θ.

nf [8] is the n × n integral matrix expressing the power basis in terms of the integral basis, and finally

nf [9] is the n × n^{2} matrix giving the multiplication table of the integral basis.

If a non monic polynomial is input, nfinit will transform it into a monic one, then reduce it (see flag = 3). It is allowed, though not very useful given the existence of nfnewprec, to input a nf or a bnf instead of a polynomial.

The special input format [x, B] is also accepted where x is a polynomial as above and B is the integer basis, as computed by nfbasis. This can be useful since nfinit uses the round 4 algorithm by default, which can be very slow in pathological cases where round 2 (nfbasis(x,2)) would succeed very quickly.

If flag = 1: does not compute the different, replace it by a dummy 0.

If flag = 2: pol is changed into another polynomial P defining the same number field, which is as simple as can easily be found using the polred algorithm, and all the subsequent computations are done using this new polynomial. In particular, the first component of the result is the modified polynomial.

If flag = 3, does a polred as in case 2, but outputs [nf , Mod(a, P)], where nf is as before and Mod(a, P) = Mod(x, pol) gives the change of variables. This is implicit when pol is not monic: first a linear change of variables is performed, to get a monic polynomial, then a polred reduction.

If flag = 4, as 2 but uses a partial polred.

If flag = 5, as 3 using a partial polred.

The library syntax is nfinit0(x, flag, prec).

3.6.86 nfisideal(nf , x): returns 1 if x is an ideal in the number field nf , 0 otherwise.

The library syntax is isideal(x).

3.6.87 nfisincl(x, y): tests whether the number field K defined by the polynomial x is conjugate to a subfield of the field L defined by y (where x and y must be in Q[X]). If they are not, the output is the number 0. If they are, the output is a vector of polynomials, each polynomial a representing an embedding of K into L, i.e. being such that y | x ∘ a.

If y is a number field (nf ), a much faster algorithm is used (factoring x over y using nffactor). Before version 2.0.14, this wasn’t guaranteed to return all the embeddings, hence was triggered by a special flag. This is no more the case.

The library syntax is nfisincl(x, y, flag).

3.6.88 nfisisom(x, y): as nfisincl, but tests for isomorphism. If either x or y is a number field, a much faster algorithm will be used.

The library syntax is nfisisom(x, y, flag).

3.6.89 nfnewprec(nf ): transforms the number field nf into the corresponding data using current (usually larger) precision. This function works as expected if nf is in fact a bnf (update bnf to current precision) but may be quite slow (many generators of principal ideals have to be computed).

The library syntax is nfnewprec(nf , prec).

3.6.90 nfkermodpr(nf , a, pr): kernel of the matrix a in Z_{K}/pr, where pr is in modpr format

(see nfmodprinit).

The library syntax is nfkermodpr(nf , a, pr).

3.6.91 nfmodprinit(nf , pr): transforms the prime ideal pr into modpr format necessary for all operations modulo pr in the number field nf . Returns a two-component vector [P, a], where P is the Hermite normal form of pr, and a is an integral element congruent to 1 modulo pr, and congruent to 0 modulo p/pr^{e}. Here p = Z ∩ pr and e is the absolute ramification index.

The library syntax is nfmodprinit(nf , pr).

3.6.92 nfsubfields(nf , {d = 0}): finds all subfields of degree d of the number field nf (all subfields if d is null or omitted). The result is a vector of subfields, each being given by [g, h], where g is an absolute equation and h expresses one of the roots of g in terms of the root x of the polynomial defining nf . This is a crude implementation by M. Olivier of an algorithm due to J. Klüners.

The library syntax is subfields(nf , d).

3.6.93 nfroots(nf , x): roots of the polynomial x in the number field nf given by nfinit without multiplicity. x has coefficients in the number field (scalar, polmod, polynomial, column vector). The main variable of nf must be of lower priority than that of x (in other words the variable number of nf must be greater than that of x). However if the coefficients of the number field occur explicitly (as polmods) as coefficients of x, the variable of these polmods must be the same as the main variable of t (see nffactor).

The library syntax is nfroots(nf , x).

3.6.94 nfrootsof1(nf ): computes the number of roots of unity w and a primitive w-th root of unity (expressed on the integral basis) belonging to the number field nf . The result is a twocomponent vector [w, z] where z is a column vector expressing a primitive w-th root of unity on the integral basis nf .zk.

The library syntax is rootsof1(nf ).

3.6.95 nfsnf(nf , x): given a torsion module x as a 3-component row vector [A, I, J] where A is a square invertible n × n matrix, I and J are two ideal lists, outputs an ideal list d_{1}, …, d_{n} which is the Smith normal form of x. In other words, x is isomorphic to Z_{K}/d_{1} ⊕ ⋯ ⊕ Z_{K}/d_{n} and d_{i} divides d_{i−1} for i ≥ 2. The link between x and [A, I, J] is as follows: if e_{i} is the canonical basis of K^{n}, I = [b_{1}, …, b_{n}] and J = [a_{1}, …, a_{n}], then x is isomorphic to

(b_{1}e_{1} ⊕ ⋯ ⊕ b_{n}e_{n})/(a_{1}A_{1} ⊕ ⋯ ⊕ a_{n}A_{n}) ,

where the A_{j} are the columns of the matrix A. Note that every finitely generated torsion module can be given in this way, and even with b_{i} = Z_{K} for all i.

The library syntax is nfsmith(nf , x).

3.6.96 nfsolvemodpr(nf , a, b, pr): solution of a · x = b in Z_{K}/pr, where a is a matrix and b a column vector, and where pr is in modpr format (see nfmodprinit).

The library syntax is nfsolvemodpr(nf , a, b, pr).

3.6.97 polcompositum(x, y, {flag = 0}): x and y being polynomials in Z[x] in the same variable, outputs a vector giving the list of all possible composita of the number fields defined by x and y, if x and y are irreducible, or of the corresponding étale algebras, if they are only squarefree. Returns an error if one of the polynomials is not squarefree.

If flag = 1, outputs a vector of 4-component vectors [z, a, b, k], where z ranges through the list of all possible compositums as above, and a (resp. b) expresses the root of x (resp. y) as a polmod in a root of z, and k is a small integer k such that a + kb is the chosen root of z.

The library syntax is polcompositum0(x, y, flag).

3.6.98 polgalois(x): Galois group of the non-constant polynomial x ∈ Q[X]. In the present version 2.0.17, x must be irreducible and the degree of x must be less than or equal to 7. On certain versions for which the data file of Galois resolvents has been installed (available in the Unix distribution as a separate package), degrees 8, 9, 10 and 11 are also implemented.

The output is a 3-component vector [n, s, k] with the following meaning: n is the cardinality of the group, s is its signature (s = 1 if the group is a subgroup of the alternating group A_{n}, s = −1 otherwise), and k is the number of the group corresponding to a given pair (n, s) (k = 1 except in 2 cases). Specifically, the groups are coded as follows, using standard notations (see GTM 138, quoted at the beginning of this section):

In degree 1: S_{1} = [1, −1, 1].
In degree 2: S_{2} = [2, −1, 1].
In degree 3: A_{3} = C_{3} = [3, 1, 1], S_{3} = [6, −1, 1].
In degree 4: C_{4} = [4, −1, 1], V_{4} = [4, 1, 1], D_{4} = [8, −1, 1], A_{4} = [12, 1, 1], S_{4} = [24, −1, 1].

In degree 5: C_{5} = [5, 1, 1], D_{5} = [10, 1, 1], M_{20} = [20, −1, 1], A_{5} = [60, 1, 1], S_{5} = [120, −1, 1].

In degree 6: C_{6} = [6, −1, 1], S_{3} = [6, −1, 2], D_{6} = [12, −1, 1], A_{4} = [12, 1, 1], G_{18} = [18, −1, 1], S_{4}^{−} = [24, −1, 1], A_{4} × C_{2} = [24, −1, 2], S_{4}^{+} = [24, 1, 1], G_{36}^{−} = [36, −1, 1], G_{36}^{+} = [36, 1, 1], S_{4} × C_{2} = [48, −1, 1], A_{5} = P SL_{2}(5) = [60, 1, 1], G_{72} = [72, −1, 1], S_{5} = P GL_{2}(5) = [120, −1, 1], A_{6} = [360, 1, 1], S_{6} = [720, −1, 1].

In degree 7: C_{7} = [7, 1, 1], D_{7} = [14, −1, 1], M_{21} = [21, 1, 1], M_{42} = [42, −1, 1], P SL_{2}(7) = P SL_{3}(2) = [168, 1, 1], A_{7} = [2520, 1, 1], S_{7} = [5040, −1, 1].

The method used is that of resolvent polynomials and is sensitive to the current precision. The precision is updated internally but, in very rare cases, a wrong result may be returned if the initial precision was not sufficient.

The library syntax is galois(x, prec).

3.6.99 polred(x, {flag = 0}, {p}): finds polynomials with reasonably small coefficients defining subfields of the number field defined by x. One of the polynomials always defines Q (hence is equal to x − 1), and another always defines the same number field as x if x is irreducible. All x accepted by nfinit are also allowed here (e.g. non-monic polynomials, nf, bnf, [x,Z K basis]).

The following binary digits of flag are significant:

1: does a partial reduction only. This means that only a suborder of the maximal order may be used.

2: gives also elements. The result is a two-column matrix, the first column giving the elements defining these subfields, the second giving the corresponding minimal polynomials.

If p is given, it is assumed that it is the two-column matrix of the factorization of the discriminant of the polynomial x.

The library syntax is polred0(x, flag, p, prec), where an omitted p is coded by gzero. Also available are polred(x, prec) and factoredpolred(x, p, prec), both corresponding to flag = 0.

3.6.100 polredabs(x, {flag = 0}): finds one of the polynomial defining the same number field as the one defined by x, and such that the sum of the squares of the modulus of the roots (i.e. the T_{2}-norm) is minimal. All x accepted by nfinit are also allowed here (e.g. non-monic polynomials,

nf, bnf, [x,Z K basis]).

The binary digits of flag mean

1: outputs a two-component row vector [P, a], where P is the default output and a is an element expressed on a root of the polynomial P, whose minimal polynomial is equal to x.

4: gives all polynomials of minimal T_{2} norm (of the two polynomials P (x) and P (−x), only one is given).

The library syntax is polredabs0(x, flag, prec).

3.6.101 polredord(x): finds polynomials with reasonably small coefficients and of the same degree as that of x defining suborders of the order defined by x. One of the polynomials always defines Q (hence is equal to (x − 1)^{n}, where n is the degree), and another always defines the same order as x if x is irreducible.

The library syntax is ordred(x).

3.6.102 poltschirnhaus(x): applies a random Tschirnhausen transformation to the polynomial x, which is assumed to be non-constant and separable, so as to obtain a new equation for the étale algebra defined by x. This is for instance useful when computing resolvents, hence is used by the

polgalois function.

The library syntax is tschirnhaus(x).

3.6.103 rnfalgtobasis(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an element of L expressed as a polynomial or polmod with polmod coefficients, expresses x on the relative integral basis.

The library syntax is rnfalgtobasis(rnf , x).

3.6.104 rnfbasis(bnf , x): given a big number field bnf as output by bnfinit, and either a polynomial x with coefficients in bnf defining a relative extension L of bnf , or a pseudo-basis x of such an extension, gives either a true bnf -basis of L if it exists, or an n + 1-element generating set of L if not, where n is the rank of L over bnf .

The library syntax is rnfbasis(bnf , x).

3.6.105 rnfbasistoalg(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an element of L expressed on the relative integral basis, computes the representation of x as a polmod with polmods coefficients.

The library syntax is rnfbasistoalg(rnf , x).

3.6.106 rnfcharpoly(nf , T, a, {v = x}): characteristic polynomial of a over nf , where a belongs to the algebra defined by T over nf , i.e. nf [X]/(T). Returns a polynomial in variable v (x by default).

The library syntax is rnfcharpoly(nf , T, a, v), where v is a variable number.

3.6.107 rnfconductor(bnf , pol): bnf being a big number field as output by bnfinit, and pol a relative polynomial defining an Abelian extension, computes the class field theory conductor of this Abelian extension. The result is a 3-component vector [conductor, rayclgp, subgroup], where conductor is the conductor of the extension given as a 2-component row vector [f_{0}, f_{∞}], rayclgp is the full ray class group corresponding to the conductor given as a 3-component vector [h,cyc,gen] as usual for a group, and subgroup is a matrix in HNF defining the subgroup of the ray class group on the given generators gen.

The library syntax is rnfconductor(rnf , pol, prec).

3.6.108 rnfdedekind(nf , pol, pr): given a number field nf as output by nfinit and a polynomial pol with coefficients in nf defining a relative extension L of nf , evaluates the relative Dedekind criterion over the order defined by a root of pol for the prime ideal pr and outputs a 3-component vector as the result. The first component is a flag equal to 1 if the enlarged order is pr-maximal and to 0 otherwise, the second component is a pseudo-basis of the enlarged order and the third component is the valuation at pr of the order discriminant.

The library syntax is rnfdedekind(nf , pol, pr).

3.6.109 rnfdet(nf , M): given a pseudomatrix M over the maximal order of nf , computes its pseudodeterminant.

The library syntax is rnfdet(nf , M).

3.6.110 rnfdisc(nf , pol): given a number field nf as output by nfinit and a polynomial pol with coefficients in nf defining a relative extension L of nf , computes the relative discriminant of L. This is a two-element row vector [D, d], where D is the relative ideal discriminant and d is the relative discriminant considered as an element of nf^{∗}/nf^{∗}^{2}. The main variable of nf must be of lower priority than that of pol.

Note: As usual, nf can be a bnf as output by nfinit.

The library syntax is rnfdiscf(bnf , pol).

3.6.111 rnfeltabstorel(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an element of L expressed as a polynomial modulo the absolute equation rnf [11][1], computes x as an element of the relative extension L/K as a polmod with polmod coefficients.

The library syntax is rnfelementabstorel(rnf , x).

3.6.112 rnfeltdown(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an element of L expressed as a polynomial or polmod with polmod coefficients, computes x as an element of K as a polmod, assuming x is in K (otherwise an error will occur). If x is given on the relative integral basis, apply rnfbasistoalg first, otherwise PARI will believe you are dealing with a vector.

The library syntax is rnfelementdown(rnf , x).

3.6.113 rnfeltreltoabs(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an element of L expressed as a polynomial or polmod with polmod coefficients, computes x as an element of the absolute extension L/Q as a polynomial modulo the absolute equation rnf [11][1]. If x is given on the relative integral basis, apply rnfbasistoalg first, otherwise PARI will believe you are dealing with a vector.

The library syntax is rnfelementreltoabs(rnf , x).

3.6.114 rnfeltup(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an element of K expressed as a polynomial or polmod, computes x as an element of the absolute extension L/Q as a polynomial modulo the absolute equation rnf [11][1]. Note that it is unnecessary to compute x as an element of the relative extension L/K (its expression would be identical to itself). If x is given on the integral basis of K, apply nfbasistoalg first, otherwise PARI will believe you are dealing with a vector.

The library syntax is rnfelementup(rnf , x).

3.6.115 rnfequation(nf , pol, {flag = 0}): given a number field nf as output by nfinit (or simply a polynomial) and a polynomial pol with coefficients in nf defining a relative extension L of nf , computes the absolute equation of L over Q.

If flag is non-zero, outputs a 3-component row vector [z, a, k], where z is the absolute equation of L over Q, as in the default behaviour, a expresses as a polmod a root β of pol in terms of a root θ of z, and k is a small integer such that θ = β + kα where α is a root of the polynomial defining the base field nf .

The main variable of nf must be of lower priority than that of pol. Note that for efficiency, this does not check whether the relative equation is irreducible over nf , but only if it is squarefree.

If it is reducible but squarefree, the result will be the absolute equation of the étale algebra defined by pol. If pol is not squarefree, an error message will be issued.

The library syntax is rnfequation0(nf , pol, flag).

3.6.116 rnfhnfbasis(bnf , x): given a big number field bnf as output by bnfinit, and either a polynomial x with coefficients in bnf defining a relative extension L of bnf , or a pseudo-basis x of such an extension, gives either a true bnf -basis of L in upper triangular Hermite normal form, if it exists, zero otherwise.

The library syntax is rnfhermitebasis(nf , x).

3.6.117 rnfidealabstorel(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an ideal of the absolute extension L/Q given in HNF (if it is not, apply idealhnf first), computes the relative pseudomatrix in HNF giving the ideal x considered as an ideal of the relative extension L/K.

The library syntax is rnfidealabstorel(rnf , x).

3.6.118 rnfidealdown(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an ideal of the absolute extension L/Q given in HNF (if it is not, apply idealhnf first), gives the ideal of K below x, i.e. the intersection of x with K. Note that, if x is given as a relative ideal (i.e. a pseudomatrix in HNF), then it is not necessary to use this function since the result is simply the first ideal of the ideal list of the pseudomatrix.

The library syntax is rnfidealdown(rnf , x).

3.6.119 rnfidealhnf(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being a relative ideal (which can be, as in the absolute case, of many different types, including of course elements), computes as a 2-component row vector the relative Hermite normal form of x, the first component being the HNF matrix (with entries on the integral basis), and the second component the ideals.

The library syntax is rnfidealhermite(rnf , x).

3.6.120 rnfidealmul(rnf , x, y): rnf being a relative number field extension L/K as output by rnfinit and x and y being ideals of the relative extension L/K given by pseudo-matrices, outputs the ideal product, again as a relative ideal.

The library syntax is rnfidealmul(rnf , x, y).

3.6.121 rnfidealnormabs(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being a relative ideal (which can be, as in the absolute case, of many different types, including of course elements), computes the norm of the ideal x considered as an ideal of the absolute extension L/Q. This is identical to idealnorm(rnfidealnormrel(rnf ,x)), only faster.

The library syntax is rnfidealnormabs(rnf , x).

3.6.122 rnfidealnormrel(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being a relative ideal (which can be, as in the absolute case, of many different types, including of course elements), computes the relative norm of x as a ideal of K in HNF.

The library syntax is rnfidealnormrel(rnf , x).

3.6.123 rnfidealreltoabs(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being a relative ideal (which can be, as in the absolute case, of many different types, including of course elements), computes the HNF matrix of the ideal x considered as an ideal of the absolute extension L/Q.

The library syntax is rnfidealreltoabs(rnf , x).

3.6.124 rnfidealtwoelt(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an ideal of the relative extension L/K given by a pseudo-matrix, gives a vector of two generators of x over Z_{L} expressed as polmods with polmod coefficients.

The library syntax is rnfidealtwoelement(rnf , x).

3.6.125 rnfidealup(rnf , x): rnf being a relative number field extension L/K as output by rnfinit and x being an ideal of K, gives the ideal xZ_{L} as an absolute ideal of L/Q (the relative ideal representation is trivial: the matrix is the identity matrix, and the ideal list starts with x, all the other ideals being Z_{K}).

The library syntax is rnfidealup(rnf , x).

3.6.126 rnfinit(nf , pol): nf being a number field in nfinit format considered as base field, and pol a polynomial defining a relative extension over nf , this computes all the necessary data to work in the relative extension. The main variable of pol must be of higher priority (i.e. lower number) than that of nf , and the coefficients of pol must be in nf .

The result is an 11-component row vector as follows (most of the components are technical), the numbering being very close to that of nfinit. In the following description, we let K be the base field defined by nf , m the degree of the base field, n the relative degree, L the large field (of relative degree n or absolute degree nm), r_{1} and r_{2} the number of real and complex places of K.

rnf [1] contains the relative polynomial pol.

rnf [2] is a row vector with r_{1} + r_{2} entries, entry j being a 2-component row vector [r_{j,1}, r_{j,2}] where r_{j,1} and r_{j,2} are the number of real and complex places of L above the j-th place of K so that r_{j,1} = 0 and r_{j,2} = n if j is a complex place, while if j is a real place we have r_{j,1} + 2r_{j,2} = n.

rnf [3] is a two-component row vector [𝔡(L/K), s] where 𝔡(L/K) is the relative ideal discriminant of L/K and s is the discriminant of L/K viewed as an element of K^{∗}/(K^{∗})^{2}, in other words it is the output of rnfdisc.

rnf [4] is the ideal index 𝔣, i.e. such that d(pol)Z_{K} = 𝔣^{2}𝔡(L/K).

rnf [5] is a vector vm with 7 entries useful for certain computations in the relative extension L/K. vm[1] is a vector of r_{1} + r_{2} matrices, the j-th matrix being an (r_{1,j} + r_{2,j}) × n matrix M_{j} representing the numerical values of the conjugates of the j-th embedding of the elements of the integral basis, where r_{i,j} is as in rnf [2]. vm[2] is a vector of r_{1} + r_{2} matrices, the j-th matrix MC_{j} being essentially the conjugate of the matrix M_{j} except that the last r_{2,j} columns are also multiplied by 2. vm[3] is a vector of r_{1} + r_{2} matrices T 2_{j}, where T 2_{j} is an n × n matrix equal to the real part of the product MC_{j} · M_{j} (which is a real positive definite matrix). vm[4] is the n × n matrix T whose entries are the relative traces of ω_{i}ω_{j} expressed as polmods in nf , where the ω_{i} are the elements of the relative integral basis. Note that the j-th embedding of T is equal to \overline{MC_{j}} · M_{j}, and in particular will be equal to T 2_{j} if r_{2,j} = 0. Note also that the relative ideal discriminant of L/K is equal to det(T) times the square of the product of the ideals in the relative pseudo-basis (in rnf [7][2]). The last 3 entries vm[5], vm[6] and vm[7] are linked to the different as in nfinit, but have not yet been implemented.

rnf [6] is a row vector with r_{1} + r_{2} entries, the j-th entry being the row vector with r_{1,j} + r_{2,j} entries of the roots of the j-th embedding of the relative polynomial pol.

rnf [7] is a two-component row vector, where the first component is the relative integral pseudo basis expressed as polynomials (in the variable of pol) with polmod coefficients in nf , and the second component is the ideal list of the pseudobasis in HNF.

rnf [8] is the inverse matrix of the integral basis matrix, with coefficients polmods in nf .

rnf [9] may be the multiplication table of the integral basis, but is not implemented at present.

rnf [10] is nf .

rnf [11] is a vector vabs with 5 entries describing the absolute extension L/Q. vabs[1] is an absolute equation. vabs[2] expresses the generator α of the number field nf as a polynomial modulo the absolute equation vabs[1]. vabs[3] is a small integer k such that, if β is an abstract root of pol and α the generator of nf , the generator whose root is vabs will be β + kα. Note that one must be very careful if k ≠ 0 when dealing simultaneously with absolute and relative quantities since the generator chosen for the absolute extension is not the same as for the relative one. If this happens, one can of course go on working, but we strongly advise to change the relative polynomial so that its root will be β + kα. Typically, the GP instruction would be

    pol = subst(pol, x, x - k*Mod(y,nf .pol))

Finally, vabs[4] is the absolute integral basis of L expressed in HNF (hence as would be output by nfinit(vabs[1])), and vabs[5] the inverse matrix of the integral basis, allowing to go from polmod to integral basis representation.

The library syntax is rnfinitalg(nf , pol, prec).

3.6.127 rnfisfree(bnf , x): given a big number field bnf as output by bnfinit, and either a polynomial x with coefficients in bnf defining a relative extension L of bnf , or a pseudo-basis x of such an extension, returns true (1) if L/bnf is free, false (0) if not.

The library syntax is rnfisfree(bnf , x), and the result is a long.

3.6.128 rnfisnorm(bnf , ext, el, {flag = 1}): similar to bnfisnorm but in the relative case. This tries to decide whether the element el in bnf is the norm of some y in ext. bnf is as output by

bnfinit.

ext is a relative extension which has to be a row vector whose components are:

ext[1]: a relative equation of the number field ext over bnf . As usual, the priority of the variable of the polynomial defining the ground field bnf (say y) must be lower than the main variable of ext[1], say x.

ext[2]: the generator y of the base field as a polynomial in x (as given by rnfequation with flag = 1).

ext[3]: is the bnfinit of the absolute extension ext/Q.

This returns a vector [a, b], where el = Norm(a) ∗ b. It looks for a solution which is an Sinteger, with S a list of places (of bnf ) containing the ramified primes, the generators of the class group of ext, as well as those primes dividing el. If ext/bnf is known to be Galois, set flag = 0 (here el is a norm iff b = 1). If flag is non zero add to S all the places above the primes which: divide flag if flag < 0, or are less than flag if flag > 0. The answer is guaranteed (i.e. el is a norm iff b = 1) under GRH, if S contains all primes less than 12 log^{2} |disc(Ext)|, where Ext is the normal closure of ext/bnf .

The library syntax is rnfisnorm(bnf , ext, x, flag, prec).

3.6.129 rnfkummer(bnr, subgroup, {deg = 0}): bnr being as output by bnrinit, finds a relative equation for the class field corresponding to the module in bnr and the given congruence subgroup. If deg is positive, outputs the list of all relative equations of degree deg contained in the ray class field defined by bnr.

(THIS PROGRAM IS STILL IN DEVELOPMENT STAGE)

The library syntax is rnfkummer(bnr, subgroup, deg, prec), where deg is a long.

3.6.130 rnflllgram(nf , pol, order): given a polynomial pol with coefficients in nf and an order order as output by rnfpseudobasis or similar, gives [[neworder], U], where neworder is a reduced order and U is the unimodular transformation matrix.

The library syntax is rnflllgram(nf , pol, order, prec).

3.6.131 rnfnormgroup(bnr, pol): bnr being a big ray class field as output by bnrinit and pol a relative polynomial defining an Abelian extension, computes the norm group (alias Artin or Takagi group) corresponding to the Abelian extension of bnf = bnr[1] defined by pol, where the module corresponding to bnr is assumed to be a multiple of the conductor (i.e. polrel defines a subextension of bnr). The result is the HNF defining the norm group on the given generators of bnr[5][3]. Note that neither the fact that pol defines an Abelian extension nor the fact that the module is a multiple of the conductor is checked. The result is undefined if the assumption is not correct.

The library syntax is rnfnormgroup(bnr, pol).

3.6.132 rnfpolred(nf , pol): relative version of polred. Given a monic polynomial pol with coefficients in nf , finds a list of relative polynomials defining some subfields, hopefully simpler and containing the original field.

The library syntax is rnfpolred(nf , pol, prec).

3.6.133 rnfpolredabs(nf , pol, {flag = 0}): relative version of polredabs. Given a monic polynomial pol with coefficients in nf , finds a simpler relative polynomial defining the same field. If flag = 1, returns [P, a] where P is the default output and a is an element expressed on a root of P whose characteristic polynomial is pol, if flag = 2, returns an absolute polynomial (same as

    rnfequation(nf ,rnfpolredabs(nf ,pol))

but faster).

Remark. In the present implementation, although this is slower than rnfpolred, it is much more efficient, the difference being more dramatic than in the absolute case. This is because the implementation of rnfpolred is based on an incomplete reduction theory of lattices over number fields (i.e. the function rnflllgram) which deserves to be improved.

The library syntax is rnfpolredabs(nf , pol, flag, prec).

3.6.134 rnfpseudobasis(nf , pol): given a number field nf as output by nfinit and a polynomial pol with coefficients in nf defining a relative extension L of nf , computes a pseudo-basis (A, I) and the relative discriminant of L. This is output as a four-element row vector [A, I, D, d], where D is the relative ideal discriminant and d is the relative discriminant considered as an element of nf^{∗}/nf^{∗}^{2}.

Note: As usual, nf can be a bnf as output by bnfinit.

The library syntax is rnfpseudobasis(nf , pol).

3.6.135 rnfsteinitz(nf , x): given a number field nf as output by nfinit and either a polynomial x with coefficients in nf defining a relative extension L of nf , or a pseudo-basis x of such an extension as output for example by rnfpseudobasis, computes another pseudo-basis (A, I) (not in HNF in general) such that all the ideals of I except perhaps the last one are equal to the ring of integers of nf , and outputs the four-component row vector [A, I, D, d] as in rnfpseudobasis. The name of this function comes from the fact that the ideal class of the last ideal of I (which is well defined) is called the Steinitz class of the module Z_{L}.

Note: nf can be a bnf as output by bnfinit.

The library syntax is rnfsteinitz(nf , x).

3.6.136 subgrouplist(bnr, {bound}, {flag = 0}): bnr being as output by bnrinit or a list of cyclic components of a finite Abelian group G, outputs the list of subgroups of G (of index bounded by bound, if not omitted). Subgroups are given as HNF left divisors of the SNF matrix corresponding to G. If flag = 0 (default) and bnr is as output by bnrinit, gives only the subgroups whose modulus is the conductor.

The library syntax is subgrouplist0(bnr, bound, flag, prec), where bound, flag and prec are long integers.

3.6.137 zetak(znf , x, {flag = 0}): znf being a number field initialized by zetakinit (not by nfinit), computes the value of the Dedekind zeta function of the number field at the complex number x. If flag = 1 computes Dedekind Λ function instead (i.e. the product of the Dedekind zeta function by its gamma and exponential factors).

The accuracy of the result depends in an essential way on the accuracy of both the zetakinit program and the current accuracy, but even so the result may be off by up to 5 or 10 decimal digits.

The library syntax is glambdak(znf , x, prec) or gzetak(znf , x, prec).

3.6.138 zetakinit(x): computes a number of initialization data concerning the number field defined by the polynomial x so as to be able to compute the Dedekind zeta and lambda functions (respectively zetak(x) and zetak(x, 1)). This function calls in particular the bnfinit program. The result is a 9-component vector v whose components are very technical and cannot really be used by the user except through the zetak function. The only component which can be used if it has not been computed already is v[1][4] which is the result of the bnfinit call.

This function is very inefficient and needs to computes millions of coefficients of the corresponding Dirichlet series if the precision is big. Unless the discriminant is small it will not be able to handle more than 9 digits of relative precision (e.g zetakinit(x^8 - 2) needs 440MB of memory at default precision).

The library syntax is initzeta(x).

3.7 Polynomials and power series.

We group here all functions which are specific to polynomials or power series. Many other functions which can be applied on these objects are described in the other sections. Also, some of the functions described here can be applied to other types.

3.7.1 O(a^b): p-adic (if a is an integer greater or equal to 2) or power series zero (in all other cases), with precision given by b.

The library syntax is ggrandocp(a, b), where b is a long.

3.7.2 deriv(x, {v}): derivative of x with respect to the main variable if v is omitted, and with respect to v otherwise. x can be any type except polmod. The derivative of a scalar type is zero, and the derivative of a vector or matrix is done componentwise. One can use x′ as a shortcut if the derivative is with respect to the main variable of x.

The library syntax is deriv(x, v), where v is a long, and an omitted v is coded as −1.

3.7.3 eval(x): replaces in x the formal variables by the values that have been assigned to them after the creation of x. This is mainly useful in GP, and not in library mode. Do not confuse this with substitution (see subst). Applying this function to a character string yields the output from the corresponding GP command, as if directly input from the keyboard (see Section 2.6.5).

The library syntax is geval(x). The more basic functions poleval(q, x), qfeval(q, x), and hqfeval(q, x) evaluate q at x, where q is respectively assumed to be a polynomial, a quadratic form (a symmetric matrix), or an Hermitian form (an Hermitian complex matrix).

3.7.4 factorpadic(pol, p, r, {flag = 0}): p-adic factorization of the polynomial pol to precision r, the result being a two-column matrix as in factor. r must be strictly larger than the p-adic valuation of the discriminant of pol for the result to make any sense. The method used is Ford-Letard’s implementation of the round 4 algorithm of Zassenhaus.

If flag = 1, use an algorithm due to Buchmann and Lenstra, which is usually less efficient.

The library syntax is factorpadic4(pol, p, r), where r is a long integer.

3.7.5 intformal(x, {v}): formal integration of x with respect to the main variable if v is omitted, with respect to the variable v otherwise. Since PARI does not know about “abstract” logarithms (they are immediately evaluated, if only to a power series), logarithmic terms in the result will yield an error. x can be of any type. When x is a rational function, it is assumed that the base ring is an integral domain of characteristic zero.

The library syntax is integ(x, v), where v is a long and an omitted v is coded as −1.

3.7.6 padicappr(pol, a): vector of p-adic roots of the polynomial pol congruent to the p-adic number a modulo p (or modulo 4 if p = 2), and with the same p-adic precision as a. The number a can be an ordinary p-adic number (type t PADIC, i.e. an element of Q_{p}) or can be an element of a finite extension of Q_{p}, in which case it is of type t POLMOD, where at least one of the coefficients of the polmod is a p-adic number. In this case, the result is the vector of roots belonging to the same extension of Q_{p} as a.

The library syntax is apprgen9(pol, a), but if a is known to be simply a p-adic number (type t PADIC), the syntax apprgen(pol, a) can be used.

3.7.7 polcoeff(x, s, {v}): coefficient of degree s of the polynomial x, with respect to the main variable if v is omitted, with respect to v otherwise.

The library syntax is polcoeff0(x, s, v), where v is a long and an omitted v is coded as −1. Also available is truecoeff(x, v).

3.7.8 poldegree(x, {v}): degree of the polynomial x in the main variable if v is omitted, in the variable v otherwise. This is to be understood as follows. When x is a polynomial or a rational function, it gives the degree of x, the degree of 0 being −1 by convention. When x is a non-zero scalar, it gives 0, and when x is a zero scalar, it gives −1. Return an error otherwise.

The library syntax is poldegree(x, v), where v and the result are longs (and an omitted v is coded as −1). Also available is degree(x), which is equivalent to poldegree(x,-1).

3.7.9 polcyclo(n, {v = x}): n-th cyclotomic polynomial, in variable v (x by default). The integer n must be positive.

The library syntax is cyclo(n, v), where n and v are long integers (v is a variable number, usually obtained through varn).

3.7.10 poldisc(pol, {v}): discriminant of the polynomial pol in the main variable is v is omitted, in v otherwise. The algorithm used is the subresultant algorithm.

The library syntax is poldisc0(x, v). Also available is discsr(x), equivalent to poldisc0(x,-

1).

3.7.11 poldiscreduced(f): reduced discriminant vector of the (integral, monic) polynomial f. This is the vector of elementary divisors of Z[α]/f′(α)Z[α], where α is a root of the polynomial f. The components of the result are all positive, and their product is equal to the absolute value of the discriminant of f.

The library syntax is reduceddiscsmith(x).

3.7.12 polinterpolate(xa, ya, {v = x}, {&e}): given the data vectors xa and ya of the same length n (xa containing the x-coordinates, and ya the corresponding y-coordinates), this function finds the interpolating polynomial passing through these points and evaluates it at v. If present, e will contain an error estimate on the returned value.

The library syntax is polint(xa, ya, v, &e), where e will contain an error estimate on the returned value.

3.7.13 polisirreducible(pol): pol being a polynomial (univariate in the present version 2.0.17), returns 1 if pol is non-constant and irreducible, 0 otherwise. Irreducibility is checked over the smallest base field over which pol seems to be defined.

The library syntax is gisirreducible(pol).

3.7.14 pollead(x, {v}): leading coefficient of the polynomial or power series x. This is computed with respect to the main variable of x if v is omitted, with respect to the variable v otherwise.

The library syntax is pollead(x, v), where v is a long and an omitted v is coded as −1. Also available is leadingcoeff(x).

3.7.15 pollegendre(n, {v = x}): creates the n^{th} Legendre polynomial, in variable v.

The library syntax is legendre(n), where x is a long.

3.7.16 polrecip(pol): reciprocal polynomial of pol, i.e. the coefficients are in reverse order. pol must be a polynomial.

The library syntax is polrecip(x).

3.7.17 polresultant(x, y, {v}, {flag = 0}): resultant of the two polynomials x and y with exact entries, with respect to the main variables of x and y if v is omitted, with respect to the variable v otherwise. The algorithm used is the subresultant algorithm by default.

If flag = 1, uses the determinant of Sylvester’s matrix instead (here x and y may have non-exact coefficients).

If flag = 2, uses Ducos’s modified subresultant algorithm. It should be much faster than the default if the coefficient ring is complicated (e.g multivariate polynomials or huge coefficients), and slightly slower otherwise.

The library syntax is polresultant0(x, y, v, flag), where v is a long and an omitted v is coded as −1. Also available are subres(x, y) (flag = 0) and resultant2(x, y) (flag = 1).

3.7.18 polroots(pol, {flag = 0}): complex roots of the polynomial pol, given as a column vector where each root is repeated according to its multiplicity. The precision is given as for transcendental functions: under GP it is kept in the variable realprecision and is transparent to the user, but it must be explicitly given as a second argument in library mode.

The algorithm used is a modification of A. Schönhage’s remarkable root-finding algorithm, due to and implemented by X. Gourdon. Barring bugs, it is guaranteed to converge and to give the roots to the required accuracy.

If flag = 1, use a variant of the Newton-Raphson method, which is not guaranteed to converge, but is rather fast. If you get the messages “too many iterations in roots” or “INTERNAL ERROR: incorrect result in roots”, use the default function (i.e. no flag or flag = 0). This used to be the default root-finding function in PARI until version 1.39.06.

The library syntax is roots(pol, prec) or rootsold(pol, prec).

3.7.19 polrootsmod(pol, p, {flag = 0}): row vector of roots modulo p of the polynomial pol. The particular non-prime value p = 4 is accepted, mainly for 2-adic computations. Multiple roots are not repeated.

If p < 100, you may try setting flag = 1, which uses a naive search. In this case, multiple roots are repeated with their order of multiplicity.

The library syntax is rootmod(pol, p) (flag = 0) or rootmod2(pol, p) (flag = 1).

3.7.20 polrootspadic(pol, p, r): row vector of p-adic roots of the polynomial pol with p-adic precision equal to r. Multiple roots are not repeated. p is assumed to be a prime.

The library syntax is rootpadic(pol, p, r), where r is a long.

3.7.21 polsturm(pol, {a}, {b}): number of real roots of the real polynomial pol in the interval ]a, b], using Sturm’s algorithm. a (resp. b) is taken to be −∞ (resp. +∞) if omitted.

The library syntax is sturmpart(pol, a, b). Use NULL to omit an argument. sturm(pol) is equivalent to sturmpart(pol,NULL,NULL). The result is a long.

3.7.22 polsubcyclo(n, d, {v = x}): gives a polynomial (in variable v) defining the sub-Abelian extension of degree d of the cyclotomic field Q(ζ_{n}), where d | ϕ(n). (Z/nZ)^{∗} has to be cyclic (i.e. n = 2, 4, p^{k} or 2p^{k} for an odd prime p).

The library syntax is subcyclo(n, d, v), where v is a variable number.

3.7.23 polsylvestermatrix(x, y): forms the Sylvester matrix corresponding to the two polynomials x and y, where the coefficients of the polynomials are put in the columns of the matrix (which is the natural direction for solving equations afterwards). The use of this matrix can be essential when dealing with polynomials with inexact entries, since polynomial Euclidean division doesn’t make much sense in this case.

The library syntax is sylvestermatrix(x, y).

3.7.24 polsym(x, n): creates the vector of the symmetric powers of the roots of the polynomial x up to power n, using Newton’s formula.

The library syntax is polsym(x).

3.7.25 poltchebi(n, {v = x}): creates the n^{th} Chebyshev polynomial, in variable v.

The library syntax is tchebi(n, v), where n and v are long integers (v is a variable number).

3.7.26 polzagier(n, m): creates Zagier’s polynomial P_{n,m} used in the functions sumalt and sumpos (with flag = 1). The exact definition can be found in a forthcoming paper. One must have m ≤ n.

The library syntax is polzagreel(n, m, prec) if the result is only wanted as a polynomial with real coefficients to the precision prec, or polzag(n, m) if the result is wanted exactly, where n and m are longs.

3.7.27 serconvol(x, y): convolution (or Hadamard product) of the two power series x and y; in other words if x = ∑ a_{k} ∗ X^{k} and y = ∑ b_{k} ∗ X^{k} then serconvol(x, y) = ∑ a_{k} ∗ b_{k} ∗ X^{k}.

The library syntax is convol(x, y).

3.7.28 serlaplace(x): x must be a power series with only non-negative exponents. If x = ∑(a_{k}/k!) ∗ X^{k} then the result is ∑ a_{k} ∗ X^{k}.

The library syntax is laplace(x).

3.7.29 serreverse(x): reverse power series (i.e. x^{−1}, not 1/x) of x. x must be a power series whose valuation is exactly equal to one.

The library syntax is recip(x).

3.7.30 subst(x, y, z): replace the simple variable y by the argument z in the “polynomial” expression x. Every type is allowed for x, but if it is not a genuine polynomial (or power series, or rational function), the substitution will be done as if the scalar components were polynomials of degree one. In particular, beware that:

    ?  subst(1, x, [1,2; 3,4])
    %1 =
    [1 0]
    [0 1]
    ?  subst(1, x, Mat([0,1]))
      *** forbidden substitution by a non square matrix

If x is a power series, z must be either a polynomial, a power series, or a rational function. y must be a simple variable name.

The library syntax is gsubst(x, v, z), where v is the number of the variable y.

3.7.31 taylor(x, y): Taylor expansion around 0 of x with respect to the simple variable y. x can be of any reasonable type, for example a rational function. The number of terms of the expansion is transparent to the user under GP, but must be given as a second argument in library mode.

The library syntax is tayl(x, y, n), where the long integer n is the desired number of terms in the expansion.

3.7.32 thue(tnf , a, {sol}): solves the equation P (x, y) = a in integers x and y, where tnf was created with thueinit(P). sol, if present, contains the solutions of Norm(x) = a modulo units of positive norm in the number field defined by P (as computed by bnfisintnorm). If tnf was computed without assuming GRH (flag = 1 in thueinit), the result is unconditional.

The library syntax is thue(tnf , a, sol), where an omitted sol is coded as NULL.

3.7.33 thueinit(P, {flag = 0}): initializes the tnf corresponding to P. It is meant to be used in conjunction with thue to solve Thue equations P (x, y) = a, where a is an integer. If flag is non-zero, certify the result unconditionnaly, Otherwise, assume GRH, this being much faster of course.

The library syntax is thueinit(P, flag, prec).

3.8 Vectors, matrices, linear algebra and sets.

Note that most linear algebra functions operating on subspaces defined by generating sets (such as mathnf, qflll, etc.) take matrices as arguments. As usual, the generating vectors are taken to be the columns of the given matrix.

3.8.1 algdep(x, k, {flag = 0}): x being real or complex, finds a polynomial of degree at most k having x as approximate root. The algorithm used is a variant of the LLL algorithm due to Hastad, Lagarias and Schnorr (STACS 1986). Note that the polynomial which is obtained is not necessarily the “correct” one (it’s not even guaranteed to be irreducible!). One can check the closeness either by a polynomial evaluation or substitution, or by computing the roots of the polynomial given by algdep. If the precision is too low, the routine may enter an infinite loop.

If flag is non-zero, use a standard LLL. flag then indicates a precision, which should be between 0.5 and 1.0 times the number of decimal digits to which x was computed.

The library syntax is algdep0(x, k, flag, prec), where k and flag are longs. Also available is algdep(x, k, prec) (flag = 0).

3.8.2 charpoly(A, {v = x}, {flag = 0}): characteristic polynomial of A with respect to the variable v, i.e. determinant of v ∗ I − A if A is a square matrix, determinant of the map “multiplication by A” if A is a scalar, in particular a polmod (e.g. charpoly(I,x)=x^2+1), error if A is of any other type. The value of flag is only significant for matrices.

If flag = 0, the method used is essentially the same as for computing the adjoint matrix, i.e. computing the traces of the powers of A.

If flag = 1, uses Lagrange interpolation which is almost always slower.

If flag = 2, uses the Hessenberg form. This is faster than the default when the coefficients are integermod a prime or real numbers, but is usually slower in other base rings.

The library syntax is charpoly0(A, v, flag), where v is the variable number. Also available are the functions caract(A, v) (flag = 1), carhess(A, v) (flag = 2), and caradj(A, v, pt) where, in this last case, pt is a GEN* which, if not equal to NULL, will receive the address of the adjoint matrix of A (see matadjoint), so both can be obtained at once.

3.8.3 concat(x, {y}): concatenation of x and y. If x or y is not a vector or matrix, it is considered as a one-dimensional vector. All types are allowed for x and y, but the sizes must be compatible. Note that matrices are concatenated horizontally, i.e. the number of rows stays the same. Using transpositions, it is easy to concatenate them vertically.

To concatenate vectors sideways (i.e. to obtain a two-row or two-column matrix), first transform the vector into a one-row or one-column matrix using the function Mat. Concatenating a row vector to a matrix having the same number of columns will add the row to the matrix (top row if the vector is x, i.e. comes first, and bottom row otherwise).

The empty matrix [;] is considered to have a number of rows compatible with any operation, in particular concatenation. (Note that this is definitely not the case for empty vectors [ ] or

[ ]~.)

If y is omitted, x has to be a row vector or a list, in which case its elements are concatenated, from left to right, using the above rules.

    ?  concat([1,2], [3,4])

%1 = [1, 2, 3, 4]
?  concat([1,2]~, [3,4]~)
%2 = [1, 2, 3, 4]~
?  concat([1,2; 3,4], [5,6]~)
%3 =
[1, 2, 5]
[3, 4, 6]
?  concat([%, [7,8]~, [1,2,3,4]])
%4 =
[1 2 5 7]
[3 4 6 8]
[1 2 3 4]

The library syntax is concat(x, y).

3.8.4 lindep(x, {flag = 0}):x being a vector with real or complex coefficients, finds a small integral linear combination among these coefficients.

If flag = 0, uses a variant of the LLL algorithm due to Hastad, Lagarias and Schnorr (STACS 1986).

If flag > 0, uses the LLL algorithm. flag is a parameter which should be between one half the number of decimal digits of precision and that number (see algdep).

If flag < 0, returns as soon as one relation has been found.

The library syntax is lindep0(x, flag, prec). Also available is lindep(x, prec) (flag = 0).

3.8.5 listcreate(n): creates an empty list of maximal length n.

This function is useless in library mode.

3.8.6 listinsert(list, x, n): inserts the object x at position n in list (which must be of type t LIST). All the remaining elements of list (from position n + 1 onwards) are shifted to the right. This and listput are the only commands which enable you to increase a list’s effective length (as long as it remains under the maximal length specified at the time of the listcreate).

This function is useless in library mode.

3.8.7 listkill(list): kill list. This deletes all elements from list and sets its effective length to 0. The maximal length is not affected.

This function is useless in library mode.

3.8.8 listput(list, x, {n}): sets the n-th element of the list list (which must be of type t LIST) equal to x. If n is omitted, or greater than the list current effective length, just appends x. This and listinsert are the only commands which enable you to increase a list’s effective length (as long as it remains under the maximal length specified at the time of the listcreate).

If you want to put an element into an occupied cell, i.e. if you don’t want to change the effective length, you can consider the list as a vector and use the usual list[n] = x construct.

This function is useless in library mode.

3.8.9 listsort(list, {flag = 0}): sorts list (which must be of type t LIST) in place. If flag is non-zero, suppresses all repeated coefficients. This is much faster than the vecsort command since no copy has to be made.

This function is useless in library mode.

3.8.10 matadjoint(x): adjoint matrix of x, i.e. the matrix y of cofactors of x, satisfying x ∗ y = det(x) ∗ Id. x must be a (non-necessarily invertible) square matrix.

The library syntax is adj(x).

3.8.11 matcompanion(x): the left companion matrix to the polynomial x.

The library syntax is assmat(x).

3.8.12 matdet(x, {flag = 0}): determinant of x. x must be a square matrix.

If flag = 0, uses Gauss-Bareiss.

If flag = 1, uses classical Gaussian elimination, which is better when the entries of the matrix are reals or integers for example, but usually much worse for more complicated entries like multivariate polynomials.

The library syntax is det(x) (flag = 0) and det2(x) (flag = 1).

3.8.13 matdetint(x): x being an m × n matrix with integer coefficients, this function computes a multiple of the determinant of the lattice generated by the columns of x if it is of rank m, and returns zero otherwise. This function can be useful in conjunction with the function mathnfmod which needs to know such a multiple. Other ways to obtain this determinant (assuming the rank is

maximal) is matdet(qflll(x,4)[2]∗x) or more simply matdet(mathnf(x)). Experiment to see

which is faster for your applications.

The library syntax is detint(x).

3.8.14 matdiagonal(x): x being a vector, creates the diagonal matrix whose diagonal entries are those of x.

The library syntax is diagonal(x).

3.8.15 mateigen(x): gives the eigenvectors of x as columns of a matrix.

The library syntax is eigen(x).

3.8.16 mathess(x): Hessenberg form of the square matrix x.

The library syntax is hess(x).

3.8.17 mathilbert(x): x being a long, creates the Hilbert matrix of order x, i.e. the matrix whose coefficient (i,j) is 1/(i + j − 1).

The library syntax is mathilbert(x).

3.8.18 mathnf(x, {flag = 0}): if x is a (not necessarily square) matrix of maximal rank, finds the upper triangular Hermite normal form of x. If the rank of x is equal to its number of rows, the result is a square matrix. In general, the columns of the result form a basis of the lattice spanned by the columns of x.

If flag = 0, uses the naive algorithm. If the Z-module generated by the columns is a lattice, it is recommanded to use mathnfmod(x, matdetint(x)) instead (much faster).

If flag = 1, uses Batut’s algorithm. Outputs a two-component row vector [H, U], where H is the upper triangular Hermite normal form of x (i.e. the default result) and U is the unimodular transformation matrix such that xU = [0|H]. If the rank of x is equal to its number of rows, H is a square matrix. In general, the columns of H form a basis of the lattice spanned by the columns of x.

If flag = 2, uses Havas’s algorithm. Outputs [H, U, P], such that H and U are as before and P is a permutation of the rows such that P applied to xU gives H. This does not work very well in present version 2.0.17.

If flag = 3, uses Batut’s algorithm, and outputs [H, U, P] as in the previous case.

If flag = 4, as in case 1 above, but uses LLL reduction along the way.

The library syntax is mathnf0(x, flag). Also available are hnf(x) (flag = 0) and hnfall(x) (flag = 1). To reduce huge (say 400 × 400 and more) relation matrices (sparse with small entries), you can use the pair hnfspec / hnfadd. Since this is rather technical and the calling interface may change, they are not documented yet. Look at the code in basemath/alglin1.c.

3.8.19 mathnfmod(x, d): if x is a (not necessarily square) matrix of maximal rank with integer entries, and d is a multiple of the (non-zero) determinant of the lattice spanned by the columns of x, finds the upper triangular Hermite normal form of x.

If the rank of x is equal to its number of rows, the result is a square matrix. In general, the columns of the result form a basis of the lattice spanned by the columns of x. This is much faster than mathnf when d is known.

The library syntax is hnfmod(x, d).

3.8.20 mathnfmodid(x, d): outputs the (upper triangular) Hermite normal form of x concatenated with d times the identity matrix.

The library syntax is hnfmodid(x, d).

3.8.21 matid(n): creates the n × n identity matrix.

The library syntax is idmat(n) where n is a long.

Related functions are gscalmat(x, n), which creates x times the identity matrix (x being a GEN and n a long), and gscalsmat(x, n) which is the same when x is a long.

3.8.22 matimage(x, {flag = 0}): gives a basis for the image of the matrix x as columns of a matrix. A priori the matrix can have entries of any type. If flag = 0, use standard Gauss pivot. If

flag = 1, use matsupplement.

The library syntax is matimage0(x, flag). Also available is image(x) (flag = 0).

3.8.23 matimagecompl(x): gives the vector of the column indices which are not extracted by the function matimage. Hence the number of components of matimagecompl(x) plus the number of columns of matimage(x) is equal to the number of columns of the matrix x.

The library syntax is imagecompl(x).

3.8.24 matindexrank(x): x being a matrix of rank r, gives two vectors y and z of length r giving a list of rows and columns respectively (starting from 1) such that the extracted matrix obtained from these two vectors using vecextract(x, y, z) is invertible.

The library syntax is indexrank(x).

3.8.25 matintersect(x, y): x and y being two matrices with the same number of rows each of whose columns are independent, finds a basis of the Q-vector space equal to the intersection of the spaces spanned by the columns of x and y respectively. See also the function idealintersect, which does the same for free Z-modules.

The library syntax is intersect(x, y).

3.8.26 matinverseimage(x, y): gives a column vector belonging to the inverse image of the column vector y by the matrix x if one exists, the empty vector otherwise. To get the complete inverse image, it suffices to add to the result any element of the kernel of x obtained for example

by matker.

The library syntax is inverseimage(x, y).

3.8.27 matisdiagonal(x): returns true (1) if x is a diagonal matrix, false (0) if not.

The library syntax is isdiagonal(x), and this returns a long integer.

3.8.28 matker(x, {flag = 0}): gives a basis for the kernel of the matrix x as columns of a matrix. A priori the matrix can have entries of any type.

If x is known to have integral entries, set flag = 1.

Note: The library function ker mod p(x, p), where x has integer entries and p is prime, which is equivalent to but many orders of magnitude faster than matker(x*Mod(1,p)) and needs much less stack space. To use it under GP, type install(ker mod p, GG) first.

The library syntax is matker0(x, flag). Also available are ker(x) (flag = 0), keri(x) (flag = 1) and ker mod p(x, p).

3.8.29 matkerint(x, {flag = 0}): gives an LLL-reduced Z-basis for the lattice equal to the kernel of the matrix x as columns of the matrix x with integer entries (rational entries are not permitted).

If flag = 0, uses a modified integer LLL algorithm.

If flag = 1, uses matrixqz(x, −2). If LLL reduction of the final result is not desired, you can

save time using matrixqz(matker(x),-2) instead.

If flag = 2, uses another modified LLL. In the present version 2.0.17, only independent rows are allowed in this case.

The library syntax is matkerint0(x, flag). Also available is kerint(x) (flag = 0).

3.8.30 matmuldiagonal(x, d): product of the matrix x by the diagonal matrix whose diagonal entries are those of the vector d. Equivalent to, but much faster than x ∗ matdiagonal(d).

The library syntax is matmuldiagonal(x, d).

3.8.31 matmultodiagonal(x, y): product of the matrices x and y knowing that the result is a diagonal matrix. Much faster than x ∗ y in that case.

The library syntax is matmultodiagonal(x, y).

3.8.32 matpascal(x, {q}): creates as a matrix the lower triangular pascal triangle of order x + 1 (i.e. with binomial coefficients up to x). If q is given, compute the q-Pascal triangle (i.e. using q-binomial coefficients).

The library syntax is matqpascal(x, q), where x is a long and q = NULL is used to omit q. Also available is matpascalx.

3.8.33 matrank(x): rank of the matrix x.

The library syntax is rank(x), and the result is a long.

3.8.34 matrixqz(x, p): x being an m × n matrix with m ≥ n with rational or integer entries, this function has varying behaviour depending on the sign of p:

If p ≥ 0, x is assumed to be of maximal rank. This function returns a matrix having only integral entries, having the same image as x, such that the GCD of all its n × n subdeterminants is equal to 1 when p is equal to 0, or not divisible by p otherwise. Here p must be a prime number (when it is non-zero). However, if the function is used when p has no small prime factors, it will either work or give the message “impossible inverse modulo” and a non-trivial divisor of p.

If p = −1, this function returns a matrix whose columns form a basis of the lattice equal to Z^{n} intersected with the lattice generated by the columns of x.

If p = −2, returns a matrix whose columns form a basis of the lattice equal to Z^{n} intersected with the Q-vector space generated by the columns of x.

The library syntax is matrixqz0(x, p).

3.8.35 matsize(x): x being a vector or matrix, returns a row vector with two components, the first being the number of rows (1 for a row vector), the second the number of columns (1 for a column vector).

The library syntax is matsize(x).

3.8.36 matsnf(X, {flag = 0}): if X is a (singular or non-singular) square matrix outputs the vector of elementary divisors of X (i.e. the diagonal of the Smith normal form of X).

The binary digits of flag mean:

1 (complete output): if set, outputs [U, V, D], where U and V are two unimodular matrices such that U × X × V is the diagonal matrix D. Otherwise output only the diagonal of D.

2 (generic input): if set, allows polynomial entries. Otherwise, assume that X has integer coefficients.

4 (cleanup): if set, cleans up the output. This means that elementary divisors equal to 1 will be deleted, i.e. outputs a shortened vector D′ instead of D. If complete output was required, returns [U′, V′, D′] so that U′XV′ = D′ holds. If this flag is set, X is allowed to be of the form D or [U, V, D] as would normally be output with the cleanup flag unset.

The library syntax is matsnf0(X, flag). Also available is smith(X) (flag = 0).

3.8.37 matsolve(x, y): x being an invertible matrix and y a column vector, finds the solution u of x ∗ u = y, using Gaussian elimination. This has the same effect as, but is a bit faster, than x^{−1} ∗ y.

The library syntax is gauss(x, y).

3.8.38 matsolvemod(m, d, y, {flag = 0}): m being any integral matrix, d a vector of positive integer moduli, and y an integral column vector, gives a small integer solution to the system of congruences ∑_{i} m_{i,j}x_{j} ≡ y_{i} (mod d_{i}) if one exists, otherwise returns zero. Shorthand notation: y (resp. d) can be given as a single integer, in which case all the y_{i} (resp. d_{i}) above are taken to be equal to y (resp. d).

If flag = 1, all solutions are returned in the form of a two-component row vector [x, u], where x is a small integer solution to the system of congruences and u is a matrix whose columns give a basis of the homogeneous system (so that all solutions can be obtained by adding x to any linear combination of columns of u). If no solution exists, returns zero.

The library syntax is matsolvemod0(m, d, y, flag). Also available are gaussmodulo(m, d, y) (flag = 0) and gaussmodulo2(m, d, y) (flag = 1).

3.8.39 matsupplement(x): assuming that the columns of the matrix x are linearly independent (if they are not, an error message is issued), finds a square invertible matrix whose first columns are the columns of x, i.e. supplement the columns of x to a basis of the whole space.

The library syntax is suppl(x).

3.8.40 mattranspose(x) or x˜ : transpose of x. This has an effect only on vectors and matrices.

The library syntax is gtrans(x).

3.8.41 qfgaussred(q): decomposition into squares of the quadratic form represented by the symmetric matrix q. The result is a matrix whose diagonal entries are the coefficients of the squares, and the non-diagonal entries represent the bilinear forms. More precisely, if (a_{ij}) denotes the output, one has

q(x) = ∑_{i} a_{ii}(x_{i} + ∑_{j} > ia_{ij}x_{j})^{2}

The library syntax is sqred(x).

3.8.42 qfjacobi(x): x being a real symmetric matrix, this gives a vector having two components: the first one is the vector of eigenvalues of x, the second is the corresponding orthogonal matrix of eigenvectors of x. The method used is Jacobi’s method for symmetric matrices.

The library syntax is jacobi(x).

3.8.43 qflll(x, {flag = 0}): LLL algorithm applied to the columns of the (not necessarily square) matrix x. The columns of x must however be of maximal rank (unless specified otherwise below). The result is a square transformation matrix T such that x·T is an LLL-reduced basis of the lattice generated by the column vectors of x.

If flag = 0 (default), the computations are done with real numbers (i.e. not with rational numbers) hence are fast but as presently programmed (version 2.0.17) are numerically unstable.

If flag = 1, it is assumed that the corresponding Gram matrix is integral. The computation is done entirely with integers and the algorithm is both accurate and quite fast. In this case, x needs not be of maximal rank.

If flag = 2, similar to case 1, except x should be an integer matrix whose columns are linearly independent. The lattice generated by the columns of x is first partially reduced before applying the LLL algorithm. [A basis is said to be partially reduced if |v_{i} ± v_{j}| ≥ |v_{i}| for any two distinct basis vectors v_{i}, v_{j}.]

This can be significantly faster than flag = 1 when one row is huge compared to the other rows.

If flag = 3, all computations are done in rational numbers. This does not incur numerical instability, but is extremely slow. This function is essentially superseded by case 1, so will soon disappear.

If flag = 4, x is assumed to have integral entries, but needs not be of maximal rank. The result is a two-component vector of matrices, the columns of the first matrix representing a basis of the integer kernel of x (not necessarily LLL-reduced) and the columns of the second matrix being an LLL-reduced Z-basis of the image of the matrix x.

If flag = 5, case as case 4, but x may have polynomial coefficients.

If flag = 7, uses an older version of case 0 above.

If flag = 8, same as case 0, where x may have polynomial coefficients.

If flag = 9, variation on case 1, using content.

The library syntax is qflll0(x, flag, prec). Also available are lll(x, prec) (flag = 0), lllint(x) (flag = 1), and lllkerim(x) (flag = 4).

3.8.44 qflllgram(x, {flag = 0}): same as qflll except that the matrix x which must now be a square symmetric real matrix is the Gram matrix of the lattice vectors, and not the coordinates of the vectors themselves. The result is again the transformation matrix T which gives (as columns) the coefficients with respect to the initial basis vectors. The flags have more or less the same meaning, but some are missing. In brief:

flag = 0: numerically unstable in the present version 2.0.17.

flag = 1: x has integer entries, the computations are all done in integers.

flag = 4: x has integer entries, gives the kernel and reduced image.

flag = 5: same as 4 for generic x.

flag = 7: an older version of case 0.

The library syntax is qflllgram0(x, flag, prec). Also available are lllgram(x, prec) (flag = 0), lllgramint(x) (flag = 1), and lllgramkerim(x) (flag = 4).

3.8.45 qfminim(x, b, m, {flag = 0}): x being a square and symmetric matrix representing a positive definite quadratic form, this function deals with the minimal vectors of x, depending on flag.

If flag = 0 (default), seeks vectors of square norm less than or equal to b (for the norm defined by x), and at most 2m of these vectors. The result is a three-component vector, the first component being the number of vectors, the second being the maximum norm found, and the last vector is a matrix whose columns are the vectors found, only one being given for each pair ±v (at most m such pairs).

If flag = 1, ignores m and returns the first vector whose norm is less than b.

In both these cases, x is assumed to have integral entries, and the function searches for the minimal non-zero vectors whenever b = 0.

If flag = 2, x can have non integral real entries, but b = 0 is now meaningless (uses Fincke-Pohst algorithm).

The library syntax is minim(x, b, m) (flag = 0), minim2(x, b, m) (flag = 1), or finally fincke pohst(x, b, m, prec) (flag = 2).

3.8.46 qfperfection(x): x being a square and symmetric matrix with integer entries representing a positive definite quadratic form, outputs the perfection rank of the form. That is, gives the rank of the family of the s symmetric matrices v_{i}v_{i}^{t}, where s is half the number of minimal vectors and the v_{i} (1 ≤ i ≤ s) are the minimal vectors.

As a side note to old-timers, this used to fail bluntly when x had more than 5000 minimal vectors. Beware that the computations can now be very lengthy when x has many minimal vectors.

The library syntax is perf(x).

3.8.47 qfsign(x): signature of the quadratic form represented by the symmetric matrix x. The result is a two-component vector.

The library syntax is signat(x).

3.8.48 setintersect(x, y): intersection of the two sets x and y.

The library syntax is setintersect(x, y).

3.8.49 setisset(x): returns true (1) if x is a set, false (0) if not. In PARI, a set is simply a row vector whose entries are strictly increasing. To convert any vector (and other objects) into a set, use the function Set.

The library syntax is setisset(x), and this returns a long.

3.8.50 setminus(x, y): difference of the two sets x and y, i.e. set of elements of x which do not belong to y.

The library syntax is setminus(x, y).

3.8.51 setsearch(x, y, {flag = 0}): searches if y belongs to the set x. If it does and flag is zero or omitted, returns the index j such that x[j] = y, otherwise returns 0. If flag is non-zero returns the index j where y should be inserted, and 0 if it already belongs to x (this is meant to be used in conjunction with listinsert).

This function works also if x is a sorted list (see listsort).

The library syntax is setsearch(x, y, flag) which returns a long integer.

3.8.52 setunion(x, y): union of the two sets x and y.

The library syntax is setunion(x, y).

3.8.53 trace(x): this applies to quite general x. If x is not a matrix, it is equal to the sum of x and its conjugate, except for polmods where it is the trace as an algebraic number.

For x a square matrix, it is the ordinary trace. If x is a non-square matrix (but not a vector), an error occurs.

The library syntax is gtrace(x).

3.8.54 vecextract(x, y, {z}): extraction of components of the vector or matrix x according to y. In case x is a matrix, its components are as usual the columns of x. The parameter y is a component specifier, which is either an integer, a string describing a range, or a vector.

If y is an integer, it is considered as a mask: the binary bits of y are read from right to left, but correspond to taking the components from left to right. For example, if y = 13 = (1101)_{2} then the components 1,3 and 4 are extracted.

If y is a vector, which must have integer entries, these entries correspond to the component numbers to be extracted, in the order specified.

If y is a string, it can be

• a single (non-zero) index giving a component number (a negative index means we start counting from the end).

• a range of the form "a..b", where a and b are indexes as above. Any of a and b can be omitted; in this case, we take as default values a = 1 and b = −1, i.e. the first and last components respectively. We then extract all components in the interval [a, b], in reverse order if b < a.

In addition, if the first character in the string is ^, the complement of the given set of indices is taken.

If z is not omitted, x must be a matrix. y is then the line specifier, and z the column specifier, where the component specifier is as explained above.

?  v = [a, b, c, d, e];
?  vecextract(v, 5) \\ mask
%1 = [a, c]
?  vecextract(v, [4, 2, 1]) \\ component list
%2 = [d, b, a]
?  vecextract(v, "2..4") \\ interval
%3 = [b, c, d]
?  vecextract(v, "-1..-3") \\ interval + reverse order
%4 = [e, d, c]
?  vecextract([1,2,3], "^2") \\ complement
%5 = [1, 3]
?  vecextract(matid(3), "2..", "..")
%6 =
[0 1 0]
[0 0 1]

The library syntax is extract(x, y) or matextract(x, y, z).

3.8.55 vecsort(x, {k}, {flag = 0}): sorts the vector x in ascending order, using the heapsort method. x must be a vector, and its components integers, reals, or fractions.

If k is present and is an integer, sorts according to the value of the k-th subcomponents of the components of x. k can also be a vector, in which case the sorting is done lexicographically according to the components listed in the vector k. For example, if k = [2, 1, 3], sorting will be done with respect to the second component, and when these are equal, with respect to the first, and when these are equal, with respect to the third.

The binary digits of flag mean:

• 1: indirect sorting of the vector x, i.e. if x is an n-component vector, returns a permutation of [1, 2, …, n] which applied to the components of x sorts x in increasing order. For example,

vecextract(x, vecsort(x,,1)) is equivalent to vecsort(x).

• 2: sorts x by ascending lexicographic order (as per the lex comparison function).

The library syntax is vecsort0(x, k, flag). To omit k, use NULL instead. You can also use the simpler functions

sort(x) (= vecsort0(x, NULL, 0)).

indexsort(x) (= vecsort0(x, NULL, 1)).

lexsort(x) (= vecsort0(x, NULL, 2)).

Also available are sindexsort and sindexlexsort which return a vector (type t VEC) of Clong integers v, where v[1] … v[n] contain the indices. Note that the resulting v is not a valid PARI object, but is in general easier to use in C programs!

3.9 Sums, products, integrals and similar functions.

Although the GP calculator is programmable, it is useful to have preprogrammed a number of loops, including sums, products, and a certain number of recursions. Also, a number of functions from numerical analysis like numerical integration and summation of series will be described here.

One of the parameters in these loops must be the control variable, hence a simple variable name. The last parameter can be any legal PARI expression, including of course expressions using loops. Since it is much easier to program directly the loops in library mode, these functions are mainly useful for GP programming. The use of these functions in library mode is a little tricky and its explanation will be mostly omitted, although the reader can try and figure it out by himself by checking the example given for the sum function. In this section we only give the library syntax, with no semantic explanation.

The letter X will always denote any simple variable name, and represents the formal parameter used in the function.

(numerical) integration: A number of Romberg-like integration methods are implemented (see intnum as opposed to intformal which we already described). The user should not require too much accuracy: 18 or 28 decimal digits is OK, but not much more. In addition, analytical cleanup of the integral must have been done: there must be no singularities in the interval or at the boundaries. In practice this can be accomplished with a simple change of variable. Furthermore, for improper integrals, where one or both of the limits of integration are plus or minus infinity, the function must decrease sufficiently rapidly at infinity. This can often be accomplished through integration by parts.

Note that infinity can be represented with essentially no loss of accuracy by 1e4000. However beware of real underflow when dealing with rapidly decreasing functions. For example, if one wants to compute the ∫_{0}^{∞} e^{−x^{2}} dx to 28 decimal digits, then one should set infinity equal to 10 for example, and certainly not to 1e4000.

The integrand may have values belonging to a vector space over the real numbers; in particular, it can be complex-valued or vector-valued.

See also the discrete summation methods below (sharing the prefix sum).

3.9.1 intnum(X = a, b, expr, {flag = 0}): numerical integration of expr (smooth in ]a, b[), with respect to X.

Set flag = 0 (or omit it altogether) when a and b are not too large, the function is smooth, and can be evaluated exactly everywhere on the interval [a, b].

If flag = 1, uses a general driver routine for doing numerical integration, making no particular assumption (slow).

flag = 2 is tailored for being used when a or b are infinite. One must have ab > 0, and in fact if for example b = +∞, then it is preferable to have a as large as possible, at least a ≥ 1.

If flag = 3, the function is allowed to be undefined (but continuous) at a or b, for example the function sin(x)/x at x = 0.

    The library syntax is intnum0(entree*e,GEN a,GEN b,char*expr,long flag,long prec).

3.9.2 matrix(m, n, {X}, {Y }, {expr = 0}): creation of the m × n matrix whose coefficients are given by the expression expr. There are two formal parameters in expr, the first one (X) corresponding to the rows, the second (Y) to the columns, and X goes from 1 to m, Y goes from 1 to n. If one of the last 3 parameters is omitted, fill the matrix with zeroes.

    The library syntax is matrice(GEN nlig,GEN ncol,entree *e1,entree *e2,char *expr).

3.9.3 prod(X = a, b, expr, {x = 1}): product of expression expr, initialized at x, the formal parameter X going from a to b. As for sum, the main purpose of the initialization parameter x is to force the type of the operations being performed. For example if it is set equal to the integer 1, operations will start being done exactly. If it is set equal to the real 1., they will be done using real numbers having the default precision. If it is set equal to the power series 1 + O(X^{k}) for a certain k, they will be done using power series of precision at most k. These are the three most common initializations.

As an extreme example, compare

    ?  prod(i=1, 100, 1-X^i); \\ this has degree 5050 !!
    time = 3,335 ms.
    ?  prod(i=1, 100, 1-X^i, 1+O(X^101))
    time = 43 ms.
    %2 = 1 - X - X^2 + X^5 + X^7 - X^12 - X^15 + X^22 + X^26 - X^35 - X^40 + X^51
    + X^57 - X^70 - X^77 + X^92 + X^100 + O(X^101)
    The library syntax is produit(entree *ep, GEN a, GEN b, char *expr, GEN x).

3.9.4 prodeuler(X = a, b, expr): product of expression expr, initialized at 1. (i.e. to a real number equal to 1 to the current realprecision), the formal parameter X ranging over the prime numbers between a and b.

    The library syntax is prodeuler(entree *ep, GEN a, GEN b, char *expr, long prec).

3.9.5 prodinf(X = a, expr, {flag = 0}): infinite product of expression expr, the formal parameter X starting at a. The evaluation stops when the relative error of the expression minus 1 is less than the default precision. The expressions must always evaluate to an element of C.

If flag = 1, do the product of the (1 + expr) instead.

The library syntax is prodinf(entree *ep, GEN a, char *expr, long prec) (flag = 0), or prodinf1 with the same arguments (flag = 1).

3.9.6 solve(X = a, b, expr): find a real root of expression expr between a and b, under the condition expr(X = a) ∗ expr(X = b) ≤ 0. This routine uses Brent’s method. This can fail miserably if expr is not defined in the whole of [a, b] (try solve(x=1, 2, tan(x)).

    The library syntax is zbrent(entree *ep, GEN a, GEN b, char *expr, long prec).

3.9.7 sum(X = a, b, expr, {x = 0}): sum of expression expr, initialized at x, the formal parameter going from a to b. As for prod, the initialization parameter x may be given to force the type of the operations being performed.

As an extreme example, compare

?  sum(i=1, 5000, 1/i); \\rational number:  denominator has 2166 digits.
time = 1,241 ms.
?  sum(i=1, 5000, 1/i, 0.)
time = 158 ms.
%2 = 9.094508852984436967261245533
The library syntax is somme(entree *ep, GEN a, GEN b, char *expr, GEN x). This is to

be used as follows: ep represents the dummy variable used in the expression expr

/* compute a^2 + . . . + b^2 */
{
  /* define the dummy variable "i" */
  entree *ep = gp variable("i");
  /* sum for a <= i <= b */
  return somme(ep, a, b, "i^2", gzero);
}

3.9.8 sumalt(X = a, expr, {flag = 0}): numerical summation of the series expr, which should be an alternating series, the formal variable X starting at a.

If flag = 0, use an algorithm of F. Villegas as modified by D. Zagier. This is much better than Euler-Van Wijngaarden’s method which was used formerly. Beware that the stopping criterion is that the term gets small enough, hence terms which are equal to 0 will create problems and should be removed.

If flag = 1, use a variant with slightly different polynomials. Sometimes faster.

Divergent alternating series can sometimes be summed by this method, as well as series which are not exactly alternating (see for example Section 2.6.3).

Important hint: a significant speed gain can be obtained by writing the (−1)^{X} which may occur in the expression as (1. - X%2*2).

The library syntax is sumalt(entree *ep, GEN a, char *expr, long flag, long prec).

3.9.9 sumdiv(n, X, expr): sum of expression expr over the positive divisors of n.

In the present version 2.0.17, n is restricted to being less than 2^{31}.

The library syntax is divsum(entree *ep, GEN num, char *expr).

3.9.10 suminf(X = a, expr): infinite sum of expression expr, the formal parameter X starting at a. The evaluation stops when the relative error of the expression is less than the default precision. The expressions must always evaluate to a complex number.

The library syntax is suminf(entree *ep, GEN a, char *expr, long prec).

3.9.11 sumpos(X = a, expr, {flag = 0}): numerical summation of the series expr, which must be a series of terms having the same sign, the formal variable X starting at a. The algorithm used is Van Wijngaarden’s trick for converting such a series into an alternating one, and is quite slow. Beware that the stopping criterion is that the term gets small enough, hence terms which are equal to 0 will create problems and should be removed.

If flag = 1, use slightly different polynomials. Sometimes faster.

    The library syntax is sumpos(entree *ep, GEN a, char *expr, long flag, long prec).

3.9.12 vector(n, {X}, {expr = 0}): creates a row vector (type t VEC) with n components whose components are the expression expr evaluated at the integer points between 1 and n. If one of the last two arguments is omitted, fill the vector with zeroes.

The library syntax is vecteur(GEN nmax, entree *ep, char *expr).

3.9.13 vectorv(n, X, expr): as vector, but returns a column vector (type t COL).

The library syntax is vvecteur(GEN nmax, entree *ep, char *expr).

3.10 Plotting functions.

Although plotting is not even a side purpose of PARI, a number of plotting functions are provided. Moreover, a lot of people felt like suggesting ideas or submitting huge patches for this section of the code. Among these, special thanks go to Klaus-Peter Nischke who suggested the recursive plotting and the forking/resizing stuff under X11, and Ilya Zakharevich who undertook a complete rewrite of the graphic code, so that most of it is now platform-independent and should be relatively easy to port or expand.

These graphic functions are either

• high-level plotting functions (all the functions starting with ploth) in which the user has little to do but explain what type of plot he wants, and whose syntax is similar to the one used in the preceding section (with somewhat more complicated flags).

• low-level plotting functions, where every drawing primitive (point, line, box, etc.) must be specified by the user. These low-level functions (called rectplot functions, sharing the prefix plot) work as follows. You have at your disposal 16 virtual windows which are filled independently, and can then be physically ORed on a single window at user-defined positions. These windows are numbered from 0 to 15, and must be initialized before being used by the function plotinit, which specifies the height and width of the virtual window (called a rectwindow in the sequel). At all times, a virtual cursor (initialized at [0, 0]) is associated to the window, and its current value can be obtained using the function plotcursor.

A number of primitive graphic objects (called rect objects) can then be drawn in these windows, using a default color associated to that window (which can be changed under X11, using the plotcolor function, black otherwise) and only the part of the object which is inside the window will be drawn, with the exception of polygons and strings which are drawn entirely (but the virtual cursor can move outside of the window). The ones sharing the prefix plotr draw relatively to the current position of the virtual cursor, the others use absolute coordinates. Those having the prefix plotrecth put in the rectwindow a large batch of rect objects corresponding to the output of the related ploth function.

Finally, the actual physical drawing is done using the function plotdraw. Note that the windows are preserved so that further drawings using the same windows at different positions or different windows can be done without extra work. If you want to erase a window (and free the corresponding memory), use the function plotkill. It is not possible to partially erase a window. Erase it completely, initialize it again and then fill it with the graphic objects that you want to keep.

In addition to initializing the window, you may want to have a scaled window to avoid unnecessary conversions. For this, use the function plotscale below. As long as this function is not called, the scaling is simply the number of pixels, the origin being at the upper left and the y-coordinates going downwards.

Note that in the present version 2.0.17 all these plotting functions (both low and high level) have been written for the X11-window system (hence also for GUI’s based on X11 such as Openwindows and Motif), and for Sunview/Suntools only, though very little code remains which is actually platform-dependent. A Macintosh, and an Atari/Gem port were provided for previous versions. These may be adapted in future releases.

Under X11/Suntools, the physical window (opened by plotdraw or any of the ploth* functions) is completely separated from GP (technically, a fork is done, and the non-graphical memory is immediately freed in the child process), which means you can go on working in the current GP session, without having to kill the window first. Under X11, this window can be closed, enlarged or reduced using the standard window manager functions. No zooming procedure is implemented though (yet).

• Finally, note that in the same way that printtex allows you to have a TEX output corresponding to printed results, the functions starting with ps allow you to have PostScript output of the plots. This will not be absolutely identical with the screen output, but will be sufficiently close. Note that you can use PostScript output even if you do not have the plotting routines enabled. The PostScript output is written in a file whose name is derived from the psfile default (./pari.ps if you did not tamper with it). Each time a new PostScript output is asked for, the PostScript output is appended to that file. Hence the user must remove this file, or change the value of psfile, first if he does not want unnecessary drawings from preceding sessions to appear. On the other hand, in this manner as many plots as desired can be kept in a single file.

None of the graphic functions are available within the PARI library, you must be under GP to use them. The reason for that is that you really should not use PARI for heavy-duty graphical work, there are much better specialized alternatives around. This whole set of routines was only meant as a convenient, but simple-minded, visual aid. If you really insist on using these in your program (we warned you), the source (plot*.c) should be readable enough for you to achieve something.

3.10.1 plot(X = a, b, expr): crude (ASCII) plot of the function represented by expression expr from a to b.

3.10.2 plotbox(w, x2, y2): let (x1, y1) be the current position of the virtual cursor. Draw in the rectwindow w the outline of the rectangle which is such that the points (x1, y1) and (x2, y2) are opposite corners. Only the part of the rectangle which is in w is drawn. The virtual cursor does not move.

3.10.3 plotclip(w): ‘clips’ the content of rectwindow w, i.e remove all parts of the drawing that would not be visible on the screen. Together with plotcopy this function enables you to draw on a scratchpad before commiting the part you’re interested in to the final picture.

3.10.4 plotcolor(w, c): set default color to c in rectwindow w. In present version 2.0.17, this is only implemented for X11 window system, and you only have the following palette to choose from:

1=black, 2=blue, 3=sienna, 4=red, 5=cornsilk, 6=grey, 7=gainsborough.

Note that it should be fairly easy for you to hardwire some more colors by tweaking the files rect.h and plotX.c. User-defined colormaps would be nice, and may be available in future versions.

3.10.5 plotcopy(w1, w2, dx, dy): copy the contents of rectwindow w1 to rectwindow w2, with offset (dx, dy).

3.10.6 plotcursor(w): give as a 2-component vector the current (scaled) position of the virtual cursor corresponding to the rectwindow w.

3.10.7 plotdraw(list): physically draw the rectwindows given in list which must be a vector whose number of components is divisible by 3. If list = [w1, x1, y1, w2, x2, y2, …], the windows w1, w2, etc. are physically placed with their upper left corner at physical position (x1, y1), (x2, y2),… respectively, and are then drawn together. Overlapping regions will thus be drawn twice, and the windows are considered transparent. Then display the whole drawing in a special window on your screen.

3.10.8 plotfile(s): set the output file for plotting output. Special filename - redirects to the same place as PARI output.

3.10.9 ploth(X = a, b, expr, {flag = 0}, {n = 0}): high precision plot of the function y = f(x) represented by the expression expr, x going from a to b. This opens a specific window (which is killed whenever you click on it), and returns a four-component vector giving the coordinates of the bounding box in the form [xmin, xmax, ymin, ymax].

Important note: Since this may involve a lot of function calls, it is advised to keep the current precision to a minimum (e.g. 9) before calling this function.

n specifies the number of reference point on the graph (0 means use the hardwired default values, that is: 1000 for general plot, 1500 for parametric plot, and 15 for recursive plot).

If no flag is given, expr is either a scalar expression f(X), in which case the plane curve y = f(X) will be drawn, or a vector [f_{1}(X), …, f_{k}(X)], and then all the curves y = f_{i}(X) will be drawn in the same window.

The binary digits of flag mean:

• 1: parametric plot. Here expr must be a vector with an even number of components. Successive pairs are then understood as the parametric coordinates of a plane curve. Each of these are then drawn.

For instance:

    ploth(X=0,2*Pi,[sin(X),cos(X)],1) will draw a circle.

ploth(X=0,2*Pi,[sin(X),cos(X)]) will draw two entwined sinusoidal curves.

    ploth(X=0,2*Pi,[X,X,sin(X),cos(X)],1) will draw a circle and the line y = x.

• 2: recursive plot. If this flag is set, only one curve can be drawn at time, i.e. expr must be either a two-component vector (for a single parametric curve, and the parametric flag has to be set), or a scalar function. The idea is to choose pairs of successive reference points, and if their middle point is not too far away from the segment joining them, draw this as a local approximation to the curve. Otherwise, add the middle point to the reference points. This is very fast, and usually more precise than usual plot. Compare the results of

              ploth(X = −1, 1, sin(1/X), 2) and ploth(X = −1, 1, sin(1/X))

for instance. But beware that if you are extremely unlucky, or choose too few reference points, you may draw some nice polygon bearing little resemblance to the original curve. For instance you should never plot recursively an odd function in a symmetric interval around 0. Try

    ploth(x = -20, 20, sin(x), 2)

to see why. Hence, it’s usually a good idea to try and plot the same curve with slightly different parameters.

• 8: do not print the x-axis.

• 16: do not print the y-axis.

• 32: do not print frame.

• 64: only plot reference points, do not join them.

3.10.10 plothraw(listx, listy, {flag = 0}): given listx and listy two vectors of equal length, plots (in high precision) the points whose (x, y)-coordinates are given in listx and listy. Automatic positioning and scaling is done, but with the same scaling factor on x and y. If flag is non-zero, join points.

3.10.11 plothsizes(): return data corresponding to the output window in the form of a 6component vector: window width and height, sizes for ticks in horizontal and vertical directions (this is intended for the gnuplot interface and is currently not significant), width and height of characters.

3.10.12 plotinit(w, x, y): initialize the rectwindow w to width x and height y, and position the virtual cursor at (0, 0). This destroys any rect objects you may have already drawn in w.

The plotting device imposes an upper bound for x and y, for instance the number of pixels for screen output. These bounds are available through the plothsizes function. The following sequence initializes in a portable way (i.e independant of the output device) a window of maximal size, accessed through coordinates in the [0, 1000] × [0, 1000] range :

    s = plothsizes();
    plotinit(0, s[1]-1, s[2]-1);
    plotscale(0, 0,1000, 0,1000);

3.10.13 plotkill(w): erase rectwindow w and free the corresponding memory. Note that if you want to use the rectwindow w again, you have to use initrect first to specify the new size. So it’s better in this case to use initrect directly as this throws away any previous work in the given rectwindow.

3.10.14 plotlines(w, X, Y, {flag = 0}): draw on the rectwindow w the polygon such that the (x,y)-coordinates of the vertices are in the vectors of equal length X and Y. For simplicity, the whole polygon is drawn, not only the part of the polygon which is inside the rectwindow. If flag is non-zero, close the polygon. In any case, the virtual cursor does not move.

X and Y are allowed to be scalars (in this case, both have to). There, a single segment will be drawn, between the virtual cursor current position and the point (X, Y). And only the part thereof which actually lies within the boundary of w. Then move the virtual cursor to (X, Y), even if it is outside the window. If you want to draw a line from (x1, y1) to (x2, y2) where (x1, y1) is not necessarily the position of the virtual cursor, use plotmove(w,x1,y1) before using this function.

3.10.15 plotlinetype(w, type): this is intended for the gnuplot interface and is currently not significant.

3.10.16 plotmove(w, x, y): move the virtual cursor of the rectwindow w to position (x, y).

3.10.17 plotpoints(w, X, Y): draw on the rectwindow w the points whose (x, y)-coordinates are in the vectors of equal length X and Y and which are inside w. The virtual cursor does not move. This is basically the same function as plothraw, but either with no scaling factor or with a scale chosen using the function plotscale.

As was the case with the plotlines function, X and Y are allowed to be (simultaneously) scalar. In this case, draw the single point (X, Y) on the rectwindow w (if it is actually inside w), and in any case move the virtual cursor to position (x, y).

3.10.18 plotpointsize(w, size): changes the “size” of following points in rectwindow w. If w = −1, change it in all rectwindows. This only works in the gnuplot interface.

3.10.19 plotpointtype(w, type): this is intended for the gnuplot interface and is currently not significant.

3.10.20 plotrbox(w, dx, dy): draw in the rectwindow w the outline of the rectangle which is such that the points (x1, y1) and (x1 + dx, y1 + dy) are opposite corners, where (x1, y1) is the current position of the cursor. Only the part of the rectangle which is in w is drawn. The virtual cursor does not move.

3.10.21 plotrecth(w, X = a, b, expr, {flag = 0}, {n = 0}): writes to rectwindow w the curve output of ploth(w, X = a, b, expr, flag, n).

3.10.22 plotrecthraw(w, data, {flag = 0}): plot graph(s) for data in rectwindow w. flag has the same significance here as in ploth, though recursive plot is no more significant. data

is a vector of vectors, each corresponding to a list a coordinates. If parametric plot is set, there must be an even number of vectors, each successive pair corresponding to a curve. Otherwise, the first one containe the x coordinates, and the other ones contain the y-coordinates of curves to plot.

3.10.23 plotrline(w, dx, dy): draw in the rectwindow w the part of the segment (x1, y1) − (x1 + dx, y1 + dy) which is inside w, where (x1, y1) is the current position of the virtual cursor, and move the virtual cursor to (x1 + dx, y1 + dy) (even if it is outside the window).

3.10.24 plotrmove(w, dx, dy): move the virtual cursor of the rectwindow w to position (x1 + dx, y1 + dy), where (x1, y1) is the initial position of the cursor (i.e. to position (dx, dy) relative to the initial cursor).

3.10.25 plotrpoint(w, dx, dy): draw the point (x1 + dx, y1 + dy) on the rectwindow w (if it is inside w), where (x1, y1) is the current position of the cursor, and in any case move the virtual cursor to position (x1 + dx, y1 + dy).

3.10.26 plotscale(w, x1, x2, y1, y2): scale the local coordinates of the rectwindow w so that x goes from x1 to x2 and y goes from y1 to y2 (x2 < x1 and y2 < y1 being allowed). Initially, after the initialization of the rectwindow w using the function plotinit, the default scaling is the graphic pixel count, and in particular the y axis is oriented downwards since the origin is at the upper left. The function plotscale allows to change all these defaults and should be used whenever functions are graphed.

3.10.27 plotstring(w, x): draw on the rectwindow w the String x (see Section 2.4), at the current position of the cursor.

3.10.28 plotterm(type): this is intended for the gnuplot interface and is currently not significant.

3.10.29 psdraw(list): same as plotdraw, except that the output is a PostScript program appended to the psfile.

3.10.30 psploth(X = a, b, expr): same as ploth, except that the output is a PostScript program appended to the psfile.

3.10.31 psplothraw(listx, listy): same as plothraw, except that the output is a PostScript program appended to the psfile.

3.11 Programming under GP.

3.11.1 Control statements.

A number of control statements are available under GP. They are simpler and have a syntax slightly different from their C counterparts, but are quite powerful enough to write any kind of program. Some of them are specific to GP, since they are made for number theorists. As usual, X will denote any simple variable name, and seq will always denote a sequence of expressions, including the empty sequence.

3.11.1.1 break({n = 1}): interrupts execution of current seq, and immediately exits from the n innermost enclosing loops, within the current function call (or the top level loop). n must be bigger than 1. If n is greater than the number of enclosing loops, all enclosing loops are exited.

3.11.1.2 for(X = a, b, seq): the formal variable X going from a to b, the seq is evaluated. Nothing is done if a > b. a and b must be in R.

3.11.1.3 fordiv(n, X, seq): the formal variable X ranging through the positive divisors of n, the sequence seq is evaluated. n must be of type integer.

3.11.1.4 forprime(X = a, b, seq): the formal variable X ranging over the prime numbers between a to b (including a and b if they are prime), the seq is evaluated. Nothing is done if a > b. Note that a and b must be in R.

3.11.1.5 forstep(X = a, b, s, seq): the formal variable X going from a to b, in increments of s, the seq is evaluated. Nothing is done if s > 0 and a > b or if s < 0 and a < b. s must be in R^{∗} or a vector of steps [s_{1}, …, s_{n}]. In the latter case, the successive steps are used in the order they appear in s.

?  forstep(x=5, 20, [2,4], print(x))
5
7
11
13
17
19

3.11.1.6 forsubgroup(H = G, {B}, seq): executes seq for each subgroup H of the abelian group G (given in SNF form or as a vector of elementary divisors), whose index is bounded by bound. The subgroups are not ordered in any obvious way, unless G is a p-group in which case Birkhoff’s algorithm produces them by decreasing index. A subgroup is given as a matrix whose columns give its generators on the implicit generators of G. For example, the following prints all subgroups of index less than 2 in G = Z/2Zg_{1} × Z/2Zg_{2} :

?  G = [2,2]; forsubgroup(H=G, 2, print(H))
[1; 1]
[1; 2]
[2; 1]
[1, 0; 1, 1]

The last one, for instance is generated by (g_{1}, g_{1} + g_{2}). This routine is intended to treat huge groups, when subgrouplist is not an option due to the sheer size of the output.

For maximal speed the subgroups have been left as produced by the algorithm. To print them in canonical form (as left divisors of G in HNF form), one can for instance use

?  G = matdiagonal([2,2]); forsubgroup(H=G, 2, print(mathnf(concat(G,H))))
[2, 1; 0, 1]
[1, 0; 0, 2]
[2, 0; 0, 1]
[1, 0; 0, 1]

Note that in this last representation, the index [G : H] is given by the determinant.

3.11.1.7 forvec(X = v, seq, {flag = 0}): v being an n-component vector (where n is arbitrary) of two-component vectors [a_{i}, b_{i}] for 1 ≤ i ≤ n, the seq is evaluated with the formal variable X[1] going from a_{1} to b_{1},…,X[n] going from a_{n} to b_{n}. The formal variable with the highest index moves the fastest. If flag = 1, generate only nondecreasing vectors X, and if flag = 2, generate only strictly increasing vectors X.

3.11.1.8 if(a, {seq1 }, {seq2 }): if a is non-zero, the expression sequence seq1 is evaluated, otherwise the expression seq2 is evaluated. Of course, seq1 or seq2 may be empty, so if (a,seq) evaluates seq if a is not equal to zero (you don’t have to write the second comma), and does nothing otherwise, whereas if (a,,seq) evaluates seq if a is equal to zero, and does nothing otherwise. You could get the same result using the ! (not) operator: if (!a,seq).

Note that the boolean operators && and || are evaluated according to operator precedence as explained in Section 2.4, but that, contrary to other operators, the evaluation of the arguments is stopped as soon as the final truth value has been determined. For instance

    if (reallydoit && longcomplicatedfunction(), . . .)

is a perfectly safe statement.

Recall that functions such as break and next operate on loops (such as forxxx, while, until). The if statement is not a loop (obviously!).

3.11.1.9 next({n = 1}): interrupts execution of current seq, resume the next iteration of the innermost enclosing loop, within the current fonction call (or top level loop). If n is specified, resume at the n-th enclosing loop. If n is bigger than the number of enclosing loops, all enclosing loops are exited.

3.11.1.10 return({x = 0}): returns from current subroutine, with result x.

3.11.1.11 until(a, seq): evaluates expression sequence seq until a is not equal to 0 (i.e. until a is true). If a is initially not equal to 0, seq is evaluated once (more generally, the condition on a is tested after execution of the seq, not before as in while).

3.11.1.12 while(a, seq): while a is non-zero evaluate the expression sequence seq. The test is made before evaluating the seq, hence in particular if a is initially equal to zero the seq will not be evaluated at all.

3.11.2 Specific functions used in GP programming.

In addition to the general PARI functions, it is necessary to have some functions which will be of use specifically for GP, though a few of these can be accessed under library mode. Before we start describing these, we recall the difference between strings and keywords (see Section 2.6.5): the latter don’t get expanded at all, and you can type them without any enclosing quotes. The former are dynamic objects, where everything outside quotes gets immediately expanded.

We need an additional notation for this chapter. An argument between braces, followed by a star, like {str}∗, means that any number of such arguments (possibly none) can be given.

3.11.2.1 addhelp(S, str): changes the help message for the symbol S. The string str is expanded on the spot and stored as the online help for S. If S is a function you have defined, its definition will still be printed before the message str. It is recommended that you document global variables and user functions in this way. Of course GP won’t protest if you don’t do it.

There’s nothing to prevent you from modifying the help of built-in PARI functions (but if you do, we’d like to hear why you needed to do it!).

3.11.2.2 alias(newkey, key): defines the keyword newkey as an alias for keyword key. key must correspond to an existing function name. This is different from the general user macros in that alias expansion takes place immediately upon execution, without having to look up any function code, and is thus much faster. A sample alias file misc/gpalias is provided with the standard distribution. Alias commands are meant to be read upon startup from the .gprc file, to cope with function names you are dissatisfied with, and should be useless in interactive usage.

3.11.2.3 allocatemem({x = 0}): this is a very special operation which allows the user to change the stack size after initialization. x must be a non-negative integer. If x! = 0, a new stack of size 16 ∗ ⌈x/16⌉ bytes will be allocated, all the PARI data on the old stack will be moved to the new one, and the old stack will be discarded. If x = 0, the size of the new stack will be twice the size of the old one.

Although it is a function, this must be the last instruction in any GP sequence. The technical reason is that this routine usually moves the stack, so objects from the current sequence might not be correct anymore. Hence, to prevent such problems, this routine terminates by a longjmp (just as an error would) and not by a return.

The library syntax is allocatemoremem(x), where x is an unsigned long, and the return type is void. GP uses a variant which ends by a longjmp.

3.11.2.4 default({key}, {val}, {flag}): sets the default corresponding to keyword key to value val. val is a string (which of course accepts numeric arguments without adverse effects, due to the expansion mechanism). See Section 2.1 for a list of available defaults, and Section 2.2 for some shortcut alternatives.

If val is omitted, prints the current value of default key. If key is omitted, prints the current values of all the defaults. If flag is set, returns the result instead of printing it.

3.11.2.5 error({str}∗): outputs its argument list (each of them interpreted as a string), then interrupts the running GP program, returning to the input prompt.

    Example: error("n = ", n, " is not squarefree !").

Note that, due to the automatic concatenation of strings, you could in fact use only one argument, just by suppressing the commas.

UNIX: 3.11.2.6 extern(str): the string str is the name of an external command (i.e. one you would type from your UNIX shell prompt). This command is immediately run and its input fed into GP, just as if read from a file.

3.11.2.7 getheap(): returns a two-component row vector giving the number of objects on the heap and the amount of memory they occupy in long words. Useful mainly for debugging purposes.

The library syntax is getheap().

3.11.2.8 getrand(): returns the current value of the random number seed. Useful mainly for debugging purposes.

The library syntax is getrand(), returns a C long.

3.11.2.9 getstack(): returns the current value of top − avma, i.e. the number of bytes used up to now on the stack. Should be equal to 0 in between commands. Useful mainly for debugging purposes.

The library syntax is getstack(), returns a C long.

3.11.2.10 gettime(): returns the time (in milliseconds) elapsed since either the last call to gettime, or to the beginning of the containing GP instruction (if inside GP), whichever came last.

The library syntax is gettime(), returns a C long.

3.11.2.11 global({list of variables}): declares the corresponding variables to be global. From now on, you will be forbidden to use them as formal parameters for function definitions or as loop indexes. This is especially useful when patching together various scripts, possibly written with different naming conventions. For instance the following situation is dangerous:

    p = 3 \\ fix characteristic
    ...
    forprime(p = 2, N, ...)
    f(p) = ...

since within the loop or within the function’s body, the true global value of p will be hidden. If the statement global(p = 3) appears at the beginning of the script, then both expressions will trigger syntax errors.

Calling global without arguments prints the list of global variables in use. In particular, eval(global) will output the values of all local variables.

3.11.2.12 input(): reads a string, interpreted as a GP expression, from the input file, usually standard input (i.e. the keyboard). If a sequence of expressions is given, the result is the result of the last expression of the sequence. When using this instruction, it is useful to prompt for the string by using the print1 function. Note that in the present version 2.19 of pari.el, when using GP under GNU Emacs (see Section 2.9) one must prompt for the string, with a string which ends with the same prompt as any of the previous ones (a "? " will do for instance).

UNIX: 3.11.2.13 install(name, code, {gpname}, {lib}): loads from dynamic library lib the function name. Assigns to it the name gpname in this GP session, with argument code code (see Section 4.9.2 for an explanation of those). If lib is omitted, uses libpari.so. If gpname is omitted, uses name.

This function is useful for adding custom functions to the GP interpreter. But it also gives you access to all (non static) functions defined in the PARI library. For instance, the function GEN addii(GEN x, GEN y) adds two PARI integers, and is not directly accessible under GP (it’s eventually called by the + operator of course):

    ?  install("addii", "GG")
    ?  addii(1, 2)
    %1 = 3

Caution: This function may not work on all systems, especially when GP has been compiled statically. In that case, the first use of an installed function will provoke a Segmentation Fault, i.e. a major internal blunder (this should never happen with a dynamically linked executable). This used to be the fate of statically linked gp on Linux and OSF1 up to and including version 2.0.3.

Hence, if you intend to use this function, please check first on some harmless example such as the one above that it works properly on your machine.

3.11.2.14 kill(x): kills the present value of the variable, alias or user-defined function x (you can only kill your own functions). The corresponding identifier can now be used to name any GP object (variable or function). This is the only way to replace a variable by a function having the same name (or the other way round), as in the following example:

?  f = 1
%1 = 1
?  f(x) = 0
  ***   unused characters: f(x)=0
                            ^----
?  kill(f)
?  f(x) = 0
?  f()
%2 = 0

When you kill a variable, all objects that used it become invalid. You can still display them, even though the killed variable will be printed in a funny way (following the same convention as used by the library function fetch var, see Section 4.6). For example:

?  a^2 + 1
%1 = a^2 + 1
?  kill(a)
?  %1
%2 = #<1>^2 + 1

If you simply want to restore a variable to its original value (monomial of degree one), use the quote operator: x = 'x.

3.11.2.15 print({str}∗): outputs its (string) arguments in raw format, ending with a newline.

3.11.2.16 print1({str}∗): outputs its (string) arguments in raw format, without ending with a newline (note that you can still embed newlines within your strings, using the \n notation !).

3.11.2.17 printp({str}∗): outputs its (string) arguments in prettyprint (beautified) format, ending with a newline.

3.11.2.18 printp1({str}∗): outputs its (string) arguments in prettyprint (beautified) format, without ending with a newline.

3.11.2.19 printtex({str}∗): outputs its (string) arguments in TEX format. This output can then be used in a TEX manuscript. The printing is done on the standard output. If you want to print it to a file you should use writetex (see there).

Another possibility is to enable the log default (see Section 2.1). You could for instance do:

default(logfile, "new.tex");
default(log, 1);
printtex(result);

(You can use the automatic string expansion/concatenation process to have dynamic file names if you wish).

3.11.2.20 quit(): exits GP.

3.11.2.21 read({str}): reads in the file whose name results from the expansion of the string str. If str is omitted, re-reads the last file that was fed into GP. The return value is the result of the last expression evaluated.

3.11.2.22 reorder({x = [ ]}): x must be a vector. If x is the empty vector, this gives the vector whose components are the existing variables in increasing order (i.e. in decreasing importance). Killed variables (see kill) will be shown as 0. If x is non-empty, it must be a permutation of variable names, and this permutation gives a new order of importance of the variables, for output only. For example, if the existing order is [x,y,z], then after reorder([z,x]) the order of importance of the variables, with respect to output, will be [z,y,x]. The internal representation is unaffected.

3.11.2.23 setrand(n): reseeds the random number generator to the value n. The initial seed is n = 1.

The library syntax is setrand(n), where n is a long. Returns n.

UNIX: 3.11.2.24 system(str): str is a string representing a system command. This command is executed, its output written to the standard output (this won’t get into your logfile), and control returns to the PARI system. This simply calls the C system command.

3.11.2.25 type(x, {t}): this is useful only under GP. If t is not present, returns the internal type number of the PARI object x. Otherwise, makes a copy of x and sets its type equal to type t, which can be either a number or, preferably since internal codes may eventually change, a symbolic name such as t FRACN (you can skip the t part here, so that FRACN by itself would also be all right). Check out existing type names with the metacommand \t.

Type changes must be used with extreme caution, or disasters may occur (SIGSEGV or SIGBUS being one’s best bet), but one instance where it can be useful is type(x,RFRACN) when x is a rational function (type t RFRAC). In this case, the created object, as well as the objects created from it, will not be reduced automatically, making the operations much faster. In fact this function is the only way to create reducible rationals (type t FRACN) or rational functions (type t RFRACN) in GP.

There is no equivalent library syntax, since the internal functions typ and settyp are available. Note that settyp does not create a copy of x, contrary to most PARI functions. It just changes the type in place (and returns nothing). typ returns a C long integer. Note also the different spellings of the internal functions (set)typ and of the GP function type*.

3.11.2.26 whatnow(key): if keyword key is the name of a function that was present in GP version 1.39.15 or lower, outputs the new function name and syntax, if it changed at all (387 out of 560 did).

3.11.2.27 write(filename, {str∗}): writes (appends) to filename the remaining arguments, and appends a newline (same output as print).

3.11.2.28 write1(filename, {str∗}): writes (appends) to filename the remaining arguments without a trailing newline (same output as print1).

3.11.2.29 writetex(filename, {str∗}): as write, in TEX format.

* This is due to the fact that type is a reserved identifier for some C compilers.

144

Programming PARI in Library Mode

4.1 Introduction: initializations, universal objects.

To be able to use PARI in library mode, you must write a C program and link it to the PARI library. See the installation guide (in Appendix A) on how to create and install the library and include files. A sample Makefile is presented in Appendix B.

Probably the best way to understand how programming is done is to work through a complete example. We will write such a program in Section 4.8. Before doing this, a few explanations are in order.

First, one must explain to the outside world what kind of objects and routines we are going to use. This is done simply with the statement

    #include <pari.h>

This file pari.h imports all the necessary constants, variables and functions, defines some important macros, and also defines the fundamental type for all PARI objects: the type GEN, which is simply a pointer to long.

Technical note: we would have liked to define a type GEN to be a pointer to itself. This unfortunately is not possible in C, except by using structures, but then the names become unwieldy. The result of this is that when we use a component of a PARI object, it will be a long, hence will need to be typecast to a GEN again if we want to avoid warnings from the compiler. This will sometimes be quite tedious, but of course is trivially done. See the discussion on typecasts in the next section.

After declaring the use of the file pari.h, the first executable statement of a main program should be to initialize the PARI system, and in particular the PARI stack which will be both a scratchboard and a repository for computed objects. This is done with a call to the function

    void pari init(long size, long maxprime).

The first argument is the number of bytes given to PARI to work with (it should not reasonably be taken below 500000), and the second is the upper limit on a precomputed prime number table. If you don’t want prime numbers, just put maxprime = 2. Be careful because lots a PARI functions need this table (certainly all the ones of interest to number theorists). If you wind up with the error message “not enough precomputed primes”, try to increase this value.

We have now at our disposal:

• a large PARI stack containing nothing. It’s a big connected chunk of memory whose size you chose when invoking pari init. All your computations are going to take place here. When doing large computations, unwanted intermediate results clutter up memory very fast so some kind of garbage collecting is needed. Most large systems do garbage collecting when the memory is getting scarce, and this slows down the performance. In PARI we have taken a different approach: you must do your own cleaning up when the intermediate results are not needed anymore. Special purpose routines have been written to do this; we will see later how (and when, if at all) you should use them.

• the following universal objects (by definition, objects which do not belong on the stack): the integers 0, 1 and 2 (respectively called gzero, gun, and gdeux), the fraction \frac{1}{2} (ghalf), the complex number i (gi). All of these are of type GEN.

In addition, space is reserved for the polynomials x_{v} (polx[v]), and the polynomials 1_{v} (polun[v]). Here, x_{v} is the name of variable number v, where 0 ≤ v ≤ MAXVARN (the exact value of which depends on your machine, at least 16383 in any case). Both polun and polx are arrays of GENs, the index being the polynomial variable number.

However, except for the ones corresponding to variables 0 and MAXVARN, these polynomials are not created upon initialization. It is the programmer’s responsibility to fill them before use. We’ll see how this is done in Section 4.6 (never through direct assignment).

• a heap which is just a linked list of permanent universal objects. For now, it contains exactly the ones listed above. You will probably very rarely use the heap yourself; and if so, only as a collection of individual copies of objects taken from the stack (called clones in the sequel). Thus you need not bother with its internal structure, which may change as PARI evolves. Some complex PARI functions may create clones for special garbage collecting purposes, usually destroying them when returning.

• a table of primes (in fact of differences between consecutive primes), called diffptr, of type byteptr (pointer to unsigned char). Its use is described in appendix C.

• access to all the built-in functions of the PARI library. These are declared to the outside world when you include pari.h, but need the above things to function properly. So if you forget the call to pari init, you will immediately get a fatal error when running your program (usually a segmentation fault).

4.2 Important technical notes.

4.2.1 Typecasts.

We have seen that, due to the non-recursiveness of the PARI types from the compiler’s point of view, many typecasts will be necessary when programming in PARI. To take an example, a vector V of dimension 2 (two components) will be represented by a chunk of memory pointed to by the GEN V. V[0] contains coded information, in particular about the type of the object, its length, etc. V[1] and V[2] contain pointers to the two components of V. Those coefficients V[i] themselves are in chunks of memory whose complexity depends on their own types, and so on. This is where typecasting will be necessary: a priori, V[i] (for i = 1, 2) is a long, but we will want to use it as a GEN. The following two constructions will be exceedingly frequent (x and V are GENs):

    V[i] = (long) x;
    x = (GEN) V[i];

Note that a typecast is not a valid lvalue (cannot be put on the left side of an assignment), so (GEN)V[i] = x would be incorrect, though some compilers may accept it.

Due to this annoyance, the PARI functions and variables that occur most frequently have analogues which are macros including the typecast. The complete list can be found in the file paricast.h (which is included by pari.h and can be found at the same place). For instance you can abbreviate:

(long) gzero     ----->  zero
(long) gun       ----->  un
(long) polx[v]   ----->  lpolx[v]
(long) gadd(x,y) ----->  ladd(x,y)

In general, replacing a leading g by an l in the name of a PARI function will typecast the result to long. Note that ldiv is an ANSI C function which is is hidden in PARI by a macro of the same name representing (long)gdiv.

The macro coeff(x, m, n) exists with exactly the meaning of x[m,n] under GP when x is a matrix. This is a purely syntactical trick to reduce the number of typecasts and thus does not create a copy of the coefficient (contrary to all the library functions). It can be put on the left side of an assignment statement, and its value, of type long integer, is a pointer to the desired coefficient object. The macro gcoeff is a synonym for (GEN) coeff, hence cannot be put on the left side of an assignment.

To retrieve the values of elements of lists of … of lists of vectors, without getting infuriated by gigantic lists of typecasts, we have the mael macros (for multidimensional array element). The syntax is maeln(x, a_{1}, …, a_{n}), where x is a GEN, the a_{i} are indexes, and n is an integer between 2 and 5 (with a standalone mael as a synonym for mael2). This stands for x[a_{1}][a_{2}] … [a_{n}] (with all the necessary typecasts), and returns a long (i.e. they are valid lvalues). The gmaeln macros are synonyms for (GEN)maeln. Note that due to the implementation of matrix types in PARI (i.e. as horizontal lists of vertical vectors), coeff(x,y) is actually completely equivalent to mael(y,x). It is suggested that you use coeff in matrix context, and mael otherwise.

4.2.2 Variations on basic functions. In the library syntax descriptions in Chapter 3, we have only given the basic names of the functions. For example gadd(x, y) assumes that x and y are PARI objects (of type GEN), and creates the result x + y on the PARI stack. For most of the basic operators and functions, many other variants are available. We give some examples for gadd, but the same is true for all the basic operators, as well as for some simple common functions (the complete list is given in Chapter 5):

    GEN gaddgs(GEN x, long y)
    GEN gaddsg(long x, GEN y)

In the following three, z is a preexisting GEN and the result of the corresponding operation is put into z. The size of the PARI stack does not change:

    void gaddz(GEN x, GEN y, GEN z)
    void gaddgsz(GEN x, long y, GEN z)
    void gaddsgz(GEN x, GEN y, GEN z)

There are also low level functions which are special cases of the above:

GEN addii(GEN x, GEN y): here x and y are GENs of type t INT (this is not checked).

GEN addrr(GEN x, GEN y): here x and y are GEN reals (type t REAL).

There also exist functions addir, addri, mpadd (whose two arguments can be of type integer or real), addis (to add a t INT and a long) and so on.

All these functions can of course be called by the user but we feel that the few microseconds lost in calling more general functions (in this case gadd) are compensated by the fact that one needs to remember a much smaller number of functions, and also because there is a hidden danger here: the types of the objects that you use, if they are themselves results of a previous computation, are not completely predetermined. For instance the multiplication of a type real t REAL by a type integer t INT usually gives a result of type real, except when the integer is 0, in which case according to the PARI philosophy the result is the exact integer 0. Hence if afterwards you call a function which specifically needs a real type argument, you are going to be in trouble.

If you really want to use these functions, their names are self-explanatory once you know that i stands for a PARI integer, r for a PARI real, mp for i or r, s for an ordinary signed long, whereas z (as a suffix) means that the result is not created on the PARI stack but assigned to a preexisting GEN object passed as an extra argument.

For completeness, Chapter 5 gives a description of all these low-level functions.

Please note that in the present version 2.0.17 the names of the functions are not always consistent. This will be changed. Hence anyone programming in PARI must be aware that the names of almost all functions that he uses might be subject to change. If the need arises (i.e. if there really are people out there who delve into the innards of PARI), updated versions with no name changes will be released.

4.2.3 Portability: 32-bit / 64-bit architectures.

PARI supports both 32-bit and 64-bit based machines, but not simultaneously! The library will have been compiled assuming a given architecture (a priori following a guess by the Configure program, see Appendix A), and some of the header files you include (through pari.h) will have been modified to match the library.

Portable macros are defined to bypass most machine dependencies. If you want your programs to run identically on 32-bit and 64-bit machines, you will have to use these, and not the corresponding numeric values, whenever the precise size of your long integers might matter. Here are the most important ones:

64-bit  32-bit

BITS IN LONG       64        32

LONG IS 64BIT  defined  undefined
DEFAULTPREC       3         4  (≈ 19 decimal digits, see formula below)
MEDDEFAULTPREC    4         6  (≈ 38 decimal digits)
BIGDEFAULTPREC    5         8  (≈ 57 decimal digits)

For instance, suppose you call a transcendental function, such as

    GEN gexp(GEN x, long prec).

The last argument prec is only used if x is an exact object (otherwise the relative precision is determined by the precision of x). But since prec sets the size of the inexact result counted in (long) words (including codewords), the same value of prec will yield different results on 32-bit and 64-bit machines. Real numbers have two codewords (see Section 4.5), so the formula for computing the bit accuracy is

                    bit accuracy(prec) = (prec − 2) ∗ BITS IN LONG

(this is actually the definition of a macro). The corresponding accuracy expressed in decimal digits would be

bit accuracy(prec) ∗ log(2)/ log(10).

For example if the value of prec is 5, the corresponding accuracy for 32-bit machines is (5 − 2) ∗ log(2^{32})/ log(10) ≈ 28 decimal digits, while for 64-bit machines it is (5 − 2) ∗ log(2^{64})/ log(10) ≈ 57 decimal digits.

Thus, you must take care to change the prec parameter you are supplying according to the bit size, either using the default precisions given by the various DEFAULTPRECs, or by using conditional constructs of the form:

    #ifndef LONG IS 64BIT
      prec = 4;
    #else
      prec = 6;
    #endif

which is in this case equivalent to the statement prec = MEDDEFAULTPREC;.

Note that for parity reasons, half the accuracies available on 32-bit architectures (the odd ones) have no precise equivalents on 64-bit machines.

4.3 Creation of PARI objects, assignments, conversions.

4.3.1 Creation of PARI objects. The basic function which creates a PARI object is the function cgetg whose prototype is:

    GEN cgetg(long length, long type).

Here length specifies the number of longwords to be allocated to the object, and type is the type number of the object, preferably in symbolic form (see Section 4.5 for the list of these). The precise effect of this function is as follows: it first creates on the PARI stack a chunk of memory of size length longwords, and saves the address of the chunk which it will in the end return. If the stack has been used up, a message to the effect that “the PARI stack overflows” will be printed, and an error raised. Otherwise, it sets the type and length of the PARI object. In effect, it fills correctly and completely its first codeword (z[0] or *z). Many PARI objects also have a second codeword (types t INT, t REAL, t PADIC, t POL, and t SER). In case you want to produce one of those from scratch (this should be exceedingly rare), it is your responsibility to fill this second codeword, either explicitly (using the macros described in Section 4.5), or implicitly using an assignment statement (using gaffect).

Note that the argument length is predetermined for a number of types: 3 for types t INTMOD,

t FRAC, t FRACN, t COMPLEX, t POLMOD, t RFRAC and t RFRACN, 4 for type t QUAD and t QFI, and

5 for type t PADIC and t QFR. However for the sake of efficiency, no checking is done in the function cgetg, so disasters will occur if you give an incorrect length.

Notes: 1) The main use of this function is to prepare for later assigments (see Section 4.3.2). Most of the time you will use GEN objects as they are created and returned by PARI functions. In this case you don’t need to use cgetg to create space to hold them.

2) For the creation of leaves, i.e. integers or reals, which is very common,

    GEN cgeti(long length)
    GEN cgetr(long length)
should be used instead of cgetg(length, t INT) and cgetg(length, t REAL) respectively.

3) The macros lgetg, lgeti, lgetr are predefined as (long)cgetg, (long)cgeti, (long)cgetr, respectively.

Examples: 1) z = cgeti(DEFAULTPREC) and cgetg(DEFAULTPREC, t INT) create an integer ob-

ject whose “precision” is bit accuracy(DEFAULTPREC) = 64. This means z can hold rational integers of absolute value less than 2^{64}. Note that in both cases, the second codeword will not be filled. Of course we could use numerical values, e.g. cgeti(4), but this would have different meanings on different machines as bit accuracy(4) equals 64 on 32-bit machines, but 128 on 64-bit machines.

2) The following creates a type “complex” object whose real and imaginary parts can hold real numbers of precision bit accuracy(MEDDEFAULTPREC) = 96 bits:

    z = cgetg(3, t COMPLEX);
    z[1] = lgetr(MEDDEFAULTPREC);
    z[2] = lgetr(MEDDEFAULTPREC);

3) To create a matrix object for 4 × 3 matrices:

    z = cgetg(4, t MAT);
    for(i=1; i<4; i++) z[i] = lgetg(5, t COL);

If one wishes to create space for the matrix elements themselves, one has to follow this with a double loop to fill each column vector.

These last two examples illustrate the fact that since PARI types are recursive, all the branches of the tree must be created. The function cgetg creates only the “root”, and other calls to cgetg must be made to produce the whole tree. For matrices, a common mistake is to think that z = cgetg(4, t MAT) (for example) will create the root of the matrix: one needs also to create the column vectors of the matrix (obviously, since we specified only one dimension in the first cgetg!). This is because a matrix is really just a row vector of column vectors (hence a priori not a basic type), but it has been given a special type number so that operations with matrices become possible.

4.3.2 Assignments. Firstly, if x and y are both declared as GEN (i.e. pointers to something), the ordinary C assignment y = x makes perfect sense: we are just moving a pointer around. However, physically modifying either x or y (for instance, x[1] = 0) will also change the other one, which is usually not desirable.

Very important note: Using the functions described in this paragraph is very inefficient and often awkward: one of the gerepile functions (see Section 4.4) should be preferred. See the paragraph end for some exceptions to this rule.

The general PARI assignment function is the function gaffect with the following syntax:

    void gaffect(GEN x, GEN y)

Its effect is to assign the PARI object x into the preexisting object y. This copies the whole structure of x into y so many conditions must be met for the assignment to be possible. For instance it is allowed to assign an integer into a real number, but the converse is forbidden. For that, you must use the truncation or rounding function of your choice (see section 3.2). It can also happen that y is not large enough or does not have the proper tree structure to receive the object x. As an extreme example, assume y is the zero integer with length equal to 2. Then all assignments of a non-zero integer into y will result in an error message since y is not large enough to accommodate a non-zero integer. In general common sense will tell you what is possible, keeping in mind the PARI philosophy which says that if it makes sense it is legal. For instance, the assignment of an imprecise object into a precise one does not make sense. However, a change in precision of imprecise objects is allowed.

All functions ending in “z” such as gaddz (see Section 4.2.2) implicitly use this function. In fact what they exactly do is record avma (see Section 4.4), perform the required operation, gaffect the result to the last operand, then restore the initial avma.

You can assign ordinary C long integers into a PARI object (not necessarily of type t INT). Use the function gaffsg with the following syntax:

    void gaffsg(long s, GEN y)

Note: due to the requirements mentionned above, it’s usually a bad idea to use gaffect statements. Two exceptions:

• for simple objects (e.g. leaves) whose size is controlled, they can be easier to use than gerepile, and about as efficient.

• to coerce an inexact object to a given precision. For instance

    gaffect(x, (tmp=cgetr(3))); x = tmp;

at the beginning of a routine where precision can be kept to a minimum (otherwise the precision of x will be used in all subsequent computations, which will be a disaster if x is known to thousands of digits).

4.3.3 Copy. It is also very useful to copy a PARI object, not just by moving around a pointer as in the y = x example, but by creating a copy of the whole tree structure, without pre-allocating a possibly complicated y to use with gaffect. The function which does this is called gcopy, with the predefined macro lcopy as a synonym for (long)gcopy. Its syntax is:

    GEN gcopy(GEN x)

and the effect is to create a new copy of x on the PARI stack. Beware that universal objects which occur in specific components of certain types (mainly moduli for types t INTMOD and t PADIC) are not copied, as they are assumed to be permanent. In this case, gcopy only copies the pointer. Use GEN forcecopy(GEN x) if you want a complete copy.

Please be sure at this point that you really understand the difference between y = x, y = gcopy(x), and gaffect(x,y): this will save you from many “obvious” mistakes later on.

4.3.4 Clones. Sometimes, it may be more efficient to create a permanent copy of a PARI object. This will not be created on the stack but on the heap. The function which does this is called gclone, with the predefined macro lclone as a synonym for (long)gclone. Its syntax is:

    GEN gclone(GEN x)

A clone can be removed from the heap (thus destroyed) using

    void gunclone(GEN x)

No PARI object should keep references to a clone which has been destroyed. If you want to copy a clone back to the stack then delete it, use forcecopy and not gcopy, otherwise some components might not be copied (moduli of t INTMODs and t POLMODs for instance).

4.3.5 Conversions. The following functions convert C objects to PARI objects (creating them on the stack as usual):

GEN stoi(long s): C long integer (“small”) to PARI integer (t INT)

GEN dbltor(double s): C double to PARI real (t REAL). The accuracy of the result is 19 decimal digits, i.e. a type t REAL of length DEFAULTPREC, although on 32-bit machines only 16 of them will be significant.

We also have the converse functions:

long itos(GEN x): x must be of type t INT,

double rtodbl(GEN x): x must be of type t REAL,

as well as the more general ones:

    long gtolong(GEN x),
    double gtodouble(GEN x).

4.4 Garbage collection.

4.4.1 Why and how.

As we have seen, the pari init routine allocates a big range of addresses (the stack) that are going to be used throughout. Recall that all PARI objects are pointers. But for universal objects, they will all point at some part of the stack.

The stack starts at the address bot and ends just before top. This means that the quantity

                             (top − bot) / sizeof(long)

is equal to the size argument of pari init. The PARI stack also has a “current stack pointer” called avma, which stands for available memory address. These three variables are global (declared for you by pari.h). For historical reasons they are of type long and not of type GEN as would seem more natural.

The stack is oriented upside-down: the more recent an object, the closer to bot. Accordingly, initially avma = top, and avma gets decremented as new objects are created. As its name indicates, avma always points just after the first free address on the stack, and (GEN)avma is always (a pointer to) the latest created object. When avma reaches bot, the stack overflows, aborting all computations, and an error message is issued. To avoid this you will need to clean up the stack from time to time, when some bunch of intermediate objects will not be needed anymore. This is called “garbage collecting.”

We are now going to describe briefly how this is done. We will see many concrete examples in the next subsection.

• First, PARI routines will do their own garbage collecting, which means that whenever a documented function from the library returns, only its result(s) will have been added to the stack (non-documented ones may not do this, for greater speed). In particular, a PARI function that does not return a GEN does not clutter the stack. Thus, if your computation is small enough (i.e. you call few PARI routines, or most of them return long integers), then you don’t need to do any garbage collecting. This will probably be the case in many of your subroutines. Of course the objects that were on the stack before the function call are left alone. Except for the ones listed below, PARI functions only collect their own garbage.

• It may happen that you need to preserve some but not all objects that were created after a certain point — for instance, if the final result you need is not a GEN, or if some search proved futile. Then, it is enough to record the value of avma just before the first garbage is created, and restore it upon exit:

    long ltop = avma; /* record initial avma */
    garbage ...
    avma = ltop; /* restore it */

All objects created in the garbage zone will eventually be overwritten: they should not be accessed anymore once avma has been restored.

• If you want to destroy (i.e. give back the memory occupied by) the latest PARI object on the stack (e.g. the latest one obtained from a function call), and the above method is not available (because the initial value of avma has been lost), just use the function

    void cgiv(GEN z)

where z is the object you want to give back.

• Unfortunately life is not so simple, and sometimes you will want to give back accumulated garbage during a computation without losing recent data. For this you need the gerepile function (or one of its variants described hereafter):

    GEN gerepile(long ltop, long lbot, GEN q)

This function cleans up the stack between ltop and lbot, where lbot < ltop, and returns the updated object q. This means:

1) we translate (copy) all the objects in the interval [avma, lbot[, so that its right extremity abuts the address ltop. Graphically

           bot          avma   lbot          ltop     top

end of stack |------------[++++++[-/-/-/-/-/-/-|++++++++| beginning of stack

free memory

becomes:

           bot                        avma   ltop     top

end of stack |--------------------------[++++++[++++++++| beginning of stack

free memory where ++ denote significant objects, -- the unused part of the stack, and -/- the garbage we remove.

2) The function then inspects all the PARI objects between avma and lbot (i.e. the ones that we want to keep and that have been translated) and looks at every component of such an object which is not a codeword. Each such component is a pointer to an object whose address is either

— between avma and lbot, in which case it will be suitably updated,

— larger than or equal to ltop, in which case it will not change, or

— between lbot and ltop in which case gerepile will scream an error message at you (“significant pointers lost in gerepile”).

3) avma is updated (we add ltop − lbot to the old value).

4) We return the (possibly updated) object q: if q initially pointed between avma and lbot, we return the translated address, as in 2). If not, the original address is still valid (and we return it!).

As stated above, no component of the remaining objects (in particular q) should belong to the erased segment [lbot, ltop[, and this is checked within gerepile. But beware as well that the addresses of all the objects in the translated zone will have changed after a call to gerepile: every pointer you may have kept around elsewhere, outside the stack objects, which previously pointed into the zone below ltop must be discarded. If you need to recover more than one object, use one of the gerepilemany functions below.

As a consequence of the preceding explanation, we must now state the most important law about programming in PARI:

If a given PARI object is to be relocated by gerepile then, apart from universal objects, the chunks of memory used by its components should be in consecutive memory locations. All GENs created by documented PARI function are guaranteed to satisfy this.

This is because the gerepile function knows only about two connected zones: the garbage that will be erased (between lbot and ltop) and the significant pointers that will be copied and updated. If there is garbage interspersed with your objects, disasters will occur when we try to update them and consider the corresponding “pointers”. So be very wary when you allow objects to become disconnected. Have a look at the examples, it’s not as complicated as it seems.

In practice this is achieved by the following programming idiom:

    ltop=avma; garbage(); lbot=avma; q=anything();
    return gerepile(ltop, lbot, q); /* returns the updated q */

Beware that

    ltop=avma; garbage();
    return gerepile(ltop, avma, anything())

might work, but should be frowned upon. We can’t predict whether avma is going to be evaluated after or before the call to anything(): it depends on the compiler. If we are out of luck, it will be after the call, so the result will belong to the garbage zone and the gerepile statement becomes equivalent to avma = ltop. Thus we would return a pointer to random garbage.

• A simple variant is

    GEN gerepileupto(long ltop, GEN q)

which cleans the stack between ltop and the connected object q and returns q updated. For this to work, q must have been created before all its components, otherwise they would belong to the garbage zone! Documented PARI functions guarantee this. If you stumble upon one that does not, consider it a bug worth reporting.

• To cope with complicated cases where many objects have to be preserved, you can use

    void gerepilemany(long ltop, GEN *gptr[], long n)

which cleans up the most recent part of the stack (between ltop and avma). All the GENs pointed at by the elements of the array gptr (of length n) are updated. A copy is done just before the cleaning to preserve them, so they don’t need to be connected before the call. This is the most robust of the gerepile functions (the less prone to user error), but also the slowest.

• More efficient, but trickier to use is

    void gerepilemanysp(long ltop, long lbot, GEN *gptr[], long n)

which cleans the stack between lbot and ltop and updates the GENs pointed at by the elements of gptr without doing any copying. This is subject to the same restrictions as gerepile, the only difference being that more than one address gets updated.

4.4.2 Examples.

Let x and y be two preexisting PARI objects and suppose that we want to compute x^{2} + y^{2}. This can trivially be done using the following program (we skip the necessary declarations; everything in sight is a GEN):

    p1 = gsqr(x); p2 = gsqr(y); z = gadd(p1,p2);

The GEN z indeed points at the desired quantity. However, consider the stack: it contains as unnecessary garbage p1 and p2. More precisely it contains (in this order) z, p2, p1 (recall that, since the stack grows downward from the top, the most recent object comes first). We need a way to get rid of this garbage (in this case it causes no harm except that it occupies memory space, but in other cases it could disconnect other PARI objects and this is dangerous).

It would not have been possible to get rid of p1, p2 before z is computed, since they are used in the final operation. We cannot record avma before p1 is computed and restore it later, since this would destroy z as well. It is not possible either to use the function cgiv since p1 and p2 are not at the bottom of the stack and we don’t want to give back z.

But using gerepile, we can give back the memory locations corresponding to p1, p2, and move the object z upwards so that no space is lost. Specifically:

    ltop = avma; /* remember the current address of the top of the stack */
    p1 = gsqr(x); p2 = gsqr(y);
    lbot = avma; /* keep the address of the bottom of the garbage pile */
    z = gadd(p1, p2); /* z is now the last object on the stack */
    z = gerepile(ltop, lbot, z); /* garbage collecting */

Of course, the last two instructions could also have been written more simply:

    z = gerepile(ltop, lbot, gadd(p1,p2));

In fact gerepileupto is even simpler to use, because the result of gadd will be the last object on the stack and gadd is guaranteed to return an object suitable for gerepileupto:

    ltop = avma;
    z = gerepileupto(ltop, gadd(gsqr(x), gsqr(y)));

As you can see, in simple conditions the use of gerepile is not really difficult. However make sure you understand exactly what has happened before you go on (use the figure from the preceding section).

Important remark: as we will see presently it is often necessary to do several gerepiles during a computation. However, the fewer the better. The only condition for gerepile to work is that the garbage be connected. If the computation can be arranged so that there is a minimal number of connected pieces of garbage, then it should be done that way.

For example suppose we want to write a function of two GEN variables x and y which creates the vector [x^{2} + y, y^{2} + x]. Without garbage collecting, one would write:

    p1 = gsqr(x); p2 = gadd(p1, y);
    p3 = gsqr(y); p4 = gadd(p3, x);
    z = cgetg(3,t VEC); z[1] = (long)p2; z[2] = (long)p4;

This leaves a dirty stack containing (in this order) z, p4, p3, p2, p1. The garbage here consists of p1 and p3, which are separated by p2. But if we compute p3 before p2 then the garbage becomes connected, and we get the following program with garbage collecting:

    ltop = avma; p1 = gsqr(x); p3 = gsqr(y); lbot = avma;
    z = cgetg(3,t VEC); z[1] = ladd(p1,y); z[2] = ladd(p3,x);
    z = gerepile(ltop,lbot,z);

Finishing by z = gerepileupto(ltop, z) would be ok as well. But when you have the choice, it’s usually clearer to brace the garbage between ltop / lbot pairs.

Beware that

    ltop = avma; p1 = gadd(gsqr(x), y); p3 = gadd(gsqr(y), x);
    z = cgetg(3,t VEC); z[1] = (long)p1; z[2] = (long)p3
    z = gerepileupto(ltop,z); /* WRONG !!! */

would be a disaster since p1 and p3 would be created before z, so the call to gerepileupto would overwrite them, leaving z[1] and z[2] pointing at random data!

We next want to write a program to compute the product of two complex numbers x and y, using a method which takes only 3 multiplications instead of 4. Let z = x∗y, and set x = x_{r} +i∗x_{i} and similarly for y and z. The well-known trick is to compute p_{1} = x_{r} ∗ y_{r}, p_{2} = x_{i} ∗ y_{i}, p_{3} = (x_{r} + x_{i}) ∗ (y_{r} + y_{i}), and then we have z_{r} = p_{1} − p_{2}, z_{i} = p_{3} − (p_{1} + p_{2}). The program is essentially as follows:

    ltop = avma; p1 = gmul(x[1],y[1]); p2 = gmul(x[2],y[2]);
    p3 = gmul(gadd(x[1],x[2]), gadd(y[1],y[2]));
    p4 = gadd(p1,p2); lbot = avma;
    z = cgetg(3,t COMPLEX); z[1] = lsub(p1,p2); z[2] = lsub(p3,p4);
    z = gerepile(ltop,lbot,z);

“Essentially,” because for instance x[1] is a long and not a GEN, so we need to insert many annoying

typecasts: p1 = gmul((GEN)x[1], (GEN)y[1]) and so on.

Let us now look at a less trivial example where more than one gerepile is needed in practice (at the expense of efficiency, one can always use only one using gcopy; see below). Suppose that we want to write a function which multiplies a line vector by a matrix (such a function is of course already part of gmul, but let’s ignore this for a moment). Then the most natural way is to do a cgetg of the result immediately, and then a gerepile for each coefficient of the result vector to get rid of the garbage which has accumulated while this particular coefficient was computed. We leave the details to the reader, who can look at the answer in the file basemath/gen1.c, in the function gmul, case t VEC times case t MAT. It would theoretically be possible to have a single connected piece of garbage, but it would be a much less natural and unnecessarily complicated program.

Let us now take an example which is probably the least trivial way of using gerepile, but is unfortunately sometimes necessary. Although it is not an infrequent occurrence, we will not give a specific example but a general one: suppose that we want to do a computation (usually inside a larger function) producing more than one PARI object as a result, say two for instance. Then even if we set up the work properly, before cleaning up we will have a stack which has the desired results z1, z2 (say), and then connected garbage from lbot to ltop. If we write

    z1 = gerepile(ltop,lbot,z1);

then the stack will be cleaned, the pointers fixed up, but we will have lost the address of z2. This is where we need one of the gerepilemany functions: we declare

    GEN *gptr[2]; /* Array of pointers to GENs */
    gptr[0] = &z1; gptr[1] = &z2;

and now the call gerepilemany(ltop, gptr, 2) copies z1 and z2 to new locations, cleans the stack from ltop to the old avma, and updates the pointers z1 and z2. Here we don’t assume anything about the stack: the garbage can be disconnected and z1, z2 need not be at the bottom of the stack. If all of these assumptions are in fact satisfied, then we can call gerepilemanysp instead, which will be faster since we don’t need the initial copy.

Another important usage is “random” garbage collection during loops whose size requirements we cannot control in advance:

      long ltop = avma, limit = (avma+bot)/2;
      GEN x, y;
      while (...)

{

        garbage(); x = anything();
        garbage(); y = anything()
        garbage();
        if (avma < limit) /* memory is running low (half spent since entry) */

{

          GEN *gptr[2];
          gptr[0] = &x; gptr[1] = &y;
          gerepilemany(ltop, gptr, 2);

}

}

Here we assume that only x and y are needed from one iteration to the next. As it would be too costly to call gerepile once for each iteration, we only do it when it seems to have become necessary. Of course, when the need arises, you can use bigger gptr arrays: in the PARI library source code, we once needed to preserve up to 10 objects at a time (in a variant of the LLL algorithm)!

Technical note: the statement limit = (avma+bot)/2 is dangerous since the addition can overflow, which would result in limit being negative. This will prevent garbage collection in the loop. To avoid this problem, we provide a robust macro stack lim(avma,n), which denotes an address where 2^{n−1}/(2^{n−1} + 1) of the total stack space is exhausted (1/2 for n = 1, 2/3 for n = 2). Hence, the above snippet should be written as

      long ltop = avma, limit = stack lim(avma,1);

…

4.4.3 Some hints and tricks. Even for the authors, the use of gerepile was not evident at first. Hence we give some indications on how to avoid most problems connected with garbage collecting:

First, although it looks complicated, gerepile has turned out to be a very flexible and fast garbage collector, which compares very favorably with much more sophisticated methods used in other systems. A few tests that we have done indicate that the price paid for using gerepile, when properly used, is usually around 1 or 2 percents of the total time, which is quite acceptable.

Secondly, in many cases, in particular when the tree structure and the size of the PARI objects which will appear in a computation are under control, one can avoid gerepile altogether by creating sufficiently large objects at the beginning (using cgetg), and then using assignment statements and operations ending with z (such as gaddz). Coming back to our first example, note that if we know that x and y are of type real and of length less than or equal to 5, we can program without using

gerepile at all:
    z = cgetr(5); ltop = avma;
    p1 = gsqr(x); p2 = gsqr(y); gaddz(p1,p2,z);
    avma = ltop;

This practice will usually be slower than a craftily used gerepile though, and is certainly more cumbersome to use. As a rule, assignment statements should generally be avoided.

Thirdly, the philosophy of gerepile is the following: keep the value of the stack pointer avma at the beginning, and just before the last operation. Afterwards, it would be too late since the lower end address of the garbage zone would have been lost. Of course you can always use gerepileupto, but you will have to assume that the object was created before its components.

Finally, if everything seems hopeless, at the expense of speed you can do the following: after saving the value of avma in ltop, perform your computation as you wish, in any order, leaving a messy stack. Let z be your result. Then write the following:

    z = gerepileupto(ltop, gcopy(z));

The trick is to force a copy of z to be created at the bottom of the stack, hence all the rest including the initial z becomes connected garbage. If you need to keep more than one result, use gerepilemany (of which the above is just a special case).

If you followed us this far, congratulations, and rejoice: the rest is much easier.

4.5 Implementation of the PARI types.

Although it is a little tedious, we now go through each type and explain its implementation. Let z be a GEN, pointing at a PARI object. In the following paragraphs, we will constantly mix two points of view: on the one hand, z will be treated as the C pointer it is (in the context of program fragments like z[1]), on the other, as PARI’s handle on (the internal representation of) some mathematical entity, so we will shamelessly write z ≠ 0 to indicate that the value thus represented is nonzero (in which case the pointer z certainly will be non-NULL). We offer no apologies for this style. In fact, you had better feel comfortable juggling both views simultaneously in your mind if you want to write correct PARI programs.

Common to all the types is the first codeword z[0], which we don’t have to worry about since this is taken care of by cgetg. Its precise structure will depend on the machine you are using, but it always contain the following data: the internal type number associated to the symbolic type name, the length of the root in longwords, and a technical bit which indicates whether the object is a clone (see below) or not. This last one is used by GP for internal garbage collecting, you won’t have to worry about it.

These data can be handled through the following macros:

long typ(GEN z) returns the type number of z.

void settyp(GEN z, long n) sets the type number of z to n (you should not have to use this function if you use cgetg).

long lg(GEN z) returns the length (in longwords) of the root of z.

long setlg(GEN z, long l) sets the length of z to l (you should not have to use this function if you use cgetg; however, see an advanced example in Section 4.8).

(If you know enough about PARI to need to access the “clone” bit, then you’ll be able to find usage hints in the code (esp. killbloc() and matrix block()). It is technical after all.)

These macros are written in such a way that you don’t need to worry about type casts when using them: i.e. if z is a GEN, typ(z[2]) will be accepted by your compiler, without your having to explicitly type typ((GEN)z[2]). Note that for the sake of efficiency, none of the codeword-handling macros check the types of their arguments even when there are stringent restrictions on their use.

The possible second codeword is used differently by the different types, and we will describe it as we now consider each of them in turn:

4.5.1 Type t INT (integers): this type has a second codeword z[1] which contains the following information:

the sign of z: coded as 1, 0 or −1 if z > 0, z = 0, z < 0 respectively.

the effective length of z, i.e. the total number of significant longwords. This means the following: apart from the integer 0, every integer is “normalized”, meaning that the first mantissa longword (i.e. z[2]) is non-zero. However, the integer may have been created with a longer length. Hence the “length” which is in z[0] can be larger than the “effective length” which is in z[1]. Accessing z[i] for i larger than or equal to the effective length will yield random results.

This information is handled using the following macros:

long signe(GEN z) returns the sign of z. void setsigne(GEN z, long s) sets the sign of z to s.

long lgefint(GEN z) returns the effective length of z.

void setlgefint(GEN z, long l) sets the effective length of z to l.

The integer 0 can be recognized either by its sign being 0, or by its effective length being equal to 2. When z ≠ 0, the word z[2] exists and is non-zero, and the absolute value of z is (z[2],z[3],…,z[lgefint(z)-1]) in base 2^BITS IN LONG, where as usual in this notation z[2] is the highest order longword.

The following further macros are available:

long mpodd(GEN x) which is 1 if x is odd, and 0 otherwise.

long mod2(GEN x), mod4(x), and so on up to mod64(x), which give the residue class of x modulo the corresponding power of 2, for positive x (you will obtain weird results if you use these on the integer 0 or on negative numbers).

These macros directly access the binary data and are thus much faster than the generic modulo functions. Besides, they return long integers instead of GENs, so they don’t clutter up the stack.

4.5.2 Type t REAL (real numbers): this type has a second codeword z[1] which also encodes its sign (obtained or set using the same functions as for the integers), and a biased binary exponent (i.e. the actual exponent value plus some constant bias, actually a power of 2, whose value is given by HIGHEXPOBIT). This exponent can be handled using the following macros:

long expo(GEN z) returns the true (unbiased) exponent of z. This is defined even when z is equal to zero, see Section 1.2.6.3.

void setexpo(GEN z, long e) sets the exponent of z to e, of course after adding the bias.

Note the functions:

long gexpo(GEN z) which tries to return an exponent for z, even if z is not a real number.

long gsigne(GEN z) which returns a sign for z, even when z is neither real nor integer (a rational number for instance).

The real zero is characterized by having its sign equal to 0. However, usually the first mantissa word z[2] is defined and equal to 0. This fact must never be used to recognize a real 0. If z is not equal to 0, the first mantissa word z[2] is normalized, i.e. its most significant bit is 1. The mantissa is (z[2],z[3],…,z[lg(z]-1]) in base 2^BITS IN LONG. Here, z[2] is the most significant longword, and the mantissa takes values between 1 (included) and 2 (excluded). Thus, assume that sizeof(long) is 32 and the length is 4, the real number 3.5 is represented as z[0] (encoding

type = t REAL, lg = 4), z[1] (encoding sign = 1, expo = 1), z[2] = 0xe0000000, z[3] = 0x0.

4.5.3 Type t INTMOD (integermods): z[1] points to the modulus, and z[2] at the number representing the class z. Both are separate GEN objects, and both must be of type integer. In principle z[1] > 0 and 0 ≤ z[2] < z[1], but this rule does not have to be strictly obeyed by the user. Any integermod obtained as the result of a PARI function call will satisfy these conditions.

It is good practice to keep the modulus object on the heap, so that new integermods resulting from operations can point at this common object, instead of carrying along their own copies of it on the stack. The library functions implement this practice almost by default.

4.5.4 Types t FRAC and t FRACN (rational numbers): z[1] points to the numerator, and z[2] to the denominator. Both must be of type integer. In principle z[2] > 0, but this rule does not have to be strictly obeyed. Note that a type t FRACN rational number can be converted to irreducible form using the function GEN gred(GEN x).

4.5.5 Type t COMPLEX (complex numbers): z[1] points to the real part, and z[2] to the imaginary part. A priori z[1] and z[2] can be of any type, but only certain types are useful and make sense.

4.5.6 Type t PADIC (p-adic numbers): this type has a second codeword [1] which contains the following information: the p-adic precision (the exponent of p modulo which the p-adic unit corresponding to z is defined if z is not 0), i.e. one less than the number of significant p-adic digits, and the biased exponent of z (the bias being equal to HIGHVALPBIT here). This information can be handled using the following functions:

long precp(GEN z) returns the p-adic precision of z.

void setprecp(GEN z, long l) sets the p-adic precision of z to l.

long valp(GEN z) returns the p-adic valuation of z (i.e. the unbiased exponent). This is defined even if z is equal to 0, see Section 1.2.6.3.

void setvalp(GEN z, long e) sets the p-adic valuation of z to e.

In addition to this codeword, z[2] points to the prime p, z[3] points to p^{precp(z)}, and z[4] points to an integer representing the p-adic unit associated to z modulo z[3] (or points to zero if z is zero). To summarize, we have the equality:

z = p^{valp(z)} ∗ (z[4] + O(z[3])) = p^{valp(z)} ∗ (z[4] + O(p^{precp(z)}))

4.5.7 Type t QUAD (quadratic numbers): z[1] points to the polynomial defining the quadratic field, z[2] to the “real part” and z[3] to the “imaginary part”, which are to be taken as the coefficients of z with respect to the “canonical” basis (1, w), see Section 1.2.3. Complex numbers are a particular case of quadratics but deserve a separate type.

4.5.8 Type t POLMOD (polmods): exactly as for integermods, z[1] points to the modulus, and z[2] to a polynomial representing the class of z. Both must be of type polynomial. However, one must obey the rules explained in Chapter 2 concerning the hierarchical structure of the variables of a polymod.

4.5.9 Type t POL (polynomials): this type has a second codeword which is analogous to the one for integers. It contains a “sign”: 0 if the polynomial is equal to 0, and 1 if not (see however the important remark below), a variable number (e.g. 0 for x, 1 for y, etc…), and an effective length.

These data can be handled with the following macros:

signe and setsigne as for reals and integers.

long lgef(GEN z) returns the effective length of z.

void setlgef(GEN z, long l) sets the effective length of z to l. long varn(GEN z) returns the variable number of the object z.

void setvarn(GEN z, long v) sets the variable number of z to v.

Note also the function long gvar(GEN z) which tries to return a variable number for z, even if z is not a polynomial or power series. The variable number of a scalar type is set by definition equal to BIGINT.

The components z[2], z[3],… z[lgef(z)-1] point to the coefficients of the polynomial in ascending order, with z[2] being the constant term and so on. Note that the degree of the polynomial is equal to its effective length minus three. The function

long degree(GEN x) returns the degree of x with respect to its main variable even when x is not a polynomial (a rational function for instance). By convention, the degree of 0 is −1.

Important remark. A zero polynomial can be characterized by the fact that its sign is 0. However, its effective length may be equal to 2, or greater than 2. If it is greater than 2, this means that all the coefficients of the polynomial are equal to zero (as they should for a zero polynomial), but not all of these zeros are exact zeros, and more precisely the leading term z[lgef(z)-1] is not an exact zero.

4.5.10 Type t SER (power series): This type also has a second codeword, which encodes a “sign”, i.e. 0 if the power series is 0, and 1 if not, a variable number as for polynomials, and a biased exponent with a bias of HIGHVALPBIT. This information can be handled with the following functions: signe, setsigne, varn, setvarn as for polynomials, and valp, setvalp for the exponent as for p-adic numbers. Beware: do not use expo and setexpo on power series.

If the power series is non-zero, z[2], z[3],… z[lg(z)-1] point to the coefficients of z in ascending order, z[2] being the first non-zero coefficient. Note that the exponent of a power series can be negative, i.e. we are then dealing with a Laurent series (with a finite number of negative terms).

4.5.11 Types t RFRAC and t RFRACN (rational functions): z[1] points to the numerator, and z[2] on the denominator. Both must be of type polynomial. Note that a type t RFRACN rational function can be converted to irreducible form using the function gred.

4.5.12 Type t QFR (indefinite binary quadratic forms): z[1], z[2], z[3] point to the three coefficients of the form and should be of type integer. z[4] is Shanks’s distance function, and should be of type real.

4.5.13 Type t QFI (definite binary quadratic forms): z[1], z[2], z[3] point to the three coefficients of the form. All three should be of type integer.

4.5.14 Types t VEC and t COL (vectors): z[1], z[2],… z[lg(z)-1] point to the components of the vector.

4.5.15 Type t MAT (matrices): z[1], z[2],… z[lg(z)-1] point to the column vectors of z, i.e. they must be of type t COL and of the same length.

The last two were introduced for specific GP use, and you’ll be much better off using the standard malloc’ed C constructs when programming in library mode. We quote them just for completeness (advising you not to use them):

4.5.16 Type t LIST (lists): This one has a second codeword which contains an effective length (handled through lgef / setlgef). z[2],…, z[lgef(z)-1] contain the components of the list.

4.5.17 Type t STR (character strings): GSTR(z) (= (z+1)) points to the first character of the (NULL-terminated) string.

4.6 PARI variables.

4.6.1 Multivariate objects

We now consider variables and formal computations. As we have seen in Section 4.5, the codewords for types t POL and t SER encode a “variable number”. This is an integer, ranging from 0 to MAXVARN. The lower it is, the higher the variable priority. PARI does not know anything about intelligent “sparse” representation of polynomials. So a multivariate polynomial in PARI is just a polynomial (in one variable), whose coefficients are themselves (arbitrary) polynomials. All computations are then just done formally on the coefficients as if the polynomial was univariate.

In fact, the way an object will be considered in formal computations depends entirely on its “principal variable number” which is given by the function

    long gvar(GEN z)

which returns a variable number for z, even if z is not a polynomial or power series. The variable number of a scalar type is set by definition equal to BIGINT which is bigger than any legal variable number. The variable number of a recursive type which is not a polynomial or power series is the minimal variable number of its components. But for polynomials and power series only the “outermost” number counts: the representation is not symmetrical at all.

Under GP, one need not worry too much since the interpreter will define the variables as it sees them and do the right thing with the polynomials produced (however, have a look at the remark in Section 2.3.8). But in library mode, they are tricky objects if you intend to build polynomials yourself (and not just let PARI functions produce them, which is usually less efficient). For instance, it does not make sense to have a variable number occur in the components of a polynomial whose main variable has a higher number (lower priority), even though there’s nothing PARI can do to prevent you from doing it.

4.6.2 Creating variables A basic difficulty is to “create” a variable. As we have seen in Section 4.1, a plethora of objects is associated to variable number v. Here is the complete list: polun[v] and polx[v], which you can use in library mode and which represent, respectively, the monic monomials of degrees 0 and 1 in v; varentries[v], and polvar[v]. The latter two are only meaningful to GP, but they have to be set nevertheless. All of them must be properly defined before you can use a given integer as a variable number.

Initially, this is done for 0 (the variable x under GP), and MAXVARN, which is there to address the need for a “temporary” new variable, which would not be used in regular objects (created by the library). We call the latter type a “temporary variable”. The regular variables meant to be used in regular objects, are called “user variables”.

4.6.2.1 User variables: When the program starts, x is the only user variable (number 0). To define new ones, use

long fetch user var(char *s)

which inspects the user variable named s (creating it if needed), and returns its variable number.

    long v = fetch user var("y");
    GEN gy = polx[v];

This function raises an error if s is already known as a function name to the interpreter.

Caveat: it is possible to use flissexpr (see Section 4.7.1) to execute a GP command and create GP variables on the fly as needed:

    GEN gy = flissexpr("y"); /* supposedly returns polx[v], for some v */
    long v = gvar(gy);

This is dangerous, especially when programming functions that will be used under GP. The code above reads the value of y, as it is currently known by the GP interpreter (possibly creating it in the process). All is well and good if y hasn’t been tampered with in previous GP commands. But if y has been modified (e.g y = 1), then the value of gy is not what you expected it to be and corresponds instead to the current value of the GP variable (e.g gun).

4.6.2.2 Temporary variables: MAXVARN should be enough for most cases, but you can create more temporary variables using

long fetch var()

This returns a variable number which is guaranteed to be unused by the library at the time you get it and as long as you do not delete it (we’ll see how to do that shortly). This has lower number (i.e. higher priority) than any temporary variable produced so far (MAXVARN is assumed to be the first such). This call updates all the aforementioned internal arrays. In particular, after the statement v = fetch var(), you can use polun[v] and polx[v]. The variables created in this way have no identifier assigned to them though, and they will be printed as #<number>, except for MAXVARN which will be printed as #. You can assign a name to a temporary variable, after creating it, by calling the function

    void name var(long n, char *s)

after which the output machinery will use the name s to represent the variable number n. The GP parser will not recognize it by that name, however, and calling this on a variable known to GP will raise an error. Temporary variables are meant to be used as free variables, and you should never assign values or functions to them as you would do with variables under GP. For that, you need a user variable.

All objects created by fetch var are on the heap and not on the stack, thus they are not subject to standard garbage collecting (they won’t be destroyed by a gerepile or avma = ltop statement). When you don’t need a variable number anymore, you can delete it using

long delete var()

which deletes the latest temporary variable created and returns the variable number of the previous one (or simply returns 0 if you try, in vain, to delete MAXVARN). Of course you should make sure that the deleted variable does not appear anywhere in the objects you use later on. Here is an example:

    {
      long first = fetch var();
      long n1 = fetch var();
      long n2 = fetch var(); /* prepare three variables for internal use */
      ...
      /* delete all variables before leaving */
      do { num = delete var(); } while (num && num <= first);
    }

The (dangerous) statement

    while (delete var()) /* empty */;

removes all temporary variables that were in use, except MAXVARN which cannot be deleted.

4.7 Input and output.

Two important aspects have not yet been explained which are specific to library mode: input and output of PARI objects.

4.7.1 Input.

For input, PARI provides you with two powerful high level functions which enables you to input your objects as if you were under GP. In fact, the second one is essentially the GP syntactical parser, hence you can use it not only for input but for (most) computations that you can do under GP. These functions are called flisexpr and flisseq. The first one has the following syntax:

    GEN flisexpr(char *s)

Its effect is to analyze the input string s and to compute the result as in GP. However it is limited to one expression. If you want to read and evaluate a sequence of expressions, use

    GEN flisseq(char *s)

In fact these two functions start by filtering out all spaces and comments in the input string (that’s what the initial f stands for). They then call the underlying basic functions, the GP parser proper: GEN lisexpr(char *s) and GEN lisseq(char *s), which are slightly faster but which you probably don’t need.

To read a GEN from a file, you can use the simpler interface

    GEN lisGEN(FILE *file)

which reads a character string of arbitrary length from the stream file (up to the first newline character), applies flisexpr to it, and returns the resulting GEN. This way, you won’t have to worry about allocating buffers to hold the string. To interactively input an expression, use

lisGEN(stdin).

Once in a while, it may be necessary to evaluate a GP expression sequence involving a call to a function you have defined in C. This is easy using install which allows you to manipulate quite an arbitrary function (GP knows about pointers!). The syntax is

    void install(void *f, char *name, char *code)

where f is the (address of) the function (cast to the C type void*), name is the name by which you want to access your function from within your GP expressions, and code is a character string describing the function call prototype (see Section 4.9.2 for the precise description of prototype strings). In case the function returns a GEN, it should satisfy gerepileupto assumptions (see Section 4.4).

4.7.2 Output. For output, there exist essentially three different functions (with variants), corresponding to the three main GP output formats (as described in Section 2.1.14), plus three extra ones, respectively devoted to TEX output, string output, and (advanced) debugging.

• “raw” format, obtained by using the function brute with the following syntax:

    void brute(GEN obj, char x, long n);

This prints the PARI object obj in format x0.n, using the notations from Section 2.1.8. Recall that here x is either 'e', 'f' or 'g' corresponding to the three numerical output formats, and n is the number of printed significant digits, and should be set to −1 if all of them are wanted (these arguments only affect the printing of real numbers). Usually you won’t need that much flexibility, so most of the time you will get by with the function

void outbrute(GEN obj), which is equivalent to brute(x,'g',-1), or even better, with

void output(GEN obj) which is equivalent to outbrute(obj) followed by a newline and a buffer flush. This is especially nice during debugging. For instance using dbx or gdb, if obj is a GEN, typing print output(obj) will enable you to see the content of obj (provided the optimizer has not put it into a register, but it’s rarely a good idea to debug optimized code).

• “prettymatrix” format: this format is identical to the preceding one except for matrices. The relevant functions are:

    void matbrute(GEN obj, char x, long n)

void outmat(GEN obj), which is followed by a newline and a buffer flush.

• “prettyprint” format: the basic function has an additional parameter m, corresponding to the (minimum) field width used for printing integers:

    void sor(GEN obj, char x, long n, long m)

The simplified version is

void outbeaut(GEN obj) which is equivalent to sor(obj,'g',-1,0) followed by a newline and a buffer flush.

• The first extra format corresponds to the texprint function of GP, and gives a TEX output of the result. It is obtained by using:

    void texe(GEN obj, char x, long n)

• The second one is the function GENtostr which converts a PARI GEN to an ASCII string. The syntax is

char* GENtostr(GEN obj), wich returns a malloc’ed character string (which you should free after use).

• The third and final one outputs the hexadecimal tree corresponding to the GP command \x using the function

void voir(GEN obj, long nb), which will only output the first nb words corresponding to leaves (very handy when you have a look at big recursive structures). If you set this parameter to −1 all significant words will be printed. Usually this last type of output would only be used for debugging purposes.

Remark. Apart from GENtostr, all PARI output is done on the stream outfile, which by default is initialized to stdout. If you want that your output be directed to another file, you should use the function void switchout(char *name) where name is a character string giving the name of the file you are going to use. The output will be appended at the end of the file. In order to close the file, simply call switchout(NULL).

Similarly, errors are sent to the stream errfile (stderr by default), and input is done on the stream infile, which you can change using the function switchin which is analogous to switchout.

(Advanced) Remark. All output is done according to the values of the pariOut / pariErr global variables which are pointers to structs of pointer to functions. If you really intend to use these, this probably means you are rewriting GP. In that case, have a look at the code in language/es.c

(init80() or GENtostr() for instance).

4.7.3 Errors.

If you want your functions to issue error messages, you can use the general error handling routine err. The basic syntax is

    err(talker, "error message");

This will print the corresponding error message and exit the program (in library mode; go back to the GP prompt otherwise). You can also use it in the more versatile guise

    err(talker, format, ...);

where format describes the format to use to write the remaining operands, as in the printf function (however, see the next section). The simple syntax above is just a special case with a constant format and no remaining arguments.

The general syntax is

    void err(numerr,...)

where numerr is a codeword which indicates what to do with the remaining arguments and what message to print. The list of valid keywords is in language/errmessages.c together with the basic corresponding message. For instance, err(typeer,"matexp") will print the message:

    *** incorrect type in matexp.

Among the codewords are warning keywords (all those which start with the prefix warn). In that case, err does not abort the computation, just print the requested message and go on. The basic example is

    err(warner, "Strategy 1 failed.  Trying strategy 2")

which is the exact equivalent of err(talker,...) except that you certainly don’t want to stop the program at this point, just inform the user that something important has occured (in particular, this output would be suitably highlighted under GP, whereas a simple printf would not).

4.7.4 Debugging output.

The global variables DEBUGLEVEL and DEBUGMEM (corresponding to the default debug and debugmem, see Section 2.1) are used throughout the PARI code to govern the amount of diagnostic and debugging output, depending on their values. You can use them to debug your own functions, especially after having made them accessible under GP through the command install (see Section 3.11.2.13).

For debugging output, you can use printf and the standard output functions (brute or output mainly), but also some special purpose functions which embody both concepts, the main one being

    void fprintferr(char *pariformat, ...)

Now let’s define what a PARI format is. It is a character string, similar to the one printf uses, where % characters have a special meaning. It describes the format to use when printing the remaining operands. But, in addition to the standard format types, you can use %Z to denote a GEN object (we would have liked to pick %G but it was already in use!). For instance you could write:

    err(talker, "x[%d] = %Z is not invertible!", i, x[i]),

since the err function accepts PARI formats. Here i is an int, x a GEN which is not a leaf and this would insert in raw format the value of the GEN x[i].

4.7.5 Timers and timing output.

To profile your functions, you can use the PARI timer. The functions long timer() and long timer2() return the elapsed time since the last call of the same function (in milliseconds). Two different functions (identical except for their independent time-of-last-call memories!) are provided so you can have both global timing and fine tuned profiling.

You can also use void msgtimer(char *format,...), which prints prints Time, then the remaining arguments as specified by format (which is a PARI format), then the output of timer2.

4.8 A complete program.

Now that the preliminaries are out of the way, the best way to learn how to use the library mode is to work through a detailed non-trivial example of a main program. We will write a program which computes the exponential of a square matrix x. The complete listing is given in Appendix B, but each part of the program will be produced and explained here. We will use an algorithm which is not optimal but is not far from the one used for the PARI function gexp (in fact embodied in the function mpexp1). This consists in calculating the sum of the series:

e^{x/(2^{n})} = ∑_{k=0}^{∞} \frac{(x/(2^{n}))^{k}}{k!}

for a suitable positive integer n, and then computing e^{x} by repeated squarings. First, we will need to compute the L^{2}-norm of the matrix x, i.e. the quantity:

z = ∥x∥_{2} = \sqrt{∑_{i,j} x_{i,j}^{2}}.

We will then choose the integer n such that the L^{2}-norm of x/(2^{n}) is less than or equal to 1, i.e.

n = ⌈ln(z)/ln(2)⌉

if z ≥ 1, and n = 0 otherwise. Then the series will converge at least as fast as the usual one for e^{1}, and the cutoff error will be easy to estimate. In fact a larger value of n would be preferable, but this is slightly machine dependent and more complicated, and will be left to the reader.

Let us start writing our program. So as to be able to use it in other contexts, we will structure it in the following way: a main program which will do the input and output, and a function which we shall call matexp which does the real work. The main program is easy to write. It can be something like this:

    #include <pari.h>
    GEN matexp(GEN x, long prec);
    int
    main()
    {
      long d, prec = 3;
      GEN x;
      /* take a stack of a 10^6 bytes, no prime table */
      pari init(1000000,2);
      printf("precision of the computation in decimal digits:\n");
      d = itos(lisGEN(stdin));
      if (d > 0) prec = (long) (d*pariK1+3);
      printf("input your matrix in GP format:\n");
      x = matexp(lisGEN(stdin), prec);
      sor(x,'g',d,0);
      exit(0);
    }

The variable prec represents the length in longwords of the real numbers used. pariK1 is a constant (defined in paricom.h) equal to ln(10)/(ln(2) ∗ BITS IN LONG), which allows us to convert from a number of decimal digits to a number of longwords, independently of the actual bit size of your long integers. The function lisGEN reads an expression (here from standard input) and converts it to a GEN, like the GP parser itself would. This means it takes care of whitespace etc. in the input, and can do computations (e.g. matid(2) or [1,0; 0,1] are equally valid inputs).

Finally, sor is the general output routine. We have chosen to give d significant digits since this is what was asked for. Note that there is a trick hidden here: if a negative d was input, then the computation will be done in precision 3 (i.e. about 9.7 decimal digits for 32-bit machines and 19.4 for 64-bit machines) and in the function sor, giving a negative third argument outputs all the significant digits, which is entirely appropriate. Now let us attack the main course, the function

matexp:
    GEN
    matexp(GEN x, long prec)

    {
      long lx=lg(x),i,k,n,lbot, ltop = avma;
      GEN y,r,s,p1,p2;
      /* check that x is a square matrix */
      if (typ(x) != t MAT) err(talker,"this expression is not a matrix");
      if (lx == 1) return cgetg(1, t MAT);
      if (lx != lg(x[1])) err(talker,"not a square matrix");
      /* compute the L2 norm of x */
      s = gzero;
      for (i=1; i<lx; i++)
        s = gadd(s, gnorml2((GEN)x[i]));
      if (typ(s) == t REAL) setlg(s,3);
      s = gsqrt(s,3); /* we do not need much precision on s */
      /* if s<1, we are happy */
      k = expo(s);
      if (k < 0) { n = 0; p1 = x; }
      else { n = k+1; p1 = gmul2n(x,-n); setexpo(s,-1); }

Before continuing, several remarks are in order.

First, before starting this computation which will produce garbage on the stack, we have carefully saved the value of the stack pointer avma in ltop. Note that we are going to assume throughout that the garbage does not overflow the currently available stack. If it ever did, we would have several options — allocate a larger stack in the main program (for instance change 1000000 into 2000000), do some gerepileing along the way, or (if you know what you are doing)

use allocatemoremem.

Secondly, the err function is the general error handler for the PARI library. This will abort the program after printing the required message.

Thirdly, notice how we handle the special case lx = 1 (empty matrix) before accessing lx(x[1]). Doing it the other way round could produce a fatal error (a segmentation fault or a bus error, most probably). Indeed, if x is of length 1, then x[1] is not a component of x. It is just the contents of the memory cell which happens to follow the one pointed to by x, and thus has no reason to be a valid GEN. Now recall that none of the codeword handling macros do any kind of type checking (see Section 4.5), thus lg would consider x[1] as a valid address, and try to access *((GEN)x[1]) (the first codeword) which is unlikely to be a legal memory address.

In the fourth place, to compute the square of the L^{2}-norm of x we just add the squares of the L^{2}-norms of the column vectors which we obtain using the library function gnorml2. Had this function not existed, the norm computation would of course have been just as easy to write, but we would have needed a double loop.

We then take the square root of s, in precision 3 (the smallest possible). The prec argument of transcendental functions (here 3) is only taken into account when the arguments are exact objects, and thus no a priori precision can be determined from the objects themselves. To cater for this possibility, if s is of type t REAL, we use the function setlg which effectively sets the precision of s to the required value. Note that here, since we are using a numeric value for a cget function, the program will run slightly differently on 32-bit and 64-bit machines: we want to use the smallest possible bit accuracy, and this is equal to BITS IN LONG.

Note that the matrix x is allowed to have complex entries, but the function gnorml2 guarantees that s is a non-negative real number (not necessarily of type t REAL of course). If we had not known this fact, we would simply have added the instruction s = greal(s); just after the for loop.

Note also that the function gnorml2 works as desired on matrices, so we really did not need this loop at all (s = gnorml2(x) would have been enough), but we wanted to give examples of function usage. Similarly, it is of course not necessary to take the square root for testing whether the norm exceeds 1.

In the fifth place, note that we initialized the sum s to gzero, which is an exact zero. This is logical, but has some disadvantages: if all the entries of the matrix are integers (or rational numbers), the computation will take rather long, about twice as long as with real numbers of the same length. It would be better to initialize s to a real zero, using for instance the instructions:

    s = cgetr(prec+1); gaffsg(0,s);

This raises the question: which real zero does this produce (have a look at Section 1.2.6.3)? In fact, the following choice has been made: it will give you the zero with exponent equal to −BITS IN LONG times the number of longwords in the mantissa, i.e. −bit accuracy(lg(s)). Instead of the above idiom, you can also use the function GEN realzero(long prec), which simply returns a real zero

to accuracy -bit accuracy(prec).

The sixth remark here is about how to determine the approximate size of a real number. The fastest way to do this is to look at its binary exponent. Hence we need to have s actually represented as a real number, and not as an integer or a rational number. The result of transcendental functions is guaranteed to be of type t REAL, or complex with t REAL components, thus this is indeed the case after the call to gsqrt since its argument is a nonnegative (real) number.

Finally, note the use of the function gmul2n. It has the following syntax:

    GEN gmul2n(GEN x, long n);

and the effect is simply to multiply x by 2^{n}, where n can be positive or negative. This is much faster than gmul or gmulgs.

There is another function gshift with exactly the same syntax. When n is non-negative, the effects of these two functions are the same. However, when n is negative, gshift acts like a right shift of -n, hence does not noramlly perform an exact division on integers. The function gshift is the PARI analogue of the C or GP operators << and >>.

We now come to the heart of the function. We have a GEN p1 which points to a certain matrix of which we want to take the exponential. We will want to transform this matrix into a matrix with real (or complex of real) entries before starting the computation. To do this, we simply multiply by the real number 1 in precision prec + 1 (to be on the side of safety). To sum the series, we will use three variables: a variable p2 which at stage k will contain p1^{k}/k!, a variable y which will contain ∑_{i=0}^{k} p1^{i}/i!, and a variable r which will contain the size estimate s^{k}/k!. Note that we do not use Horner’s rule. This is simply because we are lazy and do not want to compute in advance the number of terms that we need. We leave this modification (and many other improvements!) to the reader. The program continues as follows:

      /* initializations before the loop */
      r = cgetr(prec+1); gaffsg(1,r); p1 = gmul(r,p1);

y = gscalmat(r,lx-1); /* creates scalar matrix with r on diagonal */
p2 = p1; r = s; k = 1;
y = gadd(y,p2);
/* now the main loop */
while (expo(r) >= -BITS IN LONG*(prec-1))
{
  k++; p2 = gdivgs(gmul(p2,p1),k);
  r = gdivgs(gmul(s,r),k); y = gadd(y,p2);
}
/* now square back n times if necessary */
if (!n) { lbot = avma; y = gcopy(y); }
else
{
  for (i=0; i<n; i++) { lbot = avma; y = gsqr(y); }
}
return gerepile(ltop,lbot,y);
}

A few remarks once again. First note the use of the function gscalmat with the following syntax:

GEN gscalmat(GEN x, long m);

The effect of this function is to create the m × m scalar matrix whose diagonal entries are x. Hence the length of the matrix including the codeword will in fact be m+1. There is a corresponding function gscalsmat which takes a long as a first argument.

If we refer to what has been said above, the main loop should be self-evident.

When we do the final squarings, according to the fundamental dogma on the use of gerepile, we keep the value of avma in lbot just before the squaring, so that if it is the last one, lbot will indeed be the bottom address of the garbage pile, and gerepile will work. Note that it takes a completely negligible time to do this in each loop compared to a matrix squaring. However, when n is initially equal to 0, no squaring has to be done, and we have our final result ready but we lost the address of the bottom of the garbage pile. Hence we use the trick of copying y again to the top of the stack. This is inefficient, but does the trick. If we wanted to avoid this using only gerepile, the best thing to do would be to put the instruction lbot=avma just before both occurrences of the instruction y=gadd(p2,y). Of course, we could also rewrite the last block as follows:

/* now square back n times */
for (i=0; i<n; i++) y = gsqr(y);
return gerepileupto(ltop,y);
}

because it does not matter to gerepileupto that we have lost the address just before the final result (note that the loop is not executed if n is 0). It is safe to use gerepileupto here as y will have been created by either gsqr or gadd, both of which are guaranteed to return suitable objects.

Remarks. As such, the program should work most of the time if x is a square matrix with real or complex entries. Indeed, since essentially the first thing that we do is to multiply by the real number 1, the program should work for integer, real, rational, complex or quadratic entries. This is in accordance with the behavior of transcendental functions.

Furthermore, since this program is intended to be only an illustrative example, it has been written a little sloppily. In particular many error checks have been omitted, and the efficiency is far from optimal. An evident improvement would be the use of gerepileupto mentioned above. Another improvement is to multiply the matrix x by the real number 1 right at the beginning, speeding up the computation of the L^{2}-norm in many cases. These improvements are included in the version given in Appendix B. Still another improvement would come from a better choice of n. If the reader takes a look at the implementation of the function mpexp1 in the file basemath/trans1.c, he can make the necessary changes himself. Finally, there exist other algorithms of a different nature to compute the exponential of a matrix.

4.9 Adding functions to PARI.

4.9.1 Nota Bene. As already mentioned, modified versions of the PARI package should NOT be spread without our prior approval. If you do modify PARI, however, it is certainly for a good reason, hence we would like to know about it, so that everyone can benefit from it. There is then a good chance that the modifications that you have made will be incorporated into the next release.

(Recall the e-mail address: pari@math.u-bordeaux.fr, or use the mailing lists).

Roughly four types of modifications can be made. The first type includes all improvements to the documentation, in a broad sense. This includes correcting typos or inacurracies of course, but also items which are not really covered in this document, e.g. if you happen to write a tutorial, or pieces of code exemplifying some fine points that you think were unduly omitted.

The second type is to expand or modify the configuration routines and skeleton files (the Configure script and anything in the config/ subdirectory) so that compilation is possible (or easier, or more efficient) on an operating system previously not catered for. This includes discovering and removing idiosyncrasies in the code that would hinder its portability.

The third type is to modify existing (mathematical) code, either to correct bugs, to add new functionalities to existing functions, or to improve their efficiency.

Finally the last type is to add new functions to PARI. We explain here how to do this, so that in particular the new function can be called from GP.

4.9.2 The calling interface from GP, parser codes. A parser code is a character string describing all the GP parser needs to know about the function prototype. It contains a sequence of the following atoms:

• Syntax requirements, used by functions like for, sum, etc.:

=  separator = required at this point (between two arguments)

• Mandatory arguments, appearing in the same order as the input arguments they describe:

    G     GEN
    &     *GEN

L  long (we implicitly identify int with long)
S  symbol (i.e. GP identifier name). Function expects a *entree

V  variable (as S, but rejects symbols associated to functions)
n  variable, expects a variable number (a long, not an *entree)
I  string containing GP code (useful for control statements, to be processed with lisexpr

        or lisseq)

r  raw input (treated as a string without quotes). Quoted args are copied as strings.

Stops at first unquoted ')' or ','. Special chars can be quoted using ’\’.

Example: aa"b\n)"c yields the string "aab\n)c"

s  expanded string. Example: Pi"x"2 yields "3.142x2".

Unquoted components can be of any PARI type (converted following current output format)

• Optional arguments:

s*  any number of strings, possibly 0 (see s)
s*p idem, setting prettyp = 1 (i.e. in beautified format)
s*t idem, in TEX format
Dxxx argument has a default value.

The format to indicate a default value (atom starts with a D) is “Dvalue,type,”, where type is the code for any mandatory atom (previous group), value is any valide GP expression which is converted according to type, and the ending comma is mandatory. For instance D0,L, stands for “this optional argument will be converted to a long, and is 0 by default”. So if the user-given argument reads 1 + 3 at this point, (long)4 is sent to the function (via itos()); and (long)0 if the argument is ommitted. The following special syntaxes are available:

DG  optional GEN, send NULL if argument omitted.
D&  optional *GEN, send NULL if argument omitted.
DV  optional *entree, send NULL if argument omitted.
DI  optional *char, send NULL if argument omitted.
Dn  optional variable number, −1 if omitted.

• Automatic arguments:

f  Fake *long. C function requires a pointer but we don’t use the resulting long
p  real precision (default realprecision)
P  series precision (default seriesprecision, global variable precdl for the library)

• Return type: GEN by default, otherwise the following can appear anywhere in the code string:

l  return long
v  return void

No more than 8 arguments can be given (syntax requirements and return types are not considered as arguments). This is currently hardcoded but can trivially be changed by modifying the definition of argvec in anal.c:identifier(). This limitation should disappear in future versions.

When the function is called under GP, the prototype is scanned and each time an atom corresponding to a mandatory argument is met, a user-given argument is read (GP outputs an error message it the argument was missing). Each time an optional atom is met, a default value is inserted if the user omits the argument. The “automatic” atoms fill in the argument list transparently, supplying the current value of the corresponding variable (or a dummy pointer).

For instance, here is how you would code the following prototypes (which don’t involve default values):

GEN name(GEN x, GEN y, long prec)   ---->  "GGp"
void name(GEN x, long y, long prec) ---->  "vGLp" (or "GvLp")
long name(GEN x)                    ---->  "Gl"

If you want more examples, GP gives you easy access to the parser codes associated to all GP functions: just type \h function. You can then compare with the C prototypes as they stand in the code.

Remark: If you need to implement complicated control statements (probably for some improved summation functions), you’ll need to know about the entree type, which is not documented. Check the comment before the function list at the end of language/init.c and the source code in language/sumiter.c. You should be able to make something of it.

4.9.3 Coding guidelines. Code your function in a file of its own, using as a guide other functions in the PARI sources. One important thing to remember is to clean the stack before exiting your main function (usually using gerepile), since otherwise successive calls to the function will clutter the stack with unnecessary garbage, and stack overflow will occur sooner. Also, if it returns a GEN and you want it to be accessible to GP, you have to make sure this GEN is suitable for gerepileupto (see Section 4.4).

If error messages are to be generated in your function, use the general error handling routine err (see Section 4.7.3). Recall that, apart from the warn variants, this function does not return but ends with a longjmp statement. As well, instead of explicit printf / fprintf statements, use the following encapsulated variants:

void pariputs(char *s): write s to the GP output stream.

void fprintferr(char *s): write s to the GP error stream (this function is in fact much more versatile, see Section 4.7.4).

Declare all public functions in paridecl.h (you want the outside world to know about them). The other ones should be declared static in your file.

Your function is now ready to be used in library mode after compilation and creation of the library. If possible, compile it as a shared library (see the Makefile coming with the matexp example in the distribution). It is however still inaccessible from GP.

4.9.4 Integration with GP as a shared module

To tell GP about your function, you must do the following. First, find a name for it. It does not have to match the one used in library mode, but consistency is nice. It has to be a valid GP identifier, i.e. use only alphabetic characters, digits and the underscore character ( ), the first character being alphabetic.

Then you have to figure out the correct parser code corresponding to the function prototype. This has been explained above (Section 4.9.2).

Now, assuming your Operating System is supported by install, simply write a GP script like the following:

install(name, code, gpname, library)

addhelp(gpname, some help text)

(see Section 3.11.2.1 and 3.11.2.13). The addhelp part is not mandatory, but very useful if you want others to use your module.

Read that file from your GP session (from your preferences file for instance, see Section 2.8), and that’s it, you can use the new function gpname under GP (and we would very much like to hear about it!).

4.9.5 Integration the hard way

If install is not available for your Operating System, it’s more complicated: you have to hardcode your function in the GP binary (or install Linux). Here’s what needs to be done:

In the definition of functions basic (file language/init.c), add your entry in exact alphabetical order by its GP name (note that digits come before letters), in a line of the form:

    {"gpname",V,(void*)libname,secno,"code"},

where

libname is the name of your function in library mode,

gpname the name that you have chosen to call it under GP,

secno is the section number of Chapter 3 in which this function would belong (type ? in GP to see the list),

V is a number between 0 and 99. Right now, there are only two significant values: zero means that it’s possible to call the function without argument, and non-zero means it needs at least one argument.

code is the parser code.

Once this has been done, in the file language/helpmessages.c add in exact alphabetical order a short message describing the effect of your function: "name(x,y,...)=short descriptive

message",

The message must be a single line, of arbitrary length. Do not use \n; the necessary newlines will be inserted by GP’s online help functions. Optional arguments should be shown between braces (see the other messages for comparison).

Now, you can recompile GP.

4.9.6 Example. A complete description could look like this:

    {
      install(bnfinit0, GD0,L,DGp, ClassGroupInit, "libpari.so")
      addhelp(ClassGroupInit, "ClassGroupInit(P,{flag=0},{data=[]}):
        compute the necessary data for ...")
    }

which means we have a function ClassGroupInit under GP, which calls the library function bnfinit0 . The function has one mandatory argument (V is non-zero), and possibly two more (two 'D' in the code), plus the current real precision. More precisely, the first argument is a GEN, the second one is converted to a long using itos (0 is passed if it is omitted), and the third one is also a GEN, but we pass NULL if no argument was supplied by the user. This matches the C prototype

(from paridecl.h):
    GEN bnfinit0(GEN P, long flag, GEN data, long prec)

This function is in fact coded in basemath/buch2.c, and will in this case be completely identical to the GP function bnfinit but GP does not need to know about this, only that it can be found somewhere in the shared library libpari.so.

Important note You see in this example that it is the function’s responsibility to correctly interpret its operands: data = NULL is interpreted by the function as an empty vector. Note that since NULL is never a valid GEN pointer, this trick always enables you to distinguish between a default value and actual input: the user could explicitly supply an empty vector!

Note If install were not available we would have to modify language/helpmessages.c, and language/init.c and recompile GP. The entry in functions basic corresponding to the function above is actually

    {
    "bnfinit", 91, (void*)bnfinit0, 6, "GD0,L,DGp"
    }

178

Technical Reference Guide for Low-Level Functions

In this chapter, we give a description all public low-level functions of the PARI system. These essentially include functions for handling all the PARI types. Higher level functions, such as arithmetic or transcendental functions, are described fully in Chapter 3 of this manual.

Many other undocumented functions can be found throughout the source code. These private functions are more efficient than the library functions that call them, but much sloppier on argument checking and damage control. Use them at your own risk!

5.1 Level 0 kernel (operations on unsigned longs).

For the non-68k versions, we need level 0 operations simulating basic operations of the 68020 processor (on which PARI was originally implemented). The type ulong is defined in the file parigen.h as unsigned long. Note that in the prototypes below a ulong is sometimes implicitly typecast to int or long.

The global ulong variables overflow (which will contain only 0 or 1) and hiremainder used to be declared in the file pariinl.h. However, for certain architectures they are no longer needed, and/or have been replaced with local variables for efficiency; and the ‘functions’ mentioned below are really chunks of assembler code which will be inlined at each invocation by the compiler. If you really need to use these lowest-level operations directly, make sure you know your way through the PARI kernel sources, and understand the architecture dependencies.

To make the following descriptions valid both for 32-bit and 64-bit machines, we will set BIL to be equal to 32 (resp. 64), an abbreviation of BITS IN LONG, which is what is actually used in the source code.

int addll(int x, int y): adds the ulongs x and y, returns the lower BIL bits and puts the carry

bit into overflow.

int addllx(int x, int y): adds overflow to the sum of the ulongs x and y, returns the lower BIL bits and puts the carry bit into overflow.

int subll(int x, int y): subtracts the ulongs x and y, returns the lower BIL bits and put the carry (borrow) bit into overflow.

int subllx(int x, int y): subtracts overflow from the difference of the ulongs x and y, returns the lower BIL bits and puts the carry (borrow) bit into overflow.

int shiftl(ulong x, ulong y): shifts the ulong x left by y bits, returns the lower BIL bits and stores the high-order BIL bits into hiremainder. We must have 1 ≤ y ≤ BIL. In particular, y must be non-zero; the caller is responsible for testing this.

int shiftlr(ulong x, ulong y): shifts the ulong x << BIL right by y bits, returns the higher BIL bits and stores the low-order BIL bits into hiremainder. We must have 1 ≤ y ≤ BIL. In particular, y must be non-zero.

int bfffo(ulong x): returns the number of leading zero bits in the ulong x (i.e. the number of bit positions by which it would have to be shifted left until its leftmost bit first becomes equal to 1, which can be between 0 and BIL − 1 for nonzero x). When x is 0, BIL is returned. int mulll(ulong x, ulong y): multiplies the ulong x by the ulong y, returns the lower BIL bits and stores the high-order BIL bits into hiremainder. int addmul(ulong x, ulong y): adds hiremainder to the product of the ulongs x and y, returns the lower BIL bits and stores the high-order BIL bits into hiremainder. int divll(ulong x, ulong y): returns the Euclidean quotient of (hiremainder << BIL) + x and the ulong divisor y and stores the remainder into hiremainder. An error occurs if the quotient cannot be represented by a ulong, i.e. if hiremainder ≥ y initially.

5.2 Level 1 kernel (operations on longs, integers and reals).

In this section as elsewhere, long denotes a BIL-bit signed C-integer, “integer” denotes a PARI multiprecise integer (type t INT), “real” denotes a PARI multiprecise real (type t REAL). Refer to Chapters 1–2 and 4 for general background. Note: Many functions consist of an elementary operation, immediately followed by an assignment statement. All such functions are obtained using macros (see the file paricom.h), hence you can easily extend the list. Below, they will be introduced like in the following example: GEN gadd[z](GEN x, GEN y[, GEN z]): followed by the explicit description of the function

    GEN gadd(GEN x, GEN y)

which creates its result on the stack, returning a GEN pointer to it, and the parts in brackets indicate that there exists also a function

    void gaddz(GEN x, GEN y, GEN z)

which assigns its result to the pre-existing object z, leaving the stack unchanged.

5.2.1 Basic unit and subunit handling functions long typ(GEN x): returns the type number of x. (The header files included through pari.h will give you access to the symbolic constants t INT etc., so you should never need to know the actual numerical values.) long lg(GEN x): returns the length of x in BIL-bit words. long lgef(GEN x): returns the effective length of the polynomial x in BIL-bit words. long lgefint(GEN x): returns the effective length of the integer x in BIL-bit words. long signe(GEN x): returns the sign (−1, 0 or 1) of x. Can be used for integers, reals, polynomials and power series (for the last two types, only 0 or 1 are possible). long gsigne(GEN x): same as signe, but also valid for rational numbers (and marginally less efficient for the other types). long expo(GEN x): returns the unbiased binary exponent of the real number x. long gexpo(GEN x): same as expo, but also valid when x is not a real number. When x is an exact 0, this returns -HIGHEXPOBIT. long expi(GEN x): returns the binary exponent of the real number equal to the integer x. This is a special case of gexpo above, covering the case where x is of type t INT. long valp(GEN x): returns the unbiased 16-bit p-adic valuation (for a p-adic) or X-adic valuation (for a power series, taken with respect to the main variable) of x.

long precp(GEN x): returns the precision of the p-adic x.

long varn(GEN x): returns the variable number of x (between 0 and MAXVARN). Should be used only for polynomials and power series.

long gvar((GEN x)): returns the main variable number when any variable at all occurs in the composite object x (the smallest variable number which occurs), and BIGINT otherwise.

void settyp(GEN x, long s): sets the type number of x to s. This should be used with extreme care since usually the type is set otherwise, and the components and further codeword fields (which are left unchanged) may not match the PARI conventions for the new type.

void setlg(GEN x, long s): sets the length of x to s. Again this should be used with extreme care since usually the length is set otherwise, and increasing the length joins previously unrelated memory words to the root node of x. This is, however, an extremely efficient way of truncating vectors or polynomials.

void setlgef(GEN x, long s): sets the effective length of x to s, where x is a polynomial. The number s must be less than or equal to the length of x.

void setlgefint(GEN x, long s): sets the effective length of the integer x to s. The number s must be less than or equal to the length of x.

void setsigne(GEN x, long s): sets the sign of x to s. If x is an integer or real, s must be equal to −1, 0 or 1, and if x is a polynomial or a power series, s must be equal to 0 or 1.

void setexpo(GEN x, long s): sets the binary exponent of the real number x to s, after adding the appropriate bias. The unbiased value s must be a 24-bit signed number.

void setvalp(GEN x, long s): sets the p-adic or X-adic valuation of x to s, if x is a p-adic or a power series, respectively.

void setprecp(GEN x, long s): sets the p-adic precision of the p-adic number x to s.

void setvarn(GEN x, long s): sets the variable number of the polynomial or power series x to s (where 0 ≤ s ≤ MAXVARN).

5.2.2 Memory allocation on the PARI stack

GEN cgetg(long n, long t): allocates memory on the PARI stack for an object of length n and type t, and initializes its first codeword.

GEN cgeti(long n): allocates memory on the PARI stack for an integer of length n, and initializes its first codeword. Identical to cgetg(n,t INT).

GEN cgetr(long n): allocates memory on the PARI stack for a real of length n, and initializes its first codeword. Identical to cgetg(n,t REAL).

void cgiv(GEN x): frees object x if it is the last created on the PARI stack (otherwise disaster occurs).

GEN gerepile(long p, long q, GEN x): general garbage collector for the PARI stack. See Section 4.4 for a detailed explanation and many examples.

5.2.3 Assignments, conversions and integer parts void mpaff(GEN x, GEN z): assigns x into z (where x and z are integers or reals). void affsz(long s, GEN z): assigns the long s into the integer or real z. void affsi(long s, GEN z): assigns the long s into the integer z. void affsr(long s, GEN z): assigns the long s into the real z. void affii(GEN x, GEN z): assigns the integer x into the integer z. void affir(GEN x, GEN z): assigns the integer x into the real z. void affrs(GEN x, long s): assigns the real x into the long s… not. This is a forbidden assignment in PARI, so an error message is issued. void affri(GEN x, GEN z): assigns the real x into the integer z… no it doesn’t. This is a forbidden assignment in PARI, so an error message is issued. void affrr(GEN x, GEN z): assigns the real x into the real z. GEN stoi(long s): creates the PARI integer corresponding to the long s. long itos(GEN x): converts the PARI integer x to a C long (if possible, otherwise an error message is issued). GEN mptrunc[z](GEN x[, GEN z]): truncates the integer or real x (not the same as the integer part if x is non-integer and negative). GEN mpent[z](GEN x[, GEN z]): true integer part of the integer or real x (i.e. the floor function). 5.2.4 Valuation and shift long vals(long s): 2-adic valuation of the long s. Returns −1 if s is equal to 0, with no error. long vali(GEN x): 2-adic valuation of the integer x. Returns −1 if s is equal to 0, with no error. GEN mpshift[z](GEN x, long n[, GEN z]): shifts the real or integer x by n. If n is positive, this is a left shift, i.e. multiplication by 2^{n}. If n is negative, it is a right shift by −n, which amounts to the truncation of the quotient of x by 2^{−n}. GEN shifts(long s, long n): converts the long s into a PARI integer and shifts the value by n. GEN shifti(GEN x, long n): shifts the integer x by n. GEN shiftr(GEN x, long n): shifts the real x by n. 5.2.5 Unary operations Let “op” be some unary operation of type GEN (*)(GEN). The names and prototypes of the low-level functions corresponding to op will be as follows. GEN mpop(GEN x): creates the result of op applied to the integer or real x. GEN ops(long s): creates the result of op applied to the long s. GEN opi(GEN x): creates the result of op applied to the integer x. GEN opr(GEN x): creates the result of op applied to the real x. GEN mpopz(GEN x, GEN z): assigns the result of applying op to the integer or real x into the integer or real z.

Remark: it has not been considered useful to include the functions void opsz(long,GEN), void

opiz(GEN,GEN) and void oprz(GEN, GEN).

The above prototype schemes apply to the following operators:

op=neg: negation (−x). The result is of the same type as x. op=abs: absolute value (|x|). The result is of the same type as x.

In addition, there exist the following special unary functions with assignment: void mpinvz(GEN x, GEN z): assigns the inverse of the integer or real x into the real z. The inverse is computed as a quotient of real numbers, not as a Euclidean division. void mpinvsr(long s, GEN z): assigns the inverse of the long s into the real z. void mpinvir(GEN x, GEN z): assigns the inverse of the integer x into the real z. void mpinvrr(GEN x, GEN z): assigns the inverse of the real x into the real z.

5.2.6 Comparison operators long mpcmp(GEN x, GEN y): compares the integer or real x to the integer or real y. The result is the sign of x − y. long cmpss(long s, long t): returns the sign of s − t. long cmpsi(long s, GEN x): compares the long s to the integer x. long cmpsr(long s, GEN x): compares the long s to the real x. long cmpis(GEN x, long s): compares the integer x to the long s. long cmpii(GEN x, GEN y): compares the integer x to the integer y. long cmpir(GEN x, GEN y): compares the integer x to the real y. long cmprs(GEN x, long s): compares the real x to the long s. long cmpri(GEN x, GEN y): compares the real x to the integer y. long cmprr(GEN x, GEN y): compares the real x to the real y.

5.2.7 Binary operations Let “op” be some operation of type GEN (*)(GEN,GEN). The names and prototypes of the low-level functions corresponding to op will be as follows. In this section, the z argument in the z-functions must be of type t INT or t REAL. GEN mpop[z](GEN x, GEN y[, GEN z]): applies op to the integer-or-reals x and y. GEN opss[z](long s, long t[, GEN z]): applies op to the longs s and t. GEN opsi[z](long s, GEN x[, GEN z]): applies op to the long s and the integer x. GEN opsr[z](long s, GEN x[, GEN z]): applies op to the long s and the real x. GEN opis[z](GEN x, long s[, GEN z]): applies op to the integer x and the long s. GEN opii[z](GEN x, GEN y[, GEN z]): applies op to the integers x and y. GEN opir[z](GEN x, GEN y[, GEN z]): applies op to the integer x and the real y.

GEN oprs[z](GEN x, long s[, GEN z]): applies op to the real x and the long s.

GEN opri[z](GEN x, GEN y[, GEN z]): applies op to the real x and the integer y.

GEN oprr[z](GEN x, GEN y[, GEN z]): applies op to the reals x and y.

Each of the above can be used with the following operators.

op=add: addition (x + y). The result is real unless both x and y are integers (or longs).

op=sub: subtraction (x - y). The result is real unless both x and y are integers (or longs).

op=mul: multiplication (x * y). The result is real unless both x and y are integers (or longs), OR if x or y is the integer or long zero.

op=div: division (x / y). In the case where x and y are both integers or longs, the result is the Euclidean quotient, where the remainder has the same sign as the dividend x. If one of x or y is real, the result is real unless x is the integer or long zero. A division-by-zero error occurs if y is equal to zero.

op=res: remainder (“x % y”). This operation is defined only when x and y are longs or integers. The result is the Euclidean remainder corresponding to div, i.e. its sign is that of the dividend x. The result is always an integer.

op=mod: remainder (x % y). This operation is defined only when x and y are longs or integers. The result is the true Euclidean remainder, i.e. non-negative and less than the absolute value of y.

5.2.8 Division with remainder: the following functions return two objects, unless specifically asked for only one of them — a quotient and a remainder. The remainder will be created on the stack, and a GEN pointer to this object will be returned through the variable whose address is passed as the r argument.

GEN dvmdss(long s, long t, GEN *r): creates the Euclidean quotient and remainder of the longs s and t. If r is not NULL or ONLY REM, this puts the remainder into *r, and returns the quotient. If r is equal to NULL, only the quotient is returned. If r is equal to ONLY REM, the remainder is returned instead of the quotient. In the generic case, the remainder is created after the quotient and can be disposed of individually with a cgiv(r). The remainder is always of the sign of the dividend s.

GEN dvmdsi(long s, GEN x, GEN *r): creates the Euclidean quotient and remainder of the long s by the integer x. Obeys the same conventions with respect to r.

GEN dvmdis(GEN x, long s, GEN *r): create the Euclidean quotient and remainder of the integer x by the long s.

GEN dvmdii(GEN x, GEN y, GEN *r): returns the Euclidean quotient of the integer x by the integer y and puts the remainder into *r. If r is equal to NULL, the remainder is not created, and if r is equal to ONLY REM, only the remainder is created and returned. In the generic case, the remainder is created after the quotient and can be disposed of individually with a cgiv(r). The remainder is always of the sign of the dividend x.

GEN truedvmdii(GEN x, GEN y, GEN *r):, as dvmdii but with a non-negative remainder.

void mpdvmdz(GEN x, GEN y, GEN z, GEN *r): assigns the Euclidean quotient of the integers x and y into the integer or real z, putting the remainder into *r (unless r is equal to NULL or ONLY REM as above). void dvmdssz(long s, long t, GEN z, GEN *r): assigns the Euclidean quotient of the longs s and t into the integer or real z, putting the remainder into *r (unless r is equal to NULL or ONLY REM as above).

void dvmdsiz(long s, GEN x, GEN z, GEN *r): assigns the Euclidean quotient of the long s and the integer x into the integer or real z, putting the remainder into *r (unless r is equal to NULL or ONLY REM as above).

void dvmdisz(GEN x, long s, GEN z, GEN *r): assigns the Euclidean quotient of the integer x and the long s into the integer or real z, putting the remainder into *r (unless r is equal to NULL or ONLY REM as above).

void dvmdiiz(GEN x, GEN y, GEN z, GEN *r): assigns the Euclidean quotient of the integers x and y into the integer or real z, putting the address of the remainder into *r (unless r is equal to NULL or ONLY REM as above).

5.2.9 Miscellaneous functions

void addsii(long s, GEN x, GEN z): assigns the sum of the long s and the integer x into the integer z (essentially identical to addsiz except that z is specifically an integer).

long divise(GEN x, GEN y): if the integer y divides the integer x, returns 1 (true), otherwise returns 0 (false).

long divisii(GEN x, long s, GEN z): assigns the Euclidean quotient of the integer x and the long s into the integer z, and returns the remainder as a long.

long mpdivis(GEN x, GEN y, GEN z): if the integer y divides the integer x, assigns the quotient to the integer z and returns 1 (true), otherwise returns 0 (false).

void mulsii(long s, GEN x, GEN z): assigns the product of the long s and the integer x into the integer z (essentially dentical to mulsiz except that z is specifically an integer).

5.3 Level 2 kernel (operations on general PARI objects).

The functions available to handle subunits are the following.

GEN compo(GEN x, long n): creates a copy of the n-th true component (i.e. not counting the codewords) of the object x.

GEN truecoeff(GEN x, long n): creates a copy of the coefficient of degree n of x if x is a scalar, polynomial or power series, and otherwise of the n-th component of x.

The remaining two are macros, NOT functions (see Section 4.2.1 for a detailed explanation):

long coeff(GEN x, long i, long j): applied to a matrix x (type t MAT), this gives the address of the coefficient at row i and column j of x.

long maeln(GEN x, long a_{1}, ..., long a_{n}): stands for x[a_{1}][a_{2}]...[a_{n}], where 2 ≤ n ≤ 5, with all the necessary typecasts.

5.3.1 Copying and conversion GEN cgetp(GEN x): creates space sufficient to hold the p-adic x, and sets the prime p and the p-adic precision to those of x, but does not copy (the p-adic unit or zero representative and the modulus of) x. GEN gcopy(GEN x): creates a new copy of the object x on the PARI stack. For permanent subobjects, only the pointer is copied. GEN forcecopy(GEN x): same as copy except that even permanent subobjects are copied onto the stack. long taille(GEN x): returns the total number of BIL-bit words occupied by the tree representing x. GEN gclone(GEN x): creates a new permanent copy of the object x on the heap. GEN greffe(GEN x, long l, int use stack): applied to a polynomial x (type t POL), creates a power series (type t SER) of length l starting with x, but without actually copying the coefficients, just the pointers. If use stack is zero, this is created through malloc, and must be freed after use. Intended for internal use only. double rtodbl(GEN x): applied to a real x (type t REAL), converts x into a C double if possible. GEN dbltor(double x): converts the C double x into a PARI real. double gtodouble(GEN x): if x is a real number (but not necessarily of type t REAL), converts x into a C double if possible. long gtolong(GEN x): if x is an integer (not a C long, but not necessarily of type t INT), converts x into a C long if possible. GEN gtopoly(GEN x, long v): converts or truncates the object x into a polynomial with main variable number v. A common application would be the conversion of coefficient vectors. GEN gtopolyrev(GEN x, long v): converts or truncates the object x into a polynomial with main variable number v, but vectors are converted in reverse order. GEN gtoser(GEN x, long v): converts the object x into a power series with main variable number v. GEN gtovec(GEN x): converts the object x into a (row) vector. GEN co8(GEN x, long l): applied to a quadratic number x (type t QUAD), converts x into a real or complex number depending on the sign of the discriminant of x, to precision l BIL-bit words. GEN gcvtop(GEN x, GEN p, long l): converts x into a p-adic number of precision l. GEN gmodulcp(GEN x, GEN y): creates the object Mod(x,y) on the PARI stack, where x and y are either both integers, and the result is an integermod (type t INTMOD), or x is a scalar or a polynomial and y a polynomial, and the result is a polymod (type t POLMOD). GEN gmodulgs(GEN x, long y): same as gmodulcp except y is a long. GEN gmodulss(long x, long y): same as gmodulcp except both x and y are longs. GEN gmodulo(GEN x, GEN y): same as gmodulcp except that the modulus y is copied onto the heap and not onto the PARI stack. long gexpo(GEN x): returns the binary exponent of x or the maximal binary exponent of the coefficients of x. Returns -HIGHEXPOBIT if x has no components or is an exact zero. long gsize(GEN x): returns 0 if x is exactly 0. Otherwise, returns gexpo(x) multiplied by log_{10}(2). This gives a crude estimate for the maximal number of decimal digits of the components of x. long gsigne(GEN x): returns the sign of x (−1, 0 or 1) when x is an integer, real or (irreducible or reducible) fraction. Raises an error for all other types. long gvar(GEN x): returns the main variable of x. If no component of x is a polynomial or power series, this returns BIGINT. int precision(GEN x): If x is of type t REAL, returns the precision of x (the length of x in BIL-bit words if x is not zero, and a reasonable quantity obtained from the exponent of x if x is numerically equal to zero). If x is of type t COMPLEX, returns the minimum of the precisions of the real and imaginary part. Otherwise, returns 0 (which stands in fact for infinite precision).

5.3.2 Comparison operators and valuations int gcmp0(GEN x): returns 1 (true) if x is equal to 0, 0 (false) otherwise. int isexactzero(GEN x): returns 1 (true) if x is exactly equal to 0, 0 (false) otherwise. Note that many PARI functions will return a pointer to gzero when they are aware that the result they return is an exact zero, so it is almost always faster to test for pointer equality first, and call isexactzero (or gcmp0) only when the first test fails. int gcmp1(GEN x): returns 1 (true) if x is equal to 1, 0 (false) otherwise. int gcmp 1(GEN x): returns 1 (true) if x is equal to −1, 0 (false) otherwise. long gcmp(GEN x, GEN y): comparison of x with y (returns the sign of x − y). long gcmpsg(long s, GEN x): comparison of the long s with x. long gcmpgs(GEN x, long s): comparison of x with the long s. long lexcmp(GEN x, GEN y): comparison of x with y for the lexicographic ordering. long gegal(GEN x, GEN y): returns 1 (true) if x is equal to y, 0 otherwise. long gegalsg(long s, GEN x): returns 1 (true) if the long s is equal to x, 0 otherwise. long gegalgs(GEN x, long s): returns 1 (true) if x is equal to the long s, 0 otherwise. long iscomplex(GEN x): returns 1 (true) if x is a complex number (of component types embeddable into the reals) but is not itself real, 0 if x is a real (not necessarily of type t REAL), or raises an error if x is not embeddable into the complex numbers. long ismonome(GEN x): returns 1 (true) if x is a non-zero monomial in its main variable, 0 otherwise. long ggval(GEN x, GEN p): returns the greatest exponent e such that p^{e} divides x, when this makes sense. long gval(GEN x, long v): returns the highest power of the variable number v dividing the polynomial x. int pvaluation(GEN x, GEN p, GEN *r): applied to non-zero integers x and p, returns the highest exponent e such that p^{e} divides x, creates the quotient x/p^{e} and returns its address in *r. In particular, if p is a prime, this returns the valuation at p of x, and *r will obtain the prime-to-p part of x.

5.3.3 Assignment statements void gaffsg(long s, GEN x): assigns the long s into the object x. void gaffect(GEN x, GEN y): assigns the object x into the object y.

5.3.4 Unary operators

GEN gneg[z](GEN x[, GEN z]): yields −x.
GEN gabs[z](GEN x[, GEN z]): yields |x|.

GEN gsqr(GEN x): creates the square of x. GEN ginv(GEN x): creates the inverse of x. GEN gfloor(GEN x): creates the floor of x, i.e. the (true) integral part. GEN gfrac(GEN x): creates the fractional part of x, i.e. x minus the floor of x. GEN gceil(GEN x): creates the ceiling of x. GEN ground(GEN x): rounds the components of x to the nearest integers. Exact half-integers are rounded towards +∞. GEN grndtoi(GEN x, long *e): same as round, but in addition puts minus the number of significant binary bits left after rounding into *e. If *e is positive, all significant bits have been lost. This kind of situation raises an error message in ground but not in grndtoi. GEN gtrunc(GEN x): truncates x. This is the (false) integer part if x is an integer (i.e. the unique integer closest to x among those between 0 and x). If x is a series, it will be truncated to a polynomial; if x is a rational function, this takes the polynomial part. GEN gcvtoi(GEN x, long *e): same as grndtoi except that rounding is replaced by truncation. GEN gred[z](GEN x[, GEN z]): reduces x to lowest terms if x is a fraction or rational function (types t FRAC, t FRACN, t RFRAC and t RFRACN), otherwise creates a copy of x. GEN content(GEN x): creates the GCD of all the components of x. GEN normalize(GEN x): applied to an unnormalized power series x (i.e. type t SER with all coefficients correctly set except that x[2] might be zero), normalizes x correctly in place. Returns x. For internal use. GEN normalizepol(GEN x): applied to an unnormalized polynomial x (i.e. type t POL with all coefficients correctly set except that x[2] might be zero), normalizes x correctly in place and returns x. For internal use.

5.3.5 Binary operators GEN gmax[z](GEN x, GEN y[, GEN z]): yields the maximum of the objects x and y if they can be compared. GEN gmaxsg[z](long s, GEN x[, GEN z]): yields the maximum of the long s and the object x. GEN gmaxgs[z](GEN x, long s[, GEN z]): yields the maximum of the object x and the long s. GEN gmin[z](GEN x, GEN y[, GEN z]): yields the minimum of the objects x and y if they can be compared. GEN gminsg[z](long s, GEN x[, GEN z]): yields the minimum of the long s and the object x. GEN gmings[z](GEN x, long s[, GEN z]): yields the minimum of the object x and the long s. GEN gadd[z](GEN x, GEN y[, GEN z]): yields the sum of the objects x and y. GEN gaddsg[z](long s, GEN x[, GEN z]): yields the sum of the long s and the object x. GEN gaddgs[z](GEN x, long s[, GEN z]): yields the sum of the object x and the long s. GEN gsub[z](GEN x, GEN y[, GEN z]): yields the difference of the objects x and y. GEN gsubgs[z](GEN x, long s[, GEN z]): yields the difference of the object x and the long s. GEN gsubsg[z](long s, GEN x[, GEN z]): yields the difference of the long s and the object x. GEN gmul[z](GEN x, GEN y[, GEN z]): yields the product of the objects x and y. GEN gmulsg[z](long s, GEN x[, GEN z]): yields the product of the long s with the object x. GEN gmulgs[z](GEN x, long s[, GEN z]): yields the product of the object x with the long s. GEN gshift[z](GEN x, long n[, GEN z]): yields the result of shifting (the components of) x left by n (if n is non-negative) or right by −n (if n is negative). Applies only to integers, reals and vectors/matrices of such. For other types, it is simply multiplication by 2^{n}. GEN gmul2n[z](GEN x, long n[, GEN z]): yields the product of x and 2^{n}. This is different from gshift when n is negative and x is of type t INT: gshift truncates, while gmul2n creates a fraction if necessary. GEN gdiv[z](GEN x, GEN y[, GEN z]): yields the quotient of the objects x and y. GEN gdivgs[z](GEN x, long s[, GEN z]): yields the quotient of the object x and the long s. GEN gdivsg[z](long s, GEN x[, GEN z]): yields the quotient of the long s and the object x. GEN gdivent[z](GEN x, GEN y[, GEN z]): yields the true Euclidean quotient of x and the integer or polynomial y. GEN gdiventsg[z](long s, GEN x[, GEN z]): yields the true Euclidean quotient of the long s by the integer x. GEN gdiventgs[z](GEN x, long s[, GEN z]): yields the true Euclidean quotient of the integer x by the long s. GEN gdiventres(GEN x, GEN y): creates a 2-component vertical vector whose components are the true Euclidean quotient and remainder of x and y. GEN gdivmod(GEN x, GEN y, GEN *r): If r is not equal to NULL or ONLY REM, creates the (false) Euclidean quotient of x and y, and puts (the address of) the remainder into *r. If r is equal to NULL, do not create the remainder, and if r is equal to ONLY REM, create and output only the remainder. The remainder is created after the quotient and can be disposed of individually with a

cgiv(r).

GEN poldivres(GEN x, GEN y, GEN *r): same as gdivmod but specifically for polynomials x and y. GEN gdeuc(GEN x, GEN y): creates the Euclidean quotient of the polynomials x and y. GEN gdivround(GEN x, GEN y): if x and y are integers, returns the quotient x/y of x and y, rounded to the nearest integer. If x/y falls exactly halfway between two consecutive integers, then it is rounded towards +∞ (as for round). If x and y are not both integers, the result is the same as that of gdivent. GEN gmod[z](GEN x, GEN y[, GEN z]): yields the true remainder of x modulo the integer or polynomial y. GEN gmodsg[z](long s, GEN x[, GEN z]): yields the true remainder of the long s modulo the integer x. GEN gmodgs[z](GEN x, long s[, GEN z]): yields the true remainder of the integer x modulo the long s. GEN gres(GEN x, GEN y): creates the Euclidean remainder of the polynomial x divided by the polynomial y. GEN ginvmod(GEN x, GEN y): creates the inverse of x modulo y when it exists. GEN gpow(GEN x, GEN y, long l): creates x^{y}. The precision l is taken into account only if y is not an integer and x is an exact object. If y is an integer, binary powering is done. Otherwise, the result is exp(y ∗ log(x)) computed to precision l. GEN ggcd(GEN x, GEN y): creates the GCD of x and y. GEN glcm(GEN x, GEN y): creates the LCM of x and y. GEN subres(GEN x, GEN y): creates the resultant of the polynomials x and y computed using the subresultant algorithm. GEN gpowgs(GEN x, long n): creates x^{n} using binary powering. GEN gsubst(GEN x, long v, GEN y): substitutes the object y into x for the variable number v. int gdivise(GEN x, GEN y): returns 1 (true) if y divides x, 0 otherwise. GEN gbezout(GEN x, GEN y, GEN *u, GEN *v): creates the GCD of x and y, and puts (the adresses of) objects u and v such that ux + vy = gcd(x, y) into *u and *v.

Appendix A: Installation Guide for the UNIX Versions

1. Required tools.

We assume that you have either an ANSI C or a C++ compiler available. If your machine does not have one (for example if you still use /bin/cc in SunOS 4.1.x), we strongly suggest that you obtain the gcc/g++ compiler from the Free Software Foundation or by anonymous ftp. As for all GNU software mentioned afterwards, you can find the most convenient site to fetch gcc at the address

    http://www.gnu.ai.mit.edu/order/ftp.html

You can certainly compile PARI with a different compiler, but the PARI kernel takes advantage of some optimizations provided by gcc if it is available. This results in about 20% speedup on most architectures*.

Important Note: The graphic routines in the present version have only been tested under X11 and gnuplot, and may not work at all if you try to compile GP with an old Suntools library (even though this is supposedly supported, nobody has tested it yet!!!).

1.1. Optional packages: The following programs and libraries are useful in conjunction with GP, but not mandatory. They’re probably already installed somewhere on your system (with the possible exception of readline, which we think is really worth a try). In any case, get them before proceeding if you want the functionalities they provide. All of them are free (though you ought to make a small donation to the FSF if you use (and like) GNU wares).

• GNU readline library. This provides line editing under GP, an automatic context-dependent completion, and an editable history of commands. Note that it is incompatible with SUN commandtools (yet another reason to dump Suntools for X Windows). A recent readline (version number at least 2.2) is preferred, but older versions should be usable.

• GNU gzip/gunzip/gzcat package enables GP to read compressed data.

• GNU emacs. GP can be run in an Emacs buffer, with all the obvious advantages if you are familiar with this editor. Note that readline is still useful in this case since it provides a much better automatic completion than is provided by Emacs GP-mode.

• perl provides extended online help (full text from Chapter 3) about functions and concepts, which can be used under GP or independently (http://www.perl.com will direct you to the nearest CPAN archive site).

• A colour-capable xterm, which enables GP to use different (user configurable) colours for its output. All xterm programs which come with current X11R6.3 distributions will satisfy this requirement. Under X11R6, you can for instance use color xterm (get the latest version at

http://www.clark.net/pub/dickey/xterm).

* One notable exception is the native AIX C compiler on IBM RS/6000 workstations, which generates fast code even without any special help from the PARI kernel sources.

2. Compiling the library and the GP calculator.

2.1. Basic configuration: First, have a look at the MACHINES file to see if anything funny applies to your architecture or operating system. Then, type

    ./Configure

in the toplevel directory. This will attempt to configure GP/PARI without outside help. Note that if you want to install the end product in some nonstandard place, you can use the --prefix option, as in

    ./Configure --prefix=/an/exotic/directory

(the default prefix is /usr/local). This phase extracts some files and creates a directory Oxxx where the object files and executables will be built. The xxx part depends on your architecture and operating system, thus you can build GP for several different machines from the same source tree (the builds are completely independent, so can be done simultaneously).

Configure will let the following environment variable override the defaults if set:

AS: Assembler.

CC: C compiler.

DLLD: Dynamic library linker.

For instance, Configure avoids gcc on some architectures due to various problems which may have been fixed in your version of the compiler. You can try

    env CC=gcc Configure

and compare the benches. Also, if you run into trouble with a recent g++, try to use g++ -

fpermissive.

2.2. Troubleshooting and fine tuning: Decide whether you agree with what Configure printed on your screen (in particular the architecture, compiler and optimization flags). If anything should have been found and was not, consider that Configure failed and follow the instructions below. Look especially for the readline and X11 libraries, and the perl and gunzip (or zcat) binaries.

In case the default Configure run fails miserably, try

    ./Configure -a

(interactive mode) and answer all the questions (there aren’t that many). Of course, Configure will still provide defaults for each answer but if you accept them all, it will fail just the same, so be wary. In any case, we would appreciate a bug report including the complete output from Configure and the file Oxxx/dft.Config.in that was produced in the process.

Note that even in interactive mode, you can’t directly tell Configure where the readline library and include files are. If they are not in a standard place, it won’t find them. Nonetheless, it first searches the distribution toplevel for a readline directory. Thus, if you just want to give readline a try (as you probably should), you can get the source and compile it there (you don’t need to install it). You can also use this feature together with a symbolic link, named readline, in the PARI toplevel directory if you have compiled the readline library somewhere else, without installing it to one of its standard locations.

Technical note: Configure can build GP on different architectures simultaneously from the same toplevel sources. Instead of the readline link alluded above, you can create readline-osname-arch, using the same naming conventions as for the Oxxx directory, e.g readline-linux-i686.

2.3. Debugging/profiling: If you also want to debug the PARI library,

    Configure -g

will create a directory Oxxx.dbg containing a special Makefile ensuring that suitably non-optimized GP and PARI library will be built there. If you want to profile GP or the library (using gprof for instance),

    Configure -pg

will create an Oxxx.prf directory where a suitable version of PARI can be built.

2.4. Compilation and tests: To compile the GP binary, simply type

    make gp

in the distribution directory. If your make program supports parallel make, you can speed up the process by going to the Oxxx directory that Configure created and doing a parallel make here (for instance make -j4 with GNU make).

2.4.1. Testing

To test the binary, type make bench. This will build a static executable (the default, built by make gp is probably dynamic) and run a series of comparative tests on those two. To test only the default binary, use make dobench which starts the bench immediately.

The static binary should be slightly faster. In any case, this should not take more than one minute (user time) on modern machines. See the file MACHINES to get an idea of how much time comparable systems need (we would appreciate a short note in the same format in case your system is not listed and you nevertheless have a working GP executable).

If a [BUG] message shows up, it probably means that something is wrong. Most probably with the installation procedure, but it may be a bug in the Pari system, in which case we would appreciate a report (including the relevant *.dif file in the Oxxx directory and the file dft.Config.in). Error messages of the form “not yet available for this architecture” are an obvious special case which should not trigger a bug report (unless you implement the functionality yourself, that is!).

Note: If when running gp-dyn, you get a message of the form

    ld.so:  warning:  libpari.so.xxx has older revision than expected xxx

(possibly followed by more errors), you already have a dynamic PARI library installed and a broken local configuration. Either remove the old library or unset the LD LIBRARY PATH environment variable. Try to disable this variable in any case if anything very wrong occurs with the gp-dyn binary (e.g Illegal Instruction on startup). It doesn’t affect gp-sta.

2.4.2. Some more testing [Optional]

You can test GP in compatibility mode with make test-compat. If you want to test the graphic routines, use make test-graphic. You will have to click on the mouse button after seeing each image (under X11; under suntools you must kill the images). There will be eight of them, probably shown twice (under X11, try to resize at least one of them as a further test).

The make bench and make test-compat runs produce a Postscript file pari.ps in Oxxx which you can send to a Postscript printer. The output should bear some similarity to the screen images.

3. Installation.

When everything looks fine, type

    make install

(You may have to do this with superuser privileges, depending on the target directories.) Beware that, if you chose the same installation directory as before in the Configure process, this will wipe out any files from version 1.39.15 and below that might already be there. Libraries and executable files from newer versions (starting with version 1.900) are not removed since they are only links to files bearing the version number (beware of that as well: if you’re an avid GP fan, don’t forget to delete the old pari libraries once in a while).

This installs in the directory chosen at Configure time the default GP executable (probably gp-dyn) under the name gp, the default PARI library (probably libpari.so), the necessary include files, the manual pages, the documentation and help scripts and emacs macros.

By default, if a dynamic library libpari.so could be built, the static library libpari.a will not be created. If you want it as well, you can use the target make install-lib-sta. You can install a statically linked gp with the target make install-bin-sta. As a rule, programs linked statically (with libpari.a) may be slightly faster (about 5% gain), but use much more disk space and take more time to compile. They are also harder to upgrade: you will have to recompile them all instead of just installing the new dynamic library. (On the other hand, there’s no risk of breaking them by installing a new pari library)

3.1. The Galois package: The default polgalois function can only compute Galois groups of polynomials of degree less or equal to 7. If you want to handle polynomials of degree bigger than 7 (and less than 11), you need to fetch a separate archive: galdata.tgz which can probably be found at the same place where you got the main PARI archive, and on the megrez ftp server in any case. Untar the archive in the datadir directory which was chosen at Configure time (it’s one of the last messages on the screen if you did not run Configure -a). You can then test the polgalois function with your favourite polynomials.

 3.2. The GPRC file: Copy misc/gprc.dft (or gprc.dos if you’re using GP.EXE) to $HOME/.gprc.

Modify it to your liking. For instance, if you’re not using an ANSI terminal, remove control characters from the prompt variable. You can also enable colors.

If desired, also copy/modify misc/gpalias somewhere and call it from the gprc file (this provides some common shortcuts to lengthy names). Finally, if you have superuser privileges and want to provide system-wide defaults, you can copy your customized .gprc file to /etc/gprc.

In older versions, gphelp was hidden in pari lib directory and wasn’t meant to be used from the shell prompt, but not anymore. If gp complains it can’t find gphelp, check whether your .gprc (or the system-wide gprc) does contain explicit paths. If so, correct them according to the current

misc/gprc.dft.

4. Getting Started.

4.1. Printable Documentation: To print the user’s guide, for which you’ll need a working (plain) TEX installation; type

    make doc

This will create, in two passes, a file doc/users.dvi containing the manual with a table of contents and an index. You must then send the users.dvi file to your favourite printer in the usual way, probably via dvips. Also included are a short tutorial (doc/tutorial.dvi) and a reference card

(doc/refcard.dvi and doc/refcard.ps) for GP.

4.2. C programming: Once all libraries and include files are installed, you can link your C programs to the PARI library. A sample makefile examples/Makefile is provided to illustrate the use of the various libraries. Type make all in the examples directory to see how they perform on the mattrans.c program, which is commented in the manual.

4.3. GP scripts: Several complete sample GP programs are also given in the examples directory, for example Shanks’s SQUFOF factoring method, the Pollard rho factoring method, the Lucas-Lehmer primality test for Mersenne numbers and a simple general class group and fundamental unit algorithm (much worse than the built-in bnfinit!). See the file examples/EXPLAIN for some explanations.

4.4. EMACS: If you want to use gp under GNU Emacs, read the file emacs/pariemacs.txt. If you are familiar with Emacs, we suggest that you do so.

4.5. The PARI Community: There are three mailing lists devoted to the PARI/GP package (run courtesy of Dan Bernstein), and most feedback should be directed to those. They are:

• pari-announce: to announce major version changes. You can’t write to this one, but you should probably subscribe.

• pari-dev: for everything related to the development of PARI, including suggestions, technical questions, bug reports or patch submissions.

• pari-users: for everything else.

To subscribe, send empty messages respectively to

                      pari-announce-subscribe@list.cr.yp.to
                        pari-users-subscribe@list.cr.yp.to
                         pari-dev-subscribe@list.cr.yp.to

If you are not a member of any of those lists and don’t want to become one, you can write to us at the address

                             pari@math.u-bordeaux.fr

At the very least, we will forward you mail to the lists above and correct faulty behaviour, if necessary. But we cannot promise you will get an individual answer.

Last but not least, PARI home page (maintained by Gerhard Nicklasch) can be found at

                http://hasse.mathematik.tu-muenchen.de/ntsw/pari/

In any case, if you like this software, we would appreciate if you could send us an email message giving us some information about yourself and what you use PARI for. Put as header of your message “new user”, so we can recognize it easily.

Good luck and enjoy!

196

Appendix B: A Sample program and Makefile

We assume that you have installed the PARI library and include files as explained in Appendix A or in the installation guide. If you chose differently any of the directory names, change them accordingly in the Makefiles.

If the program example that we have given is in the file matexp.c (say as the first of several matrix transcendental functions), then a sample Makefile might look as follows. Note that the actual file examples/Makefile is much more elaborate and you should have a look at it if you intend to use install() on custom made functions, see Section 3.11.2.13.

CC = cc
INCDIR = ../../../OpenXM/include/pari
LIBDIR = ../../../OpenXM/lib
CFLAGS = -O -I$(INCDIR) -L$(LIBDIR)
all:    matexp
matexp:    matexp.c
  $(CC) $(CFLAGS) -o matexp matexp.c -lpari -lm

We then give the listing of the program examples/matexp.c seen in detail in Section 4.8, with the slight modifications explained at the end of that section.

#include <pari.h>
GEN
matexp(GEN x, long prec)
{
long lx=lg(x),i,k,n, ltop = avma;
GEN y,r,s,p1,p2;

/* check that x is a square matrix */

if (typ(x) != t MAT) err(typeer,"matexp");
if (lx == 1) return cgetg(1, t MAT);
if (lx != lg(x[1])) err(talker,"not a square matrix");

/* convert x to real or complex of real and compute its L2 norm */

s = gzero; r = cgetr(prec+1); gaffsg(1,r); x = gmul(r,x);
for (i=1; i<lx; i++)
  s = gadd(s, gnorml2((GEN)x[i]));
if (typ(s) == t REAL) setlg(s,3);

s = gsqrt(s,3); /* we do not need much precision on s */

/* if s < 1 we are happy */

k = expo(s);
if (k < 0) { n = 0; p1 = x; }
else { n = k+1; p1 = gmul2n(x,-n); setexpo(s,-1); }

/* initializations before the loop */ y = gscalmat(r,lx-1); /* creates scalar matrix with r on diagonal */

p2 = p1; r = s; k = 1;
y = gadd(y,p2);

/* now the main loop */

while (expo(r) >= -BITS IN LONG*(prec-1))
{
  k++; p2 = gdivgs(gmul(p2,p1),k);
  r = gdivgs(gmul(s,r),k); y = gadd(y,p2);
}

/* now square back n times if necessary */

for (i=0; i<n; i++) y = gsqr(y);
return gerepileupto(ltop,y);
}
int
main()
{
long d, prec = 3;
GEN x;

/* take a stack of 10^{6} bytes, no prime table */

pari init(1000000,2);
printf("precision of the computation in decimal digits:\n");
d = itos(lisGEN(stdin));
if (d > 0) prec = (long)(d*pariK1+3);
printf("input your matrix in GP format:\n");
x = matexp(lisGEN(stdin), prec);
sor(x, 'g', d, 0);
exit(0);
}

Appendix C: Summary of Available Constants

In this appendix we give the list of predefined constants available in the PARI library. All of them are in the heap and not on the PARI stack. We start by recalling the universal objects introduced in Section 4.1:

t INT: gzero (zero), gun (un), gdeux (deux)
t FRAC: ghalf (lhalf)
t COMPLEX: gi
t POL: polun[..]  (lpolun[..]), polx[..]  (lpolx[..])

Only polynomials in the variables 0 and MAXVARN are defined initially. Use fetch var() (see Section 4.6.2.2) to create new ones. The other objects are not initialized by default:

geuler. This is Euler’s constant. To initialize it, call consteuler(prec) (see Section 3.3.2). gpi. This is the number π. To initialize it, call constpi(prec) (see Section 3.3.4). bern(i). This is the 2i-th Bernoulli number (B_{0} = 1, B_{2} = 1/6, B_{4} = −1/30, etc…). To

initialize them, use the function:

void mpbern(long n, long prec);

This creates the even numbered Bernoulli numbers up to B_{2n−2} as real numbers of precision prec. They can then be used with the macro bern(i). Note that this is not a function but simply an abbreviation, hence care must be taken that i is inside the right bounds (i.e. 0 ≤ i ≤ n − 1) before using it, since no checking is done by PARI itself.

Finally, one has access to a table of (differences of) primes through the pointer diffptr. This is used as follows: when

void pari init(long size, long maxprime);

is called, this table is initialized with the successive differences of primes up to (just a little beyond) maxprime (see Section 4.1). maxprime has to be less than 436272744, whatever memory is available. A difference of 0 means we have reached the end of the table. The following example illustrates its use.

byteptr d = diffptr;
long p = 0;
while (p <= p goal) /* run through all primes up to p goal */
{
  p += *d;
  if (*d == 0) err(primer1); /* not enough primes */
  d++;
  ...
}

Here, we use the general error handling function err (see Section 4.7.3), with the codeword primer1. This will just print the error message:

    *** not enough precomputed primes

and then abort the computations.

You can use the function initprimes from the file arith2.c to compute a new table on the fly and assign it to diffptr or to a similar variable of your own. Beware that before changing diffptr, you should really free the (malloced) precomputed table first, and then all pointers into the old table will become invalid.

PARI currently guarantees that the first 6547 primes, up to and including 65557, will be present in the table, even if you set maxnum to zero.

In addition, some single or double-precision real numbers are predefined, and their list is in

the file paricom.h.

201

Index

A

Abelian extension . . . . . . . . . . 106, 111 abs . . . . . . . . . . . . . . . . . . . . . . 57 accuracy . . . . . . . . . . . . . . . . . . . . 9 acos . . . . . . . . . . . . . . . . . . . . . 57 acosh . . . . . . . . . . . . . . . . . . . . 57 addell . . . . . . . . . . . . . . . . . . . . 73 addhelp . . . . . . . . . . . . . . 38, 139, 175 addii . . . . . . . . . . . . . . . . . . . . . 147 addir . . . . . . . . . . . . . . . . . . . . . 147 addis . . . . . . . . . . . . . . . . . . . . . 147 addprimes . . . . . . . . . . . . . . . . . . 63 addri . . . . . . . . . . . . . . . . . . . . . 147 addrr . . . . . . . . . . . . . . . . . . . . 147 adj . . . . . . . . . . . . . . . . . . . . . . 120 adjoint matrix . . . . . . . . . . . . . . . 120 agm . . . . . . . . . . . . . . . . . . . . . 57 akell . . . . . . . . . . . . . . . . . . . . . 74 algdep . . . . . . . . . . . . . . . . . . . . 118 algdep0 . . . . . . . . . . . . . . . . . . . 118 algebraic dependence . . . . . . . . . . . . 118 algtobasis . . . . . . . . . . . . . . . . . . 97 alias . . . . . . . . . . . . . . . . . . . 38, 140 allocatemem . . . . . . . . . . . . . . 18, 140 allocatemoremem . . . . . . . . . . . 140, 170 alternating series . . . . . . . . . . . . . . 131 and . . . . . . . . . . . . . . . . . . . . . 48 anell . . . . . . . . . . . . . . . . . . . . . 74 apell . . . . . . . . . . . . . . . . . . . . . 74 apell2 . . . . . . . . . . . . . . . . . . . . 74 apprgen . . . . . . . . . . . . . . . . . . . 114 apprgen9 . . . . . . . . . . . . . . . . . . 114 area . . . . . . . . . . . . . . . . . . . . . 73 arg . . . . . . . . . . . . . . . . . . . . . . 57 Artin L-function . . . . . . . . . . . . . . 91 Artin root number . . . . . . . . . . . . . 91 asin . . . . . . . . . . . . . . . . . . . . . 57 asinh . . . . . . . . . . . . . . . . . . . . . 57 assignment . . . . . . . . . . . . . . . . . 151 assmat . . . . . . . . . . . . . . . . . . . . 120 atan . . . . . . . . . . . . . . . . . . . . . 57 atanh . . . . . . . . . . . . . . . . . . . . 57 automatic simplification . . . . . . . . . . 19 available commands . . . . . . . . . . . . 21 avma . . . . . . . . . . . . . . . . . . 151, 152

B backslash character . . . . . . . . . . . . 29 base . . . . . . . . . . . . . . . . . . . . . 97 base2 . . . . . . . . . . . . . . . . . . . . 97 basistoalg . . . . . . . . . . . . . . . . . . 98 Berlekamp . . . . . . . . . . . . . . . . . . 67 bern . . . . . . . . . . . . . . . . . . . . . 199 bernfrac . . . . . . . . . . . . . . . . . . . 57 Bernoulli numbers . . . . . . . . . 57, 58, 62 bernreal . . . . . . . . . . . . . . . . . . . 58 bernvec . . . . . . . . . . . . . . . . . . . 58 besseljh . . . . . . . . . . . . . . . . . . . 58 besselk . . . . . . . . . . . . . . . . . . . . 58 bestappr . . . . . . . . . . . . . . . . . . . 63 bezout . . . . . . . . . . . . . . . . . . . . 63 bezoutres . . . . . . . . . . . . . . . . . . 63 BIGDEFAULTPREC . . . . . . . . . . . 148 BIGINT . . . . . . . . . . . . . . . . 162, 163 bigomega . . . . . . . . . . . . . . . . . . 63 bilhell . . . . . . . . . . . . . . . . . . . . 74 binaire . . . . . . . . . . . . . . . . . . . . 50 binary quadratic form . . . . . . . . . . 7, 62 binary quadratic forms . . . . . . . . . . 25 binary . . . . . . . . . . . . . . . . . . . . 50 binome . . . . . . . . . . . . . . . . . . . 63 binomial coefficient . . . . . . . . . . . . . 63 binomial . . . . . . . . . . . . . . . . . . . 63 Birch and Swinnerton-Dyer conjecture . . 75 BITS IN LONG . . . . . . . . . . . . . . 148 bittest . . . . . . . . . . . . . . . . . . . . 50 bitwise and . . . . . . . . . . . . . . . . . 48 bitwise or . . . . . . . . . . . . . . . . . . 48 bit accuracy . . . . . . . . . . . . . . . . . 148 bnf . . . . . . . . . . . . . . . . . . 35, 79, 81 bnfcertify . . . . . . . . . . . . . . . . . . 82 bnfclassgrouponly . . . . . . . . . . . . . 84 bnfclassunit . . . . . . . . . . . . . . . . . 83 bnfclassunit0 . . . . . . . . . . . . . . . . 84 bnfclgp . . . . . . . . . . . . . . . . . . . 84 bnfdecodemodule . . . . . . . . . . . . . . 84 bnfinit . . . . . . . . . . . . . . . . . . 79, 84 bnfinit0 . . . . . . . . . . . . . . . . . . . 85 bnfisintnorm . . . . . . . . . . . . . . . . 85 bnfisnorm . . . . . . . . . . . . . . . . . . 85 bnfisprincipal . . . . . . . . . . . . . . . . 86 bnfissunit . . . . . . . . . . . . . . . . . . 86 bnfisunit . . . . . . . . . . . . . . . . . . . 86 bnfmake . . . . . . . . . . . . . . . . . . . 86 bnfnarrow . . . . . . . . . . . . . . . . 70, 87 bnfreg . . . . . . . . . . . . . . . . . . . . 87 bnfsignunit . . . . . . . . . . . . . . . . . 87 bnfsunit . . . . . . . . . . . . . . . . . . . 87 bnfunit . . . . . . . . . . . . . . . . . . . 87 bnr . . . . . . . . . . . . . . . . . . . . 35, 79 bnrclass . . . . . . . . . . . . . . . . . . . 88 bnrclass0 . . . . . . . . . . . . . . . . . . 88 bnrclassno . . . . . . . . . . . . . . . . . . 89 bnrclassnolist . . . . . . . . . . . . . . . . 89 bnrconductor . . . . . . . . . . . . . . . . 89 bnrconductorofchar . . . . . . . . . . . . . 89 bnrdisc . . . . . . . . . . . . . . . . . . . 89 bnrdisc0 . . . . . . . . . . . . . . . . . . . 89 bnrdisclist . . . . . . . . . . . . . . . . . . 89 bnrdisclist0 . . . . . . . . . . . . . . . . . 90 bnrinit . . . . . . . . . . . . . . . . . . 70, 90 bnrinit0 . . . . . . . . . . . . . . . . . . . 90 bnrisconductor . . . . . . . . . . . . . . . 90 bnrisprincipal . . . . . . . . . . . . . . . . 90 bnrL1 . . . . . . . . . . . . . . . . . . . . 88 bnrrootnumber . . . . . . . . . . . . . . . 91 bnrstark . . . . . . . . . . . . . . . . . 71, 91 boolean operators . . . . . . . . . . . . . 48 brace characters . . . . . . . . . . . . . . 30 break . . . . . . . . . . . . . . . . . . . . 137 brute . . . . . . . . . . . . . . . . . . 166, 168 buchfu . . . . . . . . . . . . . . . . . . . . 87 buchimag . . . . . . . . . . . . . . . . . . 71 Buchmann . . . . . . . . . . . . . 81, 83, 113 Buchmann-McCurley . . . . . . . . . . . . 70 buchnarrow . . . . . . . . . . . . . . . . . 87 buchreal . . . . . . . . . . . . . . . . . . . 71 buffersize . . . . . . . . . . . . . . . . . . 15

C

Cantor-Zassenhaus . . . . . . . . . . . . . 66 caract . . . . . . . . . . . . . . . . . . . . 118 caradj . . . . . . . . . . . . . . . . . . . . 118 carhess . . . . . . . . . . . . . . . . . . . 118 case distinction . . . . . . . . . . . . . . . 29 ceil . . . . . . . . . . . . . . . . . . . . . . 51 centerlift . . . . . . . . . . . . . . . . . . . 51 centerlift0 . . . . . . . . . . . . . . . . . . 51 certifybuchall . . . . . . . . . . . . . . . . 82 cgetg . . . . . . . . . . . . . . . 149, 150, 158 cgeti . . . . . . . . . . . . . . . . . . . . . 150 cgetr . . . . . . . . . . . . . . . . . . . . . 150 cgiv . . . . . . . . . . . . . . . . . . . . . 153 changevar . . . . . . . . . . . . . . . . 31, 51 character string . . . . . . . . . . . . . 26, 163 character . . . . . . . . . . . . 79, 88, 89, 91 characteristic polynomial . . . . . . . . . 118 charpoly . . . . . . . . . . . . . . . . . . . 118 charpoly0 . . . . . . . . . . . . . . . . . . 118 Chebyshev . . . . . . . . . . . . . . . . . 116 chinese . . . . . . . . . . . . . . . . . . . . 64 chinois . . . . . . . . . . . . . . . . . . . . 64 classno . . . . . . . . . . . . . . . . . . . . 69 classno2 . . . . . . . . . . . . . . . . . . . 69 clgp . . . . . . . . . . . . . . . . . . . . . 81 CLISP . . . . . . . . . . . . . . . . . . . . 38 clone . . . . . . . . . . . . . . . . . . 146, 152 code words . . . . . . . . . . . . . . . . . 51 codiff . . . . . . . . . . . . . . . . . . . . 81 coeff . . . . . . . . . . . . . . . . . . . 52, 147 colors . . . . . . . . . . . . . . . . . . . . 15 column vector . . . . . . . . . . . 7, 25, 162 comparison operators . . . . . . . . . . . 48 compatible . . . . . . . . . . . . . . . . . 16 completion . . . . . . . . . . . . . . . . . 41 complex number . . . . . . . . . 7, 8, 23, 161 compo . . . . . . . . . . . . . . . . . . . . 51 component . . . . . . . . . . . . . . . . . 51 components . . . . . . . . . . . . . . . . . 51 composition . . . . . . . . . . . . . . . . . 69 compraw . . . . . . . . . . . . . . . . . . 69 compress . . . . . . . . . . . . . . . . . . 22 concat . . . . . . . . . . . . . . . 36, 118, 119 conj . . . . . . . . . . . . . . . . . . . . . 52 conjvec . . . . . . . . . . . . . . . . . . . 52 consteuler . . . . . . . . . . . . . . . . 56, 199 constpi . . . . . . . . . . . . . . . . . . 57, 199 content . . . . . . . . . . . . . . . . . . . 64 contfrac . . . . . . . . . . . . . . . . . . . 64 contfrac0 . . . . . . . . . . . . . . . . . . 64 contfracpnqn . . . . . . . . . . . . . . . . 64 continued fraction . . . . . . . . . . . . . 64 Control statements . . . . . . . . . . . . . 137 conversions . . . . . . . . . . . . . . . . . 152 convol . . . . . . . . . . . . . . . . . . . . 116 coordch . . . . . . . . . . . . . . . . . . . 74 copy . . . . . . . . . . . . . . . . . . . . . 151 core . . . . . . . . . . . . . . . . . . . . . 64 core0 . . . . . . . . . . . . . . . . . . . . . 64 core2 . . . . . . . . . . . . . . . . . . . . . 64 coredisc . . . . . . . . . . . . . . . . . . . 64 coredisc0 . . . . . . . . . . . . . . . . . . 64 coredisc2 . . . . . . . . . . . . . . . . . . 64 cos . . . . . . . . . . . . . . . . . . . . . . 58 cosh . . . . . . . . . . . . . . . . . . . . . 58 cotan . . . . . . . . . . . . . . . . . . . . 58 CPU time . . . . . . . . . . . . . . . . . . 19 creation . . . . . . . . . . . . . . . . . . . 149 cyc . . . . . . . . . . . . . . . . . . . . . . 81 cyclo . . . . . . . . . . . . . . . . . . . . . 114

D

dbltor . . . . . . . . . . . . . . . . . . . . 152 debug . . . . . . . . . . . . . . . . . . 16, 168 debugfiles . . . . . . . . . . . . . . . . . . 17 debugging . . . . . . . . . . . . . . . . . . 168 debuglevel . . . . . . . . . . . . . . . . . . 66 DEBUGLEVEL . . . . . . . . . . . . . . 168 debugmem . . . . . . . . . . . . . . . . . 17 DEBUGMEM . . . . . . . . . . . . . . . . 168 debugmem . . . . . . . . . . . . . . . . . 168 decodemodule . . . . . . . . . . . . . . . . 84 decomposition into squares . . . . . . . . 124 Dedekind . . . . . . . . 59, 91, 106, 112, 113 default precision . . . . . . . . . . . . . . . 9 default . . . . . . . . . . . . . . . . . . 38, 140 DEFAULTPREC . . . . . . . . . . . . . . 148 defaults . . . . . . . . . . . . . . . . . 14, 21 definite binary quadratic form . . . . . . 162 degree . . . . . . . . . . . . . . . . . 114, 162 delete var . . . . . . . . . . . . . . . . . . 164 denom . . . . . . . . . . . . . . . . . . . . 52 denominator . . . . . . . . . . . . . . . . 52 deriv . . . . . . . . . . . . . . . . . . . . . 113 destruction . . . . . . . . . . . . . . . . . 153 det . . . . . . . . . . . . . . . . . . . . . . 120 det2 . . . . . . . . . . . . . . . . . . . . . 120 detint . . . . . . . . . . . . . . . . . . . . 120 diagonal . . . . . . . . . . . . . . . . . . . 120 diff . . . . . . . . . . . . . . . . . . . . . . 81 difference . . . . . . . . . . . . . . . . . . 46 diffptr . . . . . . . . . . . . . . . . . 146, 199 dilog . . . . . . . . . . . . . . . . . . . . . 58 dirdiv . . . . . . . . . . . . . . . . . . . . 65 direuler . . . . . . . . . . . . . . . . . . . 65 Dirichlet series . . . . . . . . . . . . . 65, 91 dirmul . . . . . . . . . . . . . . . . . . . . 65 dirzetak . . . . . . . . . . . . . . . . . . . 91 disc . . . . . . . . . . . . . . . . . . . 73, 81 discf . . . . . . . . . . . . . . . . . . . . . 98 discsr . . . . . . . . . . . . . . . . . . . . 114 divisors . . . . . . . . . . . . . . . . . . . 65 divrem . . . . . . . . . . . . . . . . . . . . 47 divsum . . . . . . . . . . . . . . . . . . . 131 dvi . . . . . . . . . . . . . . . . . . . . . . 43

E

echo . . . . . . . . . . . . . . . . . . . . . 17 ECM . . . . . . . . . . . . . . . . . . . 62, 66 editing characters . . . . . . . . . . . . . 29 effective length . . . . . . . . . . . . 160, 161 eigen . . . . . . . . . . . . . . . . . . . . . 120 eint1 . . . . . . . . . . . . . . . . . . . . . 59 element div . . . . . . . . . . . . . . . . . 98 element divmodpr . . . . . . . . . . . . . 98 element mul . . . . . . . . . . . . . . . . . 99 element mulmodpr . . . . . . . . . . . . . 99 element pow . . . . . . . . . . . . . . . . 99 element powmodpr . . . . . . . . . . . . . 99 element reduce . . . . . . . . . . . . . . . 99 element val . . . . . . . . . . . . . . . . . 99 ell . . . . . . . . . . . . . . . . . . . . . . 35 elladd . . . . . . . . . . . . . . . . . . . . 73 ellak . . . . . . . . . . . . . . . . . . . . . 74 ellan . . . . . . . . . . . . . . . . . . . . . 74 ellap . . . . . . . . . . . . . . . . . . . . . 74 ellap0 . . . . . . . . . . . . . . . . . . . . 74 ellbil . . . . . . . . . . . . . . . . . . . . . 74 ellchangecurve . . . . . . . . . . . . . . . 74 ellchangepoint . . . . . . . . . . . . . . . 74 elleisnum . . . . . . . . . . . . . . . . . . 75 elleta . . . . . . . . . . . . . . . . . . . . . 75 ellglobalred . . . . . . . . . . . . . . . . . 75 ellheight . . . . . . . . . . . . . . . . . . . 75 ellheight0 . . . . . . . . . . . . . . . . . . 75 ellheightmatrix . . . . . . . . . . . . . . . 75 ellinit . . . . . . . . . . . . . . . . . . 73, 75 ellinit0 . . . . . . . . . . . . . . . . . . . . 76 ellisoncurve . . . . . . . . . . . . . . . . . 76 ellj . . . . . . . . . . . . . . . . . . . . . . 76 elllocalred . . . . . . . . . . . . . . . . . . 76 elllseries . . . . . . . . . . . . . . . . . . . 77 ellorder . . . . . . . . . . . . . . . . . . . 77 ellordinate . . . . . . . . . . . . . . . . . . 77 ellpointtoz . . . . . . . . . . . . . . . . . . 77 ellpow . . . . . . . . . . . . . . . . . . . . 77 ellrootno . . . . . . . . . . . . . . . . . . . 77 ellsigma . . . . . . . . . . . . . . . . . . . 77 ellsub . . . . . . . . . . . . . . . . . . . . 78 elltaniyama . . . . . . . . . . . . . . . . . 78 elltors . . . . . . . . . . . . . . . . . . . . 78 elltors0 . . . . . . . . . . . . . . . . . . . 78 ellwp . . . . . . . . . . . . . . . . . . . . . 78 ellwp0 . . . . . . . . . . . . . . . . . . . . 78 ellzeta . . . . . . . . . . . . . . . . . . . . 78 ellztopoint . . . . . . . . . . . . . . . . . . 79 Emacs . . . . . . . . . . . . . . . . . . . . 40 EMX . . . . . . . . . . . . . . . . . . . 6, 13 entree . . . . . . . . . . . . . . . . . . 35, 175 environment expansion . . . . . . . . . 15, 50 environment variable . . . . . . . . . . . . 50 erfc . . . . . . . . . . . . . . . . . . . . . 59 err . . . . . . . . . . . . . . . . 167, 168, 170 errfile . . . . . . . . . . . . . . . . . . . . 167 error . . . . . . . . . . . . . . . . 38, 140, 167 eta . . . . . . . . . . . . . . . . . . . . 59, 73 Euclid . . . . . . . . . . . . . . . . . . . . 67 Euclidean quotient . . . . . . . . . . . . . 46 Euclidean remainder . . . . . . . . . . . . 47 Euler product . . . . . . . . . . . . 65, 69, 130 Euler totient function . . . . . . . . . 62, 65 Euler . . . . . . . . . . . . . . 29, 30, 56, 131 Euler-Maclaurin . . . . . . . . . . . . . . 62 eulerphi . . . . . . . . . . . . . . . . . 63, 65 eval . . . . . . . . . . . . . . . . . . . 50, 113 exact object . . . . . . . . . . . . . . . . . . 8 exp . . . . . . . . . . . . . . . . . . . . . . 59 expo . . . . . . . . . . . . . . . . . . 160, 162 expression sequence . . . . . . . . . . . . 32 expression . . . . . . . . . . . . . . . . . . 32 extern . . . . . . . . . . . . . . . . 19, 38, 140 extract . . . . . . . . . . . . . . . . . . . . 128

F

factcantor . . . . . . . . . . . . . . . . . . 66 factmod . . . . . . . . . . . . . . . . . . . 67 factmod9 . . . . . . . . . . . . . . . . . . 66 factor . . . . . . . . . . . . . . . . . . 65, 66 factor0 . . . . . . . . . . . . . . . . . . . . 66 factorback . . . . . . . . . . . . . . . . . . 66 factorcantor . . . . . . . . . . . . . . . . . 66 factoredbase . . . . . . . . . . . . . . . . . 97 factoredpolred . . . . . . . . . . . . . . . 105 factorff . . . . . . . . . . . . . . . . . . . . 66 factorial . . . . . . . . . . . . . . . . . . . 66 factorint . . . . . . . . . . . . . . . . . 65, 66 factormod . . . . . . . . . . . . . . . . . . 67 factornf . . . . . . . . . . . . . . . . . 66, 92 factorpadic . . . . . . . . . . . . . . . . . 113 factorpadic4 . . . . . . . . . . . . . . . . . 113 fetch user var . . . . . . . . . . . . . . . . 164 fetch var . . . . . . . . . . . . . . . . . . . 164 ffinit . . . . . . . . . . . . . . . . . . . . . 92 fibo . . . . . . . . . . . . . . . . . . . . . 67 fibonacci . . . . . . . . . . . . . . . . . . . 67 field discriminant . . . . . . . . . . . . . . 98 filename . . . . . . . . . . . . . . . . . . . 15 filter . . . . . . . . . . . . . . . . . . . . . 165 fincke pohst . . . . . . . . . . . . . . . . . 126 finite field . . . . . . . . . . . . . . . . . . 24 fixed floating point format . . . . . . . . 17 flisexpr . . . . . . . . . . . . . . . . . . . 165 flisseq . . . . . . . . . . . . . . . . . . . . 165 flissexpr . . . . . . . . . . . . . . . . . . . 164 floor . . . . . . . . . . . . . . . . . . . . . 52 foo . . . . . . . . . . . . . . . . . . . . . . 45 for . . . . . . . . . . . . . . . . . . . . . . 137 forcecopy . . . . . . . . . . . . . . . 151, 152 Ford . . . . . . . . . . . . . . . . . . . 97, 113 fordiv . . . . . . . . . . . . . . . . . . . . 137 formal integration . . . . . . . . . . . . . 114 format . . . . . . . . . . . . . . . 17, 166, 168 forprime . . . . . . . . . . . . . . . . . . . 138 forstep . . . . . . . . . . . . . . . . . . . . 138 forsubgroup . . . . . . . . . . . . . . . . . 138 forvec . . . . . . . . . . . . . . . . . . . . 138 fprintferr . . . . . . . . . . . . . . . 168, 175 frac . . . . . . . . . . . . . . . . . . . . . 52 FreeBSD . . . . . . . . . . . . . . . . . . . 13 fu . . . . . . . . . . . . . . . . . . . . . . 81 fundamental units . . . . . . . . . 72, 81, 83 fundunit . . . . . . . . . . . . . . . . . . . 72 futu . . . . . . . . . . . . . . . . . . . . . 81

G

gabs . . . . . . . . . . . . . . . . . . . . . 57 gach . . . . . . . . . . . . . . . . . . . . . 57 gacos . . . . . . . . . . . . . . . . . . . . 57 gadd . . . . . . . . . . . . . . . . . . . . . 46 gaddgs . . . . . . . . . . . . . . . . . . . . 147 gaddgsz . . . . . . . . . . . . . . . . . . . 147 gaddsg . . . . . . . . . . . . . . . . . . . . 147 gaddsgz . . . . . . . . . . . . . . . . . . . 147 gaddz . . . . . . . . . . . . . . 147, 151, 158 gaffect . . . . . . . . . . . . . . . . . . . . 151 gaffsg . . . . . . . . . . . . . . . . . . . . 151 Galois . . . . . . . . . . . . . . . 85, 100, 104 galois . . . . . . . . . . . . . . . . . . . . 105 Galois . . . . . . . . . . . . . . . . . . . . 110 galoisapply . . . . . . . . . . . . . . . . . 100 galoisconj . . . . . . . . . . . . . . . . . . 101 galoisconj0 . . . . . . . . . . . . . . . . . 100 galoisconj2 . . . . . . . . . . . . . . . . . 101 galoisconj4 . . . . . . . . . . . . . . . . . 101 gamma . . . . . . . . . . . . . . . . . . . 59 gammah . . . . . . . . . . . . . . . . . . . 59 gand . . . . . . . . . . . . . . . . . . . . . 48 garbage collecting . . . . . . . . . . . . . 152 garg . . . . . . . . . . . . . . . . . . . . . 57 gash . . . . . . . . . . . . . . . . . . . . . 57 gasin . . . . . . . . . . . . . . . . . . . . . 57 gatan . . . . . . . . . . . . . . . . . . . . 57 gath . . . . . . . . . . . . . . . . . . . . . 57 gauss . . . . . . . . . . . . . . . . . . . . 124 gaussmodulo . . . . . . . . . . . . . . . . 124 gaussmodulo2 . . . . . . . . . . . . . . . . 124 gbezout . . . . . . . . . . . . . . . . . . . 63 gboundcf . . . . . . . . . . . . . . . . . . 64 gcarrecomplet . . . . . . . . . . . . . . . . 68 gcarreparfait . . . . . . . . . . . . . . . . 68 gcd . . . . . . . . . . . . . . . . . . . . . . 67 gcd0 . . . . . . . . . . . . . . . . . . . . . 67 gceil . . . . . . . . . . . . . . . . . . . . . 51 gcf . . . . . . . . . . . . . . . . . . . . . . 64 gcf2 . . . . . . . . . . . . . . . . . . . . . 64 gch . . . . . . . . . . . . . . . . . . . . . . 58 gclone . . . . . . . . . . . . . . . . . . . . 152 gcmp . . . . . . . . . . . . . . . . . . . . . 48 gcmp0 . . . . . . . . . . . . . . . . . . . . 48 gcmp1 . . . . . . . . . . . . . . . . . . . . 48 gcmp 1 . . . . . . . . . . . . . . . . . . . 48 gconj . . . . . . . . . . . . . . . . . . . . . 52 gcopy . . . . . . . . . . . . . . . . . 151, 152 gcos . . . . . . . . . . . . . . . . . . . . . 58 gcotan . . . . . . . . . . . . . . . . . . . . 58 gcvtoi . . . . . . . . . . . . . . . . . . . . 55 gdeux . . . . . . . . . . . . . . . . . . . . 146 gdiv . . . . . . . . . . . . . . . . . . . . . 46 gdivent . . . . . . . . . . . . . . . . . . . 46 gdiventres . . . . . . . . . . . . . . . . . . 47 gdivround . . . . . . . . . . . . . . . . . . 47 gegal . . . . . . . . . . . . . . . . . . . . . 48 gen (member function) . . . . . . . . . . 81 GEN . . . . . . . . . . . . . . . . . 7, 145, 168 gener . . . . . . . . . . . . . . . . . . . . . 72 generic matrix . . . . . . . . . . . . . . . 37 genrand . . . . . . . . . . . . . . . . . . . 54 GENtostr . . . . . . . . . . . . . 50, 166, 167 geq . . . . . . . . . . . . . . . . . . . . . . 48 gerepile . . . . . . . . . . . . . 151, 153, 172 gerepilemany . . . . . . . . . . . . . 155, 157 gerepilemanysp . . . . . . . . . . . . . . . 155 gerepileupto . . . . . . . . . . . . . . 154, 172 getheap . . . . . . . . . . . . . . . . . . . 140 getrand . . . . . . . . . . . . . . . . . . . 140 getstack . . . . . . . . . . . . . . . . . . . 140 gettime . . . . . . . . . . . . . . . . . . . 141 geuler . . . . . . . . . . . . . . . . . . 56, 199 geval . . . . . . . . . . . . . . . . . . . . . 113 gexp . . . . . . . . . . . . . . . . . . . 59, 168 gexpo . . . . . . . . . . . . . . . . . . . . 160 gfloor . . . . . . . . . . . . . . . . . . . . 52 gfrac . . . . . . . . . . . . . . . . . . . . . 52 ggamd . . . . . . . . . . . . . . . . . . . . 59 ggamma . . . . . . . . . . . . . . . . . . . 59 ggcd . . . . . . . . . . . . . . . . . . . . . 67 gge . . . . . . . . . . . . . . . . . . . . . . 48 ggprecision . . . . . . . . . . . . . . . . . 53 ggrandocp . . . . . . . . . . . . . . . . . . 113 ggt . . . . . . . . . . . . . . . . . . . . . . 48 ggval . . . . . . . . . . . . . . . . . . . . . 55 ghalf . . . . . . . . . . . . . . . . . . . . . 146 ghell . . . . . . . . . . . . . . . . . . . . . 75 ghell2 . . . . . . . . . . . . . . . . . . . . 75 gi . . . . . . . . . . . . . . . . . . . . . . . 146 gimag . . . . . . . . . . . . . . . . . . . . 52 gisfundamental . . . . . . . . . . . . . . . 67 gisirreducible . . . . . . . . . . . . . . . . 115 gisprime . . . . . . . . . . . . . . . . . . . 67 gispsp . . . . . . . . . . . . . . . . . . . . 67 gissquarefree . . . . . . . . . . . . . . . . 68 glambdak . . . . . . . . . . . . . . . . . . 112 glcm . . . . . . . . . . . . . . . . . . . . . 68 gle . . . . . . . . . . . . . . . . . . . . . . 48 glength . . . . . . . . . . . . . . . . . . . 52 glngamma . . . . . . . . . . . . . . . . . . 60 global . . . . . . . . . . . . . . . . 31, 34, 141 globalreduction . . . . . . . . . . . . . . . 75 glog . . . . . . . . . . . . . . . . . . . . . 60 glogagm . . . . . . . . . . . . . . . . . . . 60 glt . . . . . . . . . . . . . . . . . . . . . . 48 gmax . . . . . . . . . . . . . . . . . . . . 48 gmin . . . . . . . . . . . . . . . . . . . . . 48 gmod . . . . . . . . . . . . . . . . . . . . 47 gmodulcp . . . . . . . . . . . . . . . . . . 49 gmodulo . . . . . . . . . . . . . . . . . . . 49 gmul . . . . . . . . . . . . . . . . . . . . . 46 gmul2n . . . . . . . . . . . . . . . . . 47, 171 gne . . . . . . . . . . . . . . . . . . . . . . 48 gneg . . . . . . . . . . . . . . . . . . . . . 46 gnil . . . . . . . . . . . . . . . . . . . . . 31 gnorm . . . . . . . . . . . . . . . . . . . . 53 gnorml2 . . . . . . . . . . . . . . . . . 53, 170 gnot . . . . . . . . . . . . . . . . . . . . . 48 gor . . . . . . . . . . . . . . . . . . . . . . 48 GP . . . . . . . . . . . . . . . . . . . . . . 13 gphelp . . . . . . . . . . . . . . . . . . . . 21 gpi . . . . . . . . . . . . . . . . . . . . 57, 199 gpolvar . . . . . . . . . . . . . . . . . . . 55 gpow . . . . . . . . . . . . . . . . . . . 47, 56 gprc . . . . . . . . . . . . . . . . . . . . . 15 GPRC . . . . . . . . . . . . . . . . . . . . 39 gprc . . . . . . . . . . . . . . . . . . . . . 39 gprec . . . . . . . . . . . . . . . . . . . . . 53 gpsi . . . . . . . . . . . . . . . . . . . . . 61 greal . . . . . . . . . . . . . . . . . . . . . 54 gred . . . . . . . . . . . . . . . . . . 161, 162 GRH . . . . . . 70, 71, 81, 82, 83, 85, 111, 117 grndtoi . . . . . . . . . . . . . . . . . . . 54 ground . . . . . . . . . . . . . . . . . . . . 54 gscalmat . . . . . . . . . . . . . . . . 121, 172 gscalsmat . . . . . . . . . . . . . . . 121, 172 gsh . . . . . . . . . . . . . . . . . . . . . . 61 gshift . . . . . . . . . . . . . . . . . . 47, 171 gsigne . . . . . . . . . . . . . . . . . . 48, 160 gsin . . . . . . . . . . . . . . . . . . . . . 61 gsize . . . . . . . . . . . . . . . . . . . . . 54 gsqr . . . . . . . . . . . . . . . . . . . 46, 61 gsqrt . . . . . . . . . . . . . . . . . . . . . 61 GSTR . . . . . . . . . . . . . . . . . . . . 163 gsub . . . . . . . . . . . . . . . . . . . . . 46 gsubst . . . . . . . . . . . . . . . . . . . . 117 gsumdivk . . . . . . . . . . . . . . . . . . 72 gtan . . . . . . . . . . . . . . . . . . . . . 61 gth . . . . . . . . . . . . . . . . . . . . . . 61 gtodouble . . . . . . . . . . . . . . . . . . 152 gtolong . . . . . . . . . . . . . . . . . . . 152 gtomat . . . . . . . . . . . . . . . . . . . . 49 gtopoly . . . . . . . . . . . . . . . . . . . 49 gtopolyrev . . . . . . . . . . . . . . . . . . 49 gtoser . . . . . . . . . . . . . . . . . . . . 50 gtoset . . . . . . . . . . . . . . . . . . . . 50 gtovec . . . . . . . . . . . . . . . . . . . . 50 gtrace . . . . . . . . . . . . . . . . . . . . 127 gtrans . . . . . . . . . . . . . . . . . . . . 124 gtrunc . . . . . . . . . . . . . . . . . . . . 55 gun . . . . . . . . . . . . . . . . . . . . . 146 gvar . . . . . . . . . . . . . . . . . . 162, 163 gzero . . . . . . . . . . . . . . . . . . 146, 171 gzeta . . . . . . . . . . . . . . . . . . . . . 62 gzetak . . . . . . . . . . . . . . . . . . . . 112 gzip . . . . . . . . . . . . . . . . . . . . . 22

H

Hadamard product . . . . . . . . . . . . . 116 hashing function . . . . . . . . . . . . . . 35 hashtable . . . . . . . . . . . . . . . . . . 35 hclassno . . . . . . . . . . . . . . . . . . . 69 heap . . . . . . . . . . . . . . . . 22, 146, 199 hell . . . . . . . . . . . . . . . . . . . . . . 75 help . . . . . . . . . . . . . . . . . . . . . 17 Hermite normal form . 66, 79, 80, 93, 95, 101,

103, 108, 112, 121, 138

hess . . . . . . . . . . . . . . . . . . . . . 120 hexadecimal tree . . . . . . . . . . . . . . 166 hil . . . . . . . . . . . . . . . . . . . . . . 67 Hilbert class field . . . . . . . . . . . . . 71 Hilbert matrix . . . . . . . . . . . . . . . 120 Hilbert symbol . . . . . . . . . . . . . 67, 101 hilbert . . . . . . . . . . . . . . . . . . . . 67 histsize . . . . . . . . . . . . . . . . . . . 17 hnf . . . . . . . . . . . . . . . . . . . . . . 121 hnfall . . . . . . . . . . . . . . . . . . . . 121 hnfmod . . . . . . . . . . . . . . . . . . . 121 hnfmodid . . . . . . . . . . . . . . . . . . 121 hqfeval . . . . . . . . . . . . . . . . . . . . 113 Hurwitz class number . . . . . . . . . . . 69 hyperu . . . . . . . . . . . . . . . . . . . . 59

I

I . . . . . . . . . . . . . . . . . . . 23, 29, 56 ideal list . . . . . . . . . . . . . . . . . . . 80 ideal . . . . . . . . . . . . . . . . . . . . . 79 idealadd . . . . . . . . . . . . . . . . . . . 92 idealaddtoone . . . . . . . . . . . . . . . . 92 idealaddtoone0 . . . . . . . . . . . . . . . 92 idealappr . . . . . . . . . . . . . . . . . . 92 idealappr0 . . . . . . . . . . . . . . . . . . 92 idealchinese . . . . . . . . . . . . . . . . . 92 idealcoprime . . . . . . . . . . . . . . . . 93 idealdiv . . . . . . . . . . . . . . . . . . . 93 idealdiv0 . . . . . . . . . . . . . . . . . . 93 idealdivexact . . . . . . . . . . . . . . . . 93 idealfactor . . . . . . . . . . . . . . . . . . 93 idealhermite . . . . . . . . . . . . . . . . . 93 idealhnf . . . . . . . . . . . . . . . . . . . 93 idealhnf0 . . . . . . . . . . . . . . . . . . 93 idealintersect . . . . . . . . . . . . . . 93, 122 idealinv . . . . . . . . . . . . . . . . . 93, 102 idealinv0 . . . . . . . . . . . . . . . . . . . 93 ideallist . . . . . . . . . . . . . . . . . . . 94 ideallist0 . . . . . . . . . . . . . . . . . . . 94 ideallistarch . . . . . . . . . . . . . . . . . 94 ideallistarch0 . . . . . . . . . . . . . . . . 94 ideallllred . . . . . . . . . . . . . . . . . . 95 ideallog . . . . . . . . . . . . . . . . . . . 94 idealmin . . . . . . . . . . . . . . . . . . . 94 idealmul . . . . . . . . . . . . . . . . . . . 94 idealmulred . . . . . . . . . . . . . . . . . 94 idealnorm . . . . . . . . . . . . . . . . . . 95 idealpow . . . . . . . . . . . . . . . . . . . 95 idealpowred . . . . . . . . . . . . . . . . . 95 idealpows . . . . . . . . . . . . . . . . . . 95 idealprimedec . . . . . . . . . . . . . . . . 95 idealprincipal . . . . . . . . . . . . . . . . 95 idealred . . . . . . . . . . . . . . . . . . . 95 idealstar . . . . . . . . . . . . . . . . . . . 96 idealstar0 . . . . . . . . . . . . . . . . . . 96 idealtwoelt . . . . . . . . . . . . . . . . . 96 idealval . . . . . . . . . . . . . . . . . . . 96 ideal two elt0 . . . . . . . . . . . . . . . . 96 idele . . . . . . . . . . . . . . . . . . . . . 79 ideleprincipal . . . . . . . . . . . . . . . . 96 idmat . . . . . . . . . . . . . . . . . . . . 121 if . . . . . . . . . . . . . . . . . . . . . . . 139 imag . . . . . . . . . . . . . . . . . . . . . 52 image . . . . . . . . . . . . . . . . . . . . 121 imagecompl . . . . . . . . . . . . . . . . . 122 imprecise object . . . . . . . . . . . . . . . 8 incgam . . . . . . . . . . . . . . . . . . . . 59 incgam1 . . . . . . . . . . . . . . . . . . . 59 incgam2 . . . . . . . . . . . . . . . . . . . 59 incgam3 . . . . . . . . . . . . . . . . . . . 59 incgam4 . . . . . . . . . . . . . . . . . . . 59 incgamc . . . . . . . . . . . . . . . . . . . 59 inclusive or . . . . . . . . . . . . . . . . . 48 indefinite binary quadratic form . . . . . 162 indexrank . . . . . . . . . . . . . . . . . . 122 indexsort . . . . . . . . . . . . . . . . . . 128 infile . . . . . . . . . . . . . . . . . . . . . 167 infinite product . . . . . . . . . . . . . . . 130 infinite sum . . . . . . . . . . . . . . . . . 131 infinity . . . . . . . . . . . . . . . . . . . . 129 initell . . . . . . . . . . . . . . . . . . . . 76 initzeta . . . . . . . . . . . . . . . . . . . 113 input . . . . . . . . . . . . . . . . . . 141, 165 install . . . . . . . 38, 141, 165, 168, 175, 176 integ . . . . . . . . . . . . . . . . . . . . . 114 integer . . . . . . . . . . . . . . . . 7, 23, 159 integermod . . . . . . . . . . . . 7, 8, 23, 160 integral basis . . . . . . . . . . . . . . . . 97 internal longword format . . . . . . . . . 22 internal representation . . . . . . . . . . . 23 interpolating polynomial . . . . . . . . . . 115 intersect . . . . . . . . . . . . . . . . . . . 122 intformal . . . . . . . . . . . . . . . . . . 114 intnum . . . . . . . . . . . . . . . . . . . . 129 intnum0 . . . . . . . . . . . . . . . . . . . 129 inverseimage . . . . . . . . . . . . . . . . 122 isdiagonal . . . . . . . . . . . . . . . . . . 122 isexactzero . . . . . . . . . . . . . . . . . 48 isfundamental . . . . . . . . . . . . . . . . 67 isideal . . . . . . . . . . . . . . . . . . . . 102 isprime . . . . . . . . . . . . . . . . . . . 67 isprincipalall . . . . . . . . . . . . . . . . 86 isprincipalrayall . . . . . . . . . . . . . . . 90 ispseudoprime . . . . . . . . . . . . . . . . 67 ispsp . . . . . . . . . . . . . . . . . . . . . 67 issquare . . . . . . . . . . . . . . . . . . . 68 issquarefree . . . . . . . . . . . . . . . 62, 68 isunit . . . . . . . . . . . . . . . . . . . . 86 itos . . . . . . . . . . . . . . . . . . 152, 174

J

j . . . . . . . . . . . . . . . . . . . . . . . 73 jacobi . . . . . . . . . . . . . . . . . . . . 125 jbesselh . . . . . . . . . . . . . . . . . . . 58 jell . . . . . . . . . . . . . . . . . . . . . . 76

K

kbessel . . . . . . . . . . . . . . . . . . . . 58 kbessel2 . . . . . . . . . . . . . . . . . . . 58 ker . . . . . . . . . . . . . . . . . . . . . . 122 keri . . . . . . . . . . . . . . . . . . . . . 122 kerint . . . . . . . . . . . . . . . . . . . . 122 ker mod p . . . . . . . . . . . . . . . . . . 122 keyword . . . . . . . . . . . . . . . . . . . 36 kill . . . . . . . . . . . . . . . . . . . . . . 142 Kodaira . . . . . . . . . . . . . . . . . . . 76 Kronecker symbol . . . . . . . . . . . . . 68 kronecker . . . . . . . . . . . . . . . . . . 68

L

ladd(x,y) . . . . . . . . . . . . . . . . . . 147 laplace . . . . . . . . . . . . . . . . . . . . 117 lclone . . . . . . . . . . . . . . . . . . . . 152 lcm . . . . . . . . . . . . . . . . . . . . . . 68 lcopy . . . . . . . . . . . . . . . . . . . . . 151 leadingcoeff . . . . . . . . . . . . . . . . . 115 leaves . . . . . . . . . . . . . . . . . . . . 7, 8 Legendre polynomial . . . . . . . . . . . . 115 Legendre symbol . . . . . . . . . . . . . . 68 legendre . . . . . . . . . . . . . . . . . . . 115 length . . . . . . . . . . . . . . . . . . 52, 159 Lenstra . . . . . . . . . . . . . . . . . 66, 113 lex . . . . . . . . . . . . . . . . . . . . . . 48 lexcmp . . . . . . . . . . . . . . . . . . . . 48 lexsort . . . . . . . . . . . . . . . . . . . . 128 lg . . . . . . . . . . . . . . . . . . . . 159, 170 lgef . . . . . . . . . . . . . . . . . . . 161, 163 lgefint . . . . . . . . . . . . . . . . . . . . 160 lgetg . . . . . . . . . . . . . . . . . . . . . 150 lgeti . . . . . . . . . . . . . . . . . . . . . 150 lgetr . . . . . . . . . . . . . . . . . . . . . 150 library mode . . . . . . . . . . . . . . . . 145 LiDIA . . . . . . . . . . . . . . . . . . . . 66 lift . . . . . . . . . . . . . . . . . . . . . . 53 lift0 . . . . . . . . . . . . . . . . . . . . . 53 lindep . . . . . . . . . . . . . . . . . . . . 119 lindep0 . . . . . . . . . . . . . . . . . . . 119 line editor . . . . . . . . . . . . . . . . . . 41 linear dependence . . . . . . . . . . . . . 119 lines . . . . . . . . . . . . . . . . . . . . . 17 Linux . . . . . . . . . . . . . . . . . . 13, 176 lisexpr . . . . . . . . . . . . . . . . . . . . 165 lisGEN . . . . . . . . . . . . . . . . 165, 169 Lisp . . . . . . . . . . . . . . . . . . . . . 38 lisseq . . . . . . . . . . . . . . . . . . . . . 165 list . . . . . . . . . . . . . . . . . . . . . 7, 26 List . . . . . . . . . . . . . . . . . . . . . 49 list . . . . . . . . . . . . . . . . . . . . . . 163 listcreate . . . . . . . . . . . . . . . . . . 119 listinsert . . . . . . . . . . . . . . . . . . . 119 listkill . . . . . . . . . . . . . . . . . . . . 119 listput . . . . . . . . . . . . . . . . . . . . 119 listsort . . . . . . . . . . . . . . . . . . . . 120 LLL … 84, 95, 100, 118, 119, 121, 122, 125 lll . . . . . . . . . . . . . . . . . . . . . . 125 lllgram . . . . . . . . . . . . . . . . . . . . 126 lllgramint . . . . . . . . . . . . . . . . . . 126 lllgramkerim . . . . . . . . . . . . . . . . 126 lllint . . . . . . . . . . . . . . . . . . . . . 125 lllkerim . . . . . . . . . . . . . . . . . . . 125 lngamma . . . . . . . . . . . . . . . . . . 60 local . . . . . . . . . . . . . . . . . 31, 32, 35 localreduction . . . . . . . . . . . . . . . . 76 log . . . . . . . . . . . . . . . . . . 17, 60, 142 logfile . . . . . . . . . . . . . . . . . . 17, 142 LONG IS 64BIT . . . . . . . . . . . . . . 148 lpolx . . . . . . . . . . . . . . . . . . . . . 147 lseriesell . . . . . . . . . . . . . . . . . . . 77

M

mael . . . . . . . . . . . . . . . . . . . 52, 147 makebigbnf . . . . . . . . . . . . . . . . . 86 Mat . . . . . . . . . . . . . . . . . 26, 49, 118 matadjoint . . . . . . . . . . . . . . . . . 120 matalgtobasis . . . . . . . . . . . . . . . . 96 matbasistoalg . . . . . . . . . . . . . . . . 97 matbrute . . . . . . . . . . . . . . . . . . 166 matcompanion . . . . . . . . . . . . . . . 120 matdet . . . . . . . . . . . . . . . . . . . . 120 matdetint . . . . . . . . . . . . . . . . . . 120 matdiagonal . . . . . . . . . . . . . . . . . 120 mateigen . . . . . . . . . . . . . . . . . . 120 matextract . . . . . . . . . . . . . . . . . 128 mathell . . . . . . . . . . . . . . . . . . . 75 mathess . . . . . . . . . . . . . . . . . . . 120 mathilbert . . . . . . . . . . . . . . . . . . 120 mathnf . . . . . . . . . . . . . . . . 118, 121 mathnf0 . . . . . . . . . . . . . . . . . . . 121 mathnfmod . . . . . . . . . . . . . . . . . 121 mathnfmodid . . . . . . . . . . . . . . . . 121 matid . . . . . . . . . . . . . . . . . . . . 121 matimage . . . . . . . . . . . . . . . . . . 121 matimage0 . . . . . . . . . . . . . . . . . 121 matimagecompl . . . . . . . . . . . . . . . 122 matindexrank . . . . . . . . . . . . . . . . 122 matintersect . . . . . . . . . . . . . . . . . 122 matinverseimage . . . . . . . . . . . . . . 122 matisdiagonal . . . . . . . . . . . . . . . . 122 matker . . . . . . . . . . . . . . . . . . . . 122 matker0 . . . . . . . . . . . . . . . . . . . 122 matkerint . . . . . . . . . . . . . . . . . . 122 matkerint0 . . . . . . . . . . . . . . . . . 122 matmuldiagonal . . . . . . . . . . . . . . 123 matmultodiagonal . . . . . . . . . . . . . 123 matpascal . . . . . . . . . . . . . . . . . . 123 matqpascal . . . . . . . . . . . . . . . . . 123 matrank . . . . . . . . . . . . . . . . . . . 123 matrice . . . . . . . . . . . . . . . . . . . 130 matrix . . . . . . . . . . 7, 8, 26, 37, 130, 162 matrixqz . . . . . . . . . . . . . . . . . . . 123 matrixqz0 . . . . . . . . . . . . . . . . . . 123 matsize . . . . . . . . . . . . . . . . . . . 123 matsnf . . . . . . . . . . . . . . . . . . . . 124 matsnf0 . . . . . . . . . . . . . . . . . . . 124 matsolve . . . . . . . . . . . . . . . . . . . 124 matsolvemod . . . . . . . . . . . . . . . . 124 matsolvemod0 . . . . . . . . . . . . . . . 124 matsupplement . . . . . . . . . . . . . . . 124 mattranspose . . . . . . . . . . . . . . . . 124 max . . . . . . . . . . . . . . . . . . . . . 48 MAXVARN . . . . . . . . . . . . . . 146, 163 MEDDEFAULTPREC . . . . . . . . . . . 148 member functions . . . . . . . . . 35, 73, 81 min . . . . . . . . . . . . . . . . . . . . . 48 minideal . . . . . . . . . . . . . . . . . . . 94 minim . . . . . . . . . . . . . . . . . . . . 126 minim2 . . . . . . . . . . . . . . . . . . . 126 Mod . . . . . . . . . . . . . . . . . . . . . 49 Mod0 . . . . . . . . . . . . . . . . . . . . 49 mod2 . . . . . . . . . . . . . . . . . . . . 160 modpr . . . . . . . . . . . . . . . . . . . . 103 modreverse . . . . . . . . . . . . . . . . . 97 modulargcd . . . . . . . . . . . . . . . . . 67 module . . . . . . . . . . . . . . . . . . . 80 moebius . . . . . . . . . . . . . . . . . 62, 68 Mordell-Weil group . . . . . . . . . . . 75, 77 mpadd . . . . . . . . . . . . . . . . . . . . 147 mpbern . . . . . . . . . . . . . . . . . . . 199 mpeuler . . . . . . . . . . . . . . . . . . . 56 mpfact . . . . . . . . . . . . . . . . . . . . 66 mpfactr . . . . . . . . . . . . . . . . . . . 66 mpodd . . . . . . . . . . . . . . . . . . . . 160 mppi . . . . . . . . . . . . . . . . . . . . . 57 MPQS . . . . . . . . . . . . . . . . . . 62, 66 msgtimer . . . . . . . . . . . . . . . . . . 168 mu . . . . . . . . . . . . . . . . . . . . . . 68 multivariate polynomial . . . . . . . . . . 34

N

name var . . . . . . . . . . . . . . . . . . 164 newtonpoly . . . . . . . . . . . . . . . . . 97 next . . . . . . . . . . . . . . . . . . . . . 139 nextprime . . . . . . . . . . . . . . . . . . 68 nf . . . . . . . . . . . . . . . . . . 35, 79, 81 nfalgtobasis . . . . . . . . . . . . . . . . . 97 nfbasis . . . . . . . . . . . . . . . . . . 97, 102 nfbasis0 . . . . . . . . . . . . . . . . . . . 97 nfbasistoalg . . . . . . . . . . . . . . . . . 98 nfdetint . . . . . . . . . . . . . . . . . . . 98 nfdisc . . . . . . . . . . . . . . . . . . . . 98 nfdiscf0 . . . . . . . . . . . . . . . . . . . 98 nfdiveuc . . . . . . . . . . . . . . . . . . . 98 nfdivres . . . . . . . . . . . . . . . . . . . 98 nfeltdiv . . . . . . . . . . . . . . . . . . . 98 nfeltdiveuc . . . . . . . . . . . . . . . . . 98 nfeltdivmodpr . . . . . . . . . . . . . . . . 98 nfeltdivrem . . . . . . . . . . . . . . . . . 98 nfeltmod . . . . . . . . . . . . . . . . . . . 98 nfeltmul . . . . . . . . . . . . . . . . . . . 99 nfeltmulmodpr . . . . . . . . . . . . . . . 99 nfeltpow . . . . . . . . . . . . . . . . . . . 99 nfeltpowmodpr . . . . . . . . . . . . . . . 99 nfeltreduce . . . . . . . . . . . . . . . . . 99 nfeltreducemodpr . . . . . . . . . . . . . . 99 nfeltval . . . . . . . . . . . . . . . . . . . 99 nffactor . . . . . . . . . . . . . . . . . 99, 102 nffactormod . . . . . . . . . . . . . . . . . 100 nfgaloisapply . . . . . . . . . . . . . . . . 100 nfgaloisconj . . . . . . . . . . . . . . . . . 100 nfhermite . . . . . . . . . . . . . . . . . . 101 nfhermitemod . . . . . . . . . . . . . . . . 101 nfhilbert . . . . . . . . . . . . . . . . . . . 101 nfhnf . . . . . . . . . . . . . . . . . . . . . 101 nfhnfmod . . . . . . . . . . . . . . . . . . 101 nfinit . . . . . . . . . . . . . . . . 79, 101, 105 nfinit0 . . . . . . . . . . . . . . . . . . . . 102 nfisideal . . . . . . . . . . . . . . . . . . . 102 nfisincl . . . . . . . . . . . . . . . . . 102, 103 nfisisom . . . . . . . . . . . . . . . . . . . 103 nfkermodpr . . . . . . . . . . . . . . . . . 103 nfmod . . . . . . . . . . . . . . . . . . . . 98 nfmodprinit . . . . . . . . . . . . . 98, 99, 103 nfnewprec . . . . . . . . . . . . . . . 102, 103 nfreducemodpr2 . . . . . . . . . . . . . . 99 nfroots . . . . . . . . . . . . . . . . . . . . 103 nfrootsof1 . . . . . . . . . . . . . . . . . . 103 nfsmith . . . . . . . . . . . . . . . . . . . 104 nfsnf . . . . . . . . . . . . . . . . . . . . . 104 nfsolvemodpr . . . . . . . . . . . . . . . . 104 nfsubfields . . . . . . . . . . . . . . . . . . 103 no . . . . . . . . . . . . . . . . . . . . . . 81 norm . . . . . . . . . . . . . . . . . . . . . 53 norml2 . . . . . . . . . . . . . . . . . . . . 53 not . . . . . . . . . . . . . . . . . . . . . . 48 nucomp . . . . . . . . . . . . . . . . . . . 69 nudupl . . . . . . . . . . . . . . . . . . . . 69 numbdiv . . . . . . . . . . . . . . . . . . . 68 number field . . . . . . . . . . . . . . . . 24 numdiv . . . . . . . . . . . . . . . . . . . 68 numer . . . . . . . . . . . . . . . . . . . . 53 numerator . . . . . . . . . . . . . . . . . . 53 numerical integration . . . . . . . . . . . 129 numtoperm . . . . . . . . . . . . . . . . . 53 nupow . . . . . . . . . . . . . . . . . . . . 69 Néron-Tate height . . . . . . . . . . . . . 75

O

O . . . . . . . . . . . . . . . . . . . . . 29, 113 omega . . . . . . . . . . . . . . . . 68, 71, 73 oncurve . . . . . . . . . . . . . . . . . . . 76 operator . . . . . . . . . . . . . . . . . . . 26 or . . . . . . . . . . . . . . . . . . . . . . 48 ordell . . . . . . . . . . . . . . . . . . . . 77 order . . . . . . . . . . . . . . . . . . . . . 72 orderell . . . . . . . . . . . . . . . . . . . 77 ordred . . . . . . . . . . . . . . . . . . . . 105 outbeaut . . . . . . . . . . . . . . . . . . 166 outbrute . . . . . . . . . . . . . . . . . . . 166 outfile . . . . . . . . . . . . . . . . . . . . 167 outmat . . . . . . . . . . . . . . . . . . . 166 output formats . . . . . . . . . . . . . . . 14 output . . . . . . . . . . . . . . . 18, 166, 168

P

p-adic number . . . . . . . . . . 7, 8, 24, 161 padicappr . . . . . . . . . . . . . . . . . . 114 padicprec . . . . . . . . . . . . . . . . . . 53 parametric plot . . . . . . . . . . . . . . . 134 pari.h . . . . . . . . . . . . . . . . . . . . 145 pariErr . . . . . . . . . . . . . . . . . . . 167 pariK1 . . . . . . . . . . . . . . . . . . . . 169 pariOut . . . . . . . . . . . . . . . . . . . 167 PariPerl . . . . . . . . . . . . . . . . . . . 38 pariputs . . . . . . . . . . . . . . . . . . . 175 PariPython . . . . . . . . . . . . . . . . . 38 parisize . . . . . . . . . . . . . . . . . . . 18 pari init . . . . . . . . . . . . . . . . . . . 145 parser code . . . . . . . . . . . . . . 173, 175 pascal triangle . . . . . . . . . . . . . . . 123 path . . . . . . . . . . . . . . . . . . . . . 18 perf . . . . . . . . . . . . . . . . . . . . . 126 Perl . . . . . . . . . . . . . . . . . . . . . 38 permtonum . . . . . . . . . . . . . . . . . 53 permute . . . . . . . . . . . . . . . . . . . 53 permuteInv . . . . . . . . . . . . . . . . . 53 phi . . . . . . . . . . . . . . . . . . . . . . 65 Pi . . . . . . . . . . . . . . . . . . . . 29, 57 plot . . . . . . . . . . . . . . . . . . . . . 133 plotbox . . . . . . . . . . . . . . . . . . . 133 plotclip . . . . . . . . . . . . . . . . . . . 133 plotcolor . . . . . . . . . . . . . . . . . . . 134 plotcopy . . . . . . . . . . . . . . . . 133, 134 plotcursor . . . . . . . . . . . . . . . . . . 134 plotdraw . . . . . . . . . . . . . . . . . . . 134 plotfile . . . . . . . . . . . . . . . . . . . . 134 ploth . . . . . . . . . . . . . . . . . . . . . 134 plothraw . . . . . . . . . . . . . . . . . . . 135 plothsizes . . . . . . . . . . . . . . . . . . 135 plotinit . . . . . . . . . . . . . . . . . . . 135 plotkill . . . . . . . . . . . . . . . . . . . . 135 plotlines . . . . . . . . . . . . . . . . . . . 136 plotlinetype . . . . . . . . . . . . . . . . . 136 plotmove . . . . . . . . . . . . . . . . . . 136 plotpoints . . . . . . . . . . . . . . . . . . 136 plotpointsize . . . . . . . . . . . . . . . . 136 plotpointtype . . . . . . . . . . . . . . . . 136 plotrbox . . . . . . . . . . . . . . . . . . . 136 plotrecth . . . . . . . . . . . . . . . . . . 136 plotrecthraw . . . . . . . . . . . . . . . . 136 plotrline . . . . . . . . . . . . . . . . . . . 136 plotrmove . . . . . . . . . . . . . . . . . . 137 plotrpoint . . . . . . . . . . . . . . . . . . 137 plotscale . . . . . . . . . . . . . . . . . . . 137 plotstring . . . . . . . . . . . . . . . . 38, 137 plotterm . . . . . . . . . . . . . . . . . 38, 137 pnqn . . . . . . . . . . . . . . . . . . . . . 64 pointch . . . . . . . . . . . . . . . . . . . 74 pointell . . . . . . . . . . . . . . . . . . . 79 pointer . . . . . . . . . . . . . . . . . . . . 46 Pol . . . . . . . . . . . . . . . . . . . . . . 49 polcoeff . . . . . . . . . . . . . . . . . 51, 114 polcoeff0 . . . . . . . . . . . . . . . . . . . 114 polcompositum . . . . . . . . . . . . . . . 104 polcompositum0 . . . . . . . . . . . . . . 104 polcyclo . . . . . . . . . . . . . . . . . . . 114 poldegree . . . . . . . . . . . . . . . . . . 114 poldisc . . . . . . . . . . . . . . . . . . . . 114 poldisc0 . . . . . . . . . . . . . . . . . . . 114 poldiscreduced . . . . . . . . . . . . . . . 114 poleval . . . . . . . . . . . . . . . . . . . . 113 polfnf . . . . . . . . . . . . . . . . . . . . 92 polgalois . . . . . . . . . . . . . . . . . . . 104 polint . . . . . . . . . . . . . . . . . . . . 115 polinterpolate . . . . . . . . . . . . . . . . 115 polisirreducible . . . . . . . . . . . . . . . 115 Pollard Rho . . . . . . . . . . . . . . . 62, 66 pollead . . . . . . . . . . . . . . . . . . . 115 pollegendre . . . . . . . . . . . . . . . . . 115 polmod . . . . . . . . . . . . . . 7, 8, 24, 161 polmodrecip . . . . . . . . . . . . . . . . . 97 polrecip . . . . . . . . . . . . . . . . . . . 115 polred . . . . . . . . . . . . . . . . . . . . 105 polred0 . . . . . . . . . . . . . . . . . . . 105 polredabs . . . . . . . . . . . . . . . . . . 105 polredabs0 . . . . . . . . . . . . . . . . . 105 polredord . . . . . . . . . . . . . . . . . . 105 polresultant . . . . . . . . . . . . . . . . . 115 polresultant0 . . . . . . . . . . . . . . . . 115 Polrev . . . . . . . . . . . . . . . . . . . . 49 polroots . . . . . . . . . . . . . . . . . . . 115 polrootsmod . . . . . . . . . . . . . . . . 116 polrootspadic . . . . . . . . . . . . . . . . 116 polsturm . . . . . . . . . . . . . . . . . . 116 polsubcyclo . . . . . . . . . . . . . . . . . 116 polsylvestermatrix . . . . . . . . . . . . . 116 polsym . . . . . . . . . . . . . . . . . . . . 116 poltchebi . . . . . . . . . . . . . . . . . . 116 poltschirnhaus . . . . . . . . . . . . . . . 105 polun . . . . . . . . . . . . . . . . . 146, 163 polvar . . . . . . . . . . . . . . . . . . . . 163 polx . . . . . . . . . . . . . . . . . . 146, 163 polylog . . . . . . . . . . . . . . . . . . . 60 polylog0 . . . . . . . . . . . . . . . . . . . 60 polynomial . . . . . . . . . . . . 7, 8, 25, 161 polzag . . . . . . . . . . . . . . . . . . . . 116 polzagier . . . . . . . . . . . . . . . . . . 116 polzagreel . . . . . . . . . . . . . . . . . . 116 PostScript . . . . . . . . . . . . . . . . . . 133 powell . . . . . . . . . . . . . . . . . . . . 77 power series . . . . . . . . . . . . 7, 8, 25, 162 powering . . . . . . . . . . . . . . . . 47, 56 powraw . . . . . . . . . . . . . . . . . . . 69 precdl . . . . . . . . . . . . . . . . . . . . 56 precision . . . . . . . . . . . . . . . . . 53, 55 precision0 . . . . . . . . . . . . . . . . . . 53 precp . . . . . . . . . . . . . . . . . . . . 161 precprime . . . . . . . . . . . . . . . . . . 68 preferences file . . . . . . . . . 13, 14, 39, 175 prettymatrix format . . . . . . . . . . . . 18 prettyprint format . . . . . . . . . . . . . 18 prime . . . . . . . . . . . . . . . . . . . . 68 primeform . . . . . . . . . . . . . . . . . . 70 primelimit . . . . . . . . . . . . . . . . . . 18 primes . . . . . . . . . . . . . . . . . . . . 69 principalideal . . . . . . . . . . . . . . . . 95 principalidele . . . . . . . . . . . . . . . . 96 print . . . . . . . . . . . . . . . . . 36, 38, 142 print1 . . . . . . . . . . . . . . . . . . . . 142 printf . . . . . . . . . . . . . . . . . . . . 167 printp . . . . . . . . . . . . . . . . . . . . 142 printp1 . . . . . . . . . . . . . . . . . . . 142 printtex . . . . . . . . . . . . . . . . . . . 142 priority . . . . . . . . . . . . . . . . . . . 27 prod . . . . . . . . . . . . . . . . . . . . . 130 prodeuler . . . . . . . . . . . . . . . . . . 130 prodinf . . . . . . . . . . . . . . . . . . . 130 prodinf1 . . . . . . . . . . . . . . . . . . . 130 product . . . . . . . . . . . . . . . . . . . 46 produit . . . . . . . . . . . . . . . . . . . 130 programming . . . . . . . . . . . . . . 30, 137 prompt . . . . . . . . . . . . . . . . . . . 18 psdraw . . . . . . . . . . . . . . . . . . . . 137 pseudo-basis . . . . . . . . . . . . . . . . 80 pseudo-matrix . . . . . . . . . . . . . . . 80 psfile . . . . . . . . . . . . . . . . . . . 19, 133 psi . . . . . . . . . . . . . . . . . . . . . . 61 psploth . . . . . . . . . . . . . . . . . . . 137 psplothraw . . . . . . . . . . . . . . . . . 137 Python . . . . . . . . . . . . . . . . . . . 38

Q

Qfb . . . . . . . . . . . . . . . . . . . . . 62 Qfb0 . . . . . . . . . . . . . . . . . . . . . 62 qfbclassno . . . . . . . . . . . . . . . . . . 69 qfbclassno0 . . . . . . . . . . . . . . . . . 69 qfbcompraw . . . . . . . . . . . . . . . . . 69 qfbhclassno . . . . . . . . . . . . . . . . . 69 qfbnucomp . . . . . . . . . . . . . . . . . 69 qfbnupow . . . . . . . . . . . . . . . . . . 69 qfbpowraw . . . . . . . . . . . . . . . . . 69 qfbprimeform . . . . . . . . . . . . . . . . 70 qfbred . . . . . . . . . . . . . . . . . . . . 70 qfbred0 . . . . . . . . . . . . . . . . . . . 70 qfeval . . . . . . . . . . . . . . . . . . . . 113 qfgaussred . . . . . . . . . . . . . . . . . . 124 qfi . . . . . . . . . . . . . . . . . . . . . . 62 qfjacobi . . . . . . . . . . . . . . . . . . . 125 qflll . . . . . . . . . . . . . . . . . . 118, 125 qflll0 . . . . . . . . . . . . . . . . . . . . . 125 qflllgram . . . . . . . . . . . . . . . . . . . 126 qflllgram0 . . . . . . . . . . . . . . . . . . 126 qfminim . . . . . . . . . . . . . . . . . . . 126 qfperfection . . . . . . . . . . . . . . . . . 126 qfr . . . . . . . . . . . . . . . . . . . . . . 62 qfsign . . . . . . . . . . . . . . . . . . . . 126 quadclassunit . . . . . . . . . . . . . . . . 70 quadclassunit0 . . . . . . . . . . . . . . . 71 quaddisc . . . . . . . . . . . . . . . . . . . 71 quadgen . . . . . . . . . . . . . . . . . 24, 71 quadhilbert . . . . . . . . . . . . . . . . . 71 quadpoly . . . . . . . . . . . . . . . . . . 71 quadpoly0 . . . . . . . . . . . . . . . . . . 71 quadratic number . . . . . . . . 7, 8, 24, 161 quadray . . . . . . . . . . . . . . . . . . . 71 quadregulator . . . . . . . . . . . . . . . . 72 quadunit . . . . . . . . . . . . . . . . . . 72 quit . . . . . . . . . . . . . . . . . . . 22, 142 quote . . . . . . . . . . . . . . . . . . . . 142 quotient . . . . . . . . . . . . . . . . . . . 46

R

Rabin-Miller . . . . . . . . . . . . . . . . 62 racine . . . . . . . . . . . . . . . . . . . . 72 random . . . . . . . . . . . . . . . . . . . 54 rank . . . . . . . . . . . . . . . . . . . . . 123 rational function . . . . . . . . . . 7, 25, 162 rational number . . . . . . . . . 7, 8, 23, 161 raw format . . . . . . . . . . . . . . . . . 18 rayclassno . . . . . . . . . . . . . . . . . . 89 rayclassnolist . . . . . . . . . . . . . . . . 89 read . . . . . . . . . . . . . . . . . 22, 38, 143 readline . . . . . . . . . . . . . . . . . . . 41 real number . . . . . . . . . . . . . 7, 23, 160 real . . . . . . . . . . . . . . . . . . . . . 54 realprecision . . . . . . . . . . . . . . . . . 19 realzero . . . . . . . . . . . . . . . . . . . 171 recip . . . . . . . . . . . . . . . . . . . . . 117 recursive plot . . . . . . . . . . . . . . . . 135 recursiveness . . . . . . . . . . . . . . . . . 6 redimag . . . . . . . . . . . . . . . . . . . 70 redreal . . . . . . . . . . . . . . . . . . . . 70 redrealnod . . . . . . . . . . . . . . . . . . 70 reduceddiscsmith . . . . . . . . . . . . . . 114 reduction . . . . . . . . . . . . . . . . 69, 70 reference card . . . . . . . . . . . . . . . . 21 reg . . . . . . . . . . . . . . . . . . . . . . 81 regula . . . . . . . . . . . . . . . . . . . . 72 regulator . . . . . . . . . . . . . . . . . . 87 relative matrix . . . . . . . . . . . . . . . 80 removeprimes . . . . . . . . . . . . . . . . 72 reorder . . . . . . . . . . . . . . . 31, 51, 143 resultant2 . . . . . . . . . . . . . . . . . . 115 return . . . . . . . . . . . . . . . . . . . . 139 rhoreal . . . . . . . . . . . . . . . . . . . . 70 rhorealnod . . . . . . . . . . . . . . . . . 70 Riemann zeta-function . . . . . . . . . 33, 62 rnf . . . . . . . . . . . . . . . . . . . . . . 79 rnfalgtobasis . . . . . . . . . . . . . . . . 106 rnfbasis . . . . . . . . . . . . . . . . . . . 106 rnfbasistoalg . . . . . . . . . . . . . . . . 106 rnfcharpoly . . . . . . . . . . . . . . . . . 106 rnfconductor . . . . . . . . . . . . . . . . 106 rnfdedekind . . . . . . . . . . . . . . . . . 106 rnfdet . . . . . . . . . . . . . . . . . . . . 106 rnfdisc . . . . . . . . . . . . . . . . . . . . 107 rnfdiscf . . . . . . . . . . . . . . . . . . . 107 rnfelementabstorel . . . . . . . . . . . . . 107 rnfelementdown . . . . . . . . . . . . . . . 107 rnfelementreltoabs . . . . . . . . . . . . . 107 rnfelementup . . . . . . . . . . . . . . . . 107 rnfeltabstorel . . . . . . . . . . . . . . . . 107 rnfeltdown . . . . . . . . . . . . . . . . . . 107 rnfeltreltoabs . . . . . . . . . . . . . . . . 107 rnfeltup . . . . . . . . . . . . . . . . . . . 107 rnfequation . . . . . . . . . . . . . . . . . 107 rnfequation0 . . . . . . . . . . . . . . . . 108 rnfhermitebasis . . . . . . . . . . . . . . . 108 rnfhnfbasis . . . . . . . . . . . . . . . . . 108 rnfidealabstorel . . . . . . . . . . . . . . . 108 rnfidealdown . . . . . . . . . . . . . . . . 108 rnfidealhermite . . . . . . . . . . . . . . . 108 rnfidealhnf . . . . . . . . . . . . . . . . . 108 rnfidealmul . . . . . . . . . . . . . . . . . 108 rnfidealnormabs . . . . . . . . . . . . . . . 108 rnfidealnormrel . . . . . . . . . . . . . . . 108 rnfidealreltoabs . . . . . . . . . . . . . . . 109 rnfidealtwoelement . . . . . . . . . . . . . 109 rnfidealtwoelt . . . . . . . . . . . . . . . . 109 rnfidealup . . . . . . . . . . . . . . . . . . 109 rnfinit . . . . . . . . . . . . . . . . . . . . 109 rnfinitalg . . . . . . . . . . . . . . . . . . 110 rnfisfree . . . . . . . . . . . . . . . . . . . 110 rnfisnorm . . . . . . . . . . . . . . . 110, 111 rnfkummer . . . . . . . . . . . . . . . . . 111 rnflllgram . . . . . . . . . . . . . . . . . . 111 rnfnormgroup . . . . . . . . . . . . . . . . 111 rnfpolred . . . . . . . . . . . . . . . . . . 111 rnfpolredabs . . . . . . . . . . . . . . 111, 112 rnfpseudobasis . . . . . . . . . . . . . . . 112 rnfsteinitz . . . . . . . . . . . . . . . . . . 112 rootmod . . . . . . . . . . . . . . . . . . . 116 rootmod2 . . . . . . . . . . . . . . . . . . 116 rootpadic . . . . . . . . . . . . . . . . . . 116 roots . . . . . . . . . . . . . . . . . 73, 81, 115 rootsof1 . . . . . . . . . . . . . . . . . . . 103 rootsold . . . . . . . . . . . . . . . . . . . 115 round 2 . . . . . . . . . . . . . . . . . . . 97 round 4 . . . . . . . . . . . . . . . 97, 98, 113 round . . . . . . . . . . . . . . . . . . . . 54 row vector . . . . . . . . . . . . . 7, 25, 162 rtodbl . . . . . . . . . . . . . . . . . . . . 152

S

scalar product . . . . . . . . . . . . . . . 46 scalar type . . . . . . . . . . . . . . . . . . 8 Schönhage . . . . . . . . . . . . . . . . . . 115 scientific format . . . . . . . . . . . . . . 17 secure . . . . . . . . . . . . . . . . . . . . 19 Ser . . . . . . . . . . . . . . . . . . . . . . 50 serconvol . . . . . . . . . . . . . . . . . . 116 seriesprecision . . . . . . . . . . . . . . . . 19 serlaplace . . . . . . . . . . . . . . . . . . 117 serreverse . . . . . . . . . . . . . . . . . . 117 Set . . . . . . . . . . . . . . . . . . . . . . 50 setexpo . . . . . . . . . . . . . . . . 160, 162 setintersect . . . . . . . . . . . . . . . . . 127 setisset . . . . . . . . . . . . . . . . . . . 127 setlg . . . . . . . . . . . . . . . . . . 159, 170 setlgef . . . . . . . . . . . . . . . . . 161, 163 setlgefint . . . . . . . . . . . . . . . . . . 160 setminus . . . . . . . . . . . . . . . . . . . 127 setprecp . . . . . . . . . . . . . . . . . . . 161 setrand . . . . . . . . . . . . . . . . . . . 143 setsearch . . . . . . . . . . . . . . . . . . 127 setsigne . . . . . . . . . . . . . 160, 161, 162 settyp . . . . . . . . . . . . . . . . . . . . 159 setunion . . . . . . . . . . . . . . . . . . . 127 setvalp . . . . . . . . . . . . . . . . . 161, 162 setvarn . . . . . . . . . . . . . . . . . . . 162 Shanks . . . . . . . . . . . . . . . . 62, 69, 70 shift . . . . . . . . . . . . . . . . . . . . . 47 shiftmul . . . . . . . . . . . . . . . . . . . 47 sigma . . . . . . . . . . . . . . . . . . . . 72 sign . . . . . . . . . . . . . . . . . . . 48, 81 signat . . . . . . . . . . . . . . . . . . . . 126 signe . . . . . . . . . . . . . . . 159, 161, 162 signunits . . . . . . . . . . . . . . . . . . . 87 simplefactmod . . . . . . . . . . . . . . . 67 simplify . . . . . . . . . . . . . . . . . 19, 54 sin . . . . . . . . . . . . . . . . . . . . . . 61 sindexlexsort . . . . . . . . . . . . . . . . 128 sindexsort . . . . . . . . . . . . . . . . . . 128 sinh . . . . . . . . . . . . . . . . . . . . . 61 sizebyte . . . . . . . . . . . . . . . . . . . 54 sizedigit . . . . . . . . . . . . . . . . . . . 54 smallfact . . . . . . . . . . . . . . . . . . 66 smallinitell . . . . . . . . . . . . . . . . . 76 Smith normal form …. 81, 84, 87, 96, 104,

112, 124, 138

smith . . . . . . . . . . . . . . . . . . . . 124 solve . . . . . . . . . . . . . . . . . . . . . 130 somme . . . . . . . . . . . . . . . . . . . . 131 sor . . . . . . . . . . . . . . . . . . . . . . 166 sort . . . . . . . . . . . . . . . . . . . . . 128 sqr . . . . . . . . . . . . . . . . . . . . . . 61 sqred . . . . . . . . . . . . . . . . . . . . . 124 sqrt . . . . . . . . . . . . . . . . . . . . . 61 sqrtint . . . . . . . . . . . . . . . . . . . . 72 srgcd . . . . . . . . . . . . . . . . . . . . . 67 stack . . . . . . . . . . . . . . . . 22, 145, 199 stack lim . . . . . . . . . . . . . . . . . . 158 Stark units . . . . . . . . . . . . . . . 71, 91 startup . . . . . . . . . . . . . . . . . . . 39 stderr . . . . . . . . . . . . . . . . . . . . 167 stdout . . . . . . . . . . . . . . . . . . . . 167 stoi . . . . . . . . . . . . . . . . . . . . . 152 Str . . . . . . . . . . . . . . . . . . 36, 38, 50 strftime . . . . . . . . . . . . . . . . . 15, 18 strictmatch . . . . . . . . . . . . . . . . . 19 string context . . . . . . . . . . . . . . . . 36 string . . . . . . . . . . . . . . . . . 7, 26, 36 strtoGEN . . . . . . . . . . . . . . . . . . 50 strtoGENstr . . . . . . . . . . . . . . . . . 50 sturm . . . . . . . . . . . . . . . . . . . . 116 sturmpart . . . . . . . . . . . . . . . . . . 116 subcyclo . . . . . . . . . . . . . . . . . . . 116 subell . . . . . . . . . . . . . . . . . . . . 78 subfields . . . . . . . . . . . . . . . . . . . 103 subgroup . . . . . . . . . . . . . . . . . . 138 subgrouplist . . . . . . . . . . . . . . 112, 138 subgrouplist0 . . . . . . . . . . . . . . . . 112 subres . . . . . . . . . . . . . . . . . . . . 115 subresext . . . . . . . . . . . . . . . . . . 63 subresultant algorithm . . . . . . 67, 114, 115 subst . . . . . . . . . . . . . . . . . . . . . 117 sum . . . . . . . . . . . . . . . . 46, 129, 131 sumalt . . . . . . . . . . . . . . . . . . . . 131 sumdiv . . . . . . . . . . . . . . . . . 72, 131 suminf . . . . . . . . . . . . . . . . . . . . 131 sumpos . . . . . . . . . . . . . . . . . . . 132 suppl . . . . . . . . . . . . . . . . . . . . 124 switchin . . . . . . . . . . . . . . . . . . . 167 switchout . . . . . . . . . . . . . . . . . . 167 sylvestermatrix . . . . . . . . . . . . . . . 116 symmetric powers . . . . . . . . . . . . . 116 system . . . . . . . . . . . . . . . . 19, 38, 143

T

t2 . . . . . . . . . . . . . . . . . . . . . . 81 taille . . . . . . . . . . . . . . . . . . . . . 54 taille2 . . . . . . . . . . . . . . . . . . . . 54 talker . . . . . . . . . . . . . . . . . . . . 167 Tamagawa number . . . . . . . . . . . 75, 76 tan . . . . . . . . . . . . . . . . . . . . . . 61 tanh . . . . . . . . . . . . . . . . . . . . . 61 taniyama . . . . . . . . . . . . . . . . . . 78 Taniyama-Weil . . . . . . . . . . . . . . . 74 Tate . . . . . . . . . . . . . . . . . . . . . 73 tate . . . . . . . . . . . . . . . . . . . . . 73 tayl . . . . . . . . . . . . . . . . . . . . . 117 Taylor series . . . . . . . . . . . . . . . . 46 taylor . . . . . . . . . . . . . . . . . . . . 117 tchebi . . . . . . . . . . . . . . . . . . . . 116 teich . . . . . . . . . . . . . . . . . . . . . 61 teichmuller . . . . . . . . . . . . . . . . . 61 texe . . . . . . . . . . . . . . . . . . . . . 166 texprint . . . . . . . . . . . . . . . . . . . 166 theta . . . . . . . . . . . . . . . . . . . . . 61 thetanullk . . . . . . . . . . . . . . . . . . 61 thue . . . . . . . . . . . . . . . . . . . . . 117 thueinit . . . . . . . . . . . . . . . . . . . 117 time expansion . . . . . . . . . . . . . . . 15 timer . . . . . . . . . . . . . . . . . . . 19, 168 timer2 . . . . . . . . . . . . . . . . . . . . 168 trace . . . . . . . . . . . . . . . . . . . . . 127 truecoeff . . . . . . . . . . . . . . . . . 51, 114 truncate . . . . . . . . . . . . . . . . . . . 55 tschirnhaus . . . . . . . . . . . . . . . . . 105 tu . . . . . . . . . . . . . . . . . . . . . . 81 tufu . . . . . . . . . . . . . . . . . . . . . 81 tutorial . . . . . . . . . . . . . . . . . . . 21 typ . . . . . . . . . . . . . . . . . . . . . . 159 type number . . . . . . . . . . . . . . . . 159 type . . . . . . . . . . . . . . . . . . . 38, 143 typecast . . . . . . . . . . . . . . . . . . . 146 types . . . . . . . . . . . . . . . . . . . . . . 6 t COL . . . . . . . . . . . . . . . . . . . 7, 25 t COMPLEX . . . . . . . . . . . . . . . 7, 23 t FRAC . . . . . . . . . . . . . . . . . . 7, 23 t FRACN . . . . . . . . . . . . . . . . . 7, 23 t INT . . . . . . . . . . . . . . . . . . . 7, 23 t INTMOD . . . . . . . . . . . . . . . . 7, 23 t LIST . . . . . . . . . . . . . . . . . . . 7, 26 t MAT . . . . . . . . . . . . . . . . . . . 7, 26 t PADIC . . . . . . . . . . . . . . . . . 7, 24 t POL . . . . . . . . . . . . . . . . . . . 7, 25 t POLMOD . . . . . . . . . . . . . . . . 7, 24 t QFI . . . . . . . . . . . . . . . . . . . 7, 25 t QFR . . . . . . . . . . . . . . . . . . . 7, 25 t QUAD . . . . . . . . . . . . . . . . . . 7, 24 t REAL . . . . . . . . . . . . . . . . . . 7, 23 t RFRAC . . . . . . . . . . . . . . . . . 7, 25 t RFRACN . . . . . . . . . . . . . . . . 7, 25 t SER . . . . . . . . . . . . . . . . . . . 7, 25 t STR . . . . . . . . . . . . . . . . . . . 7, 26 t VEC . . . . . . . . . . . . . . . . . . . 7, 25

U

ulong . . . . . . . . . . . . . . . . . . . . 179 un . . . . . . . . . . . . . . . . . . . . . . 147 universal object . . . . . . . . . . . . . . 199 until . . . . . . . . . . . . . . . . . . . . . 139 user defined functions . . . . . . . . . . . 32

V

valp . . . . . . . . . . . . . . . . . . 161, 162 valuation . . . . . . . . . . . . . . . . . . 55 varentries . . . . . . . . . . . . . . . . . . 163 variable (special) . . . . . . . . . . . . . . 29 variable (temporary) . . . . . . . . . . . . 164 variable (user) . . . . . . . . . . . . 163, 164 variable number . . . . . . . . 162, 163, 174 variable . . . . . . . . 24, 25, 27, 30, 55, 146 varn . . . . . . . . . . . . . . . . . . . . . 162 Vec . . . . . . . . . . . . . . . . . . . . 25, 50 vecbezout . . . . . . . . . . . . . . . . . . 63 vecbezoutres . . . . . . . . . . . . . . . . 63 veceint1 . . . . . . . . . . . . . . . . . . . 59 vecextract . . . . . . . . . . . . . . . 122, 127 vecmax . . . . . . . . . . . . . . . . . . . 48 vecmin . . . . . . . . . . . . . . . . . . . . 48 vecsort . . . . . . . . . . . . . . . . . . . . 128 vecsort0 . . . . . . . . . . . . . . . . . . . 128 vecteur . . . . . . . . . . . . . . . . . . . 132 vector . . . . . . . . . . . . . . . . 8, 132, 162 vectorv . . . . . . . . . . . . . . . . . . . 132 version number . . . . . . . . . . . . . . . 22 Vi . . . . . . . . . . . . . . . . . . . . . . 42 voir . . . . . . . . . . . . . . . . . . . . . 167 vvecteur . . . . . . . . . . . . . . . . . . . 132

W

w . . . . . . . . . . . . . . . . . . . . . . . 73 weber . . . . . . . . . . . . . . . . . . . . 62 weber0 . . . . . . . . . . . . . . . . . . . . 62 Weierstrass ℘-function . . . . . . . . . . . 79 Weierstrass equation . . . . . . . . . . . . 73 Weil curve . . . . . . . . . . . . . . . . . 78 weipell . . . . . . . . . . . . . . . . . . . . 78 wf . . . . . . . . . . . . . . . . . . . . . . 62 wf1 . . . . . . . . . . . . . . . . . . . . . . 62 wf2 . . . . . . . . . . . . . . . . . . . . . . 62 whatnow . . . . . . . . . . . . . . . . 38, 143 while . . . . . . . . . . . . . . . . . . . . . 139 write . . . . . . . . . . . . . . . . . 23, 38, 143 write1 . . . . . . . . . . . . . . . . . . . . 143 writetex . . . . . . . . . . . . . . . . . . . 143

X

x[,n] . . . . . . . . . . . . . . . . . . . . . 52 x[m,n] . . . . . . . . . . . . . . . . . . . . 52 x[m,] . . . . . . . . . . . . . . . . . . . . . 52 x[n] . . . . . . . . . . . . . . . . . . . . . 51

Z Zassenhaus . . . . . . . . . . . . . . . 66, 113 zbrent . . . . . . . . . . . . . . . . . . . . 130 zell . . . . . . . . . . . . . . . . . . . . . . 77 zero . . . . . . . . . . . . . . . . . . . 9, 147 zeta function . . . . . . . . . . . . . . . . 33 zeta . . . . . . . . . . . . . . . . . . . . . 62 zetak . . . . . . . . . . . . . . . . . . . . . 112 zetakinit . . . . . . . . . . . . . . . . . . . 113 zideallog . . . . . . . . . . . . . . . . . . . 94 zk . . . . . . . . . . . . . . . . . . . . . . 81 zkst . . . . . . . . . . . . . . . . . . . . . 81 znlog . . . . . . . . . . . . . . . . . . . . . 72 znorder . . . . . . . . . . . . . . . . . . . 72 znprimroot . . . . . . . . . . . . . . . . . 72 znstar . . . . . . . . . . . . . . . . . . . . 72